A Bibliography of Books and Other Publications about the Ada Programming Language and Its

History

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

12 November 2025 Version 2.12

Title word cross-reference

+ [Tex82]. $10.95 [Wim83a]. $1000M [Ano84b]. 129 [Ano93a]. $29.95 [Por01]. 3 [EW91, HL93]. $32.95 [Ano98]. 653 [CH97]. $7.90 [You82b]. $75.00 [Wol08]. ^{1} [TS85]. ^{3} [Ano89d]. _{Ada} [Bro80]. A^{3} [Alb05]. : = [Tex82]. N [RSC93].

* [Bie85a].

-2 [Dub85]. -3 [Dub85]. -D [HL93]. -Dimensional [EW91]. -point [RSC93].

0 [Ano83b, Ano83c, Ano84d, Aug95, Her85, Mee92, Mer84, Ped88, Per89, Tug84]. 0-07-011589-3 [Her87]. 0-13-004078-9 [Ped88]. 0-13-030834-X [Aug95]. 0-13-729756-4 [Mer84]. 0-13-730010-7 [Ano83c, Tug84]. 0-13-816174-7 [Pay93]. 0-262-19242-X [Sec88]. 0-402-45630-3 [Mee92]. 0-521-30033-9 [Mea87]. 0-521-65939-6 [Por01]. 0-85312-416-7 [Ano83b]. 0-89433-239-2 [Ano84d]. 0-946536-15-5 [Her85].

1 [Bel91, Dub85, II94, Per89, TDBP01, TDB^{+}06]. 1/0 [Per89]. 10 [Ano87d]. 10.75 [Wal83]. 100 [Ano93e]. 1003.1q [EGC02]. 1003.5 [IEE99a]. 1003.5-1992 [IEE92a]. 1003.5-1999 [IEE99a]. 1003.5b [IEE96, IEE99b]. 1003.5b-1996 [IEE96]. 1003.5b-1999 [IEE99b]. 10th [VW05]. 11 [SHLR80]. 11/40 [GBO87]. 11/780 [SHLR80]. 112 [Ano93b]. 11th [ACM94a, Ano94, vK92, PH06]. 12.50 [Wic84a]. 121 [Ano90a]. 121-131 [Ano90a]. 12th [AK07, Gau93b]. 131 [Ano90a]. 13719-3 [ISO98a]. 13th [KV08]. 14.20 [Wal84a]. 14519 [IEE99b]. 14th [KK09]. 15 [Wic84b]. 15.95 [Bud88]. 152 [Ano82f]. 154 [All84]. 15th [RV10]. 16.10 [Alb85]. 16.50 [You82b]. 161 [Ano87u]. 162 [ECM97]. 163 [Ano86h]. 16th [RV11]. 17 [Ano86f]. 1750A [Ano86a]. 17th [BP12]. 18.95 [Ano85c]. 1815A [UA83a, UA83c, Ada83, Ame95a, Uni83]. 1815A-1983 [Ame95a]. 1947 [Ano48]. 1980 [Kat82, Whi81b]. 1981 [Ano82a]. 1982 [ACM82]. 1983 [Ame95a, Ano87s, UA83a, UA83c]. 1984 [ACM84, NB84]. 1984. [Nie86]. 1985 [BF85]. 1986 [Ano88c]. 1987 [Bar87c, Taf87]. 1988 [Hei88]. 1989 [Alv89]. 1990 [CW91]. 1992 [Mee92]. 1993/Amd [ISO93]. 1994 [ISO94b]. 1995 [Ame95b, Ame95a, ISO95a, ISO95c]. 1995/Amd [ISO07]. 1995/Cor [ISO01]. 1996 [ACM96]. 1998 [ISO98a, ISO98b, ISO98c]. 1999 [IEE99a, ISO99a, ISO99b]. 19th [GST01]. 1st [Ada82, Dia11].

2 [Ano84d, Ano86e, Ano86f, Ano86g, Ano86c, Ano87l, Ano87o, Ano88a, Bie85a, Col84, GH93, Gre86, Sch86b, SH89, Sch88, ST86, Sou90, WS84, Wol08, Ano86f, Ano86e]. 20 [Ano89e]. 2005 [Bar08, BW07, GS10]. 2012 [Bar14, RCM12]. 20th [CW91]. 213 [Ano87m]. 2167A [Wal91]. 24 [Ano88a]. 244 [Mos86]. 26 [Ano86c]. 271pp [Alb85]. 28-July [Ano93f]. 2nd [ACM93a].

3 [FM87, Kin93, Mea88, Mee92]. 3-540-18008-7 [Mea88]. 30-October [Ano01]. 307 [Ano87i]. 309 [Ano87t]. 31st [Ada82]. 32 [Ano87r]. 320 [Wic88]. 341pp [Ano85c]. 343 [Ano87h]. 359pp [Bud88]. 364 [Pyl88]. 371 [Ano87v]. 38 [Ano87d]. 3rd [Wol08].

4 [II94, Mer84]. 40 [GBO87]. 400 [Kro98a]. 41 [Ano84c, Ano87p]. 432 [PCH^{+}82a, PCH^{+}82b]. 45 [Ano87l]. 459 [Ano87q]. 46 [Ano86g]. 479 [vdL84]. 48 [Ano82d]. 49 [Ano87f].

5 [Her85]. 51 [Ano87j, Ano87n]. 52 [Ano86e]. 534 [Ano93c]. 54 [Ano82c]. 55 [Ano87o]. 563 [Ano87k]. 58 [Ano86d]. 5th [PK00].

62 [Ano82e]. 66 [Ano93d]. 68000/Unix [Gar86]. 69.00 [Ano97a]. 6th [CS01].

7 [Ano83b, Ano83c, Mea88, Tug84]. 73 [Ano87e]. 74 [Ano82g]. 75th [JL11]. 780 [SHLR80]. 7th [BS02].

8 [Bus96]. 8.95 [You82a]. 80 [Ano82b, Ano93a]. ’80s [Ano82a]. ’81 [Ano81b]. 83 [MB96]. ’86 [KCGO86]. 8651-3 [ISO88]. 8652 [Ame95b, Ame95a, II94]. 8652-1995 [Ame95b, Ame95a]. 8652/1995 [T^{+}00, TDBP01, TDB^{+}06]. ’89 [IEE89]. 8th [Ano90b, RS03].

9 [Ped88]. 90 [Ano86b, Bus96]. 90-5199-142-8 [Bus96]. ’91 [ACM91a, ACM91b]. ’91/Summer [ACM91b]. ’92 [Chr91]. ’93 [ACM93b, ACM93c, Gau93b]. ’94 [ACM94b]. 95 [AR96, Ano95a, Ano97b, AH^{+}97, Bal97, Bar96, Bar97, Bei97, BLB96, Bre96, Bro96a, Bro97, Bun96, BW96, BW01, BW04, Car97, CS98, CXZY02, CU96, CK96, CSM96, CP96, CHR^{+}02, De 96, DAA96, FT96, FK96a, Fel97, FK99, GSX99, Gli96, GGP97, HP98, Hol96, HP97, H^{+}98, Int96, JPMAB00, Joh97, KP96a, Kun98, Mad96, MWR98, Mol96, Moo95, NF96, OC96, REC96, RW00, Ska97, Ska02, Smy97, SD98, TD95, Taf96a, Taf96b, TD97, Tok01, Wat97, WB96, WB97, WBP97, WJS^{+}00, Ano97a]. 9593-3 [ISO90]. ’96 [ACM96, Str96]. 9638-3 [ISO94c]. ’97 [HB97]. 978 [Wol08]. 978-0-521-86697-2 [Wol08]. ’99 [GdlP99, HD99, Ano87g]. 9th [LS04]. 9X [Ano95b, BG95, CB94, GMB93, GTG92, Nar91, Plö92, dVdV95].

A. [Ano93e, vdL84]. Abbott [BYY87]. Absolute [ZRdlP01]. Abstract [Bel91, Fel84, Ano87s, Car96, CB94, GZ87, HT96, Hil94, LAH94, NM91, Shu89b]. Abstraction [Hil83, FHT86, GH93, ØK99]. Academia [Ano93g, Ano97c]. Academic [Her85, Mos86, Ano95b]. accelerator [LDD^{+}94]. Access [SC94, JT98]. Accessing [MS02a]. Accurate [Tan90]. Acedemia [Ano96]. Achieving [CH97, Hei96, SC97]. ACM [ACM80, ACM93a, ACM94b, Ano01, Ano02, Ano03, Ano04, Ano05, Ano06, Bee94, Swa11, Gic09, Ske82, Sof85, Whi81a]. ACM-SIGPLAN [ACM80]. Acquiring [Ard87, Ano87i]. action [Jon89, NJ05, Rom98]. Actions [MWR98, RRS^{+}97, WB97]. ActiveX [Kro98b]. ADA [ACM94a, Alb85, Ano82f, Ano85c, Ano86d, Ano87s, Ano94, Ano04, BS02, BU84, Mos86, You82a, You82b, ACM80, ACM82, ACM91b, ACM93c, ACM94b, ACM96, AK07, Alv89, Ano81c, Ano82d, Ano84c, Ano85b, Ano86e, Ano86f, Ano86g, Ano86b, Ano86c, Ano87l, Ano87j, Ano87o, Ano87q, Ano87i, Ano87d, Ano87k, Ano87t, Ano87w, Ano88a, Ano88c, Ano88d, Ano89e, Ano90b, Ano93f, Ano95b, Ano95c, Ano97a, Ano11, Ano02, Ano03, Ano05, Asp98, BF85, Bar87c, Bar03, Bee94, Bee97, Boo89, BP12, Chr91, CS01, Fre82, Gau93b, GB94, Gic09, GdlP99, GWA91, HM87, HB97, Hei88, Hoo92, IEE86b, IEE92a, IEE96, IEE99a, IEE99b, ISO93, ISO95b, ISO99a, ISO99c, ISO00, I^{+}86, Kar90, vK92, KV08, KK09, LS82, Lee92, LS04, MMH88, Mac83, MO90, Mea87]. Ada [Mea88, MH87, NS87a, NS87b, Obe88, Obe94, Pay93, PH06, PK00, Pyl81, RV10, RV11, RS03, Sec88, Ska95, Smy97, Sof85, Str96, Swa11, TDBP01, Taf87, TG80, TS85, Tel84, Tou94, Tou96, Uni85, VW05, Wic84a, Wic84b, Wol08, YLT93, Zal92, ZT86, ACM94a, Ame95a, All84, Ano80a, Ano81b, Ano81a, Ano82h, Ano83a, Ano84b, Ano84e, Ano85a, Ano85d, Ano86a, Ano86i, Ano86k, Ano87b, Ano87c, Ano89c, Ano89d, Ano93d, Ano93g, Ano95d, Ano96, Ano97c, Bar82, Bar87b, Bar88, Bar89, BM82, Bel91, Bie85a, BBP^{+}84, Bou80, BMO92, Bro89c, Bur88, BW95, BW98, CCD90, CCD91, CCD93, CCS87, Coh86, Col84, CW90, Cul91, DSK90, DMM88, DG87, Dun82, Fel84, FHK88, GR80, GW90, GD84, GG82, G^{+}83, Gra88, HP89, ISO90, ISO94a, Jan80, Kem87]. ADA [Le 84, Le 85, Lig90, May82, McG83, Mos86, Mun88, Mur91, PM07, PEGR80, PZ86, Rel89b, Rel89a, Rom96, San81, Shu89b, SC82, Tes81, Thé90, Twi83, U. 82, U^{+}82, Weg79, Whi81a, Whi81b, WS83, Wil87, vdB80, vdL84, ACM87, ACM91a, ACM93b, ACM94b, Ame83, Ame95b, AR96, AH85, AE92, AGG^{+}80, AB88, Alv89, AI85, AS92, Ano79a, Ano79b, Ano80b, Ano82a, Ano83d, Ano83f, Ano83g, Ano84a, Ano86j, Ano87a, Ano87x, Ano89b, Ano89a, Ano90c, Ano91, Ano92, Ano93b, Ano93a, Ano93e, Ano95a, Ano95b, Ano97b, Ano22, Ard87, AH^{+}97, A^{+}85, BM91, BJS93, BBCS96, Bak83, Bak88, BG95, Bal97, Bar96, BF85, Bar87a, Bar94, Bar97, Bar08, Bar14, BM86, Bas87, BMM96, BB95, Bau91, BBJL92, BBB^{+}92, Bei97, Bel97, Bel80, BBH80, BA98, BA09]. Ada [Bie85b, BB98a, BAP87, Bis90, Bis85, BØ80, BST98, BLB96, BB98b, Blu88, BHM^{+}82, Boo83, Boo87, Bor95, Boy87, Bra00, BBG^{+}81, Bra84, Bre96, Bre80, Bri84, BTM89, Bro89a, Bro89d, Bro89b, Bro80, Bro92, BEE92, Bro96a, Bro97, BW03a, BW03b, Bro05, Bro84, Bro96b, BFC00, Bru84, BIM93, BDG90, Bun96, Bur85, BR86, BLW87, BW96, BW01, BW04, BW07, BK87, CC86, Car97, CS98, CCO11, Cel96, CDF^{+}83, Cha85, CKK87, CHR86, Che92, CT94a, CXYZ02, CXZY02, CU91, CU96, CK96, Chr91, CL90, CWW80, CS85, CMM85, Coh81, Col93, Con86, Con88, CB94, CDC97, Cor96, CH80, CVL84, CSM96, Cra00, CWG^{+}06, CP96, Cul97, Cur91, CS91, CHR^{+}02, DX99, DS92, DPCC96, DR96, Daw88, De 96, DAA96, DT91, DD87, DAG^{+}88, DFS^{+}80, DHGR92, Dil90b]. Ada [Dil90a, Dil91, Dil93, DM87, DRF97, DPC95, DG80, DG82, Dru82, DSd92, DBF92, DH80, DBDS93, ECM97, EOAm94, EOM95, EL87, ERB12, EW91, EST86, Eme95, EMN98, Erd02, EHMO91, EGC02, Eva84, EP85, EMB^{+}99, Fag00, FMP12, FT96, FM89, FLP90, Fel90, FK93, FK96a, Fel97, FK99, FG84, FKR86, FSO89, FM87, Fra97, FWH84, Fus90, GN93, GN97, Gal91, GM89, GKB86, GVIV12, GKPT96, Gar86, Gau93b, Gau93a, GBO87, GSX99, Geh82, Geh83, Geh84a, Geh84b, GC84, Geh87, GR88, Geh89, Ger84, GMB93, GTB91, Gil86, Gli96, GTG92, GBdlHQCGB98, GGP97, GRGG98, Goo80, Gre86, GS10, Gro92, GMP90, GMAA97, GS85, HP83, HP98, HRGG98, HD99, Hei88, HL85, Hem90, HL83, HL01, Hil88, Hil94, Hil92, HLRS80, Hol83, HSWZ94]. Ada [Hol96, HP97, Hoo85, HW89, Hug91, HvKT87, Hum92, Hun85, Hus90, HW87, H^{+}98, IEE92b, ISO88, II94, ISO94b, ISO94c, ISO95a, ISO95c, ISO96, ISO98a, ISO98b, ISO98c, ISO99b, ISO01, ISO07, ISO12, Ich79, IKBW^{+}79, Int96, Jac85, JM83, JS90, JYCM94, JKC89, JPMAB00, Jin92, JpJ90, Joh97, Jon86, Jon89, JSV97, KSB89, KB91, vJK87, vK92, Kat82, Kat84, KP96a, KP96b, Kin93, KU87, KBL80, Kro98a, Kro98b, KT96, KRS01, Kun98, KP90, LH83, Lam83, Lam02, Lam03, LM92, LL86, LN93, Lau96, Le 82, Led81, LS82, Led83, LCS91, Lev89, LvdGvK89, LRT91, Li95, LC89, LM84, LXC03, LX04, LAH94, Lof93, LZLX04, LP80, LvLS84, L^{+}87, Lun89, Lun90, Lun91, Lun92, Lut98, Lyo87, LF90, Mac80, Mad96, Mag17, Mah81, MZGT85, Man92, MDPM08, MR91]. Ada [MD92, May83, McC92, MCD^{+}94, MSH11, MA89, McG82, MGDH02, MG91, MB96, MMHS87, MP90, MAAG96, MGM^{+}02, Mit83a, Mit83b, Mit83c, Mit83d, Mit87, MWR98, Mof81, Mol96, Moo95, MH97, Mor81, Mos90, MSS89, Nai89, Nar91, NF96, NB84, Nic80, NS87c, NS88, NC90, NU89, OB80, OBM96, OCM^{+}84, Orm86, OC96, Owe87, Owe89, PV12, PCBE96, Per89, Plö92, PCH^{+}82a, PCH^{+}82b, Pri84, PW92, Pyl85, RZP^{+}88, Rad90, RSC93, Rai92, Ram89, Rap98, REC96, Ree85, Rey85, RF96, RAH^{+}01, RH01, RH02, REMC81, Rog84, RW00, Rom98, Roo89, Ros92, RLHS80, Ros85, Ros91, Ros96, Rub82, RCM12, Sag87, Sai85, Sam86, SvA^{+}98, San94, San89a, San89b, Sav81, Sca94, Sca91, Sch82, SR85a, SR85b, Sch86a, Sch86b, SH89, SKL88, ST86, Sen92]. Ada [SS22, SC88, Sha88, SMBT90, STMD96, SC94, SHLR80, Shu89a, Shu88, Sil92a, Sil81, Sil91, Ska88, Ska94a, Ska95, Ska97, Ska02, Ska94b, Ske82, SW83, Sma96, SMB83, Sme85, Smy97, SMD95, SG91, SKW^{+}86, Sri94, ST84, S^{+}85, SM91, Stexx, SFGT81, Ste80, SD98, Taf82, TD95, Taf96a, Taf96b, TD97, T^{+}00, TDB^{+}06, Taf87, TCO91, TE87, Tan90, Tem86, TN92, TDB92, Tok01, TO98, Tom89, Ton98, Tou87, Tou94, Tou96, Uni83, UA83a, UA83c, Uni81, UA83b, U. 97, VK88, VM87, VMBK89, VKT91, Wal85, WS80, Wal84b, WW84, Wal91, WCW96, WA02, War86, Wat97, WWF87, Wea92, WF97, Weg80a, Weg80b, WHD86, WMS^{+}89, Wei03, WB96, WB97, WBP97, WJS^{+}00, Wet81, Whe81, Wic84c, Wic84b, WS84, Wil06a, Win99, Wit90, Wol91, YYB^{+}21, YT90, Yeu97]. Ada [You83, YTL^{+}95, Zal88, dVdV95, vv84, vdLN81, Ano83c, Ano84d, Ano93c, Mer84, Wal83, Ano82e, Ano82g, Ano87h, Ano87m, Ano87p, Ano87r, Ano87n, Ano87u, Ano87v, Ano88b, Ano90a, Ano98, Aug95, Aus11, Bud88, Her85, Mea87, Nie86, Pyl88, Tug83, Tug84, Wal84a, Wic88, Wim83a, Wim83b, Ano83b, Lla93, Mee92, Ano82c, Ano82b, Ano86h, Ano87e, Ano87f, Ano87g, Bus96, Her87, Ped88]. Ada-95 [GSX99]. Ada-9X [GTG92]. ADA-based [Mos86, LvLS84, JSV97, Bor95, CL90, DX99, Fag00, Owe89]. Ada-Boost [SS22]. Ada-CCM [MDPM08]. Ada-compiler [vJK87]. ADA-Europe [BS02, AK07, Alv89, Asp98, Chr91, CS01, Gau93b, GdlP99, HB97, Hei88, vK92, KV08, KK09, LS04, PH06, PK00, RV10, RV11, RS03, Str96, Taf87, VW05, Gau93b, HD99]. Ada-like [CT94a, JYCM94, Ree85]. Ada-Mentoring [Ano11]. Ada-Object [BBCS96]. Ada-programming [Ano82a]. Ada-specific [CDC97]. ADA-tree [DG87]. Ada.Real Time.Clock [ZRdlP01]. Ada/O2 [MB96]. ADA/PASCAL [Ano83a]. Ada83 [Fel93, WVC^{+}01]. ADA94 [CGS94]. Ada95 [Che97, Hei96, Kem96, MS98, NMH^{+}02, OMÁ^{+}02, SAV96, Sti98a, Sti98b, WVC^{+}01, WN97, Xu98]. Ada95-like [Che97]. Ada’97 [ACM97]. Ada9x [Fel93]. Ada constructs [Tex82]. Ada education [Tex82]. ADAM [LvLS84]. adaptation [PW92]. Adaptive [TC04, CQG^{+}13, Ano87s]. AdaTEC [ACM82]. Adding [Lam02]. Addison [Ano87v, Bud88, Pyl88, vdL84]. Addison-Wesley [Ano87v, Bud88, Pyl88, vdL84]. Addition [OC96]. Additional [BLB96]. addressing [Bis85, Ano87h]. ADDS [BR86]. administration [JpJ90]. ADTEST [GN97]. advanced [GSX99, Geh83, Geh84a, Geh89, HP89, ST84, TS85]. Advancing [IEE86a]. aerosol [Fra01]. against [Bre80]. Age [CW91]. agenda [Ada82]. aid [FW96]. AIDA [Ano87x, EOAm94, EOM95]. AIDA-87 [Ano87x]. Aided [CDC97, LC89]. aimed [Ano85a]. Air [DNM^{+}10, SvA^{+}98]. Alan [Alb85, Ano98, Mea87, Mea88, Wol08]. Albuquerque [Ano06]. Algebra [LXC03, Alb05, LX04]. Algebraic [BEPP87, LM84]. Algol [LP86, Sch82]. Algorithm [Hun85, SS22, Wei03]. Algorithms [HL93, Ste80, Ano97a, Bei97]. alliance [Bla02]. Alone [CWG^{+}06]. Alonso [Ano93e]. Already [CWG^{+}06]. alternating [WY88]. Alternative [Bel97, Sti98a, Sti98b, Ano86e]. Amd [ISO93, ISO07]. amelioration [Gra88]. Amendment [IEE96, IEE99a, ISO07, TDB^{+}06]. America [Bla02]. American [Alb05, Uni83]. Ammann [Ano93a]. Amok [Lut98]. among [Bra84, Sou90]. AMPATS [Man92]. Amsterdam [Bus96]. Analogy [Kro98b]. analyser [CVL84, EOAm94]. analyses [Eva97, Mar95]. Analysing [Hol83]. Analysis [Bas87, Bel97, BB98a, BB98b, CXZY02, CHR^{+}02, DT91, DAG^{+}88, Hil88, Kro98a, Kro98b, MGDH02, RSC93, Ros96, Sav80, SAV96, U^{+}82, WW84, Ano85b, BBWF95, BST98, Cor96, DBDS93, Eas83, KB91, Lin93, NJ05, SC88, SMBT90, STMD96, Wei03, YLT93, YTL^{+}95, Kro98b]. Analytical [CSM96]. analyzer [EOM95]. Analyzing [CU91, KRS01, AE92]. ANDF [Bun96]. Andrew [Mea88]. Andy [Ano98, Wol08]. Anglo [Ano86i]. ANNA [KBL80, L^{+}87, Fra97, San89b]. Anna/Ada [Fra97]. Annasoft [Kro98b]. Annex [Bal97]. annotating [KBL80, L^{+}87]. Annual [ACM91b, ACM93c, ACM94a, ACM94b, Ano87x, Ano90b, Ano93f, Ano94, Ano01, Ano02, Ano03, Ano04, Ano05, Ano06, Gic09, IEE89, Ada82, Swa11]. anomalous [Fra01]. Anon. [Ano87r]. ANSI [IEE99b, Ada83, Ame95b, Ame95a, Uni83, UA83a, UA83c]. ANSI/ISO/IEC [Ame95b, Ame95a]. ANSI/MIL [Ada83, Ame95a, Uni83, UA83a, UA83c]. ANSI/MIL-STD-1815A [Ada83, Uni83]. ANSI/MIL-STD-1815A-1983 [UA83a, UA83c]. Answers [Wic84c]. Antonio [IEE86a]. APE [San89a]. API [IEE92b, IEE99a, IEE99b, IEE92a, IEE96, ISO99a]. apparent [GV94]. appear [Ano81b]. Appelbe [Ano87m]. Apple [CW91]. Application [BYY87, Bro96b, Ein90, GGP97, GV94, IEE92a, IEE96, IEE99a, IEE99b, ISO99a, Lau96, RH02, Rom96, STMD96, VGdlP01, BM85, IEE92b, Zal88]. Application-Defined [RH02]. Application-Level [GGP97]. Applications [ASM88, Ano88d, CW04, GTB91, Gli96, GRGG98, GMAA97, HRGG98, IEE86b, KD08, Kro98a, LM92, MSH11, MAAG96, NMH^{+}02, PV02, RF96, RH01, Ros91, Ros96, WVC^{+}01, ACM94a, ACM94b, Aus11, Bar87b, BB95, Che92, CMM85, DH80, JPMAB00, JpJ90, Sch88, Whi89]. applied [Ano87s, DG87]. Approach [Bro84, CK96, CSM96, Cur91, CHR^{+}02, Dil91, FMP12, GBdlHQCGB98, Li95, LM84, Sca94, YYB^{+}21, dVdV95, ACD^{+}87, Ano97a, Bei97, Bis85, Car96, CQG^{+}13, Cul91, Cul97, FK96b, LAH94, Mur91, RW00, SC97, Boo89, Ano84c]. Approaches [Bau91, Lam03, CP96]. approximation [Fra01]. AppSwitch [Bra00]. April [Ano87q, Ano87i, Ano87f, IEE86b, NB84]. APSE [Obe88, Bre80, Lyo87]. arbitrary [BS90]. ArcAngelC [OC08]. archetype [Gra88]. Archetypes [PV12]. Architectural [Bis85]. architecture [GS10, HSLG92, JT98]. Architectures [Dia11, Mad96]. Arcturus [ST84]. Ardo [Ano87i]. area [Bur88, WY88]. ARINC [CH97]. Arithmetic [BEE92, Fig00, Ano82b, Vig93]. Arlington [ACM82]. Array [CPD93]. Art [EMB^{+}99, CH02]. Artaza [Ano93c]. Article [Ano82f, Ano82c, Ano82d, Ano82e, Ano82g, Ano82b, Ano84c, Ano85b, Ano86d, Ano86e, Ano86f, Ano86g, Ano86b, Ano86c, Ano87l, Ano87j, Ano87o, Ano87e, Ano87q, Ano87i, Ano87h, Ano87m, Ano87p, Ano87r, Ano87d, Ano87f, Ano87k, Ano87g, Ano87n, Ano88a, Ano90a, Ano93c]. Artifact [RCM12]. Artificial [Ano87x, Wal85]. Artificial-intelligence [Wal85]. ARTK [DHGR92]. Artlandia [Kro98b]. AS/400 [Kro98a]. Ascent [CW91]. ASIS [ISO99b, KRS01]. Aspects [RT00, Ano87t, HvKT87, Sch86c]. Aspray [CW91]. assembler [GBO87]. assembly [Ano86c]. Assessing [FG84, Alb85]. Assessment [DT91, Ros96, Ano89a, ISO99c]. assisted [FM89]. Association [USE85b, USE86b]. Assurance [IEE89, Sch88]. Astro [Sti98a, Sti98b]. Asynchronous [BW03a, BW03b, BG95]. AT&T [EST86]. ATAC [BMM96]. ATC [Gro92]. Athens [Chr91]. Atlanta [Ano90b, Ano05, USE86a, Ano04]. Atlanta/Buckhead [Ano05]. Atlas [Mar95]. ATM [Lut98]. Atomic [MWR98, RRS^{+}97, Rom98, WB97]. Attention [Ano86b]. Attractions [Rap98]. Attribute [U^{+}82, MB86]. Augarten [ZT86]. Augmenting [BLB96, CS85]. August [Ano86c]. Augusta [Mit83a, Mit83b, Mit83c, Mit83d]. Austria [BS02]. autobiografia [BV07]. autobiography [BV07]. automata [Sav81]. Automated [Luq90, BST98, Hei96, SC88]. Automatic [DHGR92, DM87, DMM88, DMM90, Fra97, Hus90, IEE86a, Kro98b, NB84, NM91, Sav80, MT82]. Automating [EMN98]. Autonomic [Dia11]. Autotestcon [IEE86a]. Autumn [USE87]. Available [Kro98b, Hal83, Wal85, Whi81a]. Avionic [Ros91]. Avoidance [LM92]. AVR32 [GS10]. Axioms [BM82, Ano82d].

B [Ano86b, Ano87m, Ano88a, ERB12, IEE86a]. Babbage [CW91]. Babel [Bro81]. Back [CW91]. Background [Sei89]. Baker [Ano87j]. Baltimore [ACM90]. Barnes [Lee92]. Barringer [Ano82d]. Barry [CW91]. Based [Bro96b, Bun96, DS92, JSV97, LXC03, MDPM08, MGDH02, PV12, RCM12, Ton98, Yeu97, BK95, Bor95, Car96, Che97, CQG^{+}13, CC94, CL90, CB96, DX99, Dil91, DBF92, Fag00, HSLG92, KB91, MO94, Mos86, Owe89, PM07, RSC93, Rey87, RRS^{+}97, SMBT90, Taf82, WCW96, Wot00, YLT93, ZRC91, ZLZ^{+}96, LvLS84, SR85a, SR85b]. Bases [KCGO86]. basic [BEPP87, ISO98b, Woo89, Rel89b, Rel89a]. Basics [CW91, Ano86b]. Bastard [RAH^{+}01]. Bath [Wal84b]. Be [CWG^{+}06, Hal83, Rad90]. Beach [IEE86b]. became [Alb05]. becomes [Wal85]. before [ABCK^{+}90]. beginning [Ska88, Ska94a, Ska97, Ska02]. Behavior [OCM^{+}84, SC88]. Beidler [Ano97a]. Belgium [CS01]. Benchmarking [Cur91, FSO89]. benefits [Ano82b]. Benjamin [Wal84a, Ano92]. Benjamin/Cummings [Wal84a]. Berlin [Mea88]. BETA [ØK99]. Betty [Lla93]. between [Bla02]. Beyond [OC96, Mof81, U. 97]. Bias [GC84]. Bibliography [Bee94, Bee97, Rog84]. Big [Wic84c]. Binding [ECM97, EMN98, EGC02, IEE92a, IEE92b, IEE96, IEE99a, IEE99b, Kro98b, ISO98a, ISO99a, ISO93]. bindings [ISO88, ISO90, ISO94c]. birds [CWW80]. Birthday [JL11]. Bit [ZT86, ZT86]. Black [Kro98b]. Blackwell [Ano88b, Wic88]. Blocking [Ger84, RRS^{+}97]. Bloomington [Ano01]. Bo [Aug95]. board [BMM96]. Boasson [All84, Mer84, Wim83b]. Body [EMB^{+}99]. Booch [Ano82f, Wal84a, Lam03]. Book [Alb85, All84, Ano81c, Ano83b, Ano83c, Ano84d, Ano85c, Ano86h, Ano87s, Ano87u, Ano87t, Ano87v, Ano88b, Ano88c, Ano97a, Ano98, Aug95, Aus11, Boo89, Bud88, Bus96, Her85, Her87, Hoo92, Lee92, Lla93, Mea87, Mea88, Mee92, Mer84, Mos86, Nie86, Pay93, Ped88, Por01, Pyl88, Sec88, Tug83, Tug84, Wal83, Wal84a, Wic84a, Wic84b, Wic88, Wim83a, Wim83b, Wol08, You82a, You82b, vdL84, MCD^{+}94]. Books [Ano84d, FSJ00, Lut98, Por01]. Boost [SS22]. början [Ska95, Ska02]. Boston [ACM80, ACM87]. Both [KP96a, KP96b]. bottleneck [ZRC91]. bound [Sri07]. Bowring [Ano87p]. Box [CWG^{+}06, RAH^{+}01]. Boy [RAH^{+}01, Alb05]. Bravo [Kun98]. Brest [KK09]. Brian [Ped88, Wal83, JL11]. Brighton [Ano95b]. Britain [Bra89]. Broker [Kro98b]. Brosgol [Ano92]. Brown [Ano84c]. Brussels [Tel84]. BT [Ano86i]. Buckhead [Ano05]. buffered [Ref90]. bug [San89b]. Bührer [Ano93a]. build [Ano92]. Building [GBdlHQCGB98, MSH11, Plö92, ZAdlP97, JPMAB00, Aus11]. built [Bro81]. Bureau [Ano48, Ano84a]. Burh [YLT93]. Burns [Ano90a, Ano98, Mea87, Mea88, Wol08]. Business [CW91]. Byme [Pay93]. Byte [Bal97]. Bytecode [Int96].

C [Alb85, Ano85c, Ano86e, Ano87k, CW91, Mer84, Mos86, Wim83b, You82a, Ano89d, BASS96, Car97, Eva84, FG84, GR88, Hei96, HSWZ94, Hug91, Joh97, MO94, PCBE96, Rel89b, Wil06a, dVdV95]. C. [All84, Mac83]. C.N.E.S [Lau96]. C36 [Ano87q]. CA [Ass83]. CACI [Kro98a]. CAI [Kro98b]. CAIS [Obe88]. calcul [d’O86]. Calculating [Ano48, Har49, Har84]. Calculation [CW91, Hor82, Hun85, Woo89, d’O86]. Calculations [DPC95, LN93]. California [Ano03, Lla93, BU84]. Cambridge [Ano98, Hoo92, Lee92, Mea87, Por01, Sec88, Wic84a, Wol08]. Cameras [ZGMK07]. Can [EMN98, Mor81, Rad90]. Canada [ACM84]. Capabilities [JSV97, Ano82f]. Capability [BEE92, Goo80]. Capri [HM87, MH87]. Capsule [ZT86]. Card [CW91]. Case [AB88, Boo89, PW92, RCM12, VGdlP01, Ano86c, Blu88, Bre80, Dun82, SW94, VKT91, Kro98a, SLM91]. casts [Mos86]. CCM [MDPM08]. CCS [HL83]. CE [Kro98b]. Celebration [Hor82]. Center [ACM90, Ano88d, IEE86a, Kro98a, Lau96]. centred [CT94b]. Century [CW91, GST01]. CGI [ISO94c]. chaining [FK96b]. Challenge [TN92]. change [Ano90a, Rus87]. Changes [RAH^{+}01, Ano93e]. channels [Ref90]. Chapman [Mee92]. characterisations [FW91]. Charles [CW91]. Charrette [RLHS80]. Checking [FMP12, Thé90]. Chicago [Alb05]. Chichester [Ano83b, Nie86, Wal83]. CHILL [All84, Mer84, Wim83b, SMB83, Sme85]. choice [Alv89, Chr91, Ano86b]. Choosing [MT82]. Circuit [CW91]. Circus [CCO11, OC08]. civil [Sav81]. Class [Rom99]. Classification [Kem96, SS22]. classifier [CB09]. classwide [KP90]. Clear [Ano02]. Client [Bre96, Hei96, Kro98a, PCBE96, TC04]. Client-Server [Bre96, Hei96, Kro98a]. client-side [TC04]. Cliffs [All84, Ano81c, Aug95, Ped88]. Clocks [GRGG98, HRGG98]. CLOG [NB84]. closed [DO02]. Club [WMS^{+}89]. cluster [TC04]. COBOL [Rel89b, Rel89a]. COCOMO [Fai07]. Code [Ano97b, Bal97, DSd92, Hei96, Int96, KRS01, PV12, RCM12, Sca94, TDB92, Bra84, CCD91, CCD93, SHLR80, YYB^{+}21]. Codes [BM91]. codesign [Tem94]. coding [AI85]. Cognitive [Sca94]. Cohen [Her87]. cohesion [Dha95]. Collingbourne [Bus96]. Collins [Ano83b]. Colorado [USE86b]. colored [SBM94]. COMAL [Ath82]. Combustion [ZGMK07]. command [Bre80]. comment [Jan80]. Comments [Gre86, Kar90, MO90, TG80, WMS^{+}89, Whi81a, Xu98, YLT93, vdB80, Mah81, Nic80]. commercial [Ano83d, Ano84b]. committee [Ano81a]. Common [ECM97, ISO98a, RAH^{+}01, Fis78, Obe88]. communicating [Ram99]. Communication [GBdlHQCGB98, FHT86, KP90, Lun90, NC90, vdB80]. communications [Bus96, HW89, Kar90, Ref90]. community [Ano87g]. Companion [Lee92, Lev89, Wic84a, Mea87, Rel89b, Rel89a]. Comparative [Bau91, Vaj86]. compared [Sch82]. Comparing [Col84, FG84, FWH84, Alb85]. Comparison [Bro97, BW03a, BW03b, Bro05, Coo96, Eva84, MG91, San81, Whi89]. Compass [Ano87u, IEE89]. compatible [CDF^{+}83]. Competitiveness [ACM91b]. Compilation [Cur91, EST86, Hoo92, TN92]. compile [LZLX04]. compile-time [LZLX04]. Compiler [ACM84, Bun96, GM89, Ton98, Ano83d, Ano86b, Bro80, CGS94, Con88, CVL84, Goo80, vJK87, RMP90, RLHS80, San89b, Smy97, Ano87i]. compiler-generator [CVL84]. Compilers [WW84, Ano84e, Ano85a, Ano89c, Rog84]. Compiling [WA02]. complex [Fra01, HNVW91, ISO98b, ISO98c]. Complexities [CL05]. Complexity [DS92, Blu88, GR80, Sha88, WCW96]. compliance [Sen92]. Component [Kro98b, MDPM08, PV12, SKW^{+}86, SLM91]. Component-Based [MDPM08, PV12]. Components [LM84, MGDH02, NF96, Boo87, Eva97, HSWZ94, Sri07, Taf87]. Compositional [GSX99, KSdR^{+}88]. Comput [Ano82a, Ano82c, Ano82d, Ano82e, Ano82g, Ano82b]. Comput. [Ano84c, Ano86b, Ano87q, Ano87g]. computable [Zen13]. Computation [Ano48, GV94, Mor81, Vig93, Zen13]. Computations [CH80, Blu88]. Computer [AFI72, Ano87s, Ard87, Bro81, CW91, CDC97, IEE86b, IEE89, ISO88, ISO90, ISO94c, LC89, Mea88, RR93, Wic84b, BV07, Bus96, Fag00, FLP90, TE87, Ano87e, CW91]. Computer-aided [LC89]. Computers [Lut98, WMS^{+}89, ABCK^{+}90, WCK85]. Computing [ABCK^{+}90, Bow53, CWG^{+}06, JL11, Ano83e, KSdR^{+}88, CW91, Dia11]. con [May83]. concept [Air85, CW91]. conception [Lig90]. concepts [Fre82, IEE86a, Sch86c]. concrete [GR80]. Concurrency [BW95, BW98, BK87, CS98, Geh82, Ghe85, Lut98, MG91, Shu88, YTL^{+}95, Ano98]. Concurrent [Bur85, BW96, BW07, CXYZ02, CXZY02, GC84, GD84, Jin92, SH89, Vaj86, FM89, FLP90, Geh84b, OZC11, Rom97, Rom00, SM91, SBM94, TCO91, WCW96, BAP87, BST98, BASS96, BK87, GR88, Wol08, Mea87]. concurrente [Sch86b]. Condensed [Ano84d]. Conference [ACM82, ACM93a, ACM93b, ACM97, AFI72, AK07, Alv89, Ano87x, Ano88d, Ano90b, Ano01, Ano02, Ano03, Ano04, Ano05, Ano06, Asp98, BF85, BS02, BP12, BU84, Chr91, CS01, Gau93b, Gic09, GdlP99, HB97, Hei88, IEE86a, IEE86b, IEE89, KCGO86, vK92, KV08, KK09, LS04, PH06, PK00, RV10, RV11, RS03, Str96, Taf87, Tel84, Ass83, USE85b, USE85a, USE86b, USE87, VW05, ACM87, Swa11, USE86a, Whi81a, Ano93g]. configuration [BIM93]. Conformance [Mad96]. Conformity [ISO99c, Weg90]. confronting [BHM^{+}82]. consensus [Plö92]. Consequences [OC96]. Considerations [Sil92a]. Consolidated [T^{+}00, TDBP01]. constrained [DO02]. constraint [Car96]. constraint-based [Car96]. Constructing [CHLY12]. Construction [ACM84, CVL84, Fel97, Aug95, San94]. constructive [SC97]. consumer [Hil92]. context [Tom89, Air85]. contrôle [Car97]. Contract [Lam02]. contraction [CKS83]. Control [BW03a, BW03b, Cel96, CW90, Kro98b, Lau96, LRT91, NMH^{+}02, OMÁ^{+}02, SOK92, Sch86a, ZGMK07, Ano82f, Ano93c, BG95, BM87, Bor95, CCO11, CC94, CKS83, GS10, LDD^{+}94, PEGR80, Ref90, RT00, San95, Sav80, SC94, TM98]. Controller [PM07, Ram87]. Controls [Kro98b]. Controversial [De 96]. Convention [ACM90, IEE86a]. conventional [Rom00]. conversation [Rom96]. Conversion [GBO87, SW83, Ano83b, SC82]. Converting [Ano97b, Gli96, Mol96, Sca94]. Converts [Int96, Wal85]. coordinated [RRS^{+}97]. Copenhagen [Tou94]. Coprocessor [BMM96, Lun91]. Copy [Kro98a]. Cor [ISO01]. CORAL [San81]. CORBA [CK96, Kro98b, NMH^{+}02]. core [YYB^{+}21]. Corner [ACM94b]. corporation [Bla02]. Correct [Ano04, Ano02, Ano03, Ano05, Eva95]. Corrigenda [NS87a, NS87b]. Corrigendum [ISO01, TDB^{+}06, T^{+}00, TDBP01]. cost [Smy97, SC97]. costs [Ano82b].

Could [WN97]. Council [Ano89a]. countess [JM83]. counting [MMHS87]. Coupling [MB96, Dha95]. course [BMO92, Gau93a, LL86, LAH94, Owe89, Sil91, SC82, Tem86, Ano83b]. courses [AH85]. courseware [FLP90]. cover [Mea88, Wic88]. Coverage [Kun98]. craft [Ada10]. CRAI [HM87, MH87]. creating [Ano86c]. Critical [CSM96, KSB89, MO90, RF96, REMC81, Ros96, ZAdlP97, Ano93a, Che92, Sri94]. Criticality [CW04]. Critics [Wic84c]. Cross [Kem96, LN93]. Cross-Classification [Kem96]. cross-section [LN93]. CS1 [MCD^{+}94]. CSPL [CT94b, Che97]. Culture [Eme95, Bra89]. Cummings [Wal84a]. Current [Bau91, McG83]. curriculum [Owe87, TE87]. Cursors [MS02b]. curves [Ano87l]. cyber [SS22]. cyber-physical [SS22]. Cycle [Bas87, Wic84a]. Cyclic [ZAdlP97].

D [Ano82e, Ano84c, Ano86d, Ano86g, Ano87d, Ano87t, Ano93d, Her85, HL93]. D-W [Ano86d]. D. [Mos86]. D2 [CG91]. dalla [BV07]. Dallas [USE85b, USE85a]. DAPSE [Boy87]. DARTS [GWA91]. Data [Bei97, Bel91, Car97, CHR^{+}02, GH93, GTG92, KCGO86, Kem96, Kro98b, LP86, Lut98, Mos90, Wei03, Ano87l, CCD90, Fel84, Fel90, FK96b, GN93, GN97, GZ87, Hil94, LAH94, Luq90, NM91, Shu89b, Sil91, Kro98a, Ano97a]. Data-Modeling [Lut98]. Data-Oriented [GTG92]. Database [Erd02, Ano86k, CDF^{+}83]. DataFAN [CHR^{+}02]. dataflow [YYB^{+}21]. David [Ped88]. Dawes [Hoo92]. DBMS [SG91, SKW^{+}86]. dead [CXYZ02]. Deadlock [Ger84, LXC03, MR91, DLP89, DBDS93, KB91, STMD96, YLT93]. Deadlocks [CU91, CU96, MSS89]. Debate [WMS^{+}89]. Debugger [LF90, BTM89]. Debuggers [Sil92a]. Debugging [HL85, TCO91, Wot00, FM89, LHF94, RFF92, San89b, Sch85]. Decade [Sma96, LC89]. December [ACM80, ACM87, ACM90, ACM96, Ada82, Ano87o, Ano88a, Ano02, Ano03, Whi81b]. Decentralized [Shu89a]. Decimal [BEE92]. declarations [ISO98b, SC94]. decomposing [HL93]. decoupled [JT98]. decoupling [CQG^{+}13]. Dedicated [JL11]. Defects [CW90, Eme95, AE92, Eva97]. defence [Ano85d, Kem87]. defense [Ano87g, Ano80b, BBG^{+}81, Wal85]. define [BG84]. Defined [RH02]. definition [BBH80, Nic80]. Definitions [BB98a, BB98b, Tok01]. del [BV07]. Delays [ZRdlP01]. Delivery [ACM94b, Ano93d]. dell’inventore [BV07]. Demonstrating [Sen92, FHK88]. demonstration [FM89, Win99]. Denmark [Tou94]. Density [Wit90]. Denver [USE86b]. Department [Ano48, Ano80b, Ano87s, BBG^{+}81, Wal85, U. 97]. Dependable [DPCC96, JL11]. Dependence [Jin92]. Dependency [CXZY02, Mos90]. Deployment [Sma96]. Derivatives [Hus90]. Des. [Ano86b, Ano87g]. Description [ISO95a, ISO95b, BØ80, OB80, Sav80]. descriptions [BYY87]. Design [ASM88, Ano79b, Ano93a, Ano95c, DT91, EMB^{+}99, GMB93, Lam02, Lam03, Lee92, LM84, ND94, Ros85, SMBT90, Thé90, Whe81, Alv89, AI85, Ano87w, BYY87, Bis85, Boo91, CCD91, CCD93, CQG^{+}13, CL90, DD87, FK93, FK96a, FK99, FM87, Gom94, Hen88, Hil83, IKBW^{+}79, I^{+}86, vJK87, LvdGvK89, MMH88, MB86, Pay93, Rey87, Rey89, Sch82, SKL88, TG80, WCK85, YTL^{+}95, ZLZ^{+}96]. Design concepts [Tex82]. Designer [Wic84c]. Designing [NS87a, NS87b, NS87c, NS88, San95]. Designs [DAA96, AE92, Wot00, YT90]. Detecting [LXC03]. Detection [CU96, LM92, MR91, MSS89, CXYZ02, DLP89, San89b]. deterministic [TCO91]. d’Évaluation [Le 82]. Developer [RAH^{+}01]. Developing [Bre96, CKK87, DPCC96, GTB91, Jac85, JSV97, MO94, Ram89, Sca91, TN92, BB91]. Development [ACM94b, AS92, Ano96, CDC97, Erd02, Jac85, Kro98a, Kro98b, KT96, OMÁ^{+}02, PV12, RCM12, Sma96, Ano87b, Ano87o, Ano87v, Ano93a, Bor95, Bud88, BR86, CC94, CJ92, Cul91, FWH84, HM87, HSLG92, MA89, MH87, ND94, Pyl88, Sei89, Wil06a, Wil06b]. developmental [Cul97]. Developments [Bis90, Tok01, Har84]. devices [Ano83e, ISO94c]. Devon [Bar87c]. Devouring [CW91]. Dewayne [vdL84]. Diab [Kro98a]. diagnosis [CHLY12, HSLG92]. Diagnostics [War86]. diagrams [CCD90, CCO11]. dialogue [BR86]. dialogues [ISO94c]. DIANA [G^{+}83, Ros85]. Diego [Ano03, BU84, Ass83]. diesel [GV94]. Difference [CW91]. Different [GTB91, TN92]. Differentiation [DM87, DMM88, DMM90]. Difficult [De 96]. diffraction [Fra01]. Diffusion [CW91]. Dig [Ano82c, Ano82d, Ano82e, Ano82g, Ano82b]. Digital [Ano48, Bow53, HL01, Pay93, Sil92b, WCK85]. Dimensional [EW91, Hil88]. DIMOS [Ano93d]. Dining [Bro96a]. direct [GV94]. direct-injection [GV94]. DIS [II94]. Discipline [Ano96]. Discourse [NJ05]. discrete [BB91, Bru84]. Discussion [BHM^{+}82, Che92]. display [NM91]. displays [Ano86c]. Dissecting [Lut98]. Distributed [Ano93d, Ano04, Bal97, Bau91, BBJL92, CK96, GVIV12, GMAA97, HP98, KP96a, KP96b, KU87, LvdGvK89, LRT91, MDPM08, MGDH02, MAAG96, MWR98, Shu89a, Sma96, SG91, TM98, USE89, VM87, VKT91, Zal92, ACD^{+}87, Ano87q, Ano87k, Ano02, Ano03, Ano05, Boy87, Car96, CDF^{+}83, CB96, DG80, FK96b, HW87, KSdR^{+}88, Lun90, Mos86, NC90, Rom00, VMBK89, ZRC91, ZLZ^{+}96, Bis90, GWA91]. distributed/concurrent [Rom00]. Distributing [BAP87, JKC89]. Distribution [BBB^{+}92, Fra01]. diverse [HT96]. diversity [Rom99]. Djavaheri [Ano86e]. dla [HP89]. DM [Ano87u, Ano97a, You82b]. Dobb [Ano86d]. Document [BBG^{+}81, Uni81, Ano80b, CCD91, CCD93]. documentation [Nic80]. documenting [LP80]. DOD [Con86, Fis78, Wal91]. DoD-STD-2167A [Wal91]. domain [DLGF05]. Données [Car97, Lig90]. Dorothy [Sec88, ZT86]. Double [Ano04, Ano05]. doublet [vv84]. doubts [Mos86]. Down [LS82]. Downes [Ano83c, Tug83, Tug84]. Dowries [Wim83a]. Dr. [Ano86d]. draft [II94]. Drago [MAAG96]. driven [Rey89, RMP90]. DSA [GVIV12]. DSP [Kro98b, Sil92b]. Dublin [USE87]. duties [Ano93b]. dynamic [BB91, BG84, EOAm94, EOM95, Kro98b, Ano90a]. Dynamically [Sri07]. Dynamically-bound [Sri07].

Early [CW91, WMS^{+}89]. EASEL [ND94]. East [Wol91]. EATMS [DNM^{+}10]. ECMA-162 [ECM97]. Economics [Lut98]. Eden [CW91]. Edinburgh [RV11]. Edited [Hoo92, Alb85, Bus96]. Edition [IEE99a, Ano98, Nie86, Wol08]. Editor [RAH^{+}01, WN97, MH97]. EDN [Ano92, Ano82f]. eds [Por01]. educating [Ano82h]. education [Fai07]. educational [BHM^{+}82]. Edvard [CW91]. effect [Dru82, JT98]. Effective [BW96]. efficiency [GS85]. Efficient [Li95, MB96, WS80, Ref90, SS22]. Efficiently [MGM^{+}02]. effort [Eva95, Fis78, Pfl91]. ego [Ano89e]. Eiffel [dVdV95]. EiffelBase [Kro98b]. eight [Fel90]. eight-year [Fel90]. Eighth [ACM91b]. Einstein [SvA^{+}98]. Eisler [CW91]. Elaboration [LM92]. Elaboration-Time [LM92]. electronic [WCK85]. Element [DMM88]. Elementary [Tan90, ISO94a, ISO98c]. eleventh [ACM94b]. ELL [CVL84]. Ellis [Ano83b, Nie86]. Embedded [GTB91, Kro98a, LF90, MD92, MSH11, RH01, Shu89a, Whe81, Ano83c, Ano87t, Aus11, Bar87b, Coo96, DG82, DH80, HT96, HvKT87, Rel89a, Sag87, Tug83, Tug84, Wim83a]. Enabling [BBCS96]. enchantress [Lla93]. Encryption [SvA^{+}98]. Encyclopedia [RR93]. End [Bro81, CW91, Lut98, Bro80]. ended [Ada82]. Energy [Bra89, CH02, Whi89]. Eng. [Ano87k]. engine [CL90]. Engineering [Ano95d, Ano04, MA89, Mit87, RR93, Wal84a, Wea92, ACM93c, AH85, Ano86g, Ano93f, Ano02, Ano03, Ano05, Boo83, BMO92, CCD90, CCD91, CCD93, Dru82, Fai07, Hug91, KP90, LL86, Lin93, LC89, McC92, Mur91, Owe89, Ree85, RT00, Sav81, Sch86c, Tom89, WS84]. Engineers [BA09, BA98]. Engines [CW91, GV94]. Englewood [All84, Ano81c, Aug95, Ped88]. enhanced [SS22]. entity [San95]. entity-life [San95]. entropy [Bra89]. Enumerations [Mof81]. Environment [CW04, Erd02, Kro98a, Lam83, Obe88, OCM^{+}84, PV02, Ros85, SMD95, SFGT81, CCD90, CT94b, Che97, CC94, DLP89, LN93, Lyo87, Mit87, NU89, RFF92, Sof85, ST84, Taf82, VK88, Weg80a, WHD86, ZRC91, vdLN81, Che92, ECM97, ISO98a, TS85, Tel84, Wic84a]. Environments [Ano88d, IEE86b, Obe94, II94, Mar95, Som89]. ENVISAT [DR96]. ENVISAT-l [DR96]. equation [Sch99]. EQUEL [Rel89b]. equipment [DSK90]. equipped [BMM96]. Error [Wit90, RRS^{+}97]. ESL [San81]. Essays [JL11, Bra89]. essentials [Cra00]. Estelle [MGK91]. Euclid [BK87]. Eurocat [DNM^{+}10]. Euromicro [Ano81b]. Europe [AK07, Alv89, Asp98, BS02, BP12, Chr91, CS01, Gau93b, GdlP99, HB97, Hei88, vK92, KV08, KK09, LS04, PH06, PK00, RV10, RV11, RS03, Str96, Taf87, Tel84, Tou94, Tou96, VW05, Gau93b, HD99]. Europe/Ada [Tel84]. Eurospace [Tou94, Tou96]. Eurospace-Ada-Europe [Tou94, Tou96]. EUUG [USE87]. evaluate [Sil91]. Evaluating [Le 84, Le 85, Her85]. Evaluation [BMM96, CH80, Hus90, REMC81, TDB92, Le 82]. Event [ERB12, BB91, Bru84, GS10]. Event-B [ERB12]. Evolution [Mur91]. Evolving [Mac80]. example [FHK88]. Examples [Weg80b, Aug95, Cra00, Jon89, San94, SH89, S^{+}85, Weg79, Ano81c]. exception [Rom97, Rom00]. Exceptions [PM07, Ano87j]. Exclusion [Bro05]. executable [BIM93, Hem90]. Execution [Dil90b, Dil91, GRGG98, HRGG98, Shu89a, VM87, Ano87q, CPD93, Dil93, GS10, TCO91, VMBK89]. Execution-based [Dil91]. Executive [RF96]. Executives [ZAdlP97, BB95]. exemples [Sch86b]. expansion [CHR86]. Exper. [Ano87i, Ano87h, Ano87m]. Experience [Ard87, GTB91, RZP^{+}88, Ton98, JS90, Sei89, YTL^{+}95, Ano87i]. experienced [HP83, vdL84]. Experiences [Bis90, BBP^{+}84, Bre96, MGK91, SKL88]. Experiment [OCM^{+}84, LL86, MGK91]. Experimental [Lun89]. Expert [War86, CHLY12, Chu96]. experts [vdL84]. exploring [Zen13]. expressions [Ano82c]. Extended [CU91, DM87]. Extending [Hol96, Rom00]. Extension [IEE96, MAAG96]. Extensions [ISO96, ISO99a, IEE99b].

F [Ano93b, Ano93c]. Fachzeitschrift [Ano11]. Facilities [Moo95, BASS96, GR88, TG80, vdLN81]. facility [BJS93]. Factories [HM87, MH87]. Failure [CW91]. Far [EMN98]. Fast [ST86, RSC93]. Faster [Bow53]. Fault [CG91, DPCC96, GMAA97, KU87, LCS91, MAAG96, Ano85b, Ano87k, Ano93e, HSLG92, RW00]. Fault-Tolerant [GMAA97, KU87, MAAG96, Ano87k]. faults [Eva95]. Features [BLB96, Bro97, BW03a, BW03b, Bro05, CS98, Dav87, GMB93, Con88, GS10, Hal83, Sri94, WHD86]. featuring [ACM94b]. February [Ano86d, BU84]. Feuer [Alb85]. Field [Ska94b]. fight [Bro81]. File [MP90]. filling [Fel93]. final [FM87, MMH88]. financial [Ada82]. Findlay [Ped88]. Finite [DMM88, HP97]. firms [Ano84b]. First [Ano83d, Lau96, OCM^{+}84, Tou94, WMS^{+}89, Ano87e, BMO92, RZP^{+}88, AFI72]. Firth [Lee92]. fixed [BW97]. FL [Mos86, USE89]. flexibility [BM87]. flexible [PW92, Ref90]. flight [RT00]. Floating [Fig00]. Floating-Point [Fig00]. Florida [IEE86b]. Flow [CHR^{+}02, LXC03, CCD90]. Footnote [NB84]. force [Alb05]. Forgotten [Aus82]. Formal [ERB12, GMP90, BBH80, BØ80, CJ92, GR80, Hem90, MdMSA93, OB80]. formally [Ram89]. format [San89a]. Fort [USE89]. FORTRAN [Bro89a, Bro89d, Bro89b, Bro89c, Gra88, Mor81, Rel89b, Rel89a, Sch82, FSO89, Gli96, Lev89, Met85, SW83, Wil87]. Forum [CWG^{+}06, Ano89b]. forward [RRS^{+}97]. Foundation [ACM91b]. Fourier [RSC93]. Fourth [IEE89]. FPGA [KD08]. från [Ska95, Ska02]. Framework [CKK87, Fig00, SKW^{+}86, CB96, JPMAB00, LZLX04, SC88, VK88, Kro98b]. France [CW91, Gau93b, KK09, RS03]. Frank [Aus11]. Frankfurt [Tou96]. Frankfurt/Main [Tou96]. Frasca [Por01]. Frasca-Spada [Por01]. Fred [Mee92]. Free [Kro98b]. Freedom [Rai92]. French [Ano86i, Car97, Cha85, Le 82, U. 82]. FRG [Ano87u]. friendly [Hen88]. Fritz [Wic84b]. Fritzson [Ano87s]. Frontières [Gau93b]. full [MMHS87]. Fully [Fig00]. Function [Tan90]. functional [Fra97]. functions [ISO94a, ISO94b, ISO98c]. fundamental [Ano92]. funding [Ano86i]. Funds [Bre96]. Fungus [CW91]. Future [Sof85, Met85]. fuzzy [CC94].

G [Ano82f, Ano85b, Ano87l, Ano87j, Ano87u, Mos86, YLT93]. Gaia [VK88]. Gaithersburg [IEE89]. GANIL [LDD^{+}94]. Ganzinger [Jan80]. Gargaro [Ano87n]. Gehani [Alb85]. Gender [Ano11]. General [Sei89, Ada82, BST98, Dil90a, RSC93]. Generating [DLGF05, FK96b]. Generation [AB88, Bel91, NB84, GN93, GN97, Hei96, YYB^{+}21, vR83]. Generator [DHGR92, CVL84, SHLR80]. Generators [DSd92]. Generic [ISO94a, Tan90, DD87, ISO94b, ISO98b, ISO98c]. genericity [DO02]. generics [Bra84, EHMO91]. Genese [CW91]. Geneva [AK07]. genuinely [BJS93]. genuinely-lazy [BJS93]. Georgia [Ano90b, Ano05, USE86a]. Germany [PK00, Tou96, Bla02]. gets [Ano93b, Twi83]. Getting [Fus90, Orm86]. Giant [Aus82]. Gielen [Ano93b]. GKS [ISO88, Air85, FHK88]. Glassman [RSC93]. global [Lun90]. Glory [CW91]. glossary [Cra00]. GNARL [GB94, OBM96]. GNAT [CGS94, Erd02, MGM^{+}02, Smy97, VGdlP01]. GNAT/ORK [VGdlP01]. GNU [CGS94, GB94, Smy97]. GNU-ADA94 [CGS94]. Go [EMN98]. Goblin [Ano93b]. Goddard [Bro89a, Bro89d, Bro89b, Bro89c]. Goldsack [Ano82c, Ano83c, Tug84, Wim83a, Tug83]. Goldstein [Ano86g]. Gonzalez [IEE86a]. Good [Ano88b, Wic88]. Goos [Ano87u]. Gordian [Kro98b]. Government [Ano93g, Ano96, Ano97c, CW91]. GPIB [Ano86g]. Graduated [Weg80b, Ano81c, Weg79]. Grady [Wal84a]. Grammar [U^{+}82, Wet81, Mac83]. grammars [Hem90]. Graphical [ISO88, ISO94c, d’O86]. Graphics [Kro98b, FM89, ISO88, ISO90, ISO94c, NM91]. graphics-assisted [FM89]. graphs [Mos90]. Great [RAH^{+}01]. Greece [Chr91]. GreenTree [Kro98b]. groff [RAH^{+}01]. groomed [Ano93b]. Ground [Lau96]. Group [ACM94b, Swa11, Kro98b]. Guards [BFC00]. Guidance [MS98]. Guide [ISO00, CBSW17, Daw88, Pyl81, Rel89b, Rel89a, S^{+}85]. guided [Gil86]. Guidelines [AH^{+}97, SMD95, Orm86].

H [All84, Ano82d, Ano84c, Her87, Mer84, Wic84b, Wim83b]. Habermann [vdL84]. Hall [Alb85, All84, Ano81c, Ano83c, Ano85c, Aug95, Mee92, Mer84, Pay93, Ped88, Tug84, Wim83a, Wim83b, You82a, CW91]. halt [Bro81]. Hampshire [Ano88d]. Hand [CWG^{+}06]. Handbook [Hor82, Kor11]. Handles [CWG^{+}06]. handling [Ano87m, Rom97]. Handook [RAH^{+}01]. Hannalei [Ano03]. Hard [Ano95c, GRGG98, HRGG98, MGDH02, Ano93e, BBWF95, ZLZ^{+}96]. hardback [Pay93, Mea87]. hardcover [Sec88]. Hardware [Kro98a, Kro98b, WA02, Sca91, Tem94]. hardware-software [Tem94]. Harness [Gli96]. Harry [Ano84d]. HaRTS [ZLZ^{+}96]. Harvard [Ano48]. Harwood [Ano82g]. hash [TC04]. heat [GV94]. Heidelberg [Ano87u]. Held [Ano90b]. Help [BM91, Ano83a]. Helping [Bur88]. Hemel [Pay93]. Hempstead [Pay93]. Henry [IEE86a]. Heppenheimer [CW91]. Herman [Aus82]. hesitant [Ano87e]. Heterogeneous [Kem96]. hiding [RC94]. Hierarchical [DAA96, ISO90, BM87, PP87]. High [Ano83e, BDR98, CW91, DPC95, Fig00, Lun91, Lut98, MS98, Sam81, Yeu97, Air85, Dav87, Eas83, Hal83, ISO00, KWK05, LHF94, Smy97, SC97, TM98, Whi89]. High-End [Lut98]. High-Level [Fig00, Sam81, Air85, Dav87, Hal83, LHF94]. high-quality [Smy97]. High-speed [Ano83e]. High-tech [CW91]. Highley [Hum92]. highly [Bor95]. Hill [Her87]. Hilton [ACM93c, ACM94b, Ano93f]. Historic [JL11]. History [ACM93a, BG96, FSJ00, HHW08, Por01]. Holiday [Ano02]. Hollerith [Aus82]. Holocaust [Bla02]. HOOD [Ano93d, Ano95c, DAA96, Hei96, MO94]. HOPL [ACM93a]. HOPL-II [ACM93a]. Horwood [Ano83b, Nie86]. Hotel [ACM96, Ano03, Ano05, IEE86b, Ano04]. Hotel-Atlanta [Ano05, Ano04]. Hotel-Atlanta/Buckhead [Ano05]. House [BFC00]. Houston [Ano02]. Houston/NASA [Ano02]. HRT [Ano95c, DAA96]. HRT-HOOD [DAA96, Ano95c]. HTML [NF96]. Hugues [Aus11]. Huijsman [Ano87t]. hybrid [Gra88, Rub82]. hybridized [SS22].

i860 [Sil92b]. IAda [DBF92]. Ian [Bud88]. iAPX [PCH^{+}82a, PCH^{+}82b]. iAPX-432 [PCH^{+}82a, PCH^{+}82b]. iAPX432 [vR83]. IBM [Ano87o, Bla02, GBO87]. IC [Kro98b]. ichbiah [Lee92]. Ideas [CW91]. Identification [ST86, GR80, Jan80]. IEC [IEE99b, TDBP01, TDB^{+}06, Ame95b, Ame95a, ISO90, ISO93, II94, ISO94a, ISO94b, ISO94c, ISO95a, ISO95c, ISO95b, ISO96, ISO98a, ISO98b, ISO98c, ISO99a, ISO99b, ISO99c, ISO00, ISO01, ISO07, ISO12, T^{+}00]. IEE [Ano82c, Ano82d, Ano82e, Ano82g, Ano82b]. IEEE [Ano86g, Ano88d, IEE99a, IEE99b, Wic84b, Ano85b, Ano87j, Ano87e, Ano87q, Ano87f, Ano87k, Ano87n, Ano87w, EGC02, Fig00]. IEEE/ANSI [IEE99b]. II [ACM93a, BG96, Mar95]. IIA [Mar95]. IKBS [Ano86i]. illustrated [Ano84d, SH89]. Illustrating [PCBE96]. Illustré [Sch86b]. Immaturity [CWG^{+}06]. Impact [Mag17, Har84]. implantation [Cha85]. Implement [SG91, MdMSA93]. Implementation [DHGR92, Fra97, KU87, Li95, OBM96, PCBE96, Ram99, RRS^{+}97, WS80, Ano87k, Bel80, BBH80, Cha85, CL90, CMM85, GZ87, GR80, vJK87, MT82, MB86, MGK91, PM07, SMBT90, TG80, vv84]. implementation-oriented [BBH80]. Implementations [ERB12, Kro98b, Bri84, Car96, CKS83]. implemented [Hal83]. Implementing [Ano93e, BG95, EP85, GMB93, GGP97, GRGG98, GS10, HRGG98, KP90, WB97, YT90, ZRdlP01, Ano82g, Ano86g, Ano87j]. implementor [Whi81b]. Implications [War86, MMH88, Tel84]. impredicative [BIM93]. improved [Bak88]. In-line [Wil87]. Including [Fra97, Geh84a, ISO98b]. Incremental [Bro84, vMAW93, Ano84c, HNVW91]. Independent [IEE99a]. index [Ano84d]. India [Ano86i]. Industrial [SMD95, DH80, Tel84, YYB^{+}21]. Industry [Ano93g, Ano96, Ano97c, Hei88]. inference [CL90]. influence [Ano87g, Fai07]. influences [GST01]. informal [BYY87]. Informatik [Ano88c]. Information [Ame95b, Ano87s, Ano89a, Aus82, Bre96, CW91, IEE92a, IEE96, IEE99a, IEE99b, ISO88, ISO90, ISO94a, ISO94c, ISO95b, ISO96, ISO99c, ISO00, ISO01, ISO07, ISO12, Ame95a, CH02, IEE92b, II94, RC94, ISO94b, ISO95a, ISO95c, ISO98a, ISO98b, ISO98c, ISO99a, ISO99b]. information-hiding [RC94]. informatique [CW91]. Infrared [ZGMK07]. INGRES [Rel89b, Rel89a]. INGRES/ [Rel89a]. INGRES/EQUEL [Rel89b]. inheritance [AR96]. initial [GKB86]. injection [GV94]. Inn [Ano88d, Ano02]. Inner [SvA^{+}98]. Innocence [CW91]. Innovation [BFC00, CWG^{+}06, CW91]. Input [Ros91, Air85, Wil87]. Input/Output [Ros91, Wil87]. instances [Bra84]. Institut [Ano88c]. instruction [BBP^{+}84, JT98, Mur91]. Instruments [Har49, Har84, Hor82]. Int. [Ano87p, Ano87r, Ano87d]. IntegrAda [NU89]. integral [Sch99]. Integrated [Som89, Ano93c]. Integrating [GVIV12, Gro92, HT96, WJS^{+}00, CP96]. Integration [TN92, Chu96, FHK88, Sag87]. Integrity [BDR98, IEE89, MS98, Yeu97, ISO00, KWK05, SC97]. Intel [Lut98, Sil92b]. Intelligence [Ano87x, Wal85]. intelligent [FW96]. intended [Rom98]. Inter [NC90]. Inter-processor [NC90]. Interactive [HL01, ISO90, RAH^{+}01, CWG^{+}06]. interconnections [BEPP87]. interest [Swa11]. interfaccia [Tes81]. Interface [IEE92a, IEE96, IEE99a, IEE99b, ISO99b, Obe88, RH02, Tes81, Bak88, IEE92b, ISO99a]. Interfaces [Cel96, IEE92a, IEE96, IEE99a, IEE99b, ISO99a, Wal84b, IEE92b, II94, IEE99a, Ano86h]. Interfacing [ISO94c, MB86, Ano86c]. Intermediate [SW83, BG84, G^{+}83]. International [AK07, Alv89, Ano85c, Ano88d, Ano01, Ano02, Ano03, Ano04, Ano05, Ano06, Asp98, BF85, Bar87c, BS02, BP12, Chr91, CS01, Gau93b, Gic09, GdlP99, HB97, Hei88, IEE86a, IEE86b, KCGO86, vK92, KV08, KK09, LS04, Mer84, Obe94, PH06, PK00, Rai92, RV10, RV11, RS03, Str96, TDB^{+}06, Taf87, Tou94, Tou96, Tug84, VW05, Wim83a, Wim83b, You82a, ACM87, Sme85, Swa11, T^{+}00, TDBP01]. Internet [Taf96a, Taf96b]. Interoperability [CK96, TO98]. interpretacja [Bie85b]. interpreter [DFS^{+}80, Rub82]. interrupt [Ano87m]. introducing [NJ05]. Introduction [Bar87b, Hil94, Pri84, Weg80b, You83, AI85, Ano81c, CB94, Geh83, Geh84a, Geh89, Led81, Led83, Nai89, Nie86, Sai85, Weg79]. invariants [MSS89]. inventor [BV07]. Invitation [Kat82, Kat84, Ano84d]. Invoked [GKPT96]. IONA [Kro98b]. IOS [Bus96]. IP [CB09]. IP-LSSVM [CB09]. IPSE [MB86]. IR [II94]. IR-MA-1363-4 [II94]. Ireland [USE87]. Iriondo [Ano93c]. ISBN [Ano83b, Ano83c, Ano84d, Aug95, Bus96, Her85, Her87, Mea87, Mea88, Mee92, Mer84, Pay93, Ped88, Por01, Sec88, Tug84, Wol08]. ISE [Kro98b]. ISO [Ame95b, Ame95a, IEE99b, TDBP01, TDB^{+}06, BFC00, II94, T^{+}00]. ISO/IEC [IEE99b, TDBP01, TDB^{+}06, ISO90, ISO93, II94, ISO94a, ISO94b, ISO94c, ISO95a, ISO95c, ISO95b, ISO96, ISO98a, ISO98b, ISO98c, ISO99a, ISO99b, ISO99c, ISO00, ISO01, ISO07, ISO12, T^{+}00]. Isolation [Dil91]. Issue [Ano82a, JT98]. Issues [Fra97, GM89, GMB93, VM87, WA02, Ano87q, Ano87n, Bar87c, BHM^{+}82, Sme85]. Italian [BV07, May83, Tes81]. Italy [HM87, KV08, MH87]. Iterative [KT96]. IV [HSWZ94].

J [Ano82c, Ano82b, Ano83b, Ano84c, Ano86d, Ano86g, Ano86h, Ano87h, Ano87m, Ano87f, Ano87k, Ano87u, Ano87t, Ano90a, Ano93b, Ano93e, CW91, Hoo92, Lee92, Mea88, Mos86, Nie86, Tug83, Wic84a, Wim83a, YLT93]. J. [Ano84c, Ano86e, Ano86f, Ano86g, Ano86c, Ano87l, Ano87o, Ano88a, Ano90a, Ano93a]. Jacobs [Ano93d]. Janice [Ano85b]. January [Ano48, Ano95b, USE85b, USE86b]. Japan [AFI72, AFI72]. Jardine [Por01]. Java [Ano97b, Bal97, Bro97, BW03a, BW03b, Bro05, BW01, BW04, CW04, Car97, EMN98, Int96, KWK05, Lam03, MH97, NMH^{+}02, WN97, Wil06a]. JavaBeans [Kro98b, Lut98]. Jazyk [Ano89e]. Jennings [CW91]. Jerome [Aus11]. Johann [CW91]. John [Ano97a, Aus11]. Joint [Tel84]. Jointly [Ano48]. Jones [Ano86d]. Journal [Bee97]. JOVIAL [Sch82]. Jr [Ano84d]. JSD [YT90]. July [ACM93c, ACM94b, Ano86c, Ano87g, Ano87n, Ano93f, Wal84b, Kat82]. July/August [Ano86c]. June [ACM84, ACM93c, ACM94a, ACM94b, AK07, Alv89, Ano82f, Ano86e, Ano86f, Ano86g, Ano86b, Ano87p, Ano87r, Ano87d, Ano93f, Asp98, BS02, BP12, Gau93b, GdlP99, HB97, Hei88, IEE89, II94, vK92, KV08, KK09, LS04, PH06, PK00, RV10, RV11, RS03, Str96, Tel84, USE86a, VW05, Ano82a]. Just [Sam86].

Karam [YLT93]. Katwijk [Ano87h, Ano87t]. Katzan [Ano84d]. KBSE [BBCS96]. Keeffe [Mos86]. Kernel [DHGR92, ISO88, RH01]. Kernels [ZRdlP01]. Key [RCM12, Hun85, WHD86]. Keynote [BBWF95]. keywords [Eas83]. KL [Kro98b]. Knight [Ano87k]. Knowledge [EMB^{+}99, HT96, ZRC91]. knowledge-based [ZRC91]. KNVVT [Ano89e]. Konover [IEE86b]. Kudos [SvA^{+}98]. Kuhn [CW91].

L [Ano82b, Ano85b, Ano86g, Ano86h, Ano87n, Ano88a, Bus96, DR96]. Laboratory [Ano48, MA89, Ano86g]. Lack [CWG^{+}06]. Lake [Ano02]. langage [Le 82, U. 82]. Language [ACM80, Ada83, Ame83, Ano79b, Ano22, Bar08, BBG^{+}81, BW96, CT94a, CH80, CS91, DX99, ECM97, GC84, IEE92a, IEE96, IEE99a, IEE99b, ISO94c, ISO95b, ISO99a, Lam02, Lee92, Mag17, OC08, Sam86, Ske82, TDB^{+}06, Tes81, Tok01, Uni83, Uni81, U. 82, WA02, WWF87, WHD86, Whe81, Ame95a, Ano80b, Ano83f, Ano83g, Ano86c, Ano87w, Ano89d, Ano91, Ano92, Ano95a, Bar94, Bar97, BYY87, BK95, BST98, BBP^{+}84, Bre80, Bro81, BG84, BR86, Coh81, Coh86, Con88, DG80, Dub85, DBF92, EL87, Ein90, EP85, Fis78, Fre82, Geh84a, GR88, GG82, G^{+}83, Hil83, IEE92b, ISO88, ISO90, ISO93, II94, ISO98a, ISO99c, ISO00, IKBW^{+}79, I^{+}86, JYCM94, JKC89, vJK87, KBL80, Le 82, Lee82, LvLS84, L^{+}87, MT82, Mah81, May83, McG83, MMHS87, Mit87, Nic80, OZC11, ØK99, Ped88]. language [Pyl81, RZP^{+}88, Rad90, Ree85, Rog84, Sil81, TD95, TD97, T^{+}00, TDBP01, TO98, Tou87, UA83a, UA83c, UA83b, Weg80a, Whi81a, Wic84b, You82b, ISO95a, TG80, ACM80, Swa11, Ano85c, Her87, You82a]. language-supported [BK95]. Languages [ACM93a, Ame95b, Coo96, CDC97, FG84, Fig00, ISO01, ISO07, Sam81, Wil06a, Wim83b, Alb85, All84, BG96, BU84, BW90, BW01, BW04, BW09, Dav87, Eas83, Eva84, FW96, Ghe85, Hal83, Hen81, II94, ISO94a, ISO94b, ISO95a, ISO95c, ISO95b, ISO98b, ISO98c, ISO99b, ISO99c, ISO00, ISO12, Mac84, Mer84, PP87, Rom99, Rom00, Rus87, San81, Sch88, Sch85, SMB83, Sou90, Vaj86, Whi89]. Large [Ano48, Cel96, DSd92, FT96, KCGO86, KT96, WVC^{+}01, DMM90, Gom94, NS87a, NS87b, NS87c, NS88, Pay93, Thé90]. Large-Scale [Ano48, Gom94]. Lauderdale [USE89]. law [CCO11]. lazy [BJS93]. learned [Bro89a, Bro89d, Bro89b, Bro89c, Tom89]. Learning [Jon86, NF96, FW96, Ano86d]. Lecture [Mea88]. lectures [WCK85]. Legacy [Gli96, Sca94, Sec88, Stexx, Ano87a, ZT86]. Lessons [Bro89a, Bro89d, Bro89b, Bro89c, Tom89]. Letter [MH97, WN97]. Letters [Eme95, EMB^{+}99, RAH^{+}01, SvA^{+}98, Bee94]. Leuven [CS01]. Level [Fig00, GGP97, Sam81, Air85, CCD91, CCD93, Dav87, Eas83, Hal83, LHF94, Tem86, TM98]. Leveson [Ano85b]. Levrat [Ano88a]. lexical [Eas83]. Libraries [Ano22, Bar08, DAA96, TDB^{+}06, Ano95a, Bar97, II94, TD95, TD97, T^{+}00, TDBP01, Taf87]. Library [GMAA97, Hus90, Kro98b, KRS01, Sch86a, Bri84, DD87, GB94]. License [Ano86j]. Licklider [CW91]. Life [BV07, Bas87, CW91, Sec88, Stexx, ZT86, Ano87a, San95, Wic84a]. like [All84, CT94a, Che97, JYCM94, Ree85]. LILA [CVL84]. limitations [Bou80]. Lindgren [CW91]. line [Wil87]. linguaggio [May83, Tes81]. Linking [NF96]. Linköping [Ano87s]. Linux [Bra00, RAH^{+}01]. Lion [Ano03]. Liskov [Lam03]. LISP [Sch82, War86]. Lister [Mea88]. literature [CH02]. Livelocks [CU91]. LJ [RAH^{+}01]. Ljudupptagning [Ska02]. local [Bur88, WY88]. local-area [WY88]. locking [DLP89]. Logic [PM07, KB91, YLT93]. logic-based [KB91, YLT93]. London [HB97, Mee92]. Long [Mee92]. look [Sri94]. lookahead [BS90]. Loops [SvA^{+}98]. Loosa [RAH^{+}01]. LOTOS [CJ92, MdMSA93, MGK91]. Louis [ACM97]. Lovelace [JM83]. low [Smy97, SC97]. low-cost [Smy97]. LR [BS90]. LSSVM [CB09]. Ltd [Ano83b, Nie86]. Luker [Ano88b, Wic88].

M [All84, Ano83b, Ano86e, Ano87p, Ano92, Ano93c, Hoo92, Lee92, Mea88, Mer84, Mos86, Wim83b, YLT93]. M. [Ano93b]. M2 [DHGR92]. MA [II94, vdL84]. Machine [Bak83, SS22]. Machinery [CW91, Ano48]. Machines [Bow53, Har49, FSO89, Har84]. Macro [vdLN81]. macros [San89a]. Madrid [Alv89]. Main [Tou96]. mainframe [GBO87]. mainframes [Wal85]. mainstream [Fus90]. Mainstreaming [Ano11]. Mainten. [Ano90a]. maintenance [vMAW93]. major [Rad90]. Making [BFC00]. Mallorca [LS04]. Management [Ano95d, Bre96, DNM^{+}10, Kro98b, MB96, REMC81, Ano89b, Ano89a, Ano90a, GS85, NM91, PCH^{+}82a, PCH^{+}82b]. manager [CDF^{+}83]. managers [Ano82g]. Managing [AS92, CW91, EST86, Lin93]. Manchester [Ano88d]. Manual [Ada83, Ano79a, Ano22, BBG^{+}81, Ich79, TDB^{+}06, Uni83, Uni81, U. 82, Ano80b, Ano83f, Ano83g, Geh84a, Kat82, Le 82, L^{+}87, TD95, TD97, T^{+}00, TDBP01, UA83a, UA83c, UA83b, You82b]. Manuel [U. 82, Le 82]. Mapping [Bak83, DAA96]. March [Ano82c, Ano82d, Ano82e, Ano82g, Ano82b, Ano87e, Ano87m, Ano90b]. Marcos [Ano93c]. Marina [Por01]. marriage [LC89]. Marriott [ACM96]. Marshall [Ano86g]. MaRTE [RH01]. Mascot [MMH88, FM87]. Mass [Sec88]. Massachusetts [ACM80, ACM87]. Master [BK87]. Master/Slave [BK87]. Mathematica [Kro98b]. Mathematical [WMS^{+}89, Har84]. Mathematics [CL05, Alb05]. matrices [HL93]. matrix [ISO98b]. maturity [Col93, Bus96]. May [Ano86e, Ano86f, Ano86g, Ano87h, Ano87k, Ano88d, Bar87c, Chr91, CS01, HM87, MH87, Taf87]. May/June [Ano86e, Ano86f, Ano86g]. Mayoh [Wal83]. McCormick [Aus11]. McDermid [Wic84a]. McGettrick [Ano82e]. McGlade [Ano87d]. McGraw [Her87]. McGraw-Hill [Her87]. McLean [ACM93c, ACM94a, ACM94b, Ano93f]. MD [ACM90, IEE89]. Means [Weg80b, Ano81c, Rad90, Weg79]. Mearns [Ano82d]. Measurement [BK95, BFC00]. Measurements [HW89, Kar90]. Measuring [MA89]. mechanical [HHW08, d’O86]. Mechanism [SG91, FHT86, Ref90, Sil81]. mechanisms [Hil83]. med [Ska95, Ska02]. Medema [All84, Mer84, Wim83b]. mediated [NJ05]. Meeting [ACM91b, ACM94b, Ada82, Whi81b]. Mellor [SAV96]. Membranes [CS91]. memorial [Kno15]. Memoriam [CW91]. memory [PCH^{+}82a, PCH^{+}82b]. Mentoring [Ano11]. Message [Kro98a, Kro98b, Ref90]. Meta [Kro98a]. Meta-CASE [Kro98a]. Method [Ano95c, BM91, DM87, BYY87, LP80, Jac85]. methodologies [FWH84]. Methodology [Ros85, WWF87, BB91, Ped88]. methods [DBDS93, Gom94, Hor82, d’O86]. Metric [Rey87, RC94]. Metric-based [Rey87]. Metrics [DS92, GKB86, Wea92, Mac84, Rey87, Rey89, RMP90, Sha88, WCW96]. metrics-driven [Rey89, RMP90]. Mexico [Ano06]. Miami [IEE86b]. Micro [Jon86, Ano86d]. microcomputer [S^{+}85]. microcomputers [GBO87, Owe87, Ano87l]. Microprocessor [Lut98, DH80, vR83]. microprocessors [Dav87]. Micros [Mit83a, Mit83b, Mit83c, Mit83d]. Microsystems [CW91]. middle [Bro80]. Middleware [Dia11, GVIV12, Kro98b]. Migrating [WVC^{+}01]. Migration [Cel96]. MIL [Ame95a]. MIL-STD-1815A [Ada83, Uni83]. MIL-STD-1815A-1983 [UA83a, UA83c]. Military [Ame83]. Mill [Lla93]. mind [HHW08]. Minimal [DRF97]. minis [Wal85]. Ministry [Kem87]. Misconception [RAH^{+}01]. mission [CB96]. Mixed [CW04, Kro98b, Ein90]. Mixed-Signal [Kro98b]. ML [TO98]. MN [Ano01]. MO [ACM97]. mode [Ano93e]. Model [EW91, FMP12, MR91, Pfl91, Ano82b, DLGF05, Dil93, HSLG92, LX04, McC92, MB86, Wot00, vv84]. model-based [HSLG92, Wot00]. Modeling [DX99, Eva95, Lut98, MZGT85, MGDH02, SBM94, San95]. Modelling [CS91, ERB12, BASS96]. Models [SAV96, Dha95, GZ87, GSX99, MG91, SM91]. Modern [CW91, Hor82, Sch85]. Modernization [Bre96, DNM^{+}10]. MODULA [All84, Ano86g, Ano86c, Ano87o, Mer84, Wim83b, Col84, Sou90, Ano86e, Ano86f, Ano86g, Ano86c, Ano87l, Ano87o, Ano88a, Bie85a, BK87, GH93, Gre86, Kin93, Pyl85, Sch86b, SH89, ST86, SMB83, WS84, Ano86f, Ano86e]. Modula-2 [Ano86g, Ano86c, Ano87o, Ano88a, Col84, Sou90, Ano86e, Ano86f, Ano86g, Ano86c, Ano87l, Ano87o, Ano88a, Bie85a, GH93, Gre86, Sch86b, SH89, ST86, WS84, Ano86f, Ano86e]. Modula-3 [Kin93]. Modular [EHMO91, GBdlHQCGB98]. Modulas [Dub85]. Modulas-1 [Dub85]. Module [ISO95a, ISO95b]. modules [BEPP87]. Molau [Ano87l]. monitor [Rey85]. Monitoring [Ger84, ZGMK07, Ano93d, Car97]. monitors [Hil92, Ano82c]. Montréal [ACM84]. Montreux [Str96]. Moore [WCK85]. Moreton [Ano82c]. Moretonhampstead [Bar87c]. Morrison [Ano87v, Bud88, Pyl88]. most [Bla02]. Mounier [CW91]. Mounier-Kuhn [CW91]. Mountain [Kro98b]. move [Ano84b]. moving [vK92]. Mudge [Ano87q]. Müller [CW91]. Multi [Kro98b, LM92, Li95, Man92, YYB^{+}21, BBH80, DLP89, Mit87, RFF92]. multi-core [YYB^{+}21]. multi-language [Mit87]. Multi-ORB [Kro98b]. multi-paradigm [RFF92]. multi-processing [BBH80]. Multi-processor [Man92]. Multi-task [YYB^{+}21]. Multi-Tasking [Li95]. Multi-Unit [LM92]. multi-user [DLP89]. Multicomputers [Geh82]. Multilingual [Hug91]. multimicroprocessor [PP87]. multiple [HT96]. multiprocessing [LvLS84]. Multiprocessor [Lun89, Lun91, USE89]. Multiprocessors [REMC81, DCM79]. MulTiSIM [CB96]. multitasking [BASS96, CCS87, Ste80]. München [Ano88c]. Munich [Hei88]. Musings [EMB^{+}99]. Mustang [Kro98a]. Mutation [DAG^{+}88]. Mutual [Bro05]. My [CW91]. Myers [Ano87e].

N [Ano87q, Ano93c]. Names [RAH^{+}01]. Naming [CWG^{+}06]. Nancy [Ano85b]. Napier [Hor82, Kno15]. Narain [Alb85]. NASA [Ano89b, Ano02, Ano89a, Bro89a, Bro89d, Bro89b, Bro89c]. NASA/Goddard [Bro89a, Bro89d, Bro89b, Bro89c]. National [Ano90b, Uni83, BFC00]. natural [BYY87, GST01]. nature [Zen13]. Naturwissenschaften [Ano11]. Nautical [Mol96]. Navy [Ano48]. Nazi [Bla02]. NCC [Ano85d]. needed [AH85]. Nerima [Ano88a]. Nesting [CWW80]. Net [Jin92, Kro98b, GSX99, MSS89, SC88, SMBT90, STMD96, SM91]. Netherlands [Bus96, vK92]. Nets [CU91, MZGT85, BASS96, SBM94, TM98]. Network [Bra00, Kro98a, Kro98b]. networks [Bur88, WY88, Woo89]. Neumann [CW91]. Neural [CS91]. News [BFC00]. Newton [DM87]. next [vR83]. Nick [Por01]. Nico [vdL84]. Nielsen [Zal92]. NJ [All84, Ano81c, Ano84d, Aug95, Ped88]. No [Ano82a, Ano82f, Ano82c, Ano82d, Ano82e, Ano82g, Ano82b, Ano84c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86b, Ano86c, Ano87l, Ano87j, Ano87o, Ano87e, Ano87q, Ano87i, Ano87h, Ano87m, Ano87f, Ano87k, Ano87g, Ano88a, Ano90a, Wal84a, Wal83]. Non [Fra97]. Non-functional [Fra97]. Nondeterminism [DS92]. Norberg [CW91]. Norman [Her87]. North [Her85]. note [Mac83, San89b, Tem94]. Notes [Mea88]. Noting [EMB^{+}99]. notion [BW90, JKC89]. Nov [Ano88c]. November [ACM97, Ano83c, Ano87o, Ano88a, Ano04, Ano05, Ano06]. November/December [Ano87o, Ano88a]. nuclear [Ano93d]. numbers [Lla93]. numeric [BDG90]. Numerical [CH80, GV94, Mor81, Sch99]. nutrition [CHLY12]. NYU [DFS^{+}80].

O2 [MB96]. OASIS [KRS01]. Object [ASM88, AS92, BBCS96, Bar96, Boo91, Bor95, Bre96, Bro97, CKK87, CK96, DX99, De 96, Hol96, KRS01, Moo95, SAV96, SG91, Sti98a, Sti98b, SD98, Ano82g, Ano97a, BB91, Bei97, BK95, CB96, CP96, JPMAB00, PP87, Rom99, Sei89, Taf82, Tou87, VK88, WJS^{+}00, Ano86f]. object-based [BK95, CB96, Taf82]. Object-Oriented [AS92, Bar96, Bre96, Bro97, CKK87, DX99, De 96, KRS01, Moo95, SAV96, SG91, Boo91, Bor95, SD98, Ano97a, Bei97, BK95, CP96, Rom99, Sei89, VK88, WJS^{+}00, Ano86f]. Object-Orientedness [Hol96]. Objects [Kem96, Ano87h, BG95, LX04, Ros92, WJS^{+}00]. objets [Lig90]. OBOSS [VGdlP01]. Observing [Nar91]. occam [MG91]. Occasion [JL11]. October [ACM82, AFI72, Ano87l, Ano01, Tou96, USE89]. Offended [RAH^{+}01]. Office [BFC00]. Ogg [RAH^{+}01]. OMG [CK96]. OMG/CORBA [CK96]. onto [Bak83]. OOD [JS90]. OODBMS [Kro93]. Open [DO02, Win99, Kor11]. operacji [Bie85b]. Operating [Shu89a, Ano84e, DSK90, Mos86, ST87, Taf82]. Operational [Lau96, HNVW91]. Operations [DM87, ISO98b]. Operator [GR80, Jan80]. Optimization [DMM88, CQG^{+}13, LZLX04]. optimizer [SKL88]. option [Lin93]. opts [Kem87]. ORACLE [FHK88]. ORB [Kro98b]. Orders [Bro81]. Ordnance [Ano48]. orientation [Ros92]. Oriented [ASM88, AS92, Bar96, Bre96, Bro97, CKK87, CK96, DX99, De 96, GTG92, KRS01, Moo95, SAV96, SG91, Sti98a, Sti98b, War86, Ano97a, BB91, Bei97, BBH80, BK95, Boo91, Bor95, CP96, JPMAB00, PP87, Rom99, Sei89, SD98, Tou87, VK88, WJS^{+}00, Ano86f]. Orientedness [Hol96]. orientée [Lig90]. Origins [CW91]. ORK [VGdlP01]. Orlando [Mos86]. orthogonal [HL93]. Osborne [Ano86e]. OSI [Kar90, CJ92, HW89]. OSI-style [Kar90, HW89]. outils [Car97]. Outmuscle [WN97]. Output [Ros91, Wil87]. overhead [HW89, Kar90]. Overloading [EL87, WS80]. Overview [CC86, CDF^{+}83, Con86, Sam81, Cra00]. Oxford [Her85, Wic88].

P [All84, Ano82b, Ano86h, Ano87j, Ano88a, Mer84, Wic88, Wim83b]. P. [Ano93d]. Package [Hil88, NB84, Tan90, Ano82c, ISO94a, ISO94b, ISO98c]. Packages [DPC95, Ros91, Fel84, GKB86, ISO98b, LP80]. Pages [Mee92, Ano84d, Ano97a, Ano98, Mea87, Wal83, Wal84a]. Palma [LS04].

Panel [BHM^{+}82]. Paper [Bar94, BBWF95]. paperback [Sec88]. Papillon [NM91]. Pappas [Ano87n]. Paradigm [BBJL92, RFF92]. Parallel [CS85, Hum92, Lun89, Lun92, MSH11, Per87, Aus11, Blu88, CG91, FSO89, Kro93, RFF92, SD98, WY88, MO94]. parameterized [Sri07]. Paris [Gau93b, GKPT96]. parser [Ree85]. parsing [BS90, vMAW93]. Part [IEE92a, IEE92b, IEE96, IEE99a, ISO88, ISO90, ISO94c, Mol96, HSWZ94, ISO98a, Sch88]. partial [Rey87, Rey89, Rey85]. Partitioned [GKPT96]. partitioning [JKC89]. PASCAL [All84, BU84, Ano86e, Ano86f, Ano86g, Ano86c, Ano87l, Ano87o, Ano88a, LP86, Mer84, Wim83b, All84, Ano83a, Col84, Rel89b, San81, AGG^{+}80, Eva84, FG84, GBO87, Mof81, Pyl85, Sch82, SMB83, Alb85, Ano87s]. PASCAL-like [All84]. Pascal/Modula [Ano88a]. Pascal/Modula-2 [Ano88a]. pass [Ano84b]. Passages [BV07]. Passaggi [BV07]. passing [Ref90]. path [Ano82c, CCS87]. Paul [Ano88b]. PC [GBO87, NU89, SvA^{+}98, Ano87o]. PCs [Ano86c]. PCTE [ECM97, ISO98a]. PDP [GBO87]. PDP-11 [GBO87]. PDP-11/40 [GBO87]. PEARL [San81]. pedagogy [MCD^{+}94]. Pennsylvania [ACM96]. Pentagon [Bro81]. Penultimate [RAH^{+}01]. Penzias [CW91]. Performance [HvKT87, Lun91, ZLZ^{+}96, BBWF95, MMH88, Ano87t]. Performance-based [ZLZ^{+}96]. Perry [vdL84]. Persch [Ano87u]. Persistence [MB96, OC96]. personal [FLP90]. Perspective [BBP^{+}84, SFGT81, Fel90, Wic84b]. Peter [Ano81c, Ano87s]. Petri [BASS96, CU91, GSX99, MZGT85, MSS89, SC88, SMBT90, STMD96, SM91, SBM94, TM98]. Petrocelli [Ano84d]. PHIGS [ISO90]. Philadelphia [ACM96]. Philosopher [BV07]. Philosophers [Bro96a]. philosophy [GST01]. Physical [Sti98a, Sti98b, SS22]. physics [Har84, Whi89]. Pickett [Hoo92]. pipeline [Fel93]. PL [Bel91, LP86, Rel89b, Rel89a, Sch82]. PL/1 [Bel91]. PL/I [Rel89b, Rel89a, Sch82]. plan [FWH84, RMP90]. Planning [FT96, Ano93a]. Plant [ZGMK07, Ano93d]. Platforms [TN92]. pleased [Ano87e]. Plotting [Ano87l]. Pluvinage [Ano93d]. PM [RMP90]. Pohlmann [Ano88c]. Point [Fig00, RSC93, RT00]. Points [De 96]. Poisson [Eva97]. policies [U. 97]. Polish [HP89]. Polling [GC84]. Polynomial [MR91]. Portable [Ard87, Tan90, Ano87i, ND94, RW00, ECM97, ISO98a]. Portal [Sch86b, SH89]. Porting [MD92, Ska94b, VGdlP01]. portions [Whi81a]. Porto [PH06]. Portugal [PH06]. poses [Ano91]. Position [Ske82, Sou90]. POSIX [EGC02, IEE99a, IEE99b, BW01, GMB93, IEE92a, IEE92b, IEE96, ISO99a, OBM96, RH02, BW04]. POSIX-Ada [RH02]. POSIX/Ada [OBM96]. potential [BBP^{+}84, Dru82]. Potsdam [PK00]. power [Ano93d]. powerful [Bla02]. Pp [Por01, Sec88, All84, Ano82f, Ano82c, Ano82d, Ano82e, Ano82g, Ano82b, Ano83b, Ano83c, Ano84c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86b, Ano86c, Ano86h, Ano87l, Ano87j, Ano87o, Ano87e, Ano87q, Ano87i, Ano87h, Ano87m, Ano87p, Ano87r, Ano87d, Ano87f, Ano87k, Ano87g, Ano87n, Ano87u, Ano87t, Ano87v, Ano88a, Ano90a, Ano93b, Ano93c, Ano93d, Ano93a, Ano93e, Aug95, Bus96, Her85, Mea88, Mos86, Pyl88, Tug84, Wic84a, Wic84b, Wic88, You82a, You82b, vdL84, Wim83a, Wim83b]. pp. [Her87, Lla93]. PQCC [Bro80]. Pract [Ano87i, Ano87h, Ano87m]. Practical [BBJL92, BS90, Bro96b, CS98, CHR^{+}02, Rom97, SMD95, Jon89, LAH94, LP80, MG91]. Practice [Car97, Ano87f, Ano87w, Ano88b, A^{+}85, Wic88]. Pragmatic [dVdV95]. Pragmatix [Kro98a]. Pratique [Car97]. Praxis [Gre86]. Precision [DPC95]. Predicting [Lun92]. Prediction [CW90]. PREDULA [RFF92]. Preemption [Nar91]. Preliminary [Ano79a, Ich79, CHR86, YTL^{+}95]. Prentice [Alb85, All84, Ano81c, Ano83c, Ano85c, Aug95, Mer84, Pay93, Ped88, Tug84, Wim83a, Wim83b, You82a]. Prentice-Hall [Alb85, All84, Ano81c, Ano83c, Ano85c, Mer84, Ped88, Tug84, You82a]. Prentice/Hall [Wim83a, Wim83b]. Preprints [ACM93a]. Preprocess [MGM^{+}02]. presentation [Hen88]. Press [Ano98, Bus96, Hoo92, Lee92, Lla93, Mea87, Mos86, Pay93, Por01, Sec88, Wic84a, Wol08]. prettyprinting [Ano87s]. Price [Mea87, Mea88, Pay93, Ped88, Wal83, Wal84a]. primary [BBP^{+}84, EP85, Rad90]. Primer [Lut98]. primitive [HL93, ISO94b]. primitives [Vaj86]. Princeton [Ano84d]. Printed [CW91]. Priority [AR96, BW90, BW97]. private [SC94]. Problem [May82, CCS87, FK93, FK96a, FK99, Hil92, Wal83]. Problems [BDG90, LM92, Wet81, Ano82f, BBP^{+}84, Rad90, Mac83]. Proc [Ano82c, Ano82d, Ano82e, Ano82g, Ano82b]. Proc.-E [Ano82c, Ano82d, Ano82e, Ano82g, Ano82b]. Procedure [LH83]. Procedures [Cel05]. Proceedings [ACM91b, AFI72, Ano48, Ano90b, IEE89, KCGO86, Sof85, Ass83, USE85b, USE85a, USE87, USE89, Alv89, Ano06, BF85, BU84, Hei88, IEE86a, Taf87, USE86a, USE86b, Whi81a, ACM80, ACM82, ACM84, ACM90, ACM91a, ACM96, ACM97, AK07, Ano87x, Ano01, Ano02, Ano03, Ano04, Ano05, Asp98, Bar87c, BP12, CS01, Dia11, Gic09, GdlP99, KV08, KK09, LS04, Obe94, PH06, PK00, RV10, RV11, RS03, Swa11, Tel84, VW05, ACM93c, ACM94b, Ano93f, BS02, Chr91, Gau93b, HM87, HB97, vK92, MH87, Str96, Tou94, Tou96, Wal84b]. Process [CT94a, EMN98, IEE89, LXC03, LX04, Ano82f, Ano93c, CCD91, CCD93, CT94b, Che97, JYCM94, PEGR80, vdB80]. process-centred [CT94b]. process-control [Ano82f]. Processes [GTB91, Ram87, Ram99]. Processing [Aus82, CS85, Ano87c, BBH80, Gal91, ISO88]. processor [Man92, NC90, Roo89, ISO99c]. Processors [Sil92b, Sag87]. produce [Wal85]. producer [Hil92]. producer/consumer [Hil92]. Product [ACM94b]. production [CCD91, CCD93]. productivity [Con88, Pfl91]. Products [Kro98a]. Professional [AH^{+}97, Daw88, Gau93a]. professionals [Mun88]. Profile [BDR98, Tok01, KWK05]. Profiling [Sch85]. Program [AB88, CW91, FW91, GMAA97, IEE92a, IEE96, IEE99a, IEE99b, ISO99a, McG82, MAAG96, OCM^{+}84, Sch86a, AI85, ACD^{+}87, Ano87w, BYY87, BM85, Bri84, FK93, FK96a, FK99, IEE92b, JKC89, PW92, San89b, Win99, Ano82e]. Programación [Bar87a]. Programmable [PM07]. Programmare [May83]. Programmation [Sch86b, U. 82]. Programmer [Bar88, ISO90, Lev89, Ano83b, Nai89, S^{+}85, SC82]. Programmers [AH^{+}97, Daw88, HP83, Joh97, Pyl81, vdL84]. Programmiermethodik [Ano87u]. Programming [ACM80, ACM93a, Ada83, Ame83, Ame95b, Ano79b, Ano80b, Ano83f, Bar82, Bar89, Bar03, Bar14, BBG^{+}81, BW96, CK96, CDC97, De 96, DG82, ECM97, FG84, Fig00, Fre82, GC84, ISO94a, ISO95b, ISO99c, ISO00, ISO01, ISO07, ISO12, KD08, KP96a, KP96b, Lam83, Lee92, Lut98, Mag17, Obe88, Per87, Rus87, Sam86, Taf96a, Taf96b, TS85, Uni83, Uni81, U. 82, UA83b, Weg79, Weg80b, WB96, WS83, Wim83b, Wol08, dVdV95, Ame95a, Alb85, All84, Ano82a, Ano83g, Ano85c, Ano86f, Ano88b, Ath82, Bar94, BG96, BM85, BMO92, Bur85, BR86, BW90, BW01, BW04, BW07, BW09, Coh81, Coo96, DG80, DBF92, Eas83, Ein90, Eva84, EP85, FLP90, Fis78, FW96, FHK88, Geh84a, Geh84b, Geh87, Ghe85, GG82, Hen81, II94, ISO98a, IKBW^{+}79, I^{+}86, Jon89, Lyo87, Mac84, Mah81]. programming [McG83, Mea87, Mer84, Mit87, Nic80, NU89, ØK99, Pyl81, RZP^{+}88, Rad90, RFF92, SH89, Sch86c, Sch85, SMB83, Sou90, ST84, SD98, Swa11, Taf82, TG80, Tel84, Tou87, UA83a, UA83c, Vaj86, WHD86, WJS^{+}00, Whi81a, Whi89, Wic84b, Wic88, You82a, You82b, ISO94b, ISO95a, ISO95c, ISO98b, ISO98c, ISO99b, May83, Ano86e, Ano81c, Ano83c, Tug83, Tug84, Wim83a]. programmirovanija [Ano89e]. Programs [Bar96, Bel97, BB98a, BAP87, BB98b, BDR98, CXZY02, CU96, DAG^{+}88, Dil90b, Dil91, FMP12, Fra97, GD84, HL85, Hol83, Jac85, Jin92, KT96, LCS91, Lun92, Mad96, MR91, VM87, WF97, Ano85b, BST98, Blu88, Car96, CWW80, Cor96, Dil90a, EOAm94, EOM95, FM89, FSO89, GN93, GMP90, GS85, Hoo85, JKC89, KSB89, KBL80, LP80, L^{+}87, MO90, Mos90, Ram89, Rey85, Rey89, Rom96, Rom97, San89a, Sen92, SM91, SBM94, TCO91, VMBK89, YYB^{+}21, YTL^{+}95, Ano87q]. progress [Wol91]. Project [Bas87, Bro96b, DSd92, FT96, Kro98b, Ano87d, CGS94, KP90, Ree85, Som89, WMS^{+}89]. Projecting [AE92]. Projects [Bau91, Sil91]. proliferation [Bro81]. promise [Ano87f]. PRONAOS [Lau96]. Proof [GD84, Ano82d, BM82, BM86]. properties [Dil90a]. proposal [BJS93]. Proposed [BBG^{+}81, Uni81, Ano80b]. Prorok [Ano93a]. prospective [Ano87e, Har84]. prospects [BBP^{+}84]. protected [BG95, LX04, WJS^{+}00]. Protection [Kro98a]. PROTOB [BB91]. Protocol [IEE99a, AR96, CJ92, DLP89, Lun90, MGK91]. prototype [MB86, NM91, ST84]. Prototypes [Bel91]. Prototyping [Dun82, REC96, LvdGvK89, Luq90, SOK92, SLM91]. provide [Ano85d]. provided [Con88]. provision [BM87]. proxies [TC04]. pseudocode [Rey87, Rey89]. psychology [GST01]. Publications [Ano88b, Bee94]. Published [Alb85, Ano85c, Bud88, Wim83a, Wim83b]. Puente [Ano93e]. Pulse [Mos86, Mos86]. Punched [CW91]. pushes [Wal85]. PVM [KP96a, KP96b]. Pyle [Ano85c, You82a]. Python [RAH^{+}01].

quadrature [DD87]. Quality [EMB^{+}99, AH^{+}97, Sch88, Smy97]. Quantitative [Dha95]. Qué [ACM84]. Queries [WMS^{+}89]. query [BST98, Kro93]. quest [SW94].

R [Ano85b, Ano86g, Ano86c, Ano87q, Ano87g, Ano87t, Ano87v, Bud88, CW91, Lee92, Pyl88, Wic84b, YLT93]. R-32 [Ano85b]. Rabdology [NR90]. Race [Fel97]. races [KSB89, MO90]. railroad [McC92]. Randell [JL11]. rapid [Ano86c, SLM91]. Rapide [Mad96]. Rasmussen [Ano87m]. Rational [Kro98a, Kro98b]. Rationale [Ano79b, Bar08, IKBW^{+}79, I^{+}86, Lee92, Ano95a, Bar97, YTL^{+}95, TG80]. ratios [CHR86]. Ravenscar [BDR98, CW04, KWK05, PV12, PV02, VGdlP01]. Re [Lin93, CH97]. Re-engineering [Lin93]. re-use [CH97]. Reaching [BB98a, BB98b]. Reactive [EW91, Ram99]. Readability [PCBE96]. Readable [Boo89]. reader [Ada10]. Reading [vdL84]. Real [ASM88, Ano95c, Ano04, Bar87c, BB95, BLB96, BW03a, BW03b, Bro05, BDR98, BW01, BW04, BW09, DPCC96, FT96, GVIV12, GTB91, GRGG98, HRGG98, Hen81, LM92, Lut98, LF90, MDPM08, MD92, MSH11, MGDH02, MS02a, Rai92, RAH^{+}01, RH01, REMC81, WMS^{+}89, Wil06b, Wol08, Zal92, ZAdlP97, ZRdlP01, Ano93b, Ano93e, Ano02, Ano03, Ano05, Aus11, BBWF95, BW90, BW07, Chu96, CMM85, Coo96, Dub85, FHK88, Gal91, Gom94, Hal83, HSLG92, HT96, ISO96, ISO98b, JM83, KSdR^{+}88, KWK05, Mac80, Mah81, NS87a, NS87b, NS87c, NS88, NC90, Roo89, Sch86c, Sch88, ST87, Thé90, Zal88, ZLZ^{+}96, Ano87m]. Real-Time [Ano95c, Ano04, Bar87c, BW03a, BW03b, Bro05, BDR98, DPCC96, FT96, GVIV12, GTB91, GRGG98, HRGG98, LM92, MDPM08, MD92, MSH11, MGDH02, REC96, RH01, Zal92, ZAdlP97, ZRdlP01, BB95, BW01, BW04, BW09, Hen81, LF90, REMC81, Wil06b, Wol08, Ano93b, Ano93e, Ano02, Ano03, Ano05, Aus11, BBWF95, BW90, BW07, Chu96, CMM85, Coo96, Dub85, Gal91, Gom94, Hal83, HSLG92, HT96, ISO96, KSdR^{+}88, KWK05, Mah81, NS87a, NS87b, NS87c, NS88, NC90, Roo89, Sch86c, Sch88, ST87, Thé90, ZLZ^{+}96, Ano87m]. Real-Tune [BLB96]. Real-World [Lut98]. realization [Ano93d]. realizatsia [Ano89e]. Realtime [DRF97, IEE96, IEE99b, Ano87c, ISO99a]. reasoning [HSLG92, Rey87, Wot00]. rebels [Bro81]. recommendations [Ano89a]. recommended [Ano87w]. Reconciling [Gal91]. Reconfigurable [LRT91]. Reconfiguration [GVIV12]. Reconnaissance [BFC00]. recording [Bar03, BW04, Ska02, Wei03]. recovery [Ano93e, RRS^{+}97]. Recycling [SvA^{+}98]. Red [Ano03]. redesignation [Ame95a]. Redirector [Kro98b]. reduce [Lun90]. reduces [Ano86b]. Reducing [ZRC91]. reduction [DBDS93, STMD96]. Reference [Ada83, Ano79a, Ano83g, Ano22, BBG^{+}81, Ich79, TDB^{+}06, Uni83, UA83c, Uni81, U. 82, You82b, Ano80b, Ano83f, Geh84a, Kat82, L^{+}87, Mac83, TD95, TD97, T^{+}00, TDBP01, UA83a, UA83b, Wet81, U. 82]. Refinement [OC08, OZC11, Rey85]. Refinements [Tok01]. Reflects [CWG^{+}06]. Regard [Sil92a]. rejuvenation [Lin93]. Related [Ano04, Gic09, Ano02, Ano03, Ano05]. Relational [Tes81]. relations [WCW96]. relationship [DLGF05]. relativistic [LN93]. relazionale [Tes81]. release [Ano82h, GV94]. released [Ano89d]. Reliab. [Ano85b]. reliability [CQG^{+}13, SW94]. reliability-based [CQG^{+}13]. Reliable [AK07, Ano04, Asp98, BP12, CS01, GdlP99, HD99, HB97, Hei96, KV08, LS04, PH06, RV10, RV11, RS03, Str96, VW05, Ano92, Ano02, Ano03, Ano05, BS02, Vig93, AK07, Asp98, BS02, BP12, CS01, GdlP99, HB97, KV08, KK09, LS04, PH06, PK00, RV10, RV11, RS03, Str96, VW05]. Remotely [GKPT96]. removing [Bou80]. Rendez [BBJL92]. Rendez-Vous [BBJL92]. Rendezvous [DS92, GR88, LXC03, Nai89, Hil92, LZLX04, WCW96, Woo89]. Rendezvouz [Ano88c]. replace [Mor81]. Replicated [PV02, WB96]. Report [Ska94b, Ton98, Ano89a, Bel80, FM87, MMH88]. reports [Ada82]. Repository [Con86]. Representation [Jin92, SW83, CH02, CPD93, HLRS80]. representations [DLGF05]. Requirement [RCM12]. Requirement-Based [RCM12]. Requirements [DHGR92, WW84, Wal91, Sch82]. Reserved [ST86]. resolution [Bel80, Rom97, Rom00]. RESOLVE [HSWZ94]. resource [DLP89, Ram87]. Resources [Ano89a, Ano90c]. restoration [RW00]. restricted [JT98]. restrictive [EL87]. Result [Eme95]. Results [Bau91, GV94, SKL88]. Retargeting [Ard87, Ano87i]. Retrieval [Fra01, SLM91]. Reusability [Ano87n]. Reusable [Hei96, LM84, Ros91, ZAdlP97, Bor95, SLM91]. Reuse [BM91, SMD95, TDB92, BK95, GW90, LAH94]. Reusing [TN92]. Reverse [CCD90, CCD91, CCD93]. Review [Alb85, All84, Ano81c, Ano82f, Ano82c, Ano82d, Ano82e, Ano82g, Ano82b, Ano83b, Ano83c, Ano84c, Ano84d, Ano85b, Ano85c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86b, Ano86c, Ano86h, Ano87l, Ano87j, Ano87o, Ano87e, Ano87q, Ano87i, Ano87h, Ano87m, Ano87p, Ano87r, Ano87d, Ano87f, Ano87k, Ano87g, Ano87n, Ano87s, Ano87u, Ano87t, Ano87v, Ano88a, Ano88b, Ano88c, Ano90a, Ano93c, Ano97a, Ano98, Aug95, Aus11, Boo89, Bud88, Bus96, Her85, Her87, Hoo92, Lee92, Lla93, Mea87, Mea88, Mee92, Mer84, Mos86, Nie86, Pay93, Ped88, Por01, Pyl88, Sec88, Tug83, Tug84, Wal83, Wal84a, Wic84a, Wic84b, Wic88, Wim83a, Wim83b, Wol08, You82a, You82b, Zal92, vdL84, BLW87, Ano87o]. Reviews [CW91, ZT86]. revised [Nie86]. Revision [Ame95a, Sch86a]. Revisiting [Mag17]. RG [Ano89e]. RG-20 [Ano89e]. Rhetorical [CWG^{+}06]. Riccardi [Ano87j]. rich [OZC11]. Rigorous [Eme95, Fig00]. Ripken [Wic84a, Jan80]. risk [Ano86b]. robot [DBF92, GG82]. robotics [Fag00]. Robots [OMÁ^{+}02]. role [ACM93c, Ano93f]. ROSE [BM91, CW91]. ROSE-Ada [BM91]. Rosen [CW91]. rotations [HL93]. routines [BDG90]. routing [TC04]. RSA [Hun85]. RSI [Kro98a]. RTS [Wil06a]. rule [CC94]. rule-based [CC94]. Rules [WS80, Xu98, Ano82d, BM82]. Run [Che92, GWA91, Hol83, Lut98, Tok01, Bak88, vv84]. Run-Time [Che92, Tok01, GWA91, Hol83, Bak88, vv84]. Runtime [GB94, GTG92, SR85a, SR85b, HLRS80].

S [Ano82c, Ano86e, Ano86c, Ano87o, Mac83, Nie86, Tug83, Wic84b, Wim83a].

Safe [DRF97, RF96, Sti98a, Sti98b, SD98]. Safety [Bro96b, IEE89, LCS91, RF96, Ros96, ZAdlP97, Ano93a, Che92, Dil90a, Sch88, Ano85b]. Safety-Critical [Ros96, ZAdlP97, Ano93a]. Saib [Wic84b]. SAMeDL [ISO95b, ISO95a]. San [Ano03, BU84, IEE86a, Ass83]. Sanden [Aug95]. Sans [Gau93b]. Santander [GdlP99]. Scale [Ano48, Gom94]. Scaling [LS82]. scene [Bar94, McG83]. Schedulability [MGDH02]. Scheduler [Hum92]. Schedulers [GGP97]. Scheduling [RH02, BW97]. schema [CMM85]. schemes [GS85, Rom96, Rom98]. Scheutz [CW91]. School [WCK85]. Science [Ano87s, Mea88, RR93, Bra89, CH02, Fag00, GST01, MMHS87, TE87]. Sciences [Por01, FSJ00]. Scientific [Ano88b, FKR86, SD98, Wic88, Vig93]. scienziato [BV07]. scopes [DO02]. Script [FHT86]. Scripting [JSV97, SvA^{+}98]. SDL [MGK91]. SE [Ano87k]. SE-13 [Ano87k]. Seattle [ACM93b]. Second [IEE86b, Obe94, Tou96, Ano98, Coh86, FW96, Her87, Nie86]. Secondary [KRS01]. Secrets [Lut98]. Section [NB84, LN93]. secure [Wal85]. Security [IEE89, SKW^{+}86, Ano91]. Selecting [Hoo92]. Selection [Fra97, Whi89]. SEMANOL [BBH80]. Semantic [ISO99b, U^{+}82]. Semantics [Yeu97, Blu88, BASS96, BG84, BIM93, KSdR^{+}88]. Semantics-Based [Yeu97]. semaphores [Hil92]. seminar [Ano95b]. Sentences [MGM^{+}02]. Separate [EST86]. September [ACM93b, Ano87l, Ano87j, Ano87s, Ano92, Ano01, IEE86a, Tou94]. September/October [Ano87l]. Sequence [NMH^{+}02]. sequential [WY88]. sequential-parallel [WY88]. Serfs [Sri07]. Series [Bus96, Lee92, Wic84a, Mea87]. Server [Bre96, GGP97, Hei96, Kro98a, SvA^{+}98, Hem90]. Server-Side [SvA^{+}98]. Services [GKPT96, OBM96]. SES [Kro98a]. Session [CG91]. sessions [BW97]. Set [DM87, DPC95, Fel97, Man92, Obe88, San89a]. SETA2 [Obe94]. SETL [DG87]. Several [Xu98]. sFr [Ano97a]. Sharing [Bra84]. Shell [CC86]. Sheraton [Ano88d]. Sheraton-Wayfarer [Ano88d]. Sherman [Ano86b]. Shlaer [SAV96]. Shlaer/Mellor [SAV96]. short [Nic80]. shortest [CCS87]. should [Ros92]. Showed [CW91]. Side [SvA^{+}98, TC04, Alb05]. SIGAda [ACM94b, Ano01, Ano02, Ano03, Ano04, Ano05, Ano06, Bee94, Gic09, ACM87, ACM91b]. SIGAda’06 [Ano06]. Signal [Kro98b, OBM96, Sil92b]. SIGPLAN [ACM80, ACM84, ACM93a, Whi81a]. Silicon [OCM^{+}84, BB95]. SIMD [CS85]. Simple [Bak83, Mac84]. Simple Com [GBdlHQCGB98]. simplifié [d’O86]. Simplifying [Bou80, d’O86]. SIMULA [BU84]. simulate [DSK90]. Simulation [BU84, HP98, Kro98a, LH83, Bru84, CB96, TM98, ZRC91, BU84, Ano87p]. Simulator [DR96, JSV97, Kro98b, Mol96, GBO87]. Simulierte [Ano88c]. Singhoff [Aus11]. single [CCS87, KP90]. singly [Ref90]. singly-buffered [Ref90]. Size [Wit90, Fra01]. Skeptical [CWG^{+}06]. Slave [BK87]. Slight [CWG^{+}06]. Small [WMS^{+}89, Eva97, Ree85, FLP90]. SMALL-Ada [FLP90]. Smart [DRF97]. Smedema [All84, Mer84, Wim83b]. Society [IEE86b, Wic84b]. Soft [Ano90a, GVIV12, Kro98b, Mea88, Wic88]. Softback [Ano83c, Ano84d]. Software [ACM91b, AK07, Ano86d, Ano87k, Ano89a, Ano95d, Ano96, Ano04, BA09, Boo83, Boo87, BP12, CKK87, CT94a, Chu96, CW90, Con86, Con88, CWG^{+}06, DS92, DPCC96, DCM79, GW90, GTB91, Gom94, HM87, HD99, IEE89, KV08, KK09, Kro98a, Kro98b, KT96, Lam03, LCS91, LRT91, LS04, LF90, MD92, MH87, PH06, RV10, RV11, San94, Sch86c, Sch88, TN92, VW05, Wal85, Wal84b, Wal91, Wea92, Weg90, WS84, Wit90, Yeu97, ACM93c, AH85, AE92, Ano86h, Ano89b, Ano93a, Ano93f, Ano02, Ano03, Ano05, Asp98, BA98, BK95, BS02, Bor95, BMO92, CG91, CH97, CS01, Dha95, Dru82, Eva97, Fai07, FK96b, FWH84, FHK88, GN93, GN97, GdlP99, GV94, HB97, Hug91, II94, JS90, JYCM94, KP90, LL86, Lin93, LC89, McC92]. software [MA89, MMHS87, Mur91, Owe89, Pfl91, PK00, Ree85, RW00, RS03, RT00, SOK92, Sca91, Sei89, SLM91, Str96, Tem94, Tom89, U. 97, ZLZ^{+}96, vMAW93, Ano87j, Ano87i, Ano87h, Ano87m, Ano87n, Kro98a, Kro98b, MA89, Ano86g, Ano87v, Aug95, Bud88, Pay93, Pyl88, Wal84a]. Software-safety [Sch88]. Sold [RAH^{+}01]. Soldier [RAH^{+}01]. Solution [CWG^{+}06, MB96, TN92, Sch99]. solutions [Rad90]. Solve [Ano82f]. solving [FK93, FK96a, FK99, May82, Wal83]. Some [De 96, FW91, Mah81, Nic80, Sme85, Sri94, WHD86]. Somerville [Pyl88]. Something [SvA^{+}98]. Sommerville [Ano87v, Bud88]. Sort [Fel97]. sound [Bar03, BW04, Ska02, Wei03]. Source [AGG^{+}80, BAP87, DSd92, Int96, RCM12, TDB92, CCS87, Kor11, TO98]. Source-to-source [AGG^{+}80]. Space [Rai92, DBDS93]. Spada [Por01]. Spain [GdlP99, LS04, RV10]. SPARC [BMM96]. sparse [CB09, DMM90]. Speaks [MH97]. Special [Ano82a, NB84, Swa11, Ano82f]. specific [CDC97, Rom96]. Specification [BW03a, BW03b, Bro05, LM84, Wal91, Ano93d, BEPP87, GR80, Sav81, ISO99b]. Specifications [Bel91, Ano82c, HNVW91, MdMSA93, OZC11, Sen92]. specify [Sil91]. Specifying [Hem90, Ano82g]. spectrum [Ano93b, Ano87f]. speed [Ano83e]. Speedup [Lun92]. spending [Ano84b]. Sponsored [Ano48]. Sporadic [GGP97]. spreads [Ano87g]. Springer [Ano86h, Ano87u, Ano97a, Mea88, You82b]. Springer-Verlag [Ano86h, Ano97a, You82b]. Springer/Compass [Ano87u]. SQL [ISO95b, ISO95a, MS02a, Rel89a]. SQL/Ada [ISO95b, ISO95a]. St [ACM97]. Stan [ZT86]. Standard [Ame83, Ano22, Bar08, BBG^{+}81, IEE99a, TDB^{+}06, Uni81, Ano80b, Ano95a, Bar97, FHK88, IEE92b, II94, Rom98, Ska95, Ska02, TD95, TD97, T^{+}00, TDBP01, Ame95a, BFC00, EGC02, Fig00, LS82, Uni83, UA83a]. Standardization [Ske82, Coh81, Sme85]. Standards [Weg90, Bar94, Lee82, McG83]. standardu [Bie85b]. started [Orm86]. State [RW00, DBDS93, OZC11]. state-rich [OZC11]. statements [CXYZ02, Per89, Wil87]. States [Ano80b, BBG^{+}81]. Static [CXZY02, DT91, MR91, Ros96, MSS89, SC88]. Station [Rai92]. Status [Boy87]. STD [Ada83, Uni83, UA83a, UA83c, Wal91, IEE99b]. Stein [Sec88, ZT86]. step [CB09, Zal88]. Steps [TS85, Ano87d]. stepwise [Rey85]. still [Ano87e, VMBK89]. Stirring [WMS^{+}89]. stochastic [Vig93, Woo89]. Stockholm [BP12, Taf87]. Stolzy [Ano85b]. storage [GS85, SLM91]. Stored [Cel05, CW91]. STRAda [BBB^{+}92]. strategic [Bla02]. strategy [DCM79, MMHS87]. Stratford [Ano83b]. Stratford-Collins [Ano83b]. Strawberry [Lla93]. streams [BJS93]. Stronger [SvA^{+}98]. Strongly [TO98, BU84]. Strongly-typed [TO98]. Structure [LXC03, MB86]. Structured [Ano95c, Ath82, LM84]. Structures [Car97, Kem96, Lig90, Ano92, Ano97a, Bei97, Boo87, Fel84, Fel90, FW91, LP86, LAH94, MP90, Sch88, Sil91, Wei03, Car97]. Structuring [ACD^{+}87, Air85]. student [Ree85]. students [All84, FW96]. Studies [Bus96, Bra89, PW92]. Study [AB88, Bau91, Boo89, NMH^{+}02, Ano86c, Blu88, CHR86, Dun82, FWH84, GKB86, Rom98, SW94, Vaj86, VKT91, YYB^{+}21]. studying [FLP90]. style [AH^{+}97, HW89, Kar90]. subsequent [FW96]. Subset [Mit83a, Mit83b, Mit83c, Mit83d, HL83, LS82]. subsets [Rad90]. subsystems [Boo87]. succeed [Orm86]. success [Ano87d]. Successfully [CP96]. suggested [TG80]. Suitability [BK87]. Suite [Kro98a, Tan90, GN97, YTL^{+}95]. Summary [Whi81b]. Summation [CS91]. Summer [ACM91b, ACM94b, USE86a]. Sun [CW91]. Sunburst [CW91]. supercomputer [Ano86i, BM85]. Supercomputing [Sti98a, Sti98b]. supervision [Ano93d]. supervisory [Ano93c]. Support [Ano86k, GTG92, Lam83, Obe88, Yeu97, Ano86i, Bis85, Boy87, Lyo87, MdMSA93, NU89, Rey89, Rom99, Roo89, SS22, Som89, Taf82, Tel84, Twi83, Wic84a]. support-vector [SS22]. supported [BK95]. Supporting [ASM88, CW04, Fig00, PCH^{+}82a, PCH^{+}82b, Shu89a, RW00]. supports [Sag87]. survey [Coo96, Ghe85]. SuSe [RAH^{+}01]. Sustainable [Dia11]. Sweden [Ano87s, Asp98, BP12]. Swedish [Ska02]. Switching [Bra00]. Switzerland [AK07, Str96]. Symbolic [BB98a, BB98b, Dil90b, Dil91, LF90, CPD93]. Symposium [ACM80, ACM84, ACM91b, ACM93c, ACM94b, Ano93f, Bow53, Obe94, Tou94, Tou96, ACM94a, Ano48, Ano94, Ano89b]. symposium-forum [Ano89b]. Symposium/Summer [ACM91b]. Synchronization [GTG92, Hil92, ØK99, KP90, Sil81]. Synchronous [BW97, YYB^{+}21]. Syntax [Xu98, Ano87s, CVL84, Hen88]. Syst [Ano87r]. Syst. [Ano87p, Ano87d]. System [BBB^{+}92, Bre96, Bun96, DNM^{+}10, EST86, GD84, GTG92, GWA91, HP98, Hol83, Hoo92, IEE92a, IEE96, IEE99a, IEE99b, ISO88, ISO90, Kro98b, Lam83, LRT91, Lun89, MZGT85, OMÁ^{+}02, SR85a, SR85b, Shu89a, Whe81, Ano86e, Ano89d, Ano93c, Ano93d, Aug95, BBWF95, Bak88, BM86, BR86, CHLY12, CMM85, DG80, GN93, IEE92b, II94, Kro93, LvdGvK89, LDD^{+}94, Mos86, PP87, PW92, Rey87, Rey89, Sav80, SS22, ST87, Taf82, Thé90, HM87, ISO99a, Kro98b, MH87, Mos86].

Systematic [WF97]. Systems [Ano95c, Ano04, Bal97, Bau91, BBJL92, Cel96, CSM96, Cur91, DPCC96, DT91, EW91, GVIV12, GTB91, GBdlHQCGB98, Gro92, HL01, IEE89, KP96a, KP96b, KU87, LM84, MDPM08, MGDH02, MS98, REC96, Sma96, USE89, War86, WB96, ZAdlP97, Zal92, Ano83c, Ano84e, Ano87o, Ano92, Ano93e, Ano02, Ano03, Ano05, BB91, BM87, BW01, BW04, BW09, Bus96, Chu96, Coo96, DMM90, DG82, GN97, Gom94, HSLG92, HT96, HNVW91, HW89, HvKT87, HW87, ISO88, Kar90, Mac80, MO94, NS87a, NS87b, NS87c, NS88, NC90, San94, San95, TM98, Tug83, Tug84, Vaj86, WY88, Wal85, Wil06b, Wim83a, Win99, Kro98a, ISO00, Ano87k, Ano87t, Pay93].

T [Ano82c, Ano82g, Ano87j, Ano87q, Ano87n, DRF97]. T-Smart [DRF97]. tâches [Cha85]. Tables [WMS^{+}89]. Tactic [OC08, OZC11]. take [Ros92]. Tales [CW91]. target [Sca91]. Targeting [Gar86]. targets [Hei96]. Task [DRF97, Hum92, Jin92, MZGT85, Mol96, Nar91, REMC81, WBP97, CMM85, YYB^{+}21]. Task-Safe [DRF97]. Tasking [BMM96, BDR98, CU91, CU96, DHGR92, Dil90b, Dil91, ERB12, Ger84, HL85, Li95, Lun91, Ano87t, BTM89, BLW87, Cor96, Dil90a, Dil93, DBDS93, GSX99, HvKT87, KP90, Lun90, Roo89, SC88, Sha88, SMBT90, STMD96, TG80, Mea88]. Tasks [LH83, Shu89a, BM82, BM86, DSK90, Hem90, Kro93, Cha85, Ano82d]. Taylor [CW91]. TCOL [Bro80]. teach [Fag00]. Teaching [CDC97, Fel90, Lam03, MCD^{+}94, Fel84, JS90, LAH94, TE87, Tom89]. TEC [Sof85, Tel84]. tech [CW91]. Tech. [Ano82c, Ano82b]. Techn. [Ano82d, Ano82e, Ano82g]. Technical [ISO01, TDB^{+}06, T^{+}00, TDBP01]. Technik [Ano11]. Techniques [Bro96b, Wat97, ISO94c, Pay93, WCK85]. Technische [Ano88c]. technological [LC89]. technologie [CW91]. Technologies [AK07, Ano04, BP12, HD99, KV08, KK09, LS04, PH06, RV10, RV11, VW05, Ano02, Ano03, Ano05, Asp98, BS02, CS01, GdlP99, HB97, PK00, RS03, Str96, Gic09, Kro98b]. Technology [Ame95b, Ano90b, BBCS96, CW91, GM89, IEE92a, IEE96, IEE99a, IEE99b, ISO01, ISO07, Kro98b, Lut98, MDPM08, RV10, VGdlP01, VW05, Ame95a, CH02, GST01, IEE92b, ISO90, II94, ISO94a, ISO94b, ISO94c, ISO95a, ISO95c, ISO95b, ISO96, ISO98a, ISO98b, ISO98c, ISO99a, ISO99b, ISO99c, ISO00, ISO12, Win99]. Teleoperated [OMÁ^{+}02]. Television [HL01]. Temporal [CS91, KB91, YLT93]. Ten [Ton98, Bar94]. Tenth [ACM93c, Ano93f]. Tercentenary [Hor82, Kno15]. Termination [WBP97]. Test [Tan90, FK96b, GN93, GN97, IEE86a]. Testing [Bar96, Car96, DAG^{+}88, IEE86a, Mad96, Wat97, WF97, Weg90]. Texas [Ano02, IEE86a, USE85b]. textbook [Mos86]. Their [CU96, BEPP87, Car96, Har84, II94, Rad90]. Theodorsen [Sch99]. theory [WCK85]. Thermal [Kro98b]. Third [Ano87x, Ano88d, Tel84]. Thought [Bow53]. Threads [GMB93]. Threat [Kin93, Ano91]. three [Ano87o]. Throughput [Woo89]. Time [ASM88, Ano95c, Ano04, Bar87c, BW03a, BW03b, Bro05, BDR98, Che92, DPCC96, FT96, GVIV12, GTB91, GRGG98, HRGG98, LM92, MDPM08, MR91, MD92, MSH11, MGDH02, Rai92, REC96, RH01, Rus87, Tok01, WMS^{+}89, Zal92, ZAdlP97, ZRdlP01, Ano93b, Ano93e, Ano02, Ano03, Ano05, Aus11, BBWF95, Bak88, BB95, BW90, BW01, BW04, BW07, BW09, Chu96, CMM85, Coo96, Dub85, FHK88, Gal91, Gom94, GWA91, GS10, Hal83, HSLG92, HT96, Hen81, Hol83, ISO96, KSdR^{+}88, KWK05, LZLX04, LF90, Mac80, Mah81, NS87a, NS87b, NS87c, NS88, NC90, REMC81, Roo89, Sch86c, Sch88, ST87, Thé90, Wil06b, Wol08, Zal88, ZLZ^{+}96, vv84, Ano87m]. Timely [GVIV12]. Timing [Cor96, VM87, GS10, Ano87q]. TM [Bro97, Hei96]. Toetenel [Ano87t]. Tokyo [AFI72]. tolerance [RW00]. Tolerant [DPCC96, GMAA97, KU87, MAAG96, Ano87k, CG91]. Tomlinson [Mos86]. tongues [Bro81]. Too [RAH^{+}01, Wic84c, EL87]. Tool [BM91, ECM97, ISO98a, Int96, Kro98a, Kro98b, MdMSA93, Man92, Ros96, Ton98, ASM88, FM89, FW96, LvdGvK89, MB86, ND94, Rey85, Rey89, SLM91, YTL^{+}95]. Toole [Lla93]. toolkit [SMBT90]. Tools [Kro98a, Kro98b, Obe94, Ros85, Sch86a, Wal84b, Yeu97, Ano86h, BYY87, Boo87, Car97, Kor11, Taf87, vMAW93, Ano86d]. Toolset [Bel97, DRF97]. Toulouse [RS03]. tour [Gil86]. TR [ISO96, ISO00]. Tracing [EGC02]. Tradeoffs [PCBE96]. traditional [CP96]. Traffic [DNM^{+}10, CC94]. training [Ano80a, Fai07]. Trans [Ano85b, Ano87q, Ano87k]. Transaction [SG91]. transactional [JPMAB00]. Transfer [BW03a, BW03b, BG95]. transfert [CW91]. Transform [RSC93]. Transformation [BBB^{+}92, Ros85, BM85, GST01]. transformations [DG87]. Transforming [OCM^{+}84]. Transition [FT96, Bro89a, Bro89d, Bro89b, Bro89c]. Transitions [Bro84, Ano84c]. Translating [HL83, SAV96, Ste80]. Translation [BAP87, Kro93, VMBK89, AGG^{+}80, Luq90, TO98, Wil87, MT82]. translator [DFS^{+}80, Smy97, Ano88a]. TransLib [JPMAB00]. Transparent [PV02]. transputer [MO94]. transputer-based [MO94]. Tree [Ano04, Ano05, SW83, DG87]. Trees [LCS91, Ano85b]. Trenches [Gre86, Bie85a]. TRI [ACM96, ACM97, ACM91a, ACM93b, Ano93g, Ano95d, Ano96, Ano97c]. TRI-Ada [ACM96, Ano93g, Ano95d, Ano96, Ano97c, ACM91a, ACM93b]. TRI-Ada’97 [ACM97]. Truncated [DM87]. Tscharmer [Ano88a]. Tune [BLB96]. Turing [CBSW17]. Tutorial [CB94, Gil86, YT90, Wic84b]. Twelfth [KCGO86]. Two [Bri84, GZ87, Lam03, CB09, GTB91]. two-step [CB09]. TX [IEE86a, USE85a]. Type [Bel91, Bel80, ISO98b]. typed [BU84, TO98]. types [Ano87h, Fel84, GZ87, HT96, Hil94, HLRS80, ISO98b, NM91, Shu89b, vv84]. Tyson [ACM94b].

U [Ano93a]. U.S. [Wal85]. UCSD [Ano88a]. Uhl [Ano87u]. UIMS [ND94]. UK [Ano87v, Ano95b, Bar87c, HB97, Lee92, Mee92, Nie86, Pyl88, RV11, VW05, Wic88, Ano85d, Twi83]. UML [OMÁ^{+}02]. Undergraduate [AH85, Mur91, Owe87, TE87]. undergraduates [Tem86]. Understandable [BFC00]. Understanding [Shu88, Shu89b, Zen13]. unification [Bro81]. Unions [HP97]. Unit [LM92, OCM^{+}84, WF97]. Unit-Testing [WF97]. United [BBG^{+}81, Ano80b]. universal [Bro81]. universe [Zen13]. Universität [Ano88c]. universities [Fel93]. University [Ano48, Ano87s, Ano98, Hoo92, Lee92, Mea87, Por01, Smy97, Wic84a, Wol08]. UNIX [EST86, Gal91, Geh87, Lam83, NB84, SR85a, SR85b, Bur88, Che97, Gar86, SHLR80]. UNIX-based [SR85a, SR85b, Che97]. upper [Tem86]. upper-level [Tem86]. Uppsala [Asp98]. Ural [Ano87o]. Urquhart [Ano87k]. USA [AFI72, All84, Aug95, Lla93, Mos86, vdL84, Ano01, Ano02, Ano03, Ano05, Ano06, IEE89, USE85b, USE86a, USE86b]. Usage [Kro98b]. Use [BBJL92, CS98, CJ92, DR96, DM87, EW91, KU87, MS98, MGM^{+}02, NF96, Ros96, WY88, Ano87k, BF85, Bar87b, Bur88, CH97, CH80, DG80, FHK88, ISO00, LL86, Sav81]. used [Rad90]. User [Bee97, Hen88, DLP89, IEE86a, Bee97]. User-friendly [Hen88]. users [All84, Ano87e, Bur88]. Using [ACM87, Ano84e, Ano04, Bru84, CKK87, CU91, CS91, DSK90, DT91, DAG^{+}88, Dil90b, DMM88, DSd92, DH80, DBDS93, Fag00, GTB91, GRGG98, Gro92, HRGG98, HL01, Jac85, LH83, LCS91, LM84, NF96, OMÁ^{+}02, Owe87, REC96, Sil91, SW83, SG91, ZGMK07, Ano85b, Ano86f, Ano93c, Ano93d, Ano97a, Ano02, Ano03, Ano05, Bei97, Blu88, BASS96, DLGF05, FK96b, GMB93, HSLG92, Hei96, Hil94, Hil92, HNVW91, Hug91, KP90, LvdGvK89, LAH94, McG82, MO94, MSS89, SS22, TM98, Wot00].

V [Tug83, Wim83a]. V60 [ST87]. VA [ACM93c, Ano93f, ACM94a]. Valencia [RV10]. Valid [Ano87r]. validated [Ano86a]. Validating [FM89]. Validation [GV94, Lee82, Sil92a, Ano85d, Goo80, Mar95, ZRC91]. Validations [Ton98]. Valley [Lla93]. variable [Bis85]. Variables [LH83, HLRS80]. VAX [Con88, Mit87, SHLR80]. VDA [Jac85]. VDM [CKK87]. vector [Blu88, ISO98b, SS22]. Venice [KV08]. Verification [Dil90b, Dil91, HNVW91, LCS91, Per89, Ram87, Ano82e, EHMO91, FHT86, GMP90, McG82]. verified [Ram89]. Verifying [Dil90a, Hoo85, OCM^{+}84, LP80]. Verlag [Ano86h, Ano97a, You82b]. Verrand [Her85]. version [Ano84d, II94, RSC93]. versus [Sil92b]. Very [KCGO86, LHF94]. VHDL [KD08, Wot00]. via [CCO11, Ref90, TO98]. Victorian [Bra89]. Video [ZGMK07]. Vienna [BS02, Jac85]. View [De 96, Gre86, Bie85a, RT00]. viii [Wal83]. VIIIA [McG83]. Virginia [ACM82, ACM94b]. Virtual [Bak83, Taf82]. Visibility [Cel96]. visual [Dil93]. Vit [Ano93a]. vita [BV07]. VLDB [KCGO86]. Voelcker [Ano87f]. Vol [Ano82a, Ano82f, Ano82c, Ano82d, Ano82e, Ano82g, Ano82b, Ano84c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86b, Ano86c, Ano87l, Ano87j, Ano87o, Ano87e, Ano87q, Ano87i, Ano87h, Ano87m, Ano87f, Ano87k, Ano87g, Ano88a, Ano90a]. volume [Bus96, Kno15]. Volz [Ano87q]. Vous [BBJL92]. vs [Bie85a, Gre86, Lam03, War86].

W [Ano82g, Ano86d, Ano87e, Ano87m, Ano87t, Aus11, Pay93]. WA [ACM93b]. Wacky [RAH^{+}01]. WADAS [ACM94b, ACM91b, ACM93c, Ano94]. Wallis [Ano82b, Ano86h]. Wand [Mos86]. Washington [ACM91b, ACM93c, ACM94b, Ano93f, Ano94]. Watt [Ped88]. Way [CW91]. Wayfarer [Ano88d]. Wearing [Hoo92]. web [TC04]. Wegmann [Ano86f]. Wegner [Ano81c]. wejścia [Bie85b]. wejścia/wejścia [Bie85b]. Wellings [Ano90a, Ano98, Mea88, Wol08]. Werner [Ano88c]. Wesley [Ano87v, Bud88, Pyl88, vdL84]. West [CW91, Alb05, Wol91]. Wetherell [Mac83]. while [San89b]. White [Kro98b]. whole [Ano93b]. Wibu [Kro98a]. Wichmann [Ped88]. width [JT98]. Wiener [Ano86c]. Wiley [Wal83]. will [Ano84b, Ano85d]. William [Ped88]. Wilson [Ano87g]. Windows [Kro98b]. Winter [Ass83, USE85b, USE85a, USE86b]. within [MB86]. without [Sca91]. Wokingham [Ano87v, Pyl88]. Women [CL05]. Woodger [Lee92]. Words [ST86]. Working [ACM94b]. Workshop [Bar87c, Dia11, HM87, MH87, USE89, Wal84b, Sof85]. Workstation [Sag87]. World [Lut98, MS02a, CW91]. Write [Cel05]. Writing [Boo89].

X [Sec88, Aug95]. xiii [Mos86]. xiv [Por01]. xix [Sec88, Wal84a]. XMDS [ACD^{+}87].

year [Ada82, Fel90]. Yearbook [Mee92, Lof93]. Years [Ton98, Bar94]. Yielding [LM84]. York [Ano97a, Ano98, Her87, Smy97, VW05, WMS^{+}89]. Young [Nie86].

Z [Sen92]. zaawansowanych [HP89]. Zandvoort [vK92]. Zeit [Ano88c].

References

Ausnit:1985:AP

[A^{+}85] Christine N. Ausnit et al. Ada in practice. Springer books on professional computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 0-387-96182-8 (New York), 3-540-96182-8 (Berlin). xv + 195 pp. LCCN QA76.73.A35 A287 1985.

Allen:1988:PGA

[AB88] Pat Allen and Alan Burns. Program generation for Ada — a case study. Software—Practice and Experience, 18(12):1125– 1138, December 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Aspray:1990:CBC

[ABCK^{+}90] William Aspray, Allan G. Bromley, Martin Campbell-Kelly, Paul E. Ceruzzi, and Michael R. Williams, editors. Computing before computers. Iowa State University Press, Ames, IA, USA, 1990. ISBN 0-8138-0047-1. ix + 266 pp. LCCN QA76

.C5834 1990. URL http://www.computerhistory.org/
collections/DocumentArchive/Documents/Books/Computing\
%20Before\%20Computers/CBC.html.

Ancona:1987:SDP

[ACD^{+}87] M. Ancona, A. Clematis, G. Dodero, V. Filippone, and V. Gianuzzi. Structuring a distributed program: the XMDS approach. Microprocessing and Microprogramming, 21(1–5):287–293, August 1987. CODEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0165607487900524.

ACM:1980:PAS

[ACM80] ACM, editor. Proceedings of the ACM-SIGPLAN Symposium on the Ada Programming Language, Boston, Massachusetts, December 9–11, 1980: The Ada programming language, volume 15(11) of ACM SIGPLAN Notices. ACM Press, New York, NY 10036, USA, November 1980. CODEN SINODQ. ISBN 0-89791-030-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.73.A35 .A82 1980. ACM order no. 82500.

ACM:1982:PAC

[ACM82] ACM, editor. Proceedings of the AdaTEC Conference on Ada, Arlington, Virginia, October 6–8, 1982. ACM Press, New York, NY 10036, USA, October 1982. ISBN 0-89791-087-7 (paperback). LCCN QA76.73.A35 A35 1982. ACM order no. 825821.

ACM:1984:PSS

[ACM84] ACM, editor. Proceedings of the SIGPLAN 84 Symposium on Compiler Construction. Montréal, Qué., Canada, 17–22 June, 1984, volume 19(6) of ACM SIGPLAN Notices. ACM Press, New York, NY 10036, USA, June 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1987:UAA

[ACM87] ACM, editor. Using Ada: ACM SIGAda international conference, Boston, Massachusetts, December 8–11, 1987. ACM Press, New York, NY 10036, USA, 1987. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987. At head of title: Ada letters. Proceedings of the 1987 ACM SIGAda International Conference on the Ada Programming Language.

ACM:1990:PDB

[ACM90] ACM, editor. Proceedings: December 3–7, 1990, Baltimore Convention Center, Baltimore, MD. ACM Press, New York, NY 10036, USA, 1990. ISBN 0-89791-409-0. LCCN QA76.73.A35.

ACM:1991:TP

[ACM91a] ACM, editor. TRI-Ada ’91 Proceedings. ACM Press, New York, NY 10036, USA, 1991. ISBN 0-89791-445-7. LCCN ????

ACM:1991:WSM

[ACM91b] ACM, editor. WADAS ’91/Summer SIGAda Meeting. Eighth Annual Washington Ada Symposium/Summer SIGAda Meeting Software: Foundation for Competitiveness. Proceedings. ACM Press, New York, NY 10036, USA, 1991. ISBN 0-89791-393-0. LCCN ????

ACM:1993:ASH

[ACM93a] ACM, editor. ACM SIGPLAN HOPL-II. 2nd ACM SIGPLAN History of Programming Languages Conference (Preprints), volume 28(3) of ACM SIGPLAN Notices. ACM Press, New York, NY 10036, USA, March 1993. CODEN SINODQ. ISBN ???? ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN ????

ACM:1993:TCS

[ACM93b] ACM, editor. TRI-Ada ’93: Conference — September 1993, Seattle, WA, TRIADA -proceedings- 1993. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-621-2. LCCN ???? ACM Order No. 825930.

ACM:1993:WTA

[ACM93c] ACM, editor. WADAS ’93. Tenth Annual Washington Ada Symposium: Ada’s role in software engineering, June 28– July 1, 1993, McLean Hilton, McLean, VA: proceedings. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-609-3. LCCN QA76.73.A16W37 1993.

ACM:1994:AAA

[ACM94a] ACM, editor. ADA in applications: 11th Annual ADA symposium — June 1994, McLean, VA. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-684-0. LCCN ????

ACM:1994:AAW

[ACM94b] ACM, editor. Ada in applications: WADAS ’94: eleventh Annual Washington Ada Symposium & Summer ACM SIGAda Meeting: featuring Working Group Product Development and Delivery: June 27 – July 1, 1994, McLean Hilton, Tyson’s Corner, McLean, Virginia: proceedings. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-684-0. LCCN ????

ACM:1996:PTA

[ACM96] ACM, editor. Proceedings / TRI-Ada ’96, Philadelphia Marriott Hotel, Philadelphia, Pennsylvania, December 3–7, 1996. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-808-8. LCCN QA 76.73 A35 T75 1996.

ACM:1997:PTA

[ACM97] ACM, editor. Proceedings of the TRI-Ada’97 Conference, November 9–13, 1997, St. Louis, MO. ACM Press, New York, NY 10036, USA, 1997. ISBN 0-89791-981-5. LCCN QA 76.73

A35 T75 1997. URL http://www.acm.org/pubs/contents/

proceedings/ada/269629/. Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

ALUL:1982:AFR

[Ada82] Ada Language UK Ltd. 1st annual and financial reports for the year ended 31st December 1981, with agenda of the annual general meeting. Ada Language UK Ltd., ????, 1982. LCCN ????

AJPO:1983:RMA

[Ada83] Ada Joint Program Office. Reference Manual for the Ada Programming Language, ANSI/MIL-STD-1815A. Washington, DC, 1983.

Adamson:2010:CR

[Ada10] Glenn Adamson, editor. The craft reader. Berg Publishers, Oxford, UK, 2010. ISBN 1-84788-304-4 (hardcover), 1-84788-303-6 (paperback). xii + 641 pp. LCCN https://www.math.utah.edu/pub/tex/bib/adabooks.bib; TT149 .C733 2010.

Agresti:1992:PSD

[AE92] W. W. Agresti and W. M. Evanco. Projecting software defects from analyzing Ada designs. IEEE Transactions on Software Engineering, 18(11):988–997, November 1992. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=177368.

AFIPS:1972:FUJ

[AFI72] AFIPS, editor. First USA–Japan Computer Conference, Proceedings: October 3–5, 1972, Tokyo, Japan. AFIPS Press, Montvale, NJ, USA, 1972. LCCN QA76 .U2 1972.

Albrecht:1980:SST

[AGG^{+}80] Paul F. Albrecht, Philip E. Garrison, Susan L. Graham, Robert H. Hyerle, Patricia Ip, and Bernd Krieg-Brückner. Source-to-source translation: Ada to Pascal and Pascal to Ada. ACM SIGPLAN Notices, 15(11):183–193, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Agrawal:1985:UCN

[AH85] Jagdish C. Agrawal and Alka R. Harriger. Undergraduate courses needed in Ada and software engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 17(1):266–281, March 1985. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ausnit-Hood:1997:AQS

[AH^{+}97] Christine Ausnit-Hood et al. Ada 95 quality and style: Guidelines for Professional Programmers, volume 1344 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNCSD9. ISBN 3-540-63823-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 292 pp. LCCN QA76.73.A35

A216 1997. URL http://link.springer-ny.com/link/
service/series/0558/tocs/t1344.htm;   http://www.
springerlink.com/content/978-3-540-63823-0;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=
0302-9743&volume=1344.

Amoroso:1985:AIP

[AI85] Serafino Amoroso and Giorgio Ingargiola. Ada: an introduction to program design and coding. Pitman Publishing Ltd., London, UK, 1985. ISBN 0-273-01818-3. ix + 353 pp. LCCN QA76.73.A35 A46 1985.

Airchinnigh:1985:CHL

[Air85] Micheal Macan Airchinnigh. The CONTEXT: a highlevel structuring concept for GKS input. Computers and Graphics, 9(3):211–220, ???? 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0097849385900482.

Abdennadher:2007:RST

[AK07] Nabil Abdennadher and Fabrice Kordon, editors. Reliable Software Technologies – Ada Europe 2007: 12th Ada-Europe International Conference on Reliable Software Technologies, Geneva, Switzerland, June 25–29, 2007. Proceedings, volume 4498 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNCSD9. ISBN 3-540-73229-2 (paperback), 3-540-73230-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.R44 A33 2007. URL http://w2.vu.edu.au/

library/EBookSearch/files/Springer_E-Books_Online.
pdf; http://www.springerlink.com/content/978-3-540-
73230-3.

Albury:1985:BRB

[Alb85] David Albury. Book review: Comparing and assessing programming languages — ADA, C, and PASCAL: edited by Alan Feuer and Narain Gehani. Published by Prentice-Hall. 271pp. £16.10. Data Processing, 27(1):50, January/February 1985. CODEN ???? ISSN 0011-684X (print), 1878-3058 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0011684X85903521.

Albert:2005:HAH

[Alb05] Nancy E. Albert. A^{3} and his algebra: how a boy from Chicago’s West Side became a force in American mathematics. iUniverse, New York, NY, USA, 2005. ISBN 0-595-32817-2. xiv + 349 pp. LCCN 01.50.

Allison:1984:BRP

[All84] Colin Allison. Book review: PASCAL-like languages for users and students: C. H. Smedema, P. Medema and M. Boasson The programming languages PASCAL, MODULA, CHILL and ADA, Prentice-Hall, Englewood Cliffs, NJ, USA (1983) £9.95 pp 154. Microprocessors and Microsystems, 8(10):539, December 1984. CODEN MIMID5. ISSN 0141-9331 (print),

1872-9436 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0141933184900620.

Alvarez:1989:ADC

[Alv89] Angel (Alvarez Rodríguez) Alvarez, editor. Ada — the design choice: proceedings of the Ada-Europe International Conference, Madrid 13–15 June 1989. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-38130-4. LCCN QA76.76.D47.

ANSI:1983:MSA

[Ame83] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Military Standard Ada Programming Language, February 17 1983. Also MIL-STD-1815A. See also [IKBW^{+}79, Ich79].

ANSI:1995:ANSa

[Ame95a] American National Standards Institute. American National Standard for information technology: programming language ADA: ANSI/ISO/IEC 8652-1995: Revision and redesignation of ANSI/MIL 1815A-1983. Number 119-1 in FIPS PUB. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, revised edition, April 10, 1995. various pp.

ANSI:1995:AIIb

[Ame95b] American National Standards Institute. ANSI/ISO/IEC 8652-1995: Information Technology — Programming Languages — Ada. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1995. ISBN ???? ???? pp. LCCN ???? US$18.00. URL http://webstore.ansi.

org/ansidocstore/product.asp?sku=ANSI%2FISO%2FIEC+
8652%2D1995.

Anonymous:1948:PSL

[Ano48] Anonymous, editor. Proceedings of a Symposium on Large-Scale Digital Calculating Machinery: Jointly Sponsored by The Navy Department Bureau of Ordnance and Harvard University at The Computation Laboratory 7–10 January 1947, volume 16 of The Annals of the Computation Laboratory of Harvard University. Harvard University Press, Cambridge, MA, USA, 1948. LCCN QA75 .S96 1947.

Anonymous:1979:PAR

[Ano79a] Anonymous. Preliminary Ada reference manual. ACM SIG-PLAN Notices, 14(6A), June 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). The final standard is [Ame83].

Anonymous:1979:RDA

[Ano79b] Anonymous. Rationale for the design of the Ada programming language. ACM SIGPLAN Notices, 14(6B), June 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). The final standard is [Ame83].

Anonymous:1980:AT

[Ano80a] Anonymous. ADA training. Microprocessors and Microsystems, 4(7):292, September 1980. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0141933180903476.

Anonymous:1980:PLA

[Ano80b] Anonymous. The Programming language Ada: reference manual: proposed standard document, United States Department of Defense, volume 106 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. ISBN 0-387-10693-6, 3-540-10693-6. x + 243 pp. LCCN QA76.73.A35

P76. URL http://www.springerlink.com/openurl.asp?
genre=issue&issn=0302-9743&volume=106.

Anonymous:1981:AC

[Ano81a] Anonymous. ADA committee. Microprocessors and Microsystems, 5(5):221, June 1981. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0141933181901411.

Anonymous:1981:AAE

[Ano81b] Anonymous. ADA to appear at Euromicro ’81. Microprocessors and Microsystems, 5(6):285, July/August 1981. CO-DEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0141933181906438.

Anonymous:1981:BRB

[Ano81c] Anonymous. Book review: Programming with Ada: An introduction by means of graduated examples: Peter Wegner (Prentice-Hall, Englewood Cliffs, NJ). Computer Languages, 6(1):62–63, ???? 1981. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0096055181900618.

Anonymous:1982:APC

[Ano82a] Anonymous. Ada-programming in the ’80s: Comput. special issue vol 14 no 6 (June 1981). Microprocessors and Microsystems, 6(1):39, January/February 1982. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0141933182903143.

Anonymous:1982:ARBf

[Ano82b] Anonymous. Article review: Ada model arithmetic: costs and benefits: Wallis, P. J. L. IEE Proc.-E Comput. Dig. Tech. Vol 129 No 2 (March 1982) pp 75–80. Microprocessors and Microsystems, 6(9):497–498, November 1982. CO-DEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0141933182905154.

Anonymous:1982:ARBb

[Ano82c] Anonymous. Article review: Ada package specifications: path expressions and monitors: Goldsack, S. J. and Moreton, T. IEE Proc.-E Comput. Dig. Tech. Vol 129 No 2 (March 1982) pp 49–54. Microprocessors and Microsystems, 6(9):497, November 1982. CODEN MIMID5. ISSN 0141-9331 (print),

1872-9436 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0141933182905117.

Anonymous:1982:ARBc

[Ano82d] Anonymous. Article review: Axioms and proof rules for Ada tasks: Barringer, H. and Mearns, I. IEE Proc.-E Comput. Dig. Techn. Vol 129 No 2 (March 1982) pp 38–48. Microprocessors and Microsystems, 6(9):497, November 1982. CO-DEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0141933182905105.

Anonymous:1982:ARBd

[Ano82e] Anonymous. Article review: Program verification and Ada: McGettrick, A. D. IEE Proc.-E Comput. Dig. Techn. Vol 129 No 2 (March 1982) pp 55–62. Microprocessors and Microsystems, 6(9):497, November 1982. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0141933182905129.

Anonymous:1982:ARBa

[Ano82f] Anonymous. Article review: Solve process-control problems with ADA’s special capabilities: Booch, G. EDN Vol 27 No 13 (23 June 1982) pp 143–152. Microprocessors and Microsystems, 6(7):390, September 1982. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0141933182904562.

Anonymous:1982:ARBe

[Ano82g] Anonymous. Article review: Specifying and implementing object managers in Ada: Harwood, W. T. IEE Proc.-E Comput. Dig. Techn. Vol 129 No 2 (March 1982) pp 70–74. Microprocessors and Microsystems, 6(9):497, November 1982. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0141933182905142.

Anonymous:1982:NUR

[Ano82h] Anonymous. New US release — educating ADA. Microprocessors and Microsystems, 6(5):263, June 1982. CO-DEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0141933182901065.

Anonymous:1983:APH

[Ano83a] Anonymous. ADA/PASCAL help. Microprocessors and Microsystems, 7(1):41, January/February 1983. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0141933183902302.

Anonymous:1983:BRBa

[Ano83b] Anonymous. Book review: ADA: a programmer’s conversion course: M. J. Stratford-Collins Ellis Horwood Ltd, Chichester, 1982, 192 pp. 28.25 / £14.50 ISBN 0-85312-416-7. Advances in Engineering Software (1978), 5(2):119, April 1983. CODEN AESODT. ISSN 0141-1195 (print), 1878-3066 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0141119583901833.

Anonymous:1983:BRBb

[Ano83c] Anonymous. Book review: Programming embedded systems with ADA: Downes and Goldsack Softback, November 1982, Prentice-Hall, pp, £10.95, ISBN 0-13-730010-7. Advances in Engineering Software (1978), 5(3):172, July 1983. CO-DEN AESODT. ISSN 0965-9978 (print), 0141-1195 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0141119583901286.

Anonymous:1983:FCA

[Ano83d] Anonymous. First commercial Ada compiler. Data Processing, 25(7):44, September 1983. CODEN ???? ISSN 0011-684X (print), 1878-3058 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0011684X83902587.

Anonymous:1983:HSC

[Ano83e] Anonymous. High-speed computing devices, volume 4 of Charles Babbage Institute reprint series for the history of computing. Tomash Publishers, Los Angeles, CA, USA, 1983. ISBN 0-938228-02-1. xli + 451 pp. LCCN TK7885 .H53 1983. New introduction by Arnold A. Cohen.

Anonymous:1983:PLA

[Ano83f] Anonymous. The Programming language Ada: reference manual, volume 155 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. ISBN 3-540-12328-8 (Berlin), 0-387-12328-8 (New York). ix + 330 pp. LCCN QA76 L471 155; QA/76/ .73/A35/P762/1983 IN; QA76.73A35. American National Standards Institute, Inc. ANSI/MIL-STD-1815 A-1983, approved 17 February 1983.

Anonymous:1983:RMA

[Ano83g] Anonymous. Reference manual for the Ada programming language. Silicon Press, 25 Beverly Road, Summit, NJ 07901, USA, 1983. ISBN 0-9615336-6-8 (paperback). various pp. LCCN QA76.73.A35 R374 1983.

Anonymous:1984:AB

[Ano84a] Anonymous. Ada bureau. Data Processing, 26(8):54, October 1984. CODEN ???? ISSN 0011-684X (print), 1878-

3058 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0011684X84902272.

Anonymous:1984:ASW

[Ano84b] Anonymous. ADA spending ‘will pass $1000M by 1986’ as commercial firms move in. Microprocessors and Microsystems, 8(5):268, June 1984. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0141933184904927.

Anonymous:1984:ARB

[Ano84c] Anonymous. Article review: Transitions to Ada: an incremental approach: Brown, D. H. J., Comput. J. Vol 27 No 1 (1984) pp 37–41. Microprocessors and Microsystems, 8(4):201–202, May 1984. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0141933184903351.

Anonymous:1984:BRB

[Ano84d] Anonymous. Book review: Invitation to ADA: Condensed version by Harry Katzan, Jr.; 1984; Petrocelli Books, Inc.; Princeton, NJ; 173 pages; softback; 14.95; illustrated, index; ISBN 0-89433-239-2. Computers and Standards, 3(2): 126, ???? 1984. CODEN COSTEZ. ISSN 0167-8051 (print),

1878-3090 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0167805184900457.

Anonymous:1984:UAC

[Ano84e] Anonymous. Using ADA for compilers and operating systems. Data Processing, 26(6):??, July/August 1984. CO-DEN ???? ISSN 0011-684X (print), 1878-3058 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0011684X84902740.

Anonymous:1985:ACA

[Ano85a] Anonymous. ADA compilers are aimed at 8086 and 80186. Microprocessors and Microsystems, 9(10):517, December 1985. CODEN MIMID5. ISSN 0141-9331 (print), 1872-

9436 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0141933185901899.

Anonymous:1985:ARB

[Ano85b] Anonymous. Article review: Safety analysis of Ada programs using fault trees: Nancy G. Leveson and Janice L. Stolzy. IEEE Trans. Reliab. R-32 (5), 479 (1983). Microelectronics and Reliability, 25(3):584, ???? 1985. CO-DEN MCRLAS. ISSN 0026-2714 (print), 1872-941X (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0026271485902409.

Anonymous:1985:BRB

[Ano85c] Anonymous. Book review: The ADA programming language: I. C. Pyle. Published by Prentice-Hall International. 341pp. £18.95. Data Processing, 27(6):51, July/August 1985. CODEN ???? ISSN 0011-684X (print), 1878-3058 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0011684X85902862.

Anonymous:1985:NWP

[Ano85d] Anonymous. NCC will provide ADA validation for UK defence. Microprocessors and Microsystems, 9(9):471, November 1985. CODEN MIMID5. ISSN 0141-9331 (print), 1872-

9436 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0141933185901127.

Anonymous:1986:AV

[Ano86a] Anonymous. ADA validated for 1750A. Microprocessors and Microsystems, 10(7):413, September 1986. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0141933186904023.

Anonymous:1986:ARBe

[Ano86b] Anonymous. Article review: Attention to basics reduces risk in Ada compiler choice: Sherman, B. Comput. Des. Vol 25 No 12 (15 June 1986) pp 87–90. Microprocessors and Microsystems, 10(8):452, October 1986. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0141933186902371.

Anonymous:1986:ARBf

[Ano86c] Anonymous. Article review: Interfacing assembly language to Modula-2 — a case study: creating rapid displays for PCs: Wiener, R. S. J. Pascal, Ada & Modula-2 Vol 5 No 4 (July/August 1986) pp 21–26. Microprocessors and Microsystems, 10(8):452, October 1986. CO-DEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0141933186902383.

Anonymous:1986:ARBa

[Ano86d] Anonymous. Article review: Learning ADA on a micro: Jones, D-W Dr. Dobb’s J. Software Tools Vol 11 No 2 (February 1986) pp 42–58. Microprocessors and Microsystems, 10(6):342, July/August 1986. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0141933186903029.

Anonymous:1986:ARBb

[Ano86e] Anonymous. Article review: Modula-2: an alternative to C for system programming: Djavaheri, M. and Osborne, S. J.

Pascal, Ada, Modula-2 Vol 5 No 3 (May/June 1986) pp 47– 52. Microprocessors and Microsystems, 10(7):399, September 1986. CODEN MIMID5. ISSN 0141-9331 (print), 1872-

9436 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0141933186903704.

Anonymous:1986:ARBc

[Ano86f] Anonymous. Article review: Object-oriented programming using Modula-2: Wegmann, A. J. Pascal, Ada, Modula-2 Vol 5 No 3 (May/June 1986) pp 5–17. Microprocessors and Microsystems, 10(7):399, September 1986. CO-DEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0141933186903753.

Anonymous:1986:ARBd

[Ano86g] Anonymous. Article review: Software engineering in Modula-2: implementing the GPIB (IEEE 488) in a laboratory: Marshall, J. L. and Goldstein, R. D., J. Pascal, Ada, Modula-2 Vol 5 No 3 (May/June 1986) pp 28–46. Microprocessors and Microsystems, 10(7):399, September 1986. CO-DEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0141933186903728.

Anonymous:1986:BRB

[Ano86h] Anonymous. Book review: Ada software tools interfaces: P. J. L. Wallis, Springer-Verlag (1984) pp 163. Computer Aided Design, 18(2):111, March 1986. CODEN CAIDA5. ISSN 0010-4485 (print), 1879-2685 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010448586901879.

Anonymous:1986:BII

[Ano86i] Anonymous. BT in India — IKBS funding — support for ADA — Anglo–French supercomputer. Data Processing, 28(2):58–63, March 1986. CODEN ???? ISSN 0011-684X (print), 1878-3058 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0011684X86900997.

Anonymous:1986:LA

[Ano86j] Anonymous. License for Ada. Computer Aided Design, 18(2): 121, March 1986. CODEN CAIDA5. ISSN 0010-4485 (print),

1879-2685 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010448586902046.

Anonymous:1986:SAD

[Ano86k] Anonymous. Support for ADA in database. Data Processing, 28(7):387, September 1986. CODEN ???? ISSN 0011-684X (print), 1878-3058 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0011684X86901735.

Anonymous:1987:ALL

[Ano87a] Anonymous. Ada, a life and a legacy. MIT Press, Cambridge, MA, USA, 1987. ISBN 0-262-19242-X (print), 0-262-19242-X. LCCN ????

Anonymous:1987:AD

[Ano87b] Anonymous. ADA development for 32000. Microprocessors and Microsystems, 11(9):467, November 1987. CO-DEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0141933187900056.

Anonymous:1987:ARP

[Ano87c] Anonymous. ADA for realtime processing. Information and Software Technology, 29(8):457, October 1987. CO-DEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0950584987903016.

Anonymous:1987:ARBk

[Ano87d] Anonymous. Article review: 10 steps to Ada project success: McGlade, D. Syst. Int. (June 1987) pp 37–38. Microprocessors and Microsystems, 11(7):393, September 1987. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0141933187905515.

Anonymous:1987:ARBd

[Ano87e] Anonymous. Article review: Ada: first users — pleased; prospective users — still hesitant: Myers, W. IEEE Computer Vol 20 No 3 (March 1987) pp 68–73. Microprocessors and Microsystems, 11(5):280, June 1987. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0141933187902778.

Anonymous:1987:ARBl

[Ano87f] Anonymous. Article review: Ada: from promise to practice?: Voelcker, J. IEEE Spectrum Vol 24 No 4 (April 1987) pp 44– 49. Microprocessors and Microsystems, 11(7):393, September 1987. CODEN MIMID5. ISSN 0141-9331 (print), 1872-

9436 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0141933187905539.

Anonymous:1987:ARBn

[Ano87g] Anonymous. Article review: Ada’s influence spreads through defense community: Wilson, R. Comput. Des. Vol 26 No 13 (July 1987) pp 91–99. Microprocessors and Microsystems, 11(8):450, October 1987. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0141933187900792.

Anonymous:1987:ARBg

[Ano87h] Anonymous. Article review: Addressing types and objects in Ada: Van Katwijk, J. Software Pract. Exper. Vol 17 No 5 (May 1987) pp 319–343. Microprocessors and Microsystems, 11(6):341, July/August 1987. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0141933187905023.

Anonymous:1987:ARBf

[Ano87i] Anonymous. Article review: Experience acquiring and retargeting a portable Ada compiler: Ardo, A. Software Pract. Exper. Vol 17 No 4 (April 1987) pp 291–307. Microprocessors and Microsystems, 11(6):340, July/August 1987. CO-DEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0141933187904972.

Anonymous:1987:ARBb

[Ano87j] Anonymous. Article review: Implementing Ada exceptions: Baker, T. P. and Riccardi, G. A. IEEE Software Vol 3 No 5 (September 1986) pp 42–51. Microprocessors and Microsystems, 11(2):114, March 1987. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0141933187902420.

Anonymous:1987:ARBm

[Ano87k] Anonymous. Article review: On the implementation and use of Ada on fault-tolerant distributed systems: Knight, J. C. and Urquhart, J. I. A. IEEE Trans. Software Eng. Vol SE-13 No 5 (May 1987) pp 553–563. Microprocessors and Microsystems, 11(7):393, September 1987. CO-DEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0141933187905503.

Anonymous:1987:ARBa

[Ano87l] Anonymous. Article review: Plotting curves and data with microcomputers: Molau, G. E. J. Pascal, Ada, Modula-2 Vol 5 No 5 (September/October 1986) pp 5–45. Microprocessors and Microsystems, 11(2):112, March 1987. CO-DEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0141933187902286.

Anonymous:1987:ARBh

[Ano87m] Anonymous. Article review: Real-time interrupt handling in Ada: Rasmussen, J. B. and Appelbe, W. Software Pract. Exper. Vol 17 No 3 (March 1987) pp 197–213. Microprocessors and Microsystems, 11(6):341, July/August 1987. CO-DEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0141933187905011.

Anonymous:1987:ARBo

[Ano87n] Anonymous. Article review: Reusability issues and Ada: Gargaro, A. and Pappas, T. L. IEEE Software (July 1987) pp 43–51. Microprocessors and Microsystems, 11(9):506, November 1987. CODEN MIMID5. ISSN 0141-9331 (print), 1872-

9436 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0141933187900391.

Anonymous:1987:ARBc

[Ano87o] Anonymous. Article review: Review of three Modula-2 development systems for the IBM PC: Ural, S. J. Pascal, Ada and Modula-2 Vol 5 No 6 (November/December 1986) pp 48–55. Microprocessors and Microsystems, 11(3):168, April 1987. CODEN MIMID5. ISSN 0141-9331 (print), 1872-

9436 (electronic). URL http://www.sciencedirect.com/
science/article/pii/014193318790161X.

Anonymous:1987:ARBi

[Ano87p] Anonymous. Article review: Simulation with Ada: Bowring, M. Syst. Int. (June 1987) pp 41. Microprocessors and Microsystems, 11(7):392, September 1987. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0141933187905497.

Anonymous:1987:ARBe

[Ano87q] Anonymous. Article review: Timing issues in the distributed execution of Ada programs: Volz, R. A. and Mudge, T. N. IEEE Trans. Comput. Vol C36 No 4 (April 1987) pp 449– 459. Microprocessors and Microsystems, 11(5):282, June 1987. CODEN MIMID5. ISSN 0141-9331 (print), 1872-

9436 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0141933187902973.

Anonymous:1987:ARBj

[Ano87r] Anonymous. Article review: Valid Ada: Anon. Syst. Int. (June 1987) pp 31–32. Microprocessors and Microsystems, 11(7):392, September 1987. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0141933187905485.

Anonymous:1987:BRBa

[Ano87s] Anonymous. Book review: Adaptive prettyprinting of abstract syntax applied to ADA and PASCAL: Peter Fritzson: Department of Computer and Information Science, Linköping University, Linköping [Sweden], September 1983. Decision Support Systems, 3(2):187, June 1987. CODEN DSSYDK. ISSN 0167-9236 (print), 1873-5797 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0167923687900820.

Anonymous:1987:BRBc

[Ano87t] Anonymous. Book review: Performance aspects of Ada tasking in embedded systems: Huijsman, R. D., van Katwijk, J. and Toetenel, W. J. pp 301–309. Microprocessors and Microsystems, 11(10):557, December 1987. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0141933187904133.

Anonymous:1987:BRBb

[Ano87u] Anonymous. Book review: Programmiermethodik mit Ada: Goos, G., Persch, G. and Uhl, J., Springer/Compass, Heidelberg, FRG (1987) DM 58 pp 161. Microprocessors and Microsystems, 11(9):506, November 1987. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0141933187900469.

Anonymous:1987:BRBd

[Ano87v] Anonymous. Book review: Software development with Ada: Sommerville, I. and Morrison, R., Addison-Wesley, Wokingham, UK (1987) £15.95 pp 371. Microprocessors and Microsystems, 11(10):562, December 1987. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0141933187904637.

Anonymous:1987:IRP

[Ano87w] Anonymous. IEEE recommended practice for Ada as a program design language, volume 990–1987 of IEEE Std. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. 15 pp. LCCN QA76.73.A22

I28 1987. URL http://ieeexplore.ieee.org/servlet/
opac?punumber=2605.

Anonymous:1987:PAT

[Ano87x] Anonymous, editor. Proceedings of AIDA-87. Third Annual Conference on Artificial Intelligence and Ada. George Mason Univ, Fairfax, VA, USA, 1987.

Anonymous:1988:ARB

[Ano88a] Anonymous. Article review: A UCSD Pascal/Modula-2 translator: Levrat, B., Nerima, L. and Tscharmer, P., J. Pascal, Ada, Modula-2 Vol 6 No 6 (November/December 1987) pp 17–24. Microprocessors and Microsystems, 12(2): 114, March 1988. CODEN MIMID5. ISSN 0141-9331 (print),

1872-9436 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0141933188900610.

Anonymous:1988:BRBa

[Ano88b] Anonymous. Book review: Good programming practice in Ada: Paul A. Luker (Blackwell Scientific Publications). Computer Languages, 13(2):107, ???? 1988. CODEN CO-LADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0096055188900148.

Anonymous:1988:BRBb

[Ano88c] Anonymous. Book review: Simulierte Zeit und das Ada Rendezvouz: Werner Pohlmann: Institut für Informatik, Technische Universität München, Nov. 1986. Decision Support Systems, 4(1):151, March 1988. CODEN DSSYDK. ISSN 0167-9236 (print), 1873-5797 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0167923688901182.

Anonymous:1988:TII

[Ano88d] Anonymous, editor. The Third International IEEE Conference on Ada Applications and Environments: May 23– 25, 1988, the Sheraton-Wayfarer Inn and Conference Center, Manchester, New Hampshire. Computer Society Press, Washington, DC, USA, 1988. ISBN 0-8186-0808-0, 0-8186-8808-4, 0-8186-4808-2. LCCN QA76.73.A35 I36 1988. URL

http://ieeexplore.ieee.org/servlet/opac?punumber=
187.

Anonymous:1989:ASMb

[Ano89a] Anonymous. Ada and Software management in NASA: assessment and recommendations: a report to the Information Resources Management Council, volume 103298 of NASA-TM. National Aeronautics and Space Administration, Goddard Space Flight Center, Greenbelt, MD, 1989. ???? pp. LCCN ????

Anonymous:1989:ASMa

[Ano89b] Anonymous, editor. Ada and software management in NASA: symposium-forum, volume 103297 of NASA-TM. National Aeronautics and Space Administration, Goddard Space Flight Center, Greenbelt, MD, USA, 1989. LCCN ???? NASA technical memorandum.

Anonymous:1989:AC

[Ano89c] Anonymous. ADA compilers. Information and Software Technology, 31(2):110, March 1989. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0950584989900979.

Anonymous:1989:CAL

[Ano89d] Anonymous. C^{3} ADA language system released. Information and Software Technology, 31(1):55, January/February 1989. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0950584989900645.

Anonymous:1989:JPA

[Ano89e] Anonymous. Jazyk programmirovanija Ada i ego realizatsia: RG-20 KNVVT. Vychislitelnyĭ etisentr AN SSSR, Moskva, Russia, 1989. 137 pp. LCCN ????

Anonymous:1990:ARB

[Ano90a] Anonymous. Article review: Dynamic change management and Ada: Burns, A. and Wellings, A. J., J. Soft. Mainten. Vol 1 No 2 (1989) pp 121-131. Information and Software Technology, 32(5):382, June 1990. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/095058499090022J.

Anonymous:1990:PAN

[Ano90b] Anonymous, editor. Proceedings of the Annual National Conference of Ada Technology (8th). Held in Atlanta, Georgia on March 5–8, 1990. U.S. Army Commun.-Electron. Command, Fort Monmouth, NJ, USA, 1990. ISBN ???? LCCN ????

Anonymous:1990:RA

[Ano90c] Anonymous. Resources in Ada. Resources in computing series. ACM Press, New York, NY 10036, USA, 1990. ISBN 0-89791-374-4. x + 259 pp. LCCN QA76.73.A35 R47 1990; QA/76/.73/A35/R48/1990 IN REF. With an introduction by Gerry Fisher.

Anonymous:1991:ALP

[Ano91] Anonymous. Ada language poses threat to security. Computer Fraud and Security Bulletin, 1991(7):1–2, July 1991. CODEN CFSBEK. ISSN 0142-0496 (print), 1878-3856 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0142049609900133.

Anonymous:1992:AFL

[Ano92] Anonymous. Ada’s fundamental language structures build reliable systems: Benjamin M. Brosgol. EDN, 153 (3 September 1990). Microelectronics and Reliability, 32 (1–2):283, January/February 1992. CODEN MCRLAS. ISSN 0026-2714 (print), 1872-941X (electronic). URL

http://www.sciencedirect.com/science/article/pii/
002627149290112X.

Anonymous:1993:DPD

[Ano93a] Anonymous. 129 design planning in the development of safety-critical software with Ada: J. Prorok, K. Bührer, U. Ammann, K. Vit pp 75–80. Control Engineering Practice, 1(2):398, April 1993. CODEN CO-EPEL. ISSN 0967-0661 (print), 1873-6939 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706619391760T.

Anonymous:1993:AGG

[Ano93b] Anonymous. Ada gets groomed for the whole spectrum of real-time duties: M. Goblin, F. J. A. Gielen, pp 107–112. Control Engineering Practice, 1(1):198, February 1993. CO-DEN COEPEL. ISSN 0967-0661 (print), 1873-6939 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/096706619392135Q.

Anonymous:1993:ARB

[Ano93c] Anonymous. Article review: An integrated system for supervisory process control using ADA: M. Marcos, F. Artaza, N. Iriondo, pp 529–534. Control Engineering Practice, 1(5):891, October 1993. CODEN COEPEL. ISSN 0967-0661 (print),

1873-6939 (electronic). URL http://www.sciencedirect.
com/science/article/pii/096706619390342O.

Anonymous:1993:DDM

[Ano93d] Anonymous. DIMOS: Distributed monitoring system from specification to delivery, the realization of a nuclear power plant supervision system using HOOD and ADA: P. Jacobs, D. Pluvinage, pp 61–66. Control Engineering Practice, 1(1):198, February 1993. CODEN CO-EPEL. ISSN 0967-0661 (print), 1873-6939 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/096706619392128Q.

Anonymous:1993:IMC

[Ano93e] Anonymous. Implementing mode changes and fault recovery for hard real-time systems in Ada: A. Alonso, J. de la Puente, pp 95–100. Control Engineering Practice, 1(1):198, February 1993. CODEN COEPEL. ISSN 0967-0661 (print), 1873-

6939 (electronic). URL http://www.sciencedirect.com/
science/article/pii/096706619392133O.

Anonymous:1993:TAW

[Ano93f] Anonymous, editor. Tenth Annual Washington Ada Symposium: Ada’s role in software engineering, June 28-July 1, 1993, McLean Hilton, McLean, VA: proceedings. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-609-3. LCCN QA76.73.A16 W37 1993; QA76.73.A16 INTERNET. URL http://DC8QA4CY3N.

search.serialssolutions.com/?V=1.0&L=DC8QA4CY3N&S=
JCs&C=TC0000753080&T=marc&tab=BOOKS.

Anonymous:1993:TAI

[Ano93g] Anonymous, editor. TRI-ADA, 93: Industry, Academia and Government Conference. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-621-2 (paperback). LCCN ???? USD 56.00 Retail Price (Publisher). URL http:

//DC8QA4CY3N.search.serialssolutions.com/?V=1.0&L=
DC8QA4CY3N&S=JCs&C=TC0000779834&T=marc&tab=BOOKS.

Anonymous:1994:WAW

[Ano94] Anonymous, editor. WADAS 94: 11th Annual Washington ADA Symposium. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-684-0 (paperback). LCCN ???? URL

http://DC8QA4CY3N.search.serialssolutions.com/?V=
1.0&L=DC8QA4CY3N&S=JCs&C=TC0000779864&T=marc&tab=
BOOKS.

Anonymous:1995:ARL

[Ano95a] Anonymous. Ada 95 rationale: the language, the standard libraries. Intermetrics, Inc., Cambridge, MA, USA, 1995. various pp. LCCN QA76.73.A35 A3252 1995.

Anonymous:1995:AUA

[Ano95b] Anonymous, editor. Ada UK Ada 9X academic seminar: — January 1994, Brighton, volume 16 (3) of Ada User Journal 1995. IOS Press, Postal Drawer 10558, Burke, VA 2209-0558, USA, 1995. ISSN 0268-652x.

Anonymous:1995:HHS

[Ano95c] Anonymous. HRT-HOOD?: a Structured Design Method for Hard Real-Time Ada Systems. Elsevier, Amsterdam, The Netherlands, 1995. ISBN 0-444-82164-3 (print), 0-444-82164-3. LCCN ????

Anonymous:1995:TAE

[Ano95d] Anonymous. TRI-ADA, 95: The Engineering and Management of Software. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-705-7 (paperback). 508 pp. LCCN ????

Anonymous:1996:TAG

[Ano96] Anonymous. TRI-ADA, 96: Government, Industry and Acedemia “Discipline Software Development with ADA”. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-808-8 (paperback). 288 pp. LCCN ???? URL http:

//DC8QA4CY3N.search.serialssolutions.com/?V=1.0&L=
DC8QA4CY3N&S=JCs&C=TC0000779840&T=marc&tab=BOOKS.

Anonymous:1997:BRDe

[Ano97a] Anonymous. Book review: Data structures and algorithms: an object-oriented approach using Ada 95: By John Beidler. Springer-Verlag, New York. (1997). 364 pages. DM 78.00, öS 569.40, sFr 69.00. Computers and Mathematics with Applications, 33(7):136, April 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0898122197845981.

Anonymous:1997:CAC

[Ano97b] Anonymous. Converting Ada 95 code to Java. Object Magazine, 7(1):68, March 1997. CODEN OBMAFO. ISSN 1055-

3614. URL http://www.intermetrics.com/.

Anonymous:1997:TAG

[Ano97c] Anonymous. TRI-ADA 97: Government, Industry and Academia. ACM Press, New York, NY 10036, USA, 1997. ISBN 0-89791-981-5. 304 pp. LCCN ???? URL http:

//DC8QA4CY3N.search.serialssolutions.com/?V=1.0&L=
DC8QA4CY3N&S=JCs&C=TC0000779842&T=marc&tab=BOOKS.

Anonymous:1998:BRCm

[Ano98] Anonymous. Book review: Concurrency in Ada (second edition): By Alan Burns and Andy Wellings. Cambridge University Press, New York. (1998). 390 pages. $32.95. Computers and Mathematics with Applications, 36(8):123, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-

7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S089812219891149X.

Anonymous:2001:PAS

[Ano01] Anonymous, editor. Proceedings: ACM SIGAda Annual International Conference (SIGAda 2001): September 30-October 4, 2001, Bloomington, MN, USA. ACM Press, New York, NY 10036, USA, 2001. ISBN 1-58113-392-8. LCCN QA76.73.A35 S485 2001; QA76.73.A35 INTERNET. URL

http://DC8QA4CY3N.search.serialssolutions.com/?V=
1.0&L=DC8QA4CY3N&S=JCs&C=TC0000752882&T=marc&tab=
BOOKS.

Anonymous:2002:PAS

[Ano02] Anonymous, editor. Proceedings of the ACM SIGAda Annual International Conference (SIGAda 2002): the engineering of correct and reliable software for real-time and distributed systems using Ada and related technologies: December 8–12, 2002, Holiday Inn Houston/NASA, Houston (at Clear Lake), Texas, USA. ACM Press, New York, NY 10036, USA, 2002. ISBN 1-58113-611-0. LCCN QA76.73.A35 S485 2002; QA76.73.A35 INTERNET. URL http://DC8QA4CY3N.

search.serialssolutions.com/?V=1.0&L=DC8QA4CY3N&S=
JCs&C=TC0000722245&T=marc&tab=BOOKS.

Anonymous:2003:PAS

[Ano03] Anonymous, editor. Proceedings of the ACM SIGAda Annual International Conference (SIGAda 2003): the engineering of correct and reliable software for real-time and distributed systems using Ada and related technologies: December 7–11, 2003, Red Lion Hannalei Hotel, San Diego, California, USA. ACM Press, New York, NY 10036, USA, 2003. ISBN 1-58113-476-2. LCCN QA76.73.A35 S485 2003; QA76.73.A35 INTERNET. URL http://DC8QA4CY3N.

search.serialssolutions.com/?V=1.0&L=DC8QA4CY3N&S=
JCs&C=TC0000722249&T=marc&tab=BOOKS.

Anonymous:2004:PAS

[Ano04] Anonymous, editor. Proceedings of the ACM SIGAda Annual International Conference (SIGAda 2004): The Engineering of Correct and Reliable Software for Real-Time and Distributed Systems Using ADA and Related Technologies: November 14–18, 2004 Double Tree Hotel-Atlanta. ACM Press, New York, NY 10036, USA, 2004. ISBN 1-58113-906-3. LCCN QA76.73.A35S485 2004; QA76.73.A35S485

2004 INTERNET. URL http://DC8QA4CY3N.search.
serialssolutions.com/?V=1.0&L=DC8QA4CY3N&S=JCs&C=
TC_011595647&T=marc&tab=BOOKS.

Anonymous:2005:PAS

[Ano05] Anonymous, editor. Proceedings of the ACM SIGAda Annual International Conference (SIGAda 2005): the engineering of correct and reliable software for real-time and distributed systems using Ada and related technologies: November 13–17, 2005, Double Tree Hotel-Atlanta/Buckhead, Atlanta, Georgia, USA. ACM Press, New York, NY 10036, USA, 2005. ISBN 1-59593-185-6. LCCN QA76.73.A35 S485 2005; QA76.73.A35 INTERNET. URL http://DC8QA4CY3N.

search.serialssolutions.com/?V=1.0&L=DC8QA4CY3N&S=
JCs&C=TC0000722263&T=marc&tab=BOOKS.

Anonymous:2006:SPA

[Ano06] Anonymous, editor. SIGAda’06: proceedings of the 2006 ACM SIGAda Annual International Conference, November 12–16, 2006, Albuquerque, New Mexico, USA. ACM Press, New York, NY 10036, USA, 2006. ISBN 1-59593-563-0. LCCN QA76.73.A35 S485 2006; QA76.73.A35 INTERNET.

URL http://DC8QA4CY3N.search.serialssolutions.com/

?V=1.0&L=DC8QA4CY3N&S=JCs&C=TC0000752976&T=marc&tab=
BOOKS.

Anonymous:2002:AMF

[Ano11] Anonymous. Ada-Mentoring: Fachzeitschrift für Mentoring und Gender Mainstreaming in Technik und Naturwissenschaften. Ada-Lovelace-Mentoring, Friedrichsdorf, Germany, 2002–2011. ISSN 1860-0522. 300 pp. LCCN ????

Anonymous:2022:ARM

[Ano22] Anonymous. Ada Reference Manual: Language and Standard Libraries. AXE Consultants, ????, 2022. xxiv + 1120

pp. URL http://www.ada-auth.org/standards/22rm/RM-
Final.pdf.

Abu-Ras:1996:PIP

[AR96] Jim Abu-Ras. Priority inheritance protocol in Ada 95. Information and Software Technology, 38(2):121–125, ???? 1996. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0950584995010564.

Ardo:1987:EAR

[Ard87] Anders Ardö. Experience acquiring and retargeting a portable Ada computer. Software—Practice and Experience, 17(4):291–307, April 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anderson:1992:MAO

[AS92] John A. Anderson and John D. Sheffler. Managing Ada object-oriented development. Lecture Notes in Computer Science, 603:20–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alrebdawi:1988:STO

[ASM88] G. Alrebdawi, J. J. Skubich, and Y. Martinez. Supporting tool for object oriented design of real time applications. Annual Review in Automatic Programming, 14 (part 1)(??):29– 33, ???? 1988. CODEN ARVAAM. ISSN 0066-4138 (print),

1878-545X (electronic). URL http://www.sciencedirect.
com/science/article/pii/0066413888900055.

Asplund:1998:RST

[Asp98] Lars Asplund, editor. Reliable software technologies — Ada-Europe: 1998 Ada-Europe International Conference on Reliable Software Technologies, Uppsala, Sweden, June 8– 12, 1998: Proceedings, volume 1411 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNCSD9. ISBN 3-540-64536-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1

L43 no.1411. URL http://link.springer-ny.com/link/
service/series/0558/tocs/t1411.htm;   http://www.
springerlink.com/content/978-3-540-64536-8;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=
0302-9743&volume=1411.

USENIX:1983:UCPb

[Ass83] USENIX Association, editor. USENIX Conference Proceedings, Winter, 1983. San Diego, CA. USENIX, Berkeley, CA, USA, Winter 1983.

Atherton:1982:SPC

[Ath82] Roy Atherton. Structured programming with COMAL, volume 17 of The Ellis Horwood series in computers and their applications. Ellis Horwood, New York, NY, USA, 1982. ISBN 0-470-27318-6 (Halsted), 0-85312-416-7 (cased), 0-85312-423-X (paperback). 266 (est.) pp. LCCN QA76.73.C26

A87 1982. URL http://catalog.hathitrust.org/api/
volumes/oclc/8345491.html.

Auguston:1995:BRB

[Aug95] Mikhail Auguston. Book review: Software system construction with examples in Ada: by Bo Sanden; Prentice Hall; Englewood Cliffs, NJ, USA; 1994; 461 pp.; 47; ISBN: 0-13-030834-X. Control Engineering Practice, 3(4):595, April 1995. CODEN COEPEL. ISSN 0967-0661 (print), 1873-

6939 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0967066195900950.

Austrian:1982:HHF

[Aus82] Geoffrey D. Austrian. Herman Hollerith — Forgotten Giant of Information Processing. Columbia University Press, New York, NY, USA, 1982. ISBN 0-231-05146-8. xvi + 418 pp. LCCN QA76.2.H64 .A97.

Ausden:2011:BRB

[Aus11] Howard Ausden. Book review: Building parallel, embedded, and real-time applications with Ada by John W. McCormick, Frank Singhoff and Jerome Hugues. ACM SIGSOFT Software Engineering Notes, 36(6):28–29, November 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ben-Ari:1998:ASE

[BA98] M. Ben-Ari. Ada for software engineers. Worldwide series in computer science. Wiley, New York, NY, USA, 1998. ISBN 0-471-97912-0 (paperback/CD-ROM). xiii + 425 pp. LCCN QA76.73.A35 B46 1998. URL http://www.loc.gov/

catdir/toc/onix05/98018158.html.

Ben-Ari:2009:ASE

[BA09] Mordechai Ben-Ari, editor. Ada for Software Engineers. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second (with Ada 2005) edition, 2009. ISBN 1-84882-313-4 (paperback), 1-84882-314-2 (e-book). xiv + 509 pp. LCCN QA76.73.A35 .B46 2009. URL http:

//site.ebrary.com/lib/ucalgary/Doc?id=10288816;
http://w2.vu.edu.au/library/EBookSearch/files/Springer_
E-Books_Online.pdf.

Baker:1983:MAS

[Bak83] T. P. Baker. Mapping Ada onto a simple virtual machine. Technical Report ADA-83-02, Florida State University, 1983.

Baker:1988:IAR

[Bak88] T. P. Baker. An improved Ada run-time system interface. The Journal of Systems and Software, 8(5):373–393, December 1988. CODEN JSSODM. ISSN 0164-1212 (print), 1873-

1228 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0164121288900295.

Balfour:1997:AJB

[Bal97] B. Balfour. Ada 95, Java byte code, and the distributed systems annex. In ACM [ACM97], pages 247–

162. ISBN 0-89791-981-5. LCCN QA 76.73 A35 T75 1997.

URL http://www.acm.org/pubs/contents/proceedings/

ada/269629/. Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Bishop:1987:DCA

[BAP87] Judy M. Bishop, Stephen R. Adams, and David J. Pritchard. Distributing Concurrent Ada programs by source translation. Software—Practice and Experience, 17(12):859–884, December 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barnes:1982:PA

[Bar82] J. G. P. (John Gilbert Presslie) Barnes. Programming in ADA. International computer science series. Addison-Wes-ley, Reading, MA, USA, 1982. ISBN 0-201-13793-3, 0-201-13792-5 (paperback). x + 340 pp. LCCN QA76.73.A35 B37. £14.95£7.95.

Barnes:1987:PA

[Bar87a] J. G. P. (John Gilbert Presslie) Barnes. Programación en Ada. Ediciones Díaz de Santos, Madrid, Spain, 1987. ???? pp. Translation of [Bar82] to Spanish by Sergio Arévalo and Angel Alvarez.

Barnes:1987:IAU

[Bar87b] John Barnes. Introduction to ADA and its use for embedded applications. Microprocessors and Microsystems, 11(5):245– 254, June 1987. CODEN MIMID5. ISSN 0141-9331 (print),

1872-9436 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0141933187902675.

Barnes:1987:PIW

[Bar87c] John Barnes, editor. Proceedings of the International Workshop on Real-Time Ada issues, Moretonhampstead, Devon, UK, 13–15 May 1987, For parts, see ACM SIGADA Ada Letters vol. 7, no. 6. ACM Press, New York, NY 10036, USA, 1987. ISBN 0-89791-240-3. LCCN QA76.73.A35 A3 v.7:6. US$14.

Barnes:1988:PA

[Bar88] J. G. P. (John Gilbert Presslie) Barnes. Programmer en ADA. Informatique, intelligence artificielle. Interéditions, Paris, France, 1988. ISBN 2-7296-0216-X. xvi + 495 pp. LCCN QA 76 .73 A35 B3712 1988.

Barnes:1989:PA

[Bar89] J. G. P. (John Gilbert Presslie) Barnes. Programming in ADA. International computer science series. Addison-Wesley, Reading, MA, USA, third edition, 1989. ISBN 0-201-17566-5. xvi + 494 pp. LCCN QA76.73.A15.

Barnes:1994:PLS

[Bar94] John Barnes. The programming language standards scene, ten years on paper 9: Ada. Computer Standards and Interfaces, 16(5–6):481–485, September 1994. CODEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0920548994900299.

Barbey:1996:TAO

[Bar96] S. Barbey. Testing Ada 95 object-oriented programs. Lecture Notes in Computer Science, 1031:406–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barnes:1997:ARL

[Bar97] J. G. P. (John Gilbert Presslie) Barnes. Ada 95 rationale: the language, the standard libraries, volume 1247 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNCSD9. ISBN 3-540-63143-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 458 pp. LCCN QA76.73.A35

A217 1997. URL http://link.springer-ny.com/link/
service/series/0558/tocs/t1247.htm;   http://www.
springerlink.com/content/978-3-540-63143-9;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=

0302-9743&volume=1247. Document to accompany the new Ada standard, officially ISO/IEC 8652: 1995.

Barnes:2003:PAS

[Bar03] John Gilbert Presslie Barnes. Programming in Ada 95 [sound recording]. TPB, Enskede, The Netherlands, 2003. 2 CD–R

(69h 30m) pp. LCCN ???? URL http://katalog.tpb.se/
wsHitList.Asp?SCode1=TN&SearchStr1=C21070.

Barnes:2008:ARL

[Bar08] John Barnes, editor. Ada 2005 Rationale: The Language, The Standard Libraries, volume 5020 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNCSD9. ISBN 3-540-79700-9 (print), 3-540-79701-7 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 267 pp. LCCN QA76.73.A35 A2173 2008; QA76.73.A35

B39 2008. URL http://springerlink.com/content/
lj72202705x4/?p=d9eb747f22c149e78b620bbcce131e0b&
pi=89;   http://w2.vu.edu.au/library/EBookSearch/
files/Springer_E-Books_Online.pdf;   http://www.springerlink.
com/content/978-3-540-79701-2.

Barnes:2014:PA

[Bar14] John Barnes. Programming in Ada 2012. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-42481-X. ???? pp. LCCN ???? URL http:

//www.cambridge.org/us/academic/subjects/computer-
science/software-engineering-and-development/programming-
ada-2012.

Baskette:1987:LCA

[Bas87] Jerry Harvey Baskette. Life cycle analysis of an Ada project. IEEE Software, 4(1):40–47, January 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Boujarwah:1996:MSM

[BASS96] Abdulaziz Boujarwah, Nadia Al-Seif, and Kassem Saleh. Modelling the semantics of multitasking facilities in Concurrent C using Petri nets. Information and Software Technology, 38(1):3–9, ???? 1996. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0950584995010394.

Baumgarten:1991:DSA

[Bau91] U. Baumgarten. Distributed systems and Ada — current projects and approaches comparative study’s results. Lecture Notes in Computer Science, 499:260–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baldassari:1991:POO

[BB91] Marco Baldassari and Giorgio Bruno. PROTOB: An object oriented methodology for developing discrete event dynamic systems. Computer Languages, 16(1):39–63, ???? 1991. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0096055191900163.

Bauer:1995:RTA

[BB95] B. Bauer and C. Bouvier. Real-time Ada applications with silicon executives. Microprocessors and Microsystems, 19(2): 83–88, ???? 1995. CODEN MIMID5. ISSN 0141-9331 (print),

1872-9436 (electronic). URL http://www.sciencedirect.
com/science/article/pii/014193319598983Q.

Biieberger:1998:SRD

[BB98a] J. Biieberger and B. Burgstaller. Symbolic reaching definitions analysis of Ada programs. Lecture Notes in Computer Science, 1411:238–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blieberger:1998:SRD

[BB98b] Johann Blieberger and Bernd Burgstaller. Symbolic reaching definitions analysis of Ada programs. Lecture Notes in Computer Science, 1411:238–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/series/0558/bibs/
1411/14110238.htm;   http://link.springer-ny.com/
link/service/series/0558/papers/1411/14110238.pdf.

Bazalgette:1992:SAT

[BBB^{+}92] G. Bazalgette, D. Bekele, C. Bernon, M. Filali, J. M. Rigaud, and A. Sayah. STRAda — an Ada transformation and distribution system. Lecture Notes in Computer Science, 603: 287–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bailes:1996:KAO

[BBCS96] P. A. Bailes, P. Burnim, M. Chapman, and E. Salzman. KBSE and Ada-object and enabling technology. Lecture Notes in Computer Science, 1031:152–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brauer:1981:PLA

[BBG^{+}81] Wilfried Brauer, Per Brinch Hansen, David Gries, C. Moler, Gerhard Seegmüller, Josef Stoer, and Niklaus Wirth, editors. The Programming Language Ada Reference Manual, Proposed Standard Document, United States Department of Defense, volume 106 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. ISBN 3-540-10693-6.

Belz:1980:MPI

[BBH80] F. C. Belz, E. K. Blum, and D. Heimbigner. A multiprocessing implementation-oriented formal definition of Ada in SEMANOL. ACM SIGPLAN Notices, 15(11):202–212, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bayassi:1992:PUA

[BBJL92] M. Bayassi, H. Bitteur, J.-F. Jezequel, and P. Legrain. A practical use of the Ada rendez-vous paradigm in distributed systems. Lecture Notes in Computer Science, 603:312–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Booker:1984:EAP

[BBP^{+}84] Don M. M. Booker, Barry Burd, Jerry Przybylski, Kevin Cogan, George Corliss, Carl Brandon, Don Yee, and Phil Goldstein. Experiences in ADA: Perspective problems and prospects for a potential primary language of instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 16(1):182, February 1984. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 15th SIGCSE Technical Symposium on Computer Science Education.

Bailey:1995:KPP

[BBWF95] C. M. Bailey, A. Burns, A. J. Wellings, and C. H. Forsyth. Keynote paper: a performance analysis of a hard real-time system. Control Engineering Practice, 3(4):447–464, April 1995. CODEN COEPEL. ISSN 0967-0661 (print), 1873-

6939 (electronic). URL http://www.sciencedirect.com/
science/article/pii/096706619500017O.

Buck:1990:PAN

[BDG90] P. D. Buck, S. L. Day, and D. Gonzalez. Problems with Ada numeric routines. In Anonymous [Ano90b], pages 195–204. ISBN ???? LCCN ????

Burns:1998:RTP

[BDR98] A. Burns, B. Dobbing, and G. Romanski. The Ravenscar tasking profile for high integrity real-time programs. Lecture Notes in Computer Science, 1411:263–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/1411/14110263.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
1411/14110263.pdf.

Brosgol:1992:ADA

[BEE92] Benjamin M. Brosgol, Robert I. Eachus, and David E. Emery. An Ada decimal arithmetic capability. CrossTalk: The Journal of Defense Software Engineering, 36, Septem-

ber 1992. URL http://www.iste.uni-stuttgart.de/
ps/AdaBasis/pal_1195/ada/ajpo/work-grp/ev-team/ev-
info/summary.txt.

Beebe:1994:BPAf

[Bee94] Nelson H. F. Beebe. A bibliography of publications in ACM SIGAda Ada Letters. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 17, 1994.

44 pp. URL https://www.math.utah.edu/pub/tex/bib/
index-table-s.html#sigada. This report is updated fre-

quently.

Beebe:1997:BAU

[Bee97] Nelson H. F. Beebe. A bibliography of Ada User and Ada User Journal. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, September 9, 1997.

18 pp. URL https://www.math.utah.edu/pub/tex/bib/
index-table-a.html#adauserj. This report is updated fre-

quently.

Beidler:1997:DSA

[Bei97] John Beidler. Data structures and algorithms: an objectoriented approach using Ada 95. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 0-387-94834-1 (hardcover), 1-4612-7312-9, 1-4612-1854-3 (e-book). xvi + 364 pp. LCCN A76.64 .B43

Belmont:1980:TRA

[Bel80] Peter A. Belmont. Type resolution in Ada: an implementation report. ACM SIGPLAN Notices, 15(11):57–61, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Belkhouche:1991:GAP

[Bel91] Boumediene Belkhouche. Generation of ADA and PL/1 prototypes from abstract data type specifications. The Journal of Systems and Software, 16(3):255–264, November 1991. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0164121291900207.

Bell:1997:ATA

[Bel97] A. E. Bell. An alternative toolset for analysis of Ada programs. Lecture Notes in Computer Science, 1251:112–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blum:1987:ASM

[BEPP87] E. K. Blum, H. Ehrig, and F. Parisi-Presicce. Algebraic specification of modules and their basic interconnections. Journal of Computer and System Sciences, 34(2–3):293–339, April/ June 1987. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-

2724 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022000087900286.

Barnes:1985:AUP

[BF85] J. G. P. (John Gilbert Presslie) Barnes and Gerald A. Fisher, Jr., editors. Ada in use: proceedings of the Ada International Conference, 1985, The Ada companion series. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-30968-9. LCCN QA76.73.A16 A25 1985. URL http://uclibs.org/

PID/21238.

Brukardt:2000:NIS

[BFC00] Randall Brukardt, Nancy A. Forbes, and David N. Card. In the news: An ISO Standard guards the Ada hen house: Innovation and the National Reconnaissance Office: Making measurement understandable. IEEE Software, 17(1):92– 96, January/February 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://dlib.

computer.org/so/books/so2000/pdf/s1092.pdf.

Bryant:1984:ILD

[BG84] Barrett R. Bryant and A. A. Grau. An intermediate language to define dynamic semantics. Computer Languages, 9(3–4):149–159, ???? 1984. CODEN CO-LADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/009605518490002X.

Baker:1995:IAP

[BG95] T. P. Baker and E. W. Giering, III. Implementing Ada 9X protected objects and asynchronous transfer of control. International Journal of Mini and Microcomputers, 17(1):26–34, ???? 1995. CODEN IJMMDE. ISSN 0702-0481.

Bergin:1996:HPL

[BG96] T. J. Bergin, Jr. and R. G. Gibson, Jr., editors. History of programming languages: II. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-201-89502-1. xvi + 864 pp. LCCN QA76.7 .H57 1996.

Booch:1982:EIC

[BHM^{+}82] Grady Booch, Hal Hart, Vance Mall, Phil Miller, and Peter Wegner. The educational issues confronting Ada (panel discussion). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 14(1):261, February 1982. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 13th SIGCSE Symposium on Computer Science Education.

Bielak:1985:AVM

[Bie85a] Richard Bielak. ADA(*) vs. Modula-2: a view from the trenches. ACM SIGPLAN Notices, 20(12):13–17, December 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See comments [Gre86].

Bielecki:1985:AIS

[Bie85b] Jan Bielecki. Ada: interpretacja standardu operacji wejścia/ wejścia. Wydawnictwa Politechniki Warszawskiej, Warszawa, Poland, 1985. 117 pp. LCCN ????

Bucci:1993:EIS

[BIM93] A. Bucci, P. Inverardi, and S. Martini. An “executable” impredicative semantics for the Ada configuration. Formal Aspects of Computing, 5(2):91–120, March 1993. CO-DEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF01211301.

Biswas:1985:ASV

[Bis85] P. Biswas. Architectural support for variable addressing in Ada — a design approach. International Journal of Computer and Information Sciences, 14(1):51–72, February 1985. CODEN IJCIAH. ISSN 0091-7036.

Bishop:1990:DAD

[Bis90] Judy M. Bishop, editor. Distributed Ada: Developments and Experiences. The Ada Companion Series. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-511-52654-7 (e-book). LCCN ????

Bailes:1993:PGL

[BJS93] Paul A. Bailes, Dan B. Johnston, and Eric J. Salzman. A proposal for a genuinely-lazy streams facility for Ada. Computer Languages, 18(1):31–55, ???? 1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/009605519390029Z.

Butler:1987:SMS

[BK87] Gregory Butler and Matthew J. Kendall. The suitability for master/slave concurrency of Concurrent Euclid, Ada and Modula. Software—Practice and Experience, 17(2):117–134, February 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bieman:1995:MLS

[BK95] James M. Bieman and Santhi Karunanithi. Measurement of language-supported reuse in object-oriented and object-based software. The Journal of Systems and Software, 30(3):271–293, September 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
016412129400138D.

Black:2002:IHS

[Bla02] Edwin Black. IBM and the Holocaust: the strategic alliance between Nazi Germany and America’s most powerful corporation. Three Rivers Press, New York, NY, USA, 2002. ISBN 0-609-80899-0 (paperback). 551 pp. LCCN HD9696.2.U64 I253 2002.

Blieberger:1996:AAA

[BLB96] J. Blieberger, R. Lieger, and B. Burgstaller. Augmenting Ada 95 with additional real-tune features. Lecture Notes in Computer Science, 1088:330–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blum:1988:SCP

[Blu88] Edward K. Blum. The semantics and complexity of parallel programs for vector computations. I. A case study using Ada. BIT (Nordisk tidskrift for informationsbehandling), 28(3):530–551, 1988. CODEN NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Burns:1987:RAT

[BLW87] Alan Burns, Andrew M. Lister, and Andrew J. Wellings. A review of Ada tasking, volume 262 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNCSD9. ISBN 3-540-18008-7 (Berlin), 0-387-18008-7 (US). ISSN 0302-9743 (print), 1611-3349

(electronic).  viii + 141 pp.  LCCN QA76.73.A35

B871 1987. URL http://link.springer-ny.com/link/
service/series/0558/tocs/t0262.htm;   http://www.
springerlink.com/content/978-0-387-18008-3;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=
0302-9743&volume=262.

Barringer:1982:APR

[BM82] H. Barringer and I. Mearns. Axioms and proof rules for ADA tasks. Proceedings of the IEEE, 129(2):38–48, 1982. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Bossavit:1985:APT

[BM85] Alain Bossavit and Bertrand Meyer. An application of program transformation to supercomputer programming. Computer Physics Communications, 37(1–3):27–38, July 1985.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046558590133X.

Barringer:1986:PSA

[BM86] H. Barringer and I. Mearns. A proof system for Ada tasks. The Computer Journal, 29(5):404–415, October 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/
content/29/5/404.full.pdf+html.

Beer:1987:PFH

[BM87] Martin D. Beer and J. David T. Martin. The provision of flexibility in hierarchical control systems. Microprocessing and Microprogramming, 21(1–5):295–299, August 1987. CO-DEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165607487900536.

Badaro:1991:RAM

[BM91] N. Badaro and Th. Moineau. ROSE-Ada: a method and a tool to help reuse of Ada codes. Lecture Notes in Computer Science, 499:304–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Battini:1996:ESB

[BMM96] F. Battini, P. L. Mantovani, and M. Mattavelli. Evaluation of a SPARC board equipped with the Ada tasking coprocessor (ATAC). Lecture Notes in Computer Science, 1031:379–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bover:1992:AFC

[BMO92] David C. C. (David Charles Clifford) Bover, Kevin J. Maciunas, and Michael J. Oudshoorn. ADA: a first course in programming and software engineering. International computer science series. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-50992-X. xxvi + 391 pp. LCCN QA76.6 .B695 1991. price unknown.

Bjorner:1980:TFD

[BØ80] Dines Bjørner and Ole Nybye Øst, editors. Towards a formal description of Ada, volume 98 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNCSD9. ISBN 0-387-10283-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 630 pp. LCCN QA76.73.A35

T68 1980. URL http://link.springer-ny.com/link/
service/series/0558/tocs/t0098.htm;   http://www.
springerlink.com/content/978-0-387-10283-2;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=
0302-9743&volume=98.

Booch:1983:SEA

[Boo83] Grady Booch. Software engineering with Ada. Benjamin/Cummings series in computing and information sciences. Benjamin/Cummings Pub. Co., Menlo Park, CA, USA, 1983. ISBN 0-8053-0600-5 (paperback). xix + 504 pp. LCCN QA76.73.A35 B66 1983; QA/76/ .73/A35/B66/1983 IN. US$12.95.

Booch:1987:SCA

[Boo87] Grady Booch. Software components with Ada: structures, tools, and subsystems. Benjamin/Cummings Pub. Co., Menlo Park, CA, USA, 1987. ISBN 0-8053-0610-2. xx + 635 pp. LCCN ????

Boone:1989:BRW

[Boo89] John Boone. Book review: Writing Readable Ada — A Case Study Approach. ACM SIGPLAN Notices, 24(11):27–28, November 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Booch:1991:OOD

[Boo91] Grady Booch. Object-oriented design. Benjamin/Cummings series in ADA and software engineering. Benjamin/Cummings Pub. Co., Menlo Park, CA, USA, 1991. ISBN 0-8053-0091-0. xix + 580 pp. LCCN ????

Boriani:1995:OOA

[Bor95] Dario V. Boriani. Object-oriented Ada-based development of highly reusable control software. ISA Transactions, 34(1):39–51, March 1995. CODEN ???? ISSN ????

URL http://www.sciencedirect.com/science/article/
pii/001905789400049R.

Boute:1980:SAR

[Bou80] Raymond T. Boute. Simplifying ADA by removing limitations. ACM SIGPLAN Notices, 15(2):17–29, February 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bowden:1953:FTT

[Bow53] Baron Bertram Vivian Bowden, editor. Faster Than Thought: a Symposium on Digital Computing Machines. Sir Isaac Pitman and Sons, Ltd., London, UK, 1953. LCCN QA76.5 .B66.

URL https://archive.org/details/FasterThanThought.

With a foreword by the Right Honourable the Earl of Halsbury.

Boyd:1987:SDD

[Boy87] S. Boyd. Status of DAPSE distributed Ada support. ACM SIGSOFT Software Engineering Notes, 12(2):56, April 1987. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Brorsson:2012:RST

[BP12] Mats Brorsson and Luís Miguel Pinho, editors. Reliable Software Technologies — Ada–Europe 2012: 17th Ada–Europe International Conference on Reliable Software Technologies, Stockholm, Sweden, June 11–15, 2012. Proceedings, volume 7308 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNCSD9. ISBN 3-642-30597-0 (paperback), 3-642-30598-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C65; QA76.7-76.73.

URL http://d-nb.info/1021849022/04;   http://www.
springerlink.com/content/978-3-642-30598-6.

Burns:1986:ADD

[BR86] A. Burns and J. Robinson. ADDS — a dialogue development system for the Ada programming language. International Journal of Man-Machine Studies, 24(2):153– 170, February 1986. CODEN IJMMBC. ISSN 0020-7373.

URL http://www.sciencedirect.com/science/article/
pii/S0020737386800463.

Bray:1984:SCA

[Bra84] Gary Bray. Sharing code among instances of Ada generics. ACM SIGPLAN Notices, 19(6):276–284, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brantlinger:1989:EES

[Bra89] Patrick Brantlinger, editor. Energy and entropy: science and culture in Victorian Britain: essays from Victorian studies. Indiana University Press, Bloomington, IN, USA, 1989. ISBN 0-253-31928-5. xxii + 352 pp. LCCN Q175.52.G7 E54 1988.

Brandon:2000:ANS

[Bra00] Ann S. Brandon. AppSwitch: Network switching with Ada from Linux. Linux Journal, 69:??, January 2000. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (elec-

tronic). URL http://noframes.linuxjournal.com/lj-
issues/issue69/3675.html.

Brender:1980:CAA

[Bre80] Ronald F. Brender. The case against Ada as an APSE command language. ACM SIGPLAN Notices, 15(10):27–34, October 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bremmon:1996:FMM

[Bre96] C. Bremmon. The funds management modernization: Experiences with developing an object-oriented, client-server management information system in Ada 95. Lecture Notes in Computer Science, 1088:427–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Briggs:1984:TIA

[Bri84] J. S. Briggs. Two implementations of the Ada program library. Software—Practice and Experience, 14(5):491–500, May 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brosgol:1980:TME

[Bro80] Benjamin M. Brosgol. TCOL_{Ada} and the “middle end” of the PQCC Ada compiler. ACM SIGPLAN Notices, 15(11):101– 112, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Broad:1981:POE

[Bro81] William J. Broad. Pentagon orders end to computer Babel: To halt a proliferation of computer tongues, the Pentagon has built a universal language; but rebels fight the unification. Science, 211(4471):31–33, January 2, 1981. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/211/
4477/31.extract.

Brown:1984:TAI

[Bro84] D. H. J. Brown. Transitions to Ada: an incremental approach. The Computer Journal, 27(1):37–41, February 1984. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/27/1/37.full.pdf+html.

Brophy:1989:LLTa

[Bro89a] Carolyn Elizabeth Brophy. Lessons learned in the transition from Ada to FORTRAN at NASA/Goddard. Software Engineering Laboratory series; SEL-89-005 NASA-TM 103311, National Aeronautics and Space Administration, Goddard Space Flight Center, Greenbelt, MD, USA, 1989. ???? pp. For sale by the National Technical Information Service.

Brophy:1989:LLTc

[Bro89b] Carolyn Elizabeth Brophy. Lessons learned in the transition to Ada from FORTRAN at NASA/Goddard. Computer science technical report series CS-TR-2305, University of Maryland, College Park, MD, USA, August 1989. vii + 90 pp. Supported in part by NASA. UMIACS-TR-89-84.

Brophy:1989:LLTd

[Bro89c] Carolyn Elizabeth Brophy. Lessons learned in the transition to ADA from FORTRAN at NASA/Goddard. NASA CR-186458, Dept. of Computer Science, University of Maryland, College Park, MD, USA, 1989. ???? pp.

Brophy:1989:LLTb

[Bro89d] Carolyn Elizabeth Brophy. Lessons learned in the transition to Ada from FORTRAN at NASA/Goddard. Thesis (m.s.), University of Maryland, College Park, MD, USA, 1989. xiii + 131 pp.

Brosgol:1992:A

[Bro92] Benjamin M. Brosgol. Ada. Communications of the ACM, 35 (11):41–42, November 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0001-0782/138845.html.

Brosgol:1996:DPA

[Bro96a] B. M. Brosgol. The dining philosophers in Ada 95. Lecture Notes in Computer Science, 1088:247–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1996:PAS

[Bro96b] R. Brown. The practical application of safety techniques on an Ada based project. Lecture Notes in Computer Science, 1031:51–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brosgol:1997:COF

[Bro97] B. M. Brosgol. A comparison of the object-oriented features of Ada 95 and Java[TM]. In ACM [ACM97], pages 213– 230. ISBN 0-89791-981-5. LCCN QA 76.73 A35 T75 1997.

URL http://www.acm.org/pubs/contents/proceedings/

ada/269629/. Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Brosgol:2005:CME

[Bro05] B. M. Brosgol. A comparison of the mutual exclusion features in Ada and the real-time specification for Java. Lecture Notes in Computer Science, 3555:129–143, 2005. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruno:1984:UAD

[Bru84] Giorgio Bruno. Using Ada for discrete event simulation. Software—Practice and Experience, 14(7):685–695, July 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bermudez:1990:PAL

[BS90] Manuel E. Bermudez and Karl M. Schimpf. Practical arbitrary lookahead LR parsing. Journal of Computer and System Sciences, 41(2):230–250, October 1990. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/002200009090037L.

Blieberger:2002:RST

[BS02] Johann Blieberger and Alfred Strohmeier, editors. Reliable software technologies — ADA-Europe 2002: 7th Ada-Europe International Conference on reliable Software Technologies, Vienna, Austria, June 17–21, 2002: proceedings, volume 2361 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNCSD9. ISBN 3-540-43784-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A16 A23 2002. URL http://link.

springer-ny.com/link/service/series/0558/tocs/t2361.
htm;   http://uclibs.org/PID/23121;   http://www.
springerlink.com/content/978-3-540-43784-0;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=
0302-9743&volume=2361.

Black:1998:QLA

[BST98] C. Black, S. M. Shatz, and S. Tu. A query language for automated general analysis of Concurrent Ada programs. International Journal of Computer Systems Science and Engineering, 13(2):83–95, March 1998. CODEN CSSEEI. ISSN 0267-6192.

Brindle:1989:DAT

[BTM89] A. F. Brindle, R. N. Taylor, and D. F. Martin. A debugger for Ada tasking. IEEE Transactions on Software Engineering, 15 (3):293–304, March 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?arnumber=21757.

Bryant:1984:SST

[BU84] Ray Bryant and Brian W. Unger, editors. Simulation in strongly typed languages, ADA, PASCAL, SIMULA: proceedings of the Conference on Simulation in Strongly Typed Languages, 2–4 February 1984, San Diego, California, volume 13, no. 2 of Simulation series; 0735-9276. Simulation Councils, Inc., La Jolla, CA, USA, 1984. LCCN QA76.9C65 C66 1984; QA/76/ .9/C65/C65/1984 IN.

Budgen:1988:BRB

[Bud88] David Budgen. Book review: Software development with Ada by Ian Sommerville and R. Morrison: Published by Addison-Wesley. 359pp. £15.95. Information and Software Technology, 30(4):260–261, May 1988. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0950584988900948.

Bundgaard:1996:ABA

[Bun96] J. Bundgaard. An ANDF based Ada 95 compiler system. Lecture Notes in Computer Science, 1031:81–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burns:1985:CPA

[Bur85] Alan Burns. Concurrent programming in Ada. The Ada companion series. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-30033-9. xiv + 241 pp. LCCN QA76.73.A16 B87 1985.

Burgess:1988:UAL

[Bur88] Clifford G. Burgess. Unix and ADA — local area networks: Helping users use Unix. Computers & industrial engineering, 15(1–4):244–248, ???? 1988. CODEN CINDDL. ISSN 0360-8352 (print), 1879-0550 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0360835288900940.

Bustamante:1996:BRB

[Bus96] Donald D. Bustamante. Book review: Ada: Towards maturity: edited by L. Collingbourne. Series: Studies in computer and communications systems, volume 6; IOS Press; Amsterdam, The Netherlands; 1993; 209 pp; 75; ISBN: 90-5199-142-8. Control Engineering Practice, 4(5):744, May 1996. CODEN COEPEL. ISSN 0967-0661 (print), 1873-6939 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967066196900246.

Babbage:2007:PDV

[BV07] Charles Babbage and Andrea Villa, editors. Passaggi dalla vita di uno scienziato: autobiografia dell’inventore del computer. (Italian) [Passages from the Life of a Philosopher: autobiography of the inventor of the computer]. Frontiere / [UTET libreria]. UTET libreria, Torino, Italia, 2007. ISBN 88-02-07715-0. xvi + 414 pp. LCCN ???? With an introduction by Vittorio Marchis.

Burns:1990:NPR

[BW90] A. Burns and A. J. Wellings. The notion of priority in real-time programming languages. Computer Languages, 15(3):153–162, ???? 1990. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
009605519090008D.

Burns:1995:CA

[BW95] Alan Burns and Andrew J. Wellings. Concurrency in ADA. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-41471-7. xvi + 396 pp. LCCN QA76.73.A35 B85

1995. URL http://www.loc.gov/catdir/description/
cam026/94023789.html;   http://www.loc.gov/catdir/
toc/cam029/94023789.html.

Burns:1996:AEC

[BW96] A. Burns and A. J. Wellings. Ada 95: An effective concurrent programming language. Lecture Notes in Computer Science, 1088:58–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burns:1997:SSF

[BW97] A. Burns and A. J. Wellings. Synchronous sessions and fixed priority scheduling. Journal of Systems Architecture, 44(2):107–118, November 1997. CODEN JSARFB. ISSN 1383-7621 (print), 1873-6165 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1383762197000027.

Burns:1998:CA

[BW98] Alan Burns and Andrew J. Wellings. Concurrency in ADA. Cambridge University Press, Cambridge, UK, second edition, 1998. ISBN 0-521-62911-X (paperback). xviii + 390 pp. LCCN QA76.73.A35 B85 1998. URL http://www.loc.gov/

catdir/description/cam029/98150778.html;   http://
www.loc.gov/catdir/toc/cam026/98150778.html.

Burns:2001:RTS

[BW01] Alan Burns and Andrew J. Wellings. Real-time systems and programming languages: Ada 95, real-time Java, and realtime POSIX. International computer science series. Addison-Wesley, Reading, MA, USA, third edition, 2001. ISBN 0-201-72988-1. xvi + 738 pp. LCCN QA76.54 .B87 2001. Revised edition of Real-time systems and their programming languages, 1990.

Brosgol:2003:CATa

[BW03a] B. M. Brosgol and A. Wellings. A comparison of the asynchronous transfer of control features in Ada and the real-time specification for Java. YCS Report 350, University of York Department of Computer Science, York, UK, 2003.

Brosgol:2003:CATb

[BW03b] B. M. Brosgol and A. Wellings. A comparison of the asynchronous transfer of control features in Ada and the real-time specification for Java. Lecture Notes in Computer Science, 2655:113–128, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burns:2004:RTS

[BW04] Alan Burns and Andy Wellings. Real-time systems and programming languages [sound recording]: Ada 95, realtime Java and real-time Posix. TPB, Enskede, The Netherlands, 2004. 2 CD–R (61h 54m) pp. LCCN

????  URL http://katalog.tpb.se/wsHitList.Asp?
SCode1=TN&SearchStr1=C30785.

Burns:2007:CRT

[BW07] Alan Burns and Andrew J. Wellings. Concurrent and realtime programming in Ada 2005. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-86697-9 (hardcover). xiv + 461 pp. LCCN QA76.73.A35 B854 2007. URL http://

www.loc.gov/catdir/enhancements/fy0803/2007281389-
b.html;   http://www.loc.gov/catdir/enhancements/
fy0803/2007281389-d.html;   http://www.loc.gov/catdir/
enhancements/fy0803/2007281389-t.html.

Burns:2009:RTS

[BW09] Alan Burns and Andy Wellings. Real-time systems and programming languages. International computer science series.

Addison-Wesley, Reading, MA, USA, fourth edition, 2009. ISBN 0-321-41745-3. xviii + 602 pp. LCCN ????

Berry:1987:APD

[BYY87] Daniel M. Berry, Nancy Yavne, and Moshe Yavne. Application of program design language tools to Abbott’s method of program design by informal natural language descriptions. The Journal of Systems and Software, 7(3):221–247, September 1987. CODEN JSSODM. ISSN 0164-1212 (print), 1873-

1228 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0164121287900446.

Carver:1996:TAD

[Car96] Richard H. Carver. Testing abstract distributed programs and their implementations: a constraint-based approach. The Journal of Systems and Software, 33(3):223–237, June 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0164121296000246.

Carrez:1997:SDJ

[Car97] Christian Carrez. Structures et Données en Java, C++, et Ada 95: Pratique et outils de contrôle. (French) [Structures and Data in Java, C++, and Ada 95: Practice and monitoring tools]. Masson Editeur, Masson, France, 1997. ISBN 2-225-83007-X (paperback). x + 367 pp. LCCN ???? 325

FF. URL http://www.masson.fr/cgi-bin/bookf.pl?1:1:
is=222583007X.

Cook:1994:TIA

[CB94] Dave Cook and Eugene Bingue. Tutorial introduction to Ada 9X (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):373, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Corbin:1996:MOB

[CB96] M. J. Corbin and G. F. Butler. MulTiSIM: An object-based distributed framework for mission simulation. Simulation Practice and Theory, 3(6):383–399, January 15, 1996. CO-DEN SPTHEH. ISSN 0928-4869 (print), 1879-1433 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0928486995000178.

Carvalho:2009:ILT

[CB09] B. P. R. Carvalho and A. P. Braga. IP-LSSVM: a twostep sparse classifier. Pattern Recognition Letters, 30 (16):1507–1515, December 1, 2009. CODEN PRLEDG. ISSN 0167-8655 (print), 1872-7344 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0167865509002086.

Copeland:2017:TG

[CBSW17] B. Jack Copeland, Jonathan Bowen, Mark Sprevak, and Robin Wilson, editors. The Turing guide. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2017. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). xv + 546 pp. LCCN QA29.T8 C67 2017.

Campbell:1986:OAS

[CC86] Lisa M. Campbell and Mark D. Campbell. An overview of the Ada [1] shell. In USENIX Association [USE86b], pages 302–313.

Chiu:1994:DEF

[CC94] Stephen Chiu and Sujeet Chand. A development environment for fuzzy rule-based traffic control. Robotics and Computer-Integrated Manufacturing, 11(3):167–176,

September 1994. CODEN ????  ISSN ????  URL

http://www.sciencedirect.com/science/article/pii/
0736584594900310.

Canfora:1990:RED

[CCD90] G. Canfora, A. Cimitile, and U. De Carlini. Reverse engineering and data flow diagrams in ADA environment. Microprocessing and Microprogramming, 30(1–5):357–364, August 1990. CODEN MMICDT. ISSN 0165-6074 (print), 1878-

7061 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016560749090267D.

Canfora:1991:REP

[CCD91] G. Canfora, A. Cimitile, and U. De Carlini. A reverse engineering process for design level document production from ADA code. Microprocessors and Microsystems, 15(10):531–542, December 1991. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0141933191900125.

Canfora:1993:REP

[CCD93] G. Canfora, A. Cimitile, and U. De Carlini. A reverse engineering process for design level document production from ADA code. Information and Software Technology, 35(1):23– 34, January 1993. CODEN ISOTE7. ISSN 0950-5849 (print),

1873-6025 (electronic). URL http://www.sciencedirect.
com/science/article/pii/095058499390026Y.

Cavalcanti:2011:CLD

[CCO11] Ana Cavalcanti, Phil Clayton, and Colin O’Halloran. From control law diagrams to Ada via Circus. Formal Aspects of Computing, 23(4):465–512, July 2011. CO-DEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.com/article/10.
1007/s00165-010-0170-3.

Carlisle:1987:AMS

[CCS87] Homer Carlisle, Albert Crawford, and Sallie Sheppard. ADA multitasking and the single source shortest path problem. Parallel Computing, 4(1):75–91, February 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0167819187900640.

Cooling:1997:CAT

[CDC97] J. E. Cooling, N. Duff, and J. Cooling. Computer aided teaching of programming languages: An Ada-specific development. Lecture Notes in Computer Science, 1251:35–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1983:OAC

[CDF^{+}83] Arvola Chan, Umeshwar Dayal, Stephen Fox, Nathan Goodman, Daniel R. Ries, and Dale Skeen. Overview of an Ada compatible distributed database manager. SIGMOD Record (ACM Special Interest Group on Management of Data), 13 (4):228–237, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Celier:1996:VCM

[Cel96] V. Celier. Visibility control and migration of interfaces in large Ada systems. Lecture Notes in Computer Science, 1088: 451–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Celko:2005:HWS

[Cel05] Joe Celko. How to Write Stored Procedures, chapter 8, pages 151–170. Elsevier, Amsterdam, The Netherlands, 2005.

URL http://www.sciencedirect.com/science/article/
pii/B978012088797250008X.

Clematis:1991:SDF

[CG91] Andrea Clematis and Vittoria Gianuzzi. Session D2: Fault tolerant parallel software. Microprocessing and Microprogramming, 32(1–5):365–372, August 1991. CODEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016560749190372Z.

Comar:1994:GPG

[CGS94] C. Comar, F. Gasperoni, and E. Schonberg. The GNAT project: a GNU-ADA94 compiler. Technique et Science Informatiques, 13(6):817–836, ???? 1994. CODEN TTSIDJ. ISSN 0752-4072, 0264-7419.

Cox:1980:ELA

[CH80] M. G. Cox and S. J. Hammarling. Evaluation of the language Ada for use in numerical computations. NPL Report DNACS 30/80, National Physical Laboratory, Teddington, UK, July 1980. ???? pp.

Cook:1997:AAS

[CH97] A. Cook and K. J. R. Hunt. ARINC 653- achieving software re-use. Microprocessors and Microsystems, 20(8):479– 483, April 1997. CODEN MIMID5. ISSN 0141-9331 (print),

1872-9436 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0141933197011137.

Clarke:2002:EIR

[CH02] Bruce Clarke and Linda Dalrymple Henderson, editors. From energy to information: representation in science and technology, art, and literature. Writing science. Stanford University Press, Stanford, CA, USA, 2002. ISBN 0-8047-4176-X, 0-8047-4210-3 (paperback). xviii + 440 pp. LCCN T15 .F78 2002. URL http://

www.loc.gov/catdir/enhancements/fy0710/2002070594-
b.html;   http://www.loc.gov/catdir/enhancements/
fy0710/2002070594-d.html;   http://www.loc.gov/catdir/
toc/fy033/2002070594.html.

Chartray:1985:ITA

[Cha85] Pierre Chartray. Une implantation des tâches de Ada. (French) [An implementation of tasks in Ada]. Maître ès sciences (m.sc.)., Université de Montréal, Montréal, QC, Canada, 1985. x + 249 pp.

Chelini:1992:DAR

[Che92] James V. Chelini. A discussion on the Ada Run-Time Environment in safety critical applications. ACM SIGSOFT Software Engineering Notes, 17(4):24–27, October 1992. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chen:1997:CAL

[Che97]  Jen-Yen Jason Chen.  CSPL: an Ada95-like, Unix-

based process environment. IEEE Transactions on Software Engineering, 23(3):171–184, March 1997. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=585504.

Chen:2012:CND

[CHLY12] Yuchuan Chen, Chien-Yeh Hsu, Li Liu, and Sherry Yang. Constructing a nutrition diagnosis expert system. Expert Systems with Applications, 39(2):2132–2156, February 1, 2012. CODEN ESAPEH. ISSN ???? URL

http://www.sciencedirect.com/science/article/pii/
S0957417411010311.

Chelini:1986:PSA

[CHR86] James V. Chelini, Edmund B. Hudson, and Stephen M. Reidy. A preliminary study of Ada expansion ratios. ACM SIGSOFT Software Engineering Notes, 11(1):35–46, January 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Christodoulakis:1991:ACA

[Chr91] Dimitrios Christodoulakis, editor. Ada, the choice for ’92: Ada-Europe International Conference, Athens, Greece, May 13–17, 1991: proceedings, volume 499 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. CO-DEN LNCSD9. ISBN 3-540-54092-X (Berlin), 0-387-54092-X (US). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A35 A24 1991. URL http://link.springer-

ny.com/link/service/series/0558/tocs/t0499.htm;
http://www.springerlink.com/content/978-0-387-54092-
4;   http://www.springerlink.com/openurl.asp?genre=
issue&issn=0302-9743&volume=499.

Czarnecki:2002:DPA

[CHR^{+}02] Krzysztof Czarnecki, Michael Himsolt, Ernst Richter, Falk Vieweg, and Alfred Rosskopf. DataFAN: a practical approach to data flow analysis for Ada 95. Lecture Notes in Computer Science, 2361:231–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2361/
23610231.htm;   http://link.springer-ny.com/link/
service/series/0558/papers/2361/23610231.pdf.

Chun:1996:SIR

[Chu96] R. K. Chun. Software integration of real-time expert systems. Control Engineering Practice, 4(1):83–88, January 1996. CODEN COEPEL. ISSN 0967-0661 (print), 1873-

6939 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0967066195002103.

Clark:1992:ULF

[CJ92] Robert G. Clark and Valerie M. Jones. Use of LOTOS in the formal development of an OSI protocol. Computer Communications, 15(2):86–92, March 1992. CODEN COCOD7. ISSN 0140-3664 (print), 1873-703X (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0140366492901282.

Choukair:1996:DOO

[CK96] Z. Choukair and Y. Kermarrec. Distributed object oriented programming and interoperability for Ada 95: An OMG/ CORBA approach. Lecture Notes in Computer Science, 1031: 217–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chedgey:1987:DAS

[CKK87] Chris Chedgey, Seamus Kearney, and Hans-Jurgen Kugler. Developing Ada software using VDM in an object-oriented framework. In USENIX Association [USE87], pages 41–58. ISBN ???? LCCN ????

Coon:1983:CCI

[CKS83] Lawrence A. Coon, John P. Kearns, and Mary Lou Soffa. The contraction of control implementations. Computer Languages, 8(1):15–25, ???? 1983. CODEN CO-LADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0096055183900036.

Cianci:1990:DIA

[CL90] Maria A. Cianci and Darrell G. Linton. The design and implementation of an Ada-based inference engine. Computers & industrial engineering, 19(1–4):107–110, ???? 1990. CO-DEN CINDDL. ISSN 0360-8352 (print), 1879-0550 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0360835290900862.

Case:2005:CWM

[CL05] Bettye Anne Case and Anne M. Leggett, editors. Complexities: Women in Mathematics. Princeton University Press, Princeton, NJ, USA, 2005. ISBN 0-691-11462-5, 0-691-17109-2, 1-4008-8016-5 (e-book). xix + 412 + 22 pp. LCCN QA27.5 .C66 2005. URL http://

www.loc.gov/catdir/enhancements/fy0668/2004048843-
b.html;   http://www.loc.gov/catdir/enhancements/
fy0668/2004048843-d.html;   http://www.loc.gov/catdir/
toc/fy0612/2004048843.html.

Cocco:1985:ATS

[CMM85] N. Cocco, D. Mandrioli, and V. Milanese. The Ada task system and real-time applications: an implementation schema. Computer Languages, 10(3–4):189–209, ???? 1985. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0096055185900165.

Cohen:1981:APL

[Coh81] Paul M. Cohen. Ada programming language standardization. The Journal of Systems and Software, 2(4):351–355, December 1981. CODEN JSSODM. ISSN 0164-1212 (print),

1873-1228 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0164121281900091.

Cohen:1986:ASL

[Coh86] Norman H. Cohen. ADA as a second language. McGraw-Hill series in software engineering and technology. McGraw-Hill, New York, NY, USA, 1986. ISBN 0-07-011589-3 (paperback). xxii + 838 pp. LCCN QA76.73.A35 C64 1986.

Collins:1984:CMP

[Col84] Steve Collins. Comparing MODULA-2 with PASCAL and ADA. Data Processing, 26(10):32–34, December 1984. CO-DEN ???? ISSN 0011-684X (print), 1878-3058 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0011684X84901795.

Collingbourne:1993:ATM

[Col93] L. (Lawrence) Collingbourne, editor. Ada: towards maturity, volume 6 of Studies in computer and communications systems, 0927-5444. IOS Press, Postal Drawer 10558, Burke, VA 2209-0558, USA, 1993. ISBN 90-5199-142-8. vii + 209 pp. LCCN QA76.73.A35 A3455 1993.

Conn:1986:ODA

[Con86] Richard Conn. Overview of DOD Ada software repository. Dr. Dobb’s Journal of Software Tools, 11(2):60–??, February 1986. CODEN DDJOEB. ISSN 1044-789X.

Conti:1988:SPF

[Con88] R. A. Conti. Software productivity features provided by the Ada language and the VAX Ada compiler. Digital Technical Journal of Digital Equipment Corporation, 1(6):51–61, February 1988. CODEN DTJOEL. ISSN 0898-901X.

Cooling:1996:LPR

[Coo96] J. E. Cooling. Languages for the programming of realtime embedded systems a survey and comparison. Microprocessors and Microsystems, 20(2):67–77, April 1996. CO-DEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/014193319501067X.

Corbett:1996:TAA

[Cor96] J. C. Corbett. Timing analysis of Ada tasking programs. IEEE Transactions on Software Engineering, 22(7):461–483, July 1996. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=538604.

Cross:1996:SIT

[CP96] James H. Cross II and Thomas M. Phillips. Successfully integrating traditional and object-oriented approaches with Ada 95. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(1):19–23, March 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Coen-Porisini:1993:ARS

[CPD93] Alberto Coen-Porisini and Flavio De Paoli. Array representation in symbolic execution. Computer Languages, 18(3):197– 216, ???? 1993. CODEN COLADA. ISSN 0096-0551 (print),

1873-6742 (electronic). URL http://www.sciencedirect.
com/science/article/pii/009605519390025V.

Chen:2013:ADA

[CQG^{+}13] Zhenzhong Chen, Haobo Qiu, Liang Gao, Liu Su, and Peigen Li. An adaptive decoupling approach for reliability-based design optimization. Computers and Structures, 117(??):58–66, February 2013. CODEN CMSTCJ. ISSN 0045-7949 (print),

1879-2243 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S004579491200301X.

Crawford:2000:AEO

[Cra00] Bard S. (Bard Stockton) Crawford. Ada essentials: overview, examples, and glossary, volume 1 of LearnAda Series. Trafford, Victoria, BC, Canada, 2000. ISBN 1-55212-371-5. xi + 227 pp. LCCN QA76.73 A16 C72 2000.

Cline:1985:AAS

[CS85] C. L. Cline and H. J. Siegel. Augmenting Ada for SIMD parallel processing. IEEE Transactions on Software Engineering, SE-11(9):970–977, September 1985. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=1702117.

Curtis:1991:MTS

[CS91] M. C. Curtis and V. J. D. Sivess. Modelling the temporal summation of neural membranes using the Ada language. Lecture Notes in Computer Science, 499:2–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carter:1998:PUA

[CS98] Jeffrey R. Carter and Bo I. Sanden. Practical use of Ada 95’s concurrency features. IEEE Concurrency, 6(4):47– 56, October/December 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/pd1998/pdf/p4047.pdf;
http://www.computer.org/concurrency/pd1998/p4047abs.
htm.

Craeynest:2001:RST

[CS01] Dirk Craeynest and Alfred Strohmeier, editors. Reliable software technologies: Ada-Europe 2001: 6th Ada-Europe International Conference on Reliable Software Technologies, Leuven, Belgium, May 14–18, 2001: Proceedings, volume 2043 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN LNCSD9. ISBN 3-540-42123-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A16 A23 2001; QA267.A1

L43 no.2043. URL http://link.springer-ny.com/link/
service/series/0558/tocs/t2043.htm;   http://www.
springerlink.com/content/978-3-540-42123-8;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=
0302-9743&volume=2043.

Craigen:1996:ACS

[CSM96] D. Craigen, M. Saaltink, and S. Michell. Ada 95 and critical systems: An analytical approach. Lecture Notes in Computer Science, 1088:171–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1994:ALS

[CT94a] Jen-Yen Chen and Chia-Ming Tu. An Ada-like software process language. The Journal of Systems and Software, 27(1):

17–25, October 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chen:1994:CPC

[CT94b] Jen-Yen Chen and Chia-Ming Tu. CSPL: a process-centred environment. Information and Software Technology, 36(1): 3–11, ???? 1994. CODEN ISOTE7. ISSN 0950-5849 (print),

1873-6025 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0950584994900035.

Cheng:1991:AAT

[CU91] J. Cheng and K. Ushijima. Analyzing Ada tasking deadlocks and livelocks using extended Petri nets. Lecture Notes in Computer Science, 499:125–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:1996:TDA

[CU96] J. Cheng and K. Ushijima. Tasking deadlocks in Ada 95 programs and their detection. Lecture Notes in Computer Science, 1088:135–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Culwin:1991:ADA

[Cul91] Fintan Culwin. ADA a development approach. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1991. ISBN 0-13-489147-3 (paperback). LCCN ????

Culwin:1997:ADA

[Cul97] Fintan Culwin. Ada: a developmental approach. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1997. ISBN 0-13-264680-3, 0-13-264680-3. xviii + 778 pp. LCCN ????

Curley:1991:ABA

[Cur91] T. Curley. An approach to benchmarking Ada compilation systems. Lecture Notes in Computer Science, 499:87–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Craeynest:1984:CES

[CVL84] Dirk Craeynest, Geert Vansteenkiste, and Johan Lewi. Construction of an ELL(1) syntax analyser for Ada with the compiler-generator LILA. ACM SIGPLAN Notices, 19(1):

36–45, January 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Compton:1990:PCA

[CW90] B. Terry Compton and Carol Withrow. Prediction and control of ADA software defects. The Journal of Systems and Software, 12(3):199–207, July 1990. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016412129090040S.

Ceruzzi:1991:RPI

[CW91] Paul Ceruzzi and Eric A. Weiss. Reviews: Penzias: Ideas and Information: Managing in a High-tech World; Rose: West of Eden: The End of Innocence at Apple Computer; Lindgren: Glory & Failure: the Difference Engines of Johann Müller, Charles Babbage, and Edvard Scheutz; Jennings: The Devouring Fungus: Tales of the Computer Age; Norberg: High Technology Calculation in the Early 20th Century: Punched Card Machinery in Business and Government; Hall and Barry: Sunburst: The Ascent of Sun Microsystems; Heppenheimer: How von Neumann Showed the Way; Aspray: Back to Basics: The Stored Program Concept; Rosen: The Origins of Modern Computing; Taylor: In Memoriam: J. C. R. Licklider, 1915–1990; Mounier-Kuhn: Genese de l’informatique en France (1945–1965): Diffusion de l’innovation et transfert de technologie; Eisler: My Life with the Printed Circuit. Annals of the History of Computing, 13(2):231–234, April/June 1991. CODEN AHCOE5.

ISSN 0164-1239. URL http://dlib.computer.org/an/
books/an1991/pdf/a2231.pdf;   http://www.computer.
org/annals/an1991/a2231abs.htm.

Cai:2004:SMC

[CW04] H. Cai and A. Wellings. Supporting mixed criticality applications in a Ravenscar–Java environment. Lecture Notes in Computer Science, 3292:278–291, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crawford:2006:FIC

[CWG^{+}06] Diane Crawford, Peter Wegner, Dina Goldin, Jean-Pierre Rosen, Curtis Rhodes, C. J. Fearnley, Peter J. Denning, Andrew McGettrick, Mike Segel, and Michael J. O’Donnell. Forum: Interactive computing is already outside the box; lack of Ada reflects software immaturity; be skeptical of rhetorical slight of hand; more to innovation than innovation alone; handles not a naming solution. Communications of the ACM, 49 (3):11–13, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Clarke:1980:NAP

[CWW80] Lori A. Clarke, Jack C. Wileden, and Alexander L. Wolf. Nesting in Ada programs is for the birds. ACM SIGPLAN Notices, 15(11):139–145, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2002:CAD

[CXYZ02] Zhenqiang Chen, Baowen Xu, Hongji Yang, and Jianjun Zhao. Concurrent Ada dead statements detection. Information and Software Technology, 44(13):733–741, October 1, 2002. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-

6025 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0950584902001064.

Chen:2002:SDA

[CXZY02] Zhenqiang Chen, Baowen Xu, Jianjun Zhao, and Hongji Yang. Static dependency analysis for concurrent Ada 95 programs. Lecture Notes in Computer Science, 2361:219–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2361/23610219.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2361/23610219.pdf.

DeLaPuente:1996:MHH

[DAA96] J. A. De La Puente, A. Alonso, and A. Alvarez. Mapping HRT-HOOD designs to Ada 95 hierarchical libraries. Lecture Notes in Computer Science, 1088:78–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demillo:1988:UMA

[DAG^{+}88] R. A. Demillo, W. F. Appelbe, D. S. Guindi, K. N. King, and W. M. Mccracken. Using mutation analysis for testing Ada programs. Technical Report SERC-TR-9-P, Software Engineering Research Centre, Utrecht, The Netherlands (??), February 3, 1988.

Davies:1987:FHL

[Dav87] Anthony C. Davies. Features of high-level languages for microprocessors. Microprocessors and Microsystems, 11(2):77– 87, March 1987. CODEN MIMID5. ISSN 0141-9331 (print),

1872-9436 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0141933187902079.

Dawes:1988:PPG

[Daw88] J. (John) Dawes. The professional programmers guide to Ada. Pitman Publishing Ltd., London, UK, 1988. ISBN 0-273-02821-9. 217 pp. LCCN QA76.73.A15.

Duri:1993:USS

[DBDS93] S. Duri, U. Buy, R. Devarapalli, and S. M. Shatz. Using state space reduction methods for deadlock analysis in Ada tasking. ACM SIGSOFT Software Engineering Notes, 18(3): 51–60, July 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Duhaut:1992:ILR

[DBF92] Dominique Duhaut, Philippe Bidaud, and Daniel Fontaine. IAda: a language for robot programming based on Ada. Robotics and Autonomous Systems, 9(4):299–304,

???? 1992. CODEN ????  ISSN ????  URL

http://www.sciencedirect.com/science/article/pii/
092188909290045Z.

Dowson:1979:SSM

[DCM79] Mark Dowson, Brian Collins, and Brian McBride. Software strategy for multiprocessors. Microprocessors and Microsystems, 3(6):263–266, July/August 1979. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0141933179901601.

Delves:1987:DGQ

[DD87] L. M. Delves and B. G. S. Doman. The design of a generic quadrature library in Ada. In Numerical integration (Halifax, N.S., 1986), volume 203 of NATO Adv. Sci. Inst. Ser. C Math. Phys. Sci., pages 307–319. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1987.

DeBondeli:1996:AVS

[De 96] P. De Bondeli. An Ada 95 view of some difficult or controversial points in object-oriented programming. Lecture Notes in Computer Science, 1088:370–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dewar:1980:NAT

[DFS^{+}80] Robert B. K. Dewar, Gerald A. Fisher, Jr., Edmond Schonberg, Robert Froehlich, Stephen Bryant, Clinton F. Goss, and Michael Burke. The NYU Ada translator and interpreter. ACM SIGPLAN Notices, 15(11):194–201, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Downes:1980:UAL

[DG80] V. A. Downes and S. J. Goldsack. The use of the Ada language for programming a distributed system. Annual Review in Automatic Programming, 10(??):39–44, ???? 1980. CO-DEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0066413882900064.

Downes:1982:PES

[DG82] Valerie A. Downes and Stephen J. Goldsack. Programming embedded systems with Ada. Prentice-Hall International, Upper Saddle River, NJ 07458, USA, 1982. ISBN 0-13-730010-7. xv + 377 pp. LCCN QA76.73.A35 D68. US$16.95.

Doberkat:1987:SAT

[DG87] Ernst-Erich Doberkat and Ulrich Gutenbeil. SETL to ADA-tree transformations applied. Information and Software Technology, 29(10):548–557, December 1987. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0950584987900887.

Duncan:1980:UAI

[DH80] A. G. Duncan and J. S. Hutchison. Using Ada for industrial embedded microprocessor applications. ACM SIGPLAN Notices, 15(11):26–35, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dhama:1995:QMC

[Dha95] Harpal Dhama. Quantitative models of cohesion and coupling in software. The Journal of Systems and Software, 29(1):65– 74, April 1995. CODEN JSSODM. ISSN 0164-1212 (print),

1873-1228 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016412129400128A.

Diaz-Herrera:1992:AMK

[DHGR92] Jorge L. Díaz-Herrera, Ronald D. Graft, and Douglas B. Rupp. ARTK: M2: a kernel for Ada tasking requirements: an implementation and an automatic generator. Software— Practice and Experience, 22(4):317–348, April 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Diaconescu:2011:PWM

[Dia11] Ada Diaconescu, editor. Proceedings of the 1st Workshop on Middleware and Architectures for Autonomic and Sustainable Computing. ACM Press, New York, NY 10036, USA, 2011. ISBN 1-4503-0847-3. LCCN ????

Dillon:1990:VGS

[Dil90a] L. K. Dillon. Verifying general safety properties of Ada tasking programs. IEEE Transactions on Software Engineering, 16(1):51–63, January 1990. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=44363.

Dillon:1990:USE

[Dil90b] Laura K. Dillon. Using symbolic execution for verification of Ada tasking programs. ACM Transactions on Programming Languages and Systems, 12(4):643–669, October 1990. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
0164-0925/96551.html.

Dillon:1991:IAS

[Dil91] Laura K. Dillon. An isolation approach to symbolic execution-based verification of Ada tasking programs. The Journal of Systems and Software, 14(3):183–198, March 1991.

CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016412129190065E.

Dillon:1993:VEM

[Dil93] Laura K. Dillon. A visual execution model for Ada tasking. ACM Transactions on Software Engineering and Methodology, 2(4):311–345, October 1993. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/journals/tosem/1993-2-4/
p311-dillon/p311-dillon.pdf;   http://www.acm.org/
pubs/citations/journals/tosem/1993-2-4/p311-dillon/

.

Diaz:2005:GDR

[DLGF05] Irene Díaz, Juan Llorens, Gonzalo Genova, and J. Miguel Fuentes. Generating domain representations using a relationship model. Information systems, 30(1):1–19, March 2005. CODEN INSYD6. ISSN 0306-4379 (print), 1873-6076 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0306437903000760.

Domenici:1989:PRL

[DLP89] Andrea Domenici, Beatrice Lazzerini, and Cosimo Antonio Prete. A protocol for resource locking and deadlock detection in a multi-user environment. Microprocessing and Microprogramming, 27(1–5):431–437, August 1989. CO-DEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165607489900872.

Dixon:1987:UEO

[DM87] Lawrence C. W. Dixon and M. Mohseninia. The use of the extended operations set of Ada with automatic differentiation and the truncated Newton method. Technical Report NOC TR176, The Numerical Optimisation Center, Hatfield Polytechnic, Hatfield, UK, April 1987. ?? pp.

Dixon:1988:FEO

[DMM88] Lawrence C. W. Dixon, Z. Maany, and M. Mohseninia. Finite element optimization in ADA using automatic differentiation. Technical Report NOC TR205, The Numerical Optimisation Center, Hatfield Polytechnic, Hatfield, UK, ???? 1988. ?? pp.

Dixon:1990:ADL

[DMM90] L. C. W. Dixon, Z. Maany, and M. Mohseninia. Automatic differentiation of large sparse systems. Journal of Economic Dynamics and Control, 14(2):299–311, May 1990. CO-DEN JEDCDH. ISSN 0165-1889 (print), 1879-1743 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016518899090023A.

DelaPeyronnie:2010:MEA

[DNM^{+}10] Jerome DelaPeyronnie, Philip H. Newcomb, Vincent Morillo, Fakkhredine Trimech, Luong Nguyen, and Mark Purtill. Modernization of the Eurocat Air Traffic Management System (EATMS). In William M. Ulrich and Philip H. Newcomb, editors, Information Systems Transformation: Architecture-Driven Modernization Case Studies, chapter 5, pages 91–131. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2010. ISBN 0-12-374913-1.

URL http://www.sciencedirect.com/science/article/
pii/B9780123749130000056.

dOcagne:1986:CSG

[d’O86] Maurice d’Ocagne. Le calcul simplifié: graphical and mechanical methods for simplifying calculation, volume 11 of Charles Babbage Institute reprint series for the history of computing. MIT Press, Cambridge, MA, USA, 1986. ISBN 0-262-15032-8. xiv + x + 167 pp. LCCN QA75 .O1513 1986. Translation from French by J. Howlett and M. R. Williams.

Duggan:2002:OCS

[DO02] Dominic Duggan and John Ophel. Open and closed scopes for constrained genericity. Theoretical Computer Science, 275(1–2):215–258, March 28, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304397501001293.

Doman:1995:SAP

[DPC95] B. G. S. Doman, C. J. Pursglove, and W. M. Coen. A set of Ada packages for high precision calculations. ACM Transactions on Mathematical Software, 21(4):416–431, December 1995. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295

(electronic). URL http://www.acm.org/pubs/citations/
journals/toms/1995-21-4/p416-doman/.

David:1996:DFT

[DPCC96] P. David, T. Planche, A. Correge, and J. F. Chane. Developing fault tolerant software in Ada for real-time dependable systems. Lecture Notes in Computer Science, 1031:21–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davies:1996:UAE

[DR96] B. Davies and D. A. Rothwell. The use of Ada for the ENVISAT-l simulator. Lecture Notes in Computer Science, 1031:170–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dobbing:1997:STS

[DRF97] B. Dobbing and M. Richard-Foy. T-Smart — task-safe, minimal Ada realtime toolset. Lecture Notes in Computer Science, 1251:244–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Druffel:1982:PEA

[Dru82] Larry E. Druffel. The potential effect of Ada on software engineering in the 1980’s. ACM SIGSOFT Software Engineering Notes, 7(3):5–11, July 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Damerla:1992:SCA

[DS92] Srinivasarao Damerla and Sol M. Shatz. Software complexity and Ada rendezvous: Metrics based on nondeterminism. The Journal of Systems and Software, 17(2):119–127, February 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-

1228 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0164121292900907.

Duell:1992:UAS

[DSd92] Rob Duell, Hugo J. Sebel, and Franklin C. A. de Wit. Using Ada source code generators in a large project. Lecture Notes in Computer Science, 603:47–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeAcetis:1990:UAT

[DSK90] Louis A. DeAcetis, Oron Schmidt, and Kumar Krishen. Using ADA tasks to simulate operating equipment. Computers in Physics, 4(5):521–??, September 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168387.

Delis:1991:DAA

[DT91] A. Delis and W. M. Thomas. Design assessment of Ada systems using static analysis. Lecture Notes in Computer Science, 499:227–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubery:1985:MRR

[Dub85] J. M. Dubery. Modulas-1, -2, and -3(?) or a real realtime language. Annual Review in Automatic Programming, 13 (part 1)(??):81–89, ???? 1985. CODEN AR-VAAM. ISSN 0066-4138 (print), 1878-545X (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0066413885904483.

Duncan:1982:PAC

[Dun82] Arthur G. Duncan. Prototyping in ADA: a case study. ACM SIGSOFT Software Engineering Notes, 7(5):54–60, December 1982. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

deVivo:1995:PAC

[dVdV95] Gabriela O. de Vivo and Marco de Vivo. A pragmatic approach to C++, Eiffel and Ada 9X programming. ACM SIG-PLAN Notices, 30(9):9–16, September 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dai:1999:ABO

[DX99] Guilan Dai and Baowen Xu. An Ada-based objectoriented modeling language. ACM SIGPLAN Notices, 34 (12):47–56, December 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/b/Baowen:Xu.html;   http:/
/ftp.informatik.rwth-aachen.de/dblp/db/indices/a-
tree/d/Dai:Guilan.html.

Eastman:1983:LAK

[Eas83] C. M. Eastman. A lexical analysis of keywords in high level programming languages. International Journal of Man-Machine Studies, 19(6):595–607, December 1983. CODEN IJMMBC. ISSN 0020-7373. URL

http://www.sciencedirect.com/science/article/pii/
S002073738380073X.

ECMA-162

[ECM97] ECMA. ECMA-162: Portable Common Tool Environment (PCTE) — Ada Programming Language Binding. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, fourth edi-

tion, December 1997. URL http://www.ecma.ch/ecma1/
STAND/ECMA-162.HTM.

EspinosaMinguet:2002:ABI

[EGC02] Agustín Espinosa Minguet, Ana García Fornes, and Alfons Crespo i Lorente. An Ada binding to the IEEE 1003.1q (POSIX Tracing) Standard. Lecture Notes in Computer Science, 2361:321–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2361/
23610321.htm;   http://link.springer-ny.com/link/
service/series/0558/papers/2361/23610321.pdf.

Ernst:1991:MVA

[EHMO91] George W. Ernst, Raymond J. Hookway, James A. Menegay, and William F. Ogden. Modular verification of Ada generics. Computer Languages, 16(3–4):259–280, ???? 1991. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0096055191900107.

Einarsson:1990:AML

[Ein90] Bo Einarsson. Application of mixed language programming. Computer Physics Communications, 61(1–2):150–162, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/001046559090113F.

Eckart:1987:OAL

[EL87] J. Dana Eckart and Richard J. LeBlanc. Overloading in the Ada language: Is it too restrictive? Computer Languages, 12(3–4):163–171, ???? 1987. CODEN CO-LADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0096055187900208.

Evans:1999:LQM

[EMB^{+}99] John Evans, David B. Moffett, Randall Brukardt, Richard Kuhn, Bill Dietrich, David Budgen, and Soheil Khajenoori. Letters: Quality musings; what about Ada?; the art of design; noting another body of knowledge. IEEE Software, 16(6):9–12, November/December 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:

//dlib.computer.org/so/books/so1999/pdf/s6009.pdf.

Emery:1995:LDR

[Eme95] David Emery. Letters: “Defects” the result of a rigorous Ada culture. IEEE Software, 12(4):7–8, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Emery:1998:AAB

[EMN98] David E. Emery, Robert F. Mathis, and Karl A. Nyberg. Automating the Ada binding process for Java — how far can we go? Lecture Notes in Computer Science, 1411:29–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/1411/14110029.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/1411/14110029.pdf.

Eassa:1994:ADA

[EOAm94] F. E. Eassa, L. J. Osterweil, and M. Z. Abdel-mageed. AIDA: a dynamic analyser for Ada programs. Information and Software Technology, 36(2):107–117, ???? 1994. CO-DEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0950584994900906.

Eassa:1995:ADA

[EOM95] F. E. Eassa, L. J. Osterweil, and M. Z. Abdel Mageed. AIDA: a dynamic analyzer for Ada programs. The Journal of Systems and Software, 31(3):239–255, December 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0164121294001014.

Evans:1985:IAP

[EP85] Howard Evans and Wayne Patterson. Implementing Ada as the primary programming language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 17 (1):255–265, March 1985. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Edmunds:2012:FMA

[ERB12] Andrew Edmunds, Abdolbaghi Rezazadeh, and Michael Butler. Formal modelling for Ada implementations: Tasking event-B. Lecture Notes in Computer Science, 7308:119–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-30598-6_9/.

Erdmann:2002:GAD

[Erd02] Michael Erdmann. GNAT Ada database development environment. Lecture Notes in Computer Science, 2361:334–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2361/23610334.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2361/23610334.pdf.

Elsesser:1986:MSC

[EST86] G. W. Elsesser, M. S. Safran, and T. Tieger. Managing separate compilation in AT&T’s UNIX Ada system. In USENIX Association [USE86b], pages 252–260.

Evans:1984:CPL

[Eva84] Arthur Evans, Jr. A comparison of programming languages Ada, Pascal, C. In Feuer and Gehani [FG84], page ?? ISBN 0-13-154840-9 (paperback), 0-13-154857-3 (hardcover). LCCN QA76.73.A35 C66 1984.

Evanco:1995:MEC

[Eva95] William M. Evanco. Modeling the effort to correct faults. The Journal of Systems and Software, 29(1):75–84, April 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-

1228 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016412129400129B.

Evanco:1997:PAD

[Eva97] William M. Evanco. Poisson analyses of defects for small software components. The Journal of Systems and Software, 38(1):27–35, July 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0164121297000630.

Elrad:1991:UAR

[EW91] T. Elrad and V. Winans. The use of Ada in reactive systems: a 3-dimensional model. Lecture Notes in Computer Science, 499:20–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fagin:2000:UAB

[Fag00] Barry Fagin. Using Ada-based robotics to teach computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):148–151, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fairley:2007:ICS

[Fai07] Richard E. (Dick) Fairley. The influence of COCOMO on software engineering education and training. The Journal of Systems and Software, 80(8):1201–1208, August 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121206002901.

Feldman:1984:ATA

[Fel84] Michael B. Feldman. Abstract types, ADA packages, and the teaching of data structures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 16 (1):183–189, February 1984. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 15th SIGCSE Technical Symposium on Computer Science Education.

Feldman:1990:TDS

[Fel90] Michael B. Feldman. Teaching data structures with Ada: an eight-year perspective. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(2):21– 29, June 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Feldman:1993:AAU

[Fel93] M. B. Feldman. Ada83 and Ada9x in the universities: filling the pipeline. In ACM [ACM93c], page 106. ISBN 0-89791-609-3. LCCN QA76.73.A16W37 1993.

Feldman:1997:ASR

[Fel97] M. B. Feldman. An Ada 95 sort race construction set. Lecture Notes in Computer Science, 1251:23–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feuer:1984:CAP

[FG84] Alan R. Feuer and Narain Gehani, editors. Comparing and Assessing Programming Languages: Ada, C, and Pascal. Prentice-Hall software series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1984. ISBN 0-13-154840-9 (paperback), 0-13-154857-3 (hardcover). xiv + 271 pp. LCCN QA76.73.A35 C66 1984.

Fritz:1988:USS

[FHK88] W. Fritz, V. Haase, and R. Kalcher. The use of standard software in real time programming: An example demonstrating the integration of ADA, ORACLE and GKS. Annual Review in Automatic Programming, 14 (part 1)(??):43–49, ???? 1988. CODEN ARVAAM. ISSN 0066-4138 (print),

1878-545X (electronic). URL http://www.sciencedirect.
com/science/article/pii/0066413888900079.

Francez:1986:SCA

[FHT86] Nissim Francez, Brent Hailpern, and Gadi Taubenfeld. Script: a communication abstraction mechanism and its verification. Science of Computer Programming, 6(??):35–88, ???? 1986. CODEN SCPGD4. ISSN 0167-6423 (print),

1872-7964 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0167642386900183.

FigueroadelCid:2000:RFF

[Fig00] Samuel Figueroa del Cid. A Rigorous Framework for Fully Supporting the IEEE Standard for Floating-Point Arithmetic in High-Level Programming Languages. Ph.D. thesis, Department of Computer Science, New York University, New York, NY, USA, January 2000. 345 pp. URL http://www.cs.

nyu.edu/csweb/Research/theses.html;   http://www.
cs.nyu.edu/csweb/Research/Theses/figueroa_sam.html;
 http://www.cs.nyu.edu/csweb/Research/Theses/figueroa_
sam.pdf;   http://www.cs.nyu.edu/csweb/Research/Theses/
figueroa_sam.ps.gz;   http://wwwlib.umi.com/dissertations/
fullcit/9956669;   http://wwwlib.umi.com/dissertations/
preview/9956669.

Fisher:1978:DCP

[Fis78] David A. Fisher. DoD’s common programming language effort. Computer, 11(3):24–33, March 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Feldman:1993:APS

[FK93] Michael B. Feldman and Elliot B. Koffman. Ada: problem solving and program design. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-52279-9. xvii + 795 pp. LCCN QA76.73.A35 F44 1993.

Feldman:1996:APS

[FK96a] Michael B. Feldman and Elliot B. Koffman. Ada 95: problem solving and program design. Addison-Wesley, Reading, MA, USA, second edition, 1996. ISBN 0-201-87009-6. xvii + 814 pp. LCCN QA76.73.A16 F44 1996.

Ferguson:1996:GTD

[FK96b] Roger Ferguson and Bogdan Korel. Generating test data for distributed software using the chaining approach. Information and Software Technology, 38(5):343–353, May 1996. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0950584995010645.

Feldman:1999:APS

[FK99] Michael B. Feldman and Elliot B. Koffman. Ada 95: problem solving and program design. Addison-Wesley, Reading, MA, USA, third edition, 1999. ISBN 0-201-36123-X (paperback). xv + 784 pp. LCCN QA76.73.A35 F43 1999. Includes CD-ROM.

Ford:1986:SA

[FKR86] B. Ford, J. Kok, and M. W. Rogers, editors. Scientific Ada. Cambridge University Press, ????, 1986. ISBN 0-521-33258-3. 388 pp.

Feldman:1990:SAP

[FLP90] Michael B. Feldman, Arthur Vargas Lopes, and Manuel Perez. SMALL-Ada: personal computer courseware for studying concurrent programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22 (1):206–211, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Foulkes:1987:AMD

[FM87] R. Foulkes and D. N. MacLean. Ada in a Mascot 3 design: final report. Yard, Glasgow, Scotland, 1987. LCCN ????

Feldman:1989:VDT

[FM89] M. B. Feldman and M. L. Moran. Validating a demonstration tool for graphics-assisted debugging of Ada concurrent programs. IEEE Transactions on Software Engineering, 15 (3):305–313, March 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?arnumber=21758.

Faria:2012:AMC

[FMP12] José Miguel Faria, João Martins, and Jorge Sousa Pinto. An approach to model checking Ada programs. Lecture Notes in Computer Science, 7308:105–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/chapter/10.1007/978-
3-642-30598-6_8/.

Franch:1997:INF

[Fra97] X. Franch. Including non-functional issues in Anna/Ada programs for automatic implementation selection. Lecture Notes in Computer Science, 1251:88–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franssens:2001:RAS

[Fra01] Ghislain R. Franssens. Retrieval of the aerosol size distribution in the complex anomalous diffraction approximation. Atmospheric Environment, 35(30):5099–5104, October 2001. CODEN ATENBP. ISSN 0004-6981 (print), 1878-2442 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1352231001003314.

Freedman:1982:PCA

[Fre82] Roy S. Freedman. Programming concepts with the Ada language. PBI series for the computer and data processing professional. Petrocelli Books, New York, NY, USA, 1982. ISBN 0-89433-190-6 (paperback). viii + 162 pp. LCCN QA76.73.A35 F73 1982.

Frasca-Spada:2000:BSH

[FSJ00] Marina Frasca-Spada and Nicholas Jardine, editors. Books and the sciences in history. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-65063-1 (hardcover), 0-521-65939-6 (paperback). xiv + 438 pp. LCCN Q125 .B623

2000. URL http://www.loc.gov/catdir/description/
cam0210/99087281.html;   http://www.loc.gov/catdir/
toc/cam021/99087281.html.

Fosdick:1989:BFA

[FSO89] Lloyd Dudley Fosdick, C. J. C. Schauble, and Kurt M. Olender. Benchmarking Fortran and Ada programs on parallel machines. Technical report CU-CS-420-89, University of Colorado, Boulder, Dept. of Computer Science, Boulder, CO, USA, January 1989. 47 pp.

Feith:1996:PTA

[FT96] R. Feith and M. Tonndorf. Planning the transition to Ada 95 for a large real-time project. Lecture Notes in Computer Science, 1088:500–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fussichen:1990:GAM

[Fus90] Kenneth Fussichen. Getting Ada into the mainstream in the 1990’s. In ACM [ACM90], page 428. ISBN 0-89791-409-0. LCCN QA76.73.A35.

Fenton:1991:PSS

[FW91] N. E. Fenton and R. W. Whitty. Program structures: Some new characterisations. Journal of Computer and System Sciences, 43(3):467–483, December 1991. CODEN JC-SSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/002200009190025Z.

Fix:1996:ITA

[FW96] Vikki Fix and Susan Wiedenbeck. An intelligent tool to aid students in learning second and subsequent programming languages. Computers and Education, 27(2):71–83, September 1996. CODEN COMEDR. ISSN 0360-1315 (print), 1873-

782X (electronic). URL http://www.sciencedirect.com/
science/article/pii/036013159600022X.

Freeman:1984:CSD

[FWH84] Peter Freeman, Anthony I. Wasserman, and Raymond C. Houghton. Comparing software development methodologies for Ada: a study plan. ACM SIGSOFT Software Engineering Notes, 9(4):22–55, July 1984. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Goos:1983:DIL

[G^{+}83] Gerhard Goos et al., editors. DIANA: an intermediate language for ADA, volume 161 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., revised edition, 1983. CODEN LNCSD9. ISBN 3-540-12695-3 (Berlin), 0-387-12695-3 (US). ISSN 0302-9743 (print), 1611-3349 (electronic). vii + 201 pp. LCCN QA76.73.A16

D5 1983. URL http://link.springer-ny.com/link/
service/series/0558/tocs/t0161.htm;   http://www.
springerlink.com/content/978-0-387-12695-1;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=
0302-9743&volume=161.

Gallmeister:1991:RUA

[Gal91] William O. Gallmeister. Reconciling UNIX, Ada, and realtime processing. Dr. Dobb’s Journal of Software Tools, 16 (6):56, 58, 60, 62, 64, June 1991. CODEN DDJOEB. ISSN 1044-789X.

Gart:1986:TAU

[Gar86] Mitchell Gart. Targeting Ada to 68000/Unix. In USENIX Association [USE86a], pages 261–274.

Gauthier:1993:APC

[Gau93a] Michel Gauthier. Ada: a professional course. MacMillan Publishing Company, New York, NY, USA, 1993. ISBN 0-333-58001-x, 0-333-58001-x. 480 pp. LCCN ????

Gauthier:1993:AEA

[Gau93b] Michel J. Gauthier, editor. Ada-Europe ’93: 12th Ada-Europe International Conference, “Ada Sans Frontières”, Paris, France, June 14–18, 1993: proceedings, volume 688 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNCSD9. ISBN 3-540-56802-6 (Berlin), 0-387-56802-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A35

A24 1993. URL http://link.springer-ny.com/link/
service/series/0558/tocs/t0688.htm;   http://www.
springerlink.com/content/978-3-540-56802-5;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=
0302-9743&volume=688.

Geiring:1994:GAR

[GB94] E. W. Geiring and T. Baker. The GNU Ada Runtime Library (GNARL). In ACM [ACM94a], pages 97–107. ISBN 0-89791-684-0. LCCN ????

Gonzalez-Barahona:1998:BMC

[GBdlHQCGB98] Jesús M. González-Barahona, Pedro de-las Heras-Quirós, José Centeno-González, and Francisco Ballesteros. Building modular communication systems in Ada: The Simple Com approach. Lecture Notes in Computer Science, 1411:225–237, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/1411/14110225.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/1411/14110225.pdf.

Gayer:1987:CPA

[GBO87] Richard Gayer, Catherine M. Beise, and G. Scott Owen. Conversion of a PDP-11/40 assembler and simulator from mainframe Pascal to Ada on IBM PC microcomputers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 19(1):378–381, February 1987. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gehani:1984:CPA

[GC84] Narain H. Gehani and T. A. Cargill. Concurrent programming in the Ada language: The polling bias. Software—Practice and Experience, 14(5):413–427, January 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gerth:1984:PSC

[GD84] Rob Gerth and Willem P. De Roever. A proof system for concurrent ADA programs. Science of Computer Programming, 4(2):159–204, August 1984. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0167642384900182.

GonzalezHarbour:1999:RST

[GdlP99] Michael González Harbour and Juan A. de la Puente, editors. Reliable software technologies — Ada-Europe ’99: 1999 Ada-Europe International Conference on Reliable Software Technologies, Santander, Spain, June 7–11, 1999: Proceedings, volume 1622 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNCSD9. ISBN 3-540-66093-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.R44 A33 1999; QA267.A1 L43 no.1622. URL http://link.springer-

ny.com/link/service/series/0558/tocs/t1622.htm;
http://uclibs.org/PID/11342;   http://www.springerlink.
com/content/978-3-540-66093-4;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=
1622.

Gehani:1982:CAM

[Geh82] N. H. Gehani. Concurrency in Ada and multicomputers. Computer Languages, 7(1):21–23, ???? 1982. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0096055182900182.

Gehani:1983:AAI

[Geh83] Narain Gehani. Ada: an advanced introduction. Prentice-Hall software series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1983. ISBN 0-13-003962-4 (paperback). xvii + 325 pp. LCCN QA76.73.A35 G43 1983.

Gehani:1984:AAI

[Geh84a] Narain Gehani. Ada, an advanced introduction: including reference manual for the Ada programming language. Prentice-Hall software series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1984. ISBN 0-13-003997-7. various pp. LCCN QA76.73.A35 G43 1984; QA76.73.A35G43 1984.

Gehani:1984:ACP

[Geh84b] Narain Gehani. Ada: concurrent programming. Prentice-Hall software series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1984. ISBN 0-13-004011-8. 261 pp. LCCN QA76.73.A35 G433 1984.

Gehani:1987:UAP

[Geh87] Narain Gehani. UNIX Ada programming. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1987. ISBN 0-13-938325-5 (paperback). xix + 310 pp. LCCN QA76.73.A35 G437 1987. Revised edition of [Geh84a].

Gehani:1989:AAI

[Geh89] Narain Gehani. Ada: an advanced introduction. Prentice-Hall software series. Prentice-Hall International, Upper Saddle River, NJ 07458, USA, second edition, 1989. ISBN 0-13-004334-6, 0-13-004359-1. xvii + 280 pp. LCCN QA76.73.A35 G43 1989.

German:1984:MDB

[Ger84] S. M. German. Monitoring for deadlock and blocking in Ada tasking. IEEE Transactions on Software Engineering, SE-10(6):764–777, November/December 1984. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=5010305.

Gini:1982:ALR

[GG82] Giuseppina Gini and Maria Gini. ADA: a language for robot programming? Computers in industry, 3(4):253–259, December 1982. CODEN CINUD4. ISSN 0166-3615 (print),

1872-6194 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0166361582900136.

GonzalezHarbour:1997:IAL

[GGP97] M. Gonzalez Harbour, J. J. Gutierrez Garcia, and J. C. Palencia Gutierrez. Implementing application-level sporadic server schedulers in Ada 95. Lecture Notes in Computer Science, 1251:125–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glaser:1993:DAM

[GH93] H. Glaser and R. Harrison. Data abstraction in Modula-2. Information and Software Technology, 35(11–12): 619–626, November/December 1993. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
095058499390076F.

Ghezzi:1985:CPL

[Ghe85] Carlo Ghezzi. Concurrency in programming languages: a survey. Parallel Computing, 2(3):229–241, November 1985. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0167819185900055.

Gicca:2009:PAS

[Gic09] Greg Editor Gicca, editor. Proceedings of the Acm Sigada Annual International Conference on Ada and Related Technologies. ACM Press, New York, NY 10036, USA, 2009. ISBN 1-60558-475-4. LCCN ???? USD 22.00 Re-

tail Price (Publisher). URL http://DC8QA4CY3N.search.
serialssolutions.com/?V=1.0&L=DC8QA4CY3N&S=JCs&C=
TC_011597804&T=marc&tab=BOOKS.

Gilpin:1986:AGT

[Gil86] Geoff Gilpin. Ada, a guided tour and tutorial. Prentice Hall Press, New York, NY, USA, 1986. ISBN 0-13-730599-0 (paperback). xiv + 450 pp. LCCN QA76.73.A35 G55 1986; QA76.73.A35G55 1986.

Gannon:1986:MAP

[GKB86] J. D. Gannon, E. E. Katz, and V. R. Basili. Metrics for Ada packages: an initial study. Communications of the ACM, 29 (7):616–623, July 1986. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0001-0782/6144.html.

Gargaro:1996:PPA

[GKPT96] A. Gargaro, Y. Kermarrec, L. Pautet, and S. Tardieu. PARIS — partitioned Ada for remotely invoked services. Lecture Notes in Computer Science, 1031:191–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gliss:1996:AHC

[Gli96] B. Gliss. An Ada 95 harness for converting legacy Fortran applications. Lecture Notes in Computer Science, 1088:413– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ganapathi:1989:IAC

[GM89] M. Ganapathi and G. O. Mendal. Issues in Ada compiler technology. Computer, 22(2):52–60, February 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Guerra:1997:ALP

[GMAA97] F. Guerra, J. Miranda, A. Alvarez, and S. Arevalo. An Ada library to program fault-tolerant distributed applications. Lecture Notes in Computer Science, 1251:230–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giering:1993:IAF

[GMB93] E. W. Giering, F. Mueller, and T. P. Baker. Implementing Ada 9X features using POSIX threads: Design issues. In ACM [ACM93b], pages 214–228. ISBN 0-89791-621-2. LCCN ???? ACM Order No. 825930.

Guaspari:1990:FVA

[GMP90] D. Guaspari, C. Marceau, and W. Polak. Formal verification of Ada programs. IEEE Transactions on Software Engineering, 16(9):1058–1075, September 1990. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=58790.

Gallagher:1993:SSG

[GN93] M. J. Gallagher and V. L. Narasimhan. A software system for the generation of test data for Ada programs. Microprocessing and Microprogramming, 38(1–5):637–644, September 1993. CODEN MMICDT. ISSN 0165-6074 (print), 1878-

7061 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016560749390204X.

Gallagher:1997:ATD

[GN97] M. J. Gallagher and V. Lakshmi Narasimhan. ADTEST: a test data generation suite for Ada software systems. IEEE Transactions on Software Engineering, 23(8):473–484, August 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=624304.

Gomaa:1994:SDM

[Gom94] Hassan Gomaa. Software design methods for the design of large-scale real-time systems. The Journal of Systems and Software, 25(2):127–146, May 1994. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0164121294900027.

Goodenough:1980:ACV

[Goo80] John B. Goodenough. The Ada compiler validation capability. ACM SIGPLAN Notices, 15(11):1–8, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ganzinger:1980:OIA

[GR80] Harald Ganzinger and Knut Ripken. Operator identification in ADA: formal specification, complexity, and concrete implementation. ACM SIGPLAN Notices, 15(2):30–42, February 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See comment [Jan80].

Gehani:1988:RFC

[GR88] N. H. Gehani and W. D. Roome. Rendezvous facilities: Concurrent C and the Ada language. IEEE Transactions on Software Engineering, 14(11):1546–1553, November 1988.

CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=9043.

Gran:1988:HAF

[Gra88] Pamela K. Gran. A hybrid archetype for FORTRAN to ADA amelioration. Thesis (m.s.), National University, Vista, CA, USA, 1988. vii + 76 pp.

Greenwood:1986:CVT

[Gre86] J. R. Greenwood. Comments on “A View from the Trenches”. Ada vs. Modula-2 vs. Praxis. ACM SIGPLAN Notices, 21(5): 45–49, May 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [Bie85a].

GonzalezHarbour:1998:IUE

[GRGG98] M. Gonzalez Harbour, M. A. Rivas, J. J. Gutierrez Garcia, and J. C. P. Gutierrez. Implementing and using execution time clocks in Ada hard real-time applications. Lecture Notes in Computer Science, 1411:90–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Groeneveld:1992:UAI

[Gro92] Miech Groeneveld. Using Ada in integrating ATC systems. Lecture Notes in Computer Science, 603:72–??, 1992. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1985:ESM

[GS85] Rajiv Gupta and Mary Lou Soffa. The efficiency of storage management schemes for Ada programs. ACM SIGPLAN Notices, 20(11):30–38, November 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gregertsen:2010:INA

[GS10] Kristoffer Nyborg Gregertsen and Amund Skavhaug. Implementing the new Ada 2005 timing event and execution time control features on the AVR32 architecture. Journal of Systems Architecture, 56(10):509–522, October 2010. CO-DEN JSARFB. ISSN 1383-7621 (print), 1873-6165 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1383762110000937.

Green:2001:TPI

[GST01] Christopher D. Green, Marlene Gay Shore, and Thomas Teo, editors. The transformation of psychology: influences of 19th century philosophy, technology, and natural science. American Psychological Association, Washington, DC, USA, 2001. ISBN 1-55798-776-9. xvii + 245 pp. LCCN BF103 .T73 2001.

Gedela:1999:CPN

[GSX99] Ravi K. Gedela, Sol M. Shatz, and Haiping Xu. Compositional Petri net models of advanced tasking in Ada-

95. Computer Languages, 25(2):55–87, July 1999. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (elec-

tronic). URL http://www.elsevier.nl/gej-ng/10/15/
18/28/27/25/abstract.html;   http://www.elsevier.
nl/gej-ng/10/15/18/28/27/25/article.pdf;   http://
www.sciencedirect.com/science/article/pii/S0096055199000144.

Gilbert:1991:EDT

[GTB91] B. Gilbert, M. Taylor, and G. Bek. Experience developing two Ada applications for embedded real-time systems using different software processes. Lecture Notes in Computer Science, 499:394–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gobin:1992:RSS

[GTG92] M. Gobin, M. Timmerman, and F. J. A. Gielen. Runtime system support for data-oriented synchronization in Ada-9X. Lecture Notes in Computer Science, 603:128–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guilain:1994:VNS

[GV94] S. Guilain and J. Vignes. Validation of numerical software results — application to the computation of apparent heat release in direct-injection diesel engines. Mathematics and Computers in Simulation, 37(1):73–92, November 21, 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0378475494900604.

Garcia-Valls:2012:IMT

[GVIV12] Marisol García-Valls and Felipe Ibáñez-Vázquez. Integrating middleware for timely reconfiguration of distributed soft real-time systems with Ada DSA. Lecture Notes in Computer Science, 7308:35–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30598-6_3/.

Gautier:1990:SRA

[GW90] R. J. Gautier and Peter J. L. Wallis, editors. Software reuse with ADA, volume 16 of IEE computing series. Peter Peregrinus on behalf of the Institution of Electrical Engineers, London, UK, 1990. ISBN 0-86341-173-8. xvi + 205 pp. LCCN IN PROCESS (ONLINE).

Gothe:1991:DAR

[GWA91] M. Carlsson Göthe, D. Wengelin, and L. Asplund. The Distributed Ada Run-time System DARTS. Software—Practice and Experience, 21(11):1249–1263, November 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gannon:1987:TIM

[GZ87] John D. Gannon and Marvin V. Zelkowitz. Two implementation models of abstract data types. Computer Languages, 12(1):21–25, ???? 1987. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0096055187900099.

Huzar:1998:A

[H^{+}98] Zbigniew Huzar et al. Ada 95. “Helion”, Gliwice, Poland, 1998. ISBN 83-86718-38-2. 372 pp. LCCN ????

Halang:1983:RTF

[Hal83] Wolfgang A. Halang. On real-time features available in highlevel languages and yet to be implemented. Microprocessing and Microprogramming, 12(2):79–87, September 1983. CO-DEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165607483901850.

Hartree:1949:CIM

[Har49] Douglas R. (Douglas Rayner) Hartree. Calculating Instruments and Machines. University of Illinois Press, Urbana, IL, USA, 1949. ix + 138 pp. LCCN QA85 .H29.

Hartree:1984:CMR

[Har84] Douglas Rayner (Douglas Rayner) Hartree. Calculating machines: recent and prospective developments and their impact on mathematical physics; and, Calculating instruments and machines, volume 6 of Charles Babbage Institute reprint series for the history of computing. MIT Press, Cambridge, MA, USA, 1984. ISBN 0-262-08147-4. xvi + 40; viii + 138 + 3 pp. LCCN QA76.95; QA85 .H3 1984. With a new introduction by Maurice V. Wilkes.

Hardy:1997:RST

[HB97] Keith Hardy and Jim Briggs, editors. Reliable software technologies — Ada-Europe ’97: 1997 Ada-Europe International Conference on Reliable Software Technologies, London, UK, June 2–6, 1997: proceedings, volume 1251 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNCSD9. ISBN 3-540-63114-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A16 A23 1997. URL http://link.springer-

ny.com/link/service/series/0558/tocs/t1251.htm;
http://uclibs.org/PID/126855;   http://www.springerlink.
com/content/978-3-540-63114-9;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=
1251.

Harbour:1999:RST

[HD99] M. G. Harbour and J. A. De la Puente. Reliable software technologies Ada-Europe ’99. Lecture Notes in Computer Science, 1622:i–xiii, 1–449, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heilbrunner:1988:AIP

[Hei88] Stephan Heilbrunner, editor. Ada in industry: proceedings of the Ada-Europe International Conference, Munich 7–9 June 1988, The Ada companion series. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-36347-0. LCCN ????

Heitz:1996:ARR

[Hei96] M. Heitz. Achieving reusable and reliable client-server code using HOOD[TM] automated code generation for Ada95 and C++ targets. Lecture Notes in Computer Science, 1031:419– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hemmendinger:1990:SAS

[Hem90] D. Hemmendinger. Specifying Ada server tasks with executable formal grammars. IEEE Transactions on Software Engineering, 16(7):741–754, July 1990. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=56100.

Henry:1981:RTP

[Hen81] Roger Henry. Real-time programming languages. International Journal of Man-Machine Studies, 14(3):355– 369, April 1981. CODEN IJMMBC. ISSN 0020-7373.

URL http://www.sciencedirect.com/science/article/
pii/S0020737381800636.

Henno:1988:UFS

[Hen88] J. Henno. User-friendly syntax: design and presentation. International Journal of Man-Machine Studies, 28(5):551– 572, May 1988. CODEN IJMMBC. ISSN 0020-7373.

URL http://www.sciencedirect.com/science/article/
pii/S0020737388800609.

Hermann:1985:BRB

[Her85] Peter Hermann. Book review: Evaluating Ada: D. Le Verrand (North Oxford Academic, Oxford, 1985) 288 pp., ISBN 0-946536-15-5. Computer Methods in Applied Mechanics and Engineering, 53(2):201, November 1985. CO-DEN CMMECC. ISSN 0045-7825 (print), 1879-2138 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0045782585900064.

Hermann:1987:BRB

[Her87] Peter Hermann. Book review: Ada as a second language: Norman H. Cohen (McGraw-Hill, New York, 1986) ISBN 0-07-011589-3, 838 pp. Computer Methods in Applied Mechanics and Engineering, 65(3):293, December 1987. CO-DEN CMMECC. ISSN 0045-7825 (print), 1879-2138 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0045782587901617.

Husbands:2008:MMH

[HHW08] Philip Husbands, Owen Holland, and Michael Wheeler, editors. The mechanical mind in history. MIT Press, Cambridge, MA, USA, 2008. ISBN 0-262-08377-9 (hardcover). viii + 458 pp. LCCN Q335 .M3956 2008.

Hilfinger:1983:AML

[Hil83] Paul N. Hilfinger. Abstraction mechanisms and language design. ACM distinguished dissertations. MIT Press, Cambridge, MA, USA, 1983. ISBN 0-262-08134-2. xi + 176 pp. LCCN QA76.73.A15.

Hilfinger:1988:APD

[Hil88] Paul N. Hilfinger. An Ada package for dimensional analysis. ACM Transactions on Programming Languages and Systems, 10(2):189–203, April 1988. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0164-0925/42346.html.

Hilzer:1992:SPC

[Hil92] Ralph C. Hilzer, Jr. Synchronization of the producer/ consumer problem using semaphores, monitors, and the Ada rendezvous. Operating Systems Review, 26(3):31–39, July 1992. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hillam:1994:IAD

[Hil94] Bruce Hillam. Introduction to abstract data types using Ada. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-124215-6. xxi + 662 pp. LCCN ????

Hennessy:1983:TSA

[HL83] M. C. B. Hennessy and W. Li. Translating a subset of Ada into CCS. In Formal description of programming concepts, II (Garmisch-Partenkirchen, 1982), pages 227–249. North-Hol-land, Amsterdam, The Netherlands, 1983.

Helmbold:1985:DAT

[HL85] David Helmbold and David Luckham. Debugging Ada tasking programs. IEEE Software, 2(2):47–57, March 1985. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Herter:1993:ADD

[HL93] Thomas Herter and Klaus Lott. Algorithms for decomposing 3-D orthogonal matrices into primitive rotations. Computers and Graphics, 17(5):517–527, September/October 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/009784939390003R.

Heraud:2001:UAI

[HL01] Pascal Héraud and Thierry Lelégard. Using Ada in interactive digital television systems. Lecture Notes in Computer Science, 2043:21–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2043/
20430021.htm;   http://link.springer-ny.com/link/
service/series/0558/papers/2043/20430021.pdf.

Hisgen:1980:RRA

[HLRS80] Andy Hisgen, David Alex Lamb, Jonathan Rosenberg, and Mark Sherman. A runtime representation for Ada variables and types. ACM SIGPLAN Notices, 15(11):82–90, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Habermann:1987:SDA

[HM87] A. Nico Habermann and U. Montanari, editors. System development and Ada: CRAI Workshop on Software Factories and Ada, Capri, Italy, May 26–30 1986 proceedings, volume 275 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNCSD9. ISBN 0-387-18341-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.S88

W661 1986. URL http://link.springer-ny.com/link/
service/series/0558/tocs/t0275.htm;   http://www.
springerlink.com/content/978-0-387-18341-1;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=
0302-9743&volume=275.

Howden:1991:VCS

[HNVW91] William E. Howden, David Nesbitt, Cheron Vail, and Bruce Wieand. Verification of complex systems using incremental operational specifications. Information Sciences, 57-58(??):427–450, September/December 1991. CODEN ISIJBC. ISSN 0020-0255 (print), 1872-6291 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0020025591900929.

Holdsworth:1983:SAA

[Hol83] David Holdsworth. A system for analysing Ada programs at run-time. Software—Practice and Experience, 13(5):407– 421, May 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Holzmueller:1996:EOO

[Hol96] B. Holzmueller. Extending the object-orientedness of Ada 95. Lecture Notes in Computer Science, 1088:357–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hookway:1985:VAP

[Hoo85] Raymond J. Hookway. Verifying Ada programs. ACM SIGSOFT Software Engineering Notes, 10(4):51–52, August 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hook:1992:BRS

[Hoo92] Audrey A. Hook. Book review: Selecting an Ada Compilation System Edited by J. Dawes, M. J. Pickett, and A. Wearing: (Cambridge University Press 1990). ACM SIGPLAN Notices, 27(1):13, January 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Horsburgh:1982:HNT

[Hor82] E. M. (Ellice Martin) Horsburgh, editor. Handbook of the Napier Tercentenary Celebration, or, Modern instruments and methods of calculation, volume 3 of Charles Babbage Institute reprint series for the history of computing. Tomash Publishers, Los Angeles, CA, USA, 1982. ISBN 0-938228-10-2. xxi + vii + 343 + xii pp. LCCN QA71 .M62 1982. With a new introduction by Michael R. Williams.

Habermann:1983:AEP

[HP83] Arie Nicolaas Habermann and Dewayne E. Perry. Ada for experienced programmers. Addison-Wesley series in computer science. Addison-Wesley, Reading, MA, USA, 1983. ISBN 0-201-11481-X (paperback). v + 479 pp. LCCN QA 76.73 A35 H33 1983.

Habermann:1989:ADZ

[HP89] A. Nico Habermann and Dewayne E. Perry. ADA dla zaawansowanych. (Polish) [ADA for the advanced]. Biblioteka Inśynierii Oprogramowania. Wydawnictwa Naukowo-Techniczne, Warszawa, Poland, 1989. ISBN 83-204-1058-4. 459 + 5 pp. LCCN ????

Holzmueller:1997:FUA

[HP97] B. Holzmueller and E. Plödereder. Finite unions for Ada 95. Lecture Notes in Computer Science, 1251:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hagenauer:1998:ADS

[HP98] Helge Hagenauer and Werner Pohlmann. Ada 95 for a distributed simulation system. Lecture Notes in Computer Science, 1411:140–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/1411/
14110140.htm;   http://link.springer-ny.com/link/
service/series/0558/papers/1411/14110140.pdf.

Harbour:1998:IUE

[HRGG98] M. González Harbour, M. Aldea Rivas, J. J. Gutiérrez García, and J. C. Palencia Gutiérrez. Implementing and using execution time clocks in Ada hard real-time applications. Lecture Notes in Computer Science, 1411:90–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/1411/14110090.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/1411/14110090.pdf.

Hall:1992:ADR

[HSLG92] Gardiner A. Hall, James Schuetzle, David LaVallee, and Uday Gupta. An architecture for the development of real-time fault diagnosis systems using model-based reasoning. Telematics and Informatics, 9(3–4):163–172, Summer– Autumn 1992. CODEN TEINEG. ISSN 0736-5853 (print),

1879-324X (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0736585305800346.

Hollingsworth:1994:PIR

[HSWZ94] Joseph E. Hollingsworth, Sethu Sreerama, Bruce W. Weide, and Sergey Zhupanov. Part IV: RESOLVE components in Ada and C++. ACM SIGSOFT Software Engineering Notes, 19(4):52–63, October 1994. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Harrison:1996:IMD

[HT96] Alan Harrison and Peter G. Thomas. Integrating multiple and diverse abstract knowledge types in real-time embedded systems. Knowledge-Based Systems, 9(7):417– 434, November 1996. CODEN ???? ISSN ???? URL

http://www.sciencedirect.com/science/article/pii/
S0950705196010520.

Hughes:1991:MSE

[Hug91] David K. Hughes. Multilingual software engineering using Ada and C. ACM SIGSOFT Software Engineering Notes, 16 (4):55–59, October 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Hummel:1992:HPA

[Hum92] Susan Flynn Hummel. A highley parallel Ada task scheduler. Lecture Notes in Computer Science, 603:275–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hunter:1985:ARK

[Hun85] D. G. N. Hunter. Algorithm 121: RSA key calculation in Ada. The Computer Journal, 28(3):343–345, 1985. CODEN CMJOA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Huss:1990:ALA

[Hus90] Ronald E. Huss. An Ada library for automatic evaluation of derivatives. Applied Mathematics and Computation, 35(2 (part II)):103–123, January 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/009630039090112G. Also appeared as Working Paper, Hughes Aircraft Company, February, 1989.

Huijsman:1987:PAA

[HvKT87] R. D. Huijsman, J. van Katwijk, and W. J. Toetenel. Performance aspects of Ada tasking in embedded systems. Microprocessing and Microprogramming, 21(1–5):301–309, August 1987. CODEN MMICDT. ISSN 0165-6074 (print), 1878-

7061 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165607487900548.

Hutcheon:1987:ADS

[HW87] A. D. Hutcheon and A. J. Wellings. Ada for distributed systems. Computer Standards and Interfaces, 6(1):71–81, ???? 1987. CODEN CSTIEZ. ISSN 0920-5489 (print), 1872-

7018 (electronic). URL http://www.sciencedirect.com/
science/article/pii/092054898790047X.

Howes:1989:MAO

[HW89] N. R. Howes and A. C. Weaver. Measurements of Ada overhead in OSI-style communications systems. IEEE Transactions on Software Engineering, 15(12):1507–1517, December 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520

(electronic). URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=58763. See comments [Kar90].

Ichbiah:1986:RDA

[I^{+}86] Jean D. Ichbiah et al. Rationale for the design of the Ada programming language. Silicon Press, 25 Beverly Road, Summit, NJ 07901, USA, 1986. ISBN 0-9615336-5-X (paperback). ix + 393 pp. LCCN ????

Ichbiah:1979:PAR

[Ich79] Jean D. Ichbiah. Preliminary Ada reference manual. ACM SIGPLAN Notices, 14(6a):1–145, June 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). The final standard is [Ame83].

IEEE:1986:ATC

[IEE86a] IEEE, editor. Advancing test concepts and the user: proceedings / Autotestcon 86, San Antonio, TX, September 8– 11, 1986, IEEE International Automatic Testing Conference, Henry B. Gonzalez Convention Center, San Antonio, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1986. CODEN AUPRDX. ISBN ???? ISSN 0734-7510. LCCN TK 7895 A8 A98 1986.

IEEE:1986:ICS

[IEE86b] IEEE, editor. IEEE Computer Society Second International Conference on Ada Applications and Environments: April 8–10, 1986, Konover Hotel, Miami Beach, Florida. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1986. ISBN 0-8186-0718-1 (paperback), 0-8186-8718-5 (hard), 0-8186-4718-3 (microfiche). LCCN QA76.73.A35 I33 1986. IEEE catalog number 86CH2281-4. IEEE Computer Society order number 718.

IEEE:1989:CPF

[IEE89] IEEE, editor. COMPASS ’89 Proceedings of the Fourth Annual Conference on Computer Assurance. Systems Integrity, Software Safety and Process Security, Gaithersburg, MD, USA, June 19–23, 1989. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. IEEE catalog number 89CH2656-7.

IEEE:1992:IIT

[IEE92a] IEEE. IEEE 1003.5-1992: Information Technology — POSIX Ada Language Interfaces — Part 1: Binding for System Application Program Interface (API). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN ???? ???? pp. LCCN ????

US$110. URL http://www.ansi.org/.

IEEE:1992:ISIa

[IEE92b] IEEE Standards Board. IEEE standard for information technology: POSIX Ada language interfaces — Part 1: Binding for system application program interface (API). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, June 18, 1992. xii + 305 pp. LCCN QA76.76.O63I445 1992. Std 1003.5-1992.

IEEE:1996:IITa

[IEE96] IEEE. IEEE 1003.5b-1996: Information Technology — POSIX Ada Language Interfaces — Part 1: Binding for System Application Program Interface (API) — Amendment 1: Realtime Extension. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN ???? ???? pp. LCCN ???? US$195. URL

http://www.ansi.org/.

IEEE:1999:EIS

[IEE99a] IEEE. 1003.5-1999 Edition IEEE Standard for Information Technology — POSIX Ada Language Interfaces — Part 1: Binding for System Application Program Interface (API) — Amendment 2: Protocol — Independent Interfaces. IEEE, New York, NY, USA, 1999. ISBN 0-7381-1539-8 (print), 0-7381-1540-1 (electronic). 890 pp. LCCN ???? US$145.00. This edition incorporates IEEE Std 1003.5-1992, IEEE Std 1003.5b-1996 and IEEE Std 1003.5c-1998.

IEEE:1999:III

[IEE99b] IEEE. 14519:1999 1(ISO/IEC) IEEE/ANSI Std 1003.5b-1999 Information Technology — POSIX Ada Language Interfaces — Binding for System Application Program Interface (API) — Realtime Extensions. IEEE, New York, NY, USA, 1999. ISBN 0-7381-1570-3. 548 pp. LCCN ????

US$95.00. URL http://standards.ieee.org/reading/
ieee/interp/1003.5.html.

ISO:1994:IID

[II94] International Organization for Standardization and International Electrotechnical Commission. ISO/IEC DIS 8652: information technology — programming languages — their environments and system software interfaces, programming language Ada, language and standard libraries, draft, version 5.0, 1 June 1994, IR-MA-1363-4. Draft international standard. Intermetrics, Inc., Cambridge, MA, USA, 1994. ISBN ???? various pp. LCCN ???? URL http://www.iso.ch/

cate/d22983.html.

Ichbiah:1979:RDA

[IKBW^{+}79] Jean D. Ichbiah, Bernd Krieg-Brueckner, Brian A. Wichmann, John G. P. Barnes, Olivier Roubine, and Jean-Claude Heliard. Rationale for the design of the Ada programming language. ACM SIGPLAN Notices, 14(6b):1–261, June 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). The final standard is [Ame83].

Intermetrics:1996:TCA

[Int96] Intermetrics. Tool converts Ada 95 source code to Java bytecode. Java Report Online, 1996. URL http://www.sigs.

com/publications/docs/jro/twij96/twij961216.html#

TOOL. Intermetrics Inc. debuted AppletMagic, a tool that converts Ada 95 source code to Java bytecode for execution by any Java-capable Web browser. AppletMagic simplifies the development of complex, high-reliability applets and can be used as a supplement or an alternative to the Java language. Ada provides compile-time advantages such as enumeration types and generic templates, as well as in, in-out, and out parameter modes. The Java execution technology contributes runtime flexibility through automatic garbage collection, dynamic linking, and platform independence.

ISO:1988:IIPc

[ISO88] ISO. ISO 8651-3:1988: Information processing systems — Computer graphics — Graphical Kernel System (GKS) language bindings — Part 3: Ada. International Organization for Standardization, Geneva, Switzerland, 1988. ISBN ???? 184 pp. LCCN ???? CHF 224. URL http://www.iso.ch/

cate/d16026.html.

ISO:1990:IIIb

[ISO90] ISO. ISO/IEC 9593-3:1990: Information technology — Computer graphics — Programmer’s Hierarchical Interactive Graphics System (PHIGS) language bindings — Part 3: ADA. International Organization for Standardization, Geneva, Switzerland, 1990. ISBN ???? 285 pp. LCCN

???? CHF 270. URL http://www.iso.ch/cate/d17368.

html. Available in English only.

ISO:1993:IIA

[ISO93] ISO. ISO/IEC 10728:1993/Amd 2:1996: Ada language binding. International Organization for Standardization, Geneva, Switzerland, 1993. ISBN ???? 20 pp. LCCN ???? CHF 13.

URL http://www.iso.ch/cate/d24342.html. Available in

English only.

ISO:1994:IIIb

[ISO94a] ISO. ISO/IEC 11430:1994: Information technology — Programming languages — Generic package of elementary functions for ADA. International Organization for Standardization, Geneva, Switzerland, 1994. ISBN ???? 50 pp. LCCN ???? CHF 136; US$92.00. URL http://webstore.ansi.

org/ansidocstore/product.asp?sku=ISO%2FIEC+11430%

3A1994; http://www.iso.ch/cate/d19370.html. Avail-

able in English only.

ISO:1994:IIIc

[ISO94b] ISO. ISO/IEC 11729:1994: Information technology — Programming languages — Generic package of primitive functions for Ada. International Organization for Standardization, Geneva, Switzerland, 1994. ISBN ???? 19 pp. LCCN ???? CHF 86; US$58.00. URL http://webstore.ansi.

org/ansidocstore/product.asp?sku=ISO%2FIEC+11729%
3A1994; http://www.iso.ch/cate/d20879.html. Avail-

able in English only.

ISO:1994:IIIg

[ISO94c] ISO. ISO/IEC 9638-3:1994: Information technology — Computer graphics — Interfacing techniques for dialogues with graphical devices (CGI) — Language bindings — Part 3: Ada. International Organization for Standardization, Geneva, Switzerland, 1994. ISBN ???? 361 pp. LCCN

???? CHF 306. URL http://www.iso.ch/cate/d17461.

html. Available in English only.

ISO:1995:IIIa

[ISO95a] ISO. ISO/IEC 12227:1995: Information technology — Programming languages — SQL/Ada Module Description Language (SAMeDL). International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? 120 pp.

LCCN ???? CHF 188. URL http://www.ansi.org/;
http://www.iso.ch/cate/d2345.html.

ISO:1995:IIIk

[ISO95b] ISO. ISO/IEC 12227:1995: Information technology — Programming languages — SQL/Ada Module Description Language (SAMeDL). International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? 120 pp. LCCN ???? CHF 188; US$128.00. URL http:

//webstore.ansi.org/ansidocstore/product.asp?sku=
ISO%2FIEC+12227%3A1995; http://www.iso.ch/cate/d2345.

html. Available in English only.

ISO:1995:IIIg

[ISO95c] ISO. ISO/IEC 8652:1995: Information technology — Programming languages — Ada. International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? 511 pp. LCCN ???? CHF 334. URL http://www.adaic.

org/standards/05rm/RM-Final.pdf;   http://www.iso.

ch/cate/d22983.html. Available in English only.

ISO:1996:IIT

[ISO96] ISO. ISO/IEC TR 11735:1996: Information technology — Extensions for real-time Ada. International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 337

pp. LCCN ???? CHF 288. URL http://www.iso.ch/cate/
d22986.html.

ISO:1998:IIIg

[ISO98a] ISO. ISO/IEC 13719-3:1998: Information technology — Portable Common Tool Environment (PCTE) — Part 3: Ada programming language binding. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 161 pp. LCCN ???? CHF 212. URL http://www.iso.ch/ cate/d30522.html. Available in English only.

ISO:1998:IIIi

[ISO98b] ISO. ISO/IEC 13813:1998: Information technology — Programming languages — Generic packages of real and complex type declarations and basic operations for Ada (including vector and matrix types). International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 62 pp. LCCN ???? CHF 152; US$105.00. URL http:

//webstore.ansi.org/ansidocstore/product.asp?sku=
ISO%2FIEC+13813%3A1998; http://www.iso.ch/cate/d28992.

html. Available in English only.

ISO:1998:IIIj

[ISO98c] ISO. ISO/IEC 13814:1998: Information technology — Programming languages — Generic package of complex elementary functions for Ada. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 36 pp. LCCN ???? CHF 122; US$84.00. URL http:

//webstore.ansi.org/ansidocstore/product.asp?sku=
ISO%2FIEC+13814%3A1998; http://www.iso.ch/cate/d22985.

html. Available in English only.

ISO:1999:IIIb

[ISO99a] ISO. ISO/IEC 14519:1999: Information technology — POSIX Ada Language Interfaces — Binding for System Application Program Interface (API) — Realtime Extensions. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 529 pp. LCCN ???? CHF 334.

URL http://www.iso.ch/cate/d29522.html. Available in

English only.

ISO:1999:IIId

[ISO99b] ISO. ISO/IEC 15291:1999: Information technology — Programming languages — Ada Semantic Interface Specification (ASIS). International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 283 pp. LCCN ???? CHF 270; US$185.00. URL http://webstore.ansi.

org/ansidocstore/product.asp?sku=ISO%2FIEC+15291%
3A1999; http://www.iso.ch/cate/d27169.html. Avail-

able in English only.

ISO:1999:IIIs

[ISO99c] ISO. ISO/IEC 18009:1999: Information technology — Programming languages — Ada: Conformity assessment of a language processor. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 25 pp. LCCN ???? CHF 98; US$68.00. URL http://webstore.ansi.

org/ansidocstore/product.asp?sku=ISO%2FIEC+18009%
3A1999; http://www.iso.ch/cate/d31051.html. Avail-

able in English only.

ISO:2000:IIT

[ISO00] ISO. ISO/IEC TR 15942:2000: Information technology — Programming languages — Guide for the use of the Ada programming language in high integrity systems. International Organization for Standardization, Geneva, Switzerland, 2000. ISBN ???? 48 pp. LCCN ???? CHF 136; US$92.00. URL http:

//webstore.ansi.org/ansidocstore/product.asp?sku=
ISO%2FIEC+TR+15942%3A2000; http://www.iso.ch/cate/

d29575.html. Available in English only.

ISO:2001:IICa

[ISO01] ISO. ISO/IEC 8652:1995/Cor 1:2001, Technical Corrigendum to Information Technology — Programming Languages — Ada. International Organization for Standardization, Geneva, Switzerland, 2001. ISBN ???? ???? pp. LCCN ????

ISO:2007:IIA

[ISO07] ISO. ISO/IEC 8652:1995/Amd 1:2007, Amendment to Information Technology — Programming Languages — Ada. International Organization for Standardization, Geneva, Switzerland, 2007. ???? pp. LCCN ???? URL http://www.iso.

org/iso/home/store/catalogue_ics/catalogue_detail_
ics.htm?csnumber=45001.

ISO:2012:IIIb

[ISO12] ISO. ISO/IEC 8652:2012 Information technology — Programming languages — Ada. International Organization for Standardization, Geneva, Switzerland, 2012. ISBN

????  832 (est.) pp. LCCN ????  URL http:

//www.ada-auth.org/standards/ada12.html;   http://
www.iso.org/iso/home/store/catalogue_ics/catalogue_
detail_ics.htm?csnumber=61507.

Jackson:1985:DAP

[Jac85] M. I. Jackson. Developing Ada programs using the Vienna Development Method (VDA). Software—Practice and Experience, 15(3):305–318, March 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Janas:1980:COI

[Jan80] Jürgen M. Janas. A comment on ‘operator identification in ADA’ by Ganzinger and Ripken. ACM SIGPLAN Notices, 15(9):39–43, September 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [GR80].

Jingde:1992:TDN

[Jin92] Cheng Jingde. Task dependence net as a representation for concurrent Ada programs. Lecture Notes in Computer Science, 603:150–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jha:1989:APP

[JKC89] R. Jha, J. M. Kamrad, and D. T. Cornhill. Ada program partitioning language: a notion for distributing Ada programs. IEEE Transactions on Software Engineering, 15(3): 271–280, March 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?arnumber=21755.

Jones:2011:DHC

[JL11] Cliff B. Jones and J. L. (John L.) Lloyd, editors. Dependable and Historic Computing: Essays Dedicated to Brian Randell on the Occasion of His 75th Birthday, volume 6875 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 3-642-24540-4, 3-642-24541-2. ISSN 0302-9743. xii +

523 pp. LCCN QA76.17. URL http://link.springer.com/
10.1007/978-3-642-24541-1.

James:1983:RAC

[JM83] Carol L. James and Duncan E. Morrill. The real Ada, countess of Lovelace. ACM SIGSOFT Software Engineering Notes, 8(1):30–31, January 1983. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Johnston:1997:ACC

[Joh97] Simon Johnston. Ada 95 for C and C++ programmers. International computer science series. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-40363-3, 0-201-40363-3. xvii + 364 pp. LCCN ????

Jones:1986:LAM

[Jon86] Do-While Jones. Learning Ada on a micro. Dr. Dobb’s Journal of Software Tools, 11(2):42–??, February 1986. CODEN DDJOEB. ISSN 1044-789X.

Jones:1989:AAP

[Jon89] Do-While Jones. Ada in action: with practical programming examples. Wiley, New York, NY, USA, 1989. ISBN 0-471-60708-8. xiv + 490 pp. LCCN ????

Johnson:1990:AAA

[JpJ90] Philip I. Johnson and Philip I. Ada primer Johnson. Ada: applications and administration. McGraw-Hill, New York, NY, USA, second edition, 1990. ISBN 0-07-032627-4. xxii + 209 pp. LCCN QA76.73.A35 J63 1990.

Jimenez-Peris:2000:TAO

[JPMAB00] R. Jiménez-Peris, M. Patiño Martínez, S. Arévalo, and F. J. Ballesteros. TransLib: an Ada 95 object oriented framework for building transactional applications. International Journal of Computer Systems Science and Engineering, 15(1):7–??, January 2000. CODEN CSSEEI. ISSN 0267-6192.

Jean:1990:ETO

[JS90] C. Jean and A. Strohmeier. An experience in teaching OOD for Ada software. ACM SIGSOFT Software Engineering Notes, 15(5):44–49, October 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jovanovic:1997:DSC

[JSV97] S. S. Jovanovic, D. M. Sotirovski, and P. Van Aswegen. Developing scripting capabilities for an Ada-based simulator. Lecture Notes in Computer Science, 1251:56–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1998:ERI

[JT98] G. P. Jones and N. P. Topham. The effect of restricted instruction issue width on an access decoupled architecture. Advances in Parallel Computing, 12(??): 665–672, ???? 1998. CODEN ???? ISSN ????

URL http://www.sciencedirect.com/science/article/
pii/S0927545298800853.

Jen-Yen:1994:ALS

[JYCM94] Chen Jen-Yen and Tu Chia-Ming. An Ada-like software process language. The Journal of Systems and Software, 27(1):17–25, October 1994. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0164121294901112.

Karam:1990:CMA

[Kar90] G. M. Karam. Comments on “Measurements of Ada overhead in OSI-style communications systems”. IEEE Transactions on Software Engineering, 16(12):1435–1439, December 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=62440. See [HW89].

Katzan:1982:IAA

[Kat82] Harry Katzan. Invitation to Ada and Ada reference manual (July 1980). PBI, New York, NY, USA, 1982. ISBN 0-89433-132-9. xi + 429 pp. LCCN QA76.73.A35 K37 1982.

Katzan:1984:IA

[Kat84] Harry Katzan. Invitation to Ada. Petrocelli invitation to series. PBI, New York, NY, USA, 1984. ISBN 0-89433-239-2 (paperback). x + 173 pp. LCCN QA76.73.A35 K36 1984. US$14.95.

Karam:1991:TLB

[KB91] G. M. Karam and R. J. A. Buhr. Temporal logic-based deadlock analysis for Ada. IEEE Transactions on Software Engineering, 17(10):1109–1125, October 1991. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?arnumber=99197. See comments [YLT93].

Krieg-Bruckner:1980:ATL

[KBL80] Bernd Krieg-Brückner and David C. Luckham. ANNA: towards a language for annotating Ada programs. ACM SIG-PLAN Notices, 15(11):128–138, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kambayashi:1986:TIC

[KCGO86] Yahiko Kambayashi, Wesley Chu, Georges Gardarin, and Setsuo Ohsuga, editors. Twelfth International Conference on Very Large Data Bases, Proceedings (VLDB ’86). Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1986. ISBN 0-934613-18-4. LCCN QA 76.9 D3 I61 1986.

Kapre:2008:PFA

[KD08] Nachiket Kapre and André DeHon. Programming FPGA applications in VHDL. In Scott Hauck and André De-Hon, editors, Reconfigurable Computing: The Theory and Practice of FPGA-Based Computation, pages 129– 153. Elsevier, Amsterdam, The Netherlands, 2008. URL

http://www.sciencedirect.com/science/article/pii/
B978012370522850011X.

Kemp:1987:MDO

[Kem87] Ray Kemp. Ministry of Defence opts for ADA. Information and Software Technology, 29(6):282, July/August 1987. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0950584987900255.

Kempe:1996:HDS

[Kem96] M. Kempe. Heterogeneous data structures and crossclassification of objects with Ada95. Lecture Notes in Computer Science, 1031:71–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

King:1993:MTA

[Kin93] K. N. King. Modula-3: a threat to Ada? In Proceedings of the Tenth Annual Washington Ada Symposium on Ada: Ada’s Role in Software Engineering, WADAS ’93, pages 147– 154. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-609-3.

Kordon:2009:RST

[KK09] Fabrice Kordon and Yvon Kermarrec, editors. Reliable Software Technologies — Ada-Europe 2009: 14th Ada-Europe International Conference, Brest, France, June 8–12, 2009. Proceedings, volume 5570 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNCSD9. ISBN 3-642-01923-4 (print), 3-642-01924-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-01924-1.

Knott:1915:NTM

[Kno15] Cargill Gilston Knott, editor. Napier tercentenary memorial volume. Published for the Royal society of Edinburgh, by Longmans, Green and Company, London, UK, 1915. xi + 441 + 1 pp. LCCN QA29.N2 K5. URL http://resolver.

library.cornell.edu/math/2013807.

Koranne:2011:HOS

[Kor11] Sandeep Koranne. Handbook of open source tools. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 1-4419-7719-8 (e-book). xxxvi + 484 pp. LCCN QA76.76.S46 K67 2011.

Kurtz:1990:ISC

[KP90] Barry L. Kurtz and Thomas H. Puckett. Implementing a single classwide project in software engineering using Ada tasking for synchronization and communication. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):6–11, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kermarrec:1996:PDSa

[KP96a] Y. Kermarrec and L. Pautet. Programming distributed systems with both Ada 95 and PVM. Lecture Notes in Computer Science, 1031:206–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kermarrec:1996:PDSb

[KP96b] Y. Kermarrec and L. Pautet. Programming distributed systems with both Ada 95 and PVM. In Toussaint [Tou96], pages 206–216. CODEN LNCSD9. ISBN 3-540-60757-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A35 I57 1995. URL http://link.

springer-ny.com/link/service/series/0558/tocs/t1031.
htm;   http://www.springerlink.com/content/978-3-
540-60757-1;   http://www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&volume=1031.

Kroha:1993:TQO

[Kro93] P. Kroha. Translation of a query in OODBMS into a system of parallel tasks. Microprocessing and Microprogramming, 37(1–5):167–170, January 1993. CODEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016560749390040R.

Kroeker:1998:NTSd

[Kro98a] Kirk L. Kroeker. New tools: Software development: Rational Software’s Ada development environment; Pragmatix Software’s Meta-CASE tool; Wibu Systems’ copy protection tools. hardware development: Diab Data’s suite of analysis tools for embedded applications; Mustang Software’s message center tool; RSI’s AS/400 client-server tool; SES’s client-server simulation tool; CACI Products’ network analysis tools. Computer, 31(9):102–104, 106, September 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814

(electronic). URL http://dlib.computer.org/co/books/
co1998/pdf/r9102.pdf.

Kroeker:1998:NTSe

[Kro98b] Kirk L. Kroeker. New tools: Software development: Rational Software’s development tools for Ada, Artlandia’s Mathematica graphics tool, ISE’s EiffelBase library available for free, Annasoft System’s Windows CE development tool. hardware development: White Mountain’s DSP tools, Analogy’s mixed-signal IC simulator, Dynamic Soft Analysis’ thermal analysis tool; net development: Framework Technologies’ project management system, CAI’s message broker, Gordian’s network redirector software; component technology: Black & White’s usage control software for multi-ORB CORBA implementations, IONA Technologies’ middleware development tools, GreenTree Technologies’ ActiveX controls, KL Group’s JavaBeans with automatic data binding. Computer, 31(10):110–114, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/co/books/co1998/pdf/
rx110.pdf.

Kuchumov:2001:OAS

[KRS01] Alexei Kuchumov, Sergey Rybin, and Alfred Strohmeier. OASIS — an ASIS secondary library for analyzing objectoriented Ada code. Lecture Notes in Computer Science, 2043: 113–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer-ny.

com/link/service/series/0558/bibs/2043/20430113.htm;
 http://link.springer-ny.com/link/service/series/
0558/papers/2043/20430113.pdf.

Karam:1989:CRA

[KSB89] G. M. Karam, C. M. Stanczyk, and G. W. Bond. Critical races in Ada programs. IEEE Transactions on Software Engineering, 15(11):1471–1480, November 1989. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?arnumber=41338. See comments [MO90].

Koymans:1988:CSR

[KSdR^{+}88] R. Koymans, R. K. Shyamasundar, W. P. de Roever, R. Gerth, and S. Arun-Kumar. Compositional semantics for real-time distributed computing. Information and Computation, 79(3):210–256, December 1988. CODEN IN-FCEC. ISSN 0890-5401 (print), 1090-2651 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/089054018890020X.

Kruchten:1996:ISD

[KT96] P. Kruchten and C. J. Thompson. Iterative software development for large Ada programs. Lecture Notes in Computer Science, 1088:101–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knight:1987:IUA

[KU87] J. C. Knight and J. I. A. Urquhart. On the implementation and use of Ada on fault-tolerant distributed systems. IEEE Transactions on Software Engineering, SE-13 (5):553–563, May 1987. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?arnumber=1702255.

Kunz:1998:BAC

[Kun98] Michelle Kunz. Bravo on Ada 95 coverage. Computers in Physics, 12(4):312–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

aip.scitation.org/doi/10.1063/1.4822631.

Kordon:2008:RST

[KV08] Fabrice Kordon and Tullio Vardanega, editors. Reliable Software Technologies – Ada-Europe 2008: 13th Ada-Europe International Conference on Reliable Software Technologies, Venice, Italy, June 16–20, 2008. Proceedings, volume 5026 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNCSD9. ISBN 3-540-68621-5 (print), 3-540-68624-X (e-book). ISSN 0302-9743 (print), 1611-3349

(electronic). LCCN ???? URL http://www.springerlink.
com/content/978-3-540-68624-8.

Kwon:2005:RJH

[KWK05] Jagun Kwon, Andy Wellings, and Steve King. Ravenscar– Java: a high integrity profile for real-time Java. Concurrency and Computation: Practice and Experience, 17(5–6): 681–714, April/May 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luckham:1987:ALA

[L^{+}87] David C. Luckham et al. ANNA, a language for annotating Ada programs: reference manual, volume 260 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. CODEN LNCSD9. ISBN 3-540-17980-1 (Berlin), 0-387-17980-1 (US). ISSN 0302-9743 (print), 1611-3349 (electronic). v + 143 pp. LCCN QA76.73.A54

A561 1987. URL http://link.springer-ny.com/link/
service/series/0558/tocs/t0260.htm;   http://www.
springerlink.com/content/978-0-387-17980-3;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=

0302-9743&volume=260. Spine title: ANNA reference manual.

Lodgher:1994:PAT

[LAH94] Akhtar Lodgher and Hisham Al-Haddad. A practical approach for teaching reuse in a data structures course using Ada (abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):407, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lamb:1983:TUS

[Lam83] J. Eli Lamb. Towards a UNIX system Ada programming support environment. In Association [Ass83], pages 143–?? Abstract only.

Lamm:2002:ADC

[Lam02] Ehud Lamm. Adding design by contract to the Ada language. Lecture Notes in Computer Science, 2361:205–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2361/23610205.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2361/23610205.pdf.

Lamm:2003:BAV

[Lam03] E. Lamm. Booch’s Ada vs. Liskov’s Java: Two approaches to teaching software design. Lecture Notes in Computer Science, 2655:102–112, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laurens:1996:PGC

[Lau96] A. Laurens. PRONAOS ground control center: First operational Ada application in C.N.E.S. Lecture Notes in Computer Science, 1031:124–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Linton:1989:CAS

[LC89] Darrell G. Linton and Maria A. Cianci. Computer-aided software engineering and Ada – the technological marriage of the decade. Computers & industrial engineering, 17(1–4):542– 545, ???? 1989. CODEN CINDDL. ISSN 0360-8352 (print),

1879-0550 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0360835289901204.

Leveson:1991:SVA

[LCS91] Nancy G. Leveson, Stephen S. Cha, and Timothy J. Shimeall. Safety verification of Ada programs using software fault trees. IEEE Software, 8(4):48–59, July 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Luong:1994:NAC

[LDD^{+}94] T. T. Luong, L. David, P. Duneau, P. De Saint Jores, E. Lécorché, E. Lemaître, P. Lermine, J. M. Loyant, C. Maugeais, F. Régnault, J. F. Rozé, A. Souf, and M. Ulrich. The new accelerator control system of GANIL. Nuclear instruments and methods in physics research. Section A, Accelerators, spectrometers, detectors and associated equipment, 352(1–2):96–102, December 15, 1994. CODEN NIMAER. ISSN 0168-9002, 0167-5087. URL

http://www.sciencedirect.com/science/article/pii/
0168900294914710.

LeVerrand:1982:LAM

[Le 82] Dominique Le Verrand. Le langage Ada: manuel d’Évaluation. (French) [The Ada language: evaluation manual]. Dunod, Paris, France, 1982. ISBN 2-04-015499-X (paperback). ii + 266 pp. LCCN ????

LeVerrand:1984:EA

[Le 84] Dominique Le Verrand. Evaluating ADA. MacMillan Publishing Company, New York, NY, USA, revised edition, 1984. ISBN 0-02-947660-7. ???? pp. LCCN QA76.73.A35 L3813 1984. Translation from French original [Le 82] by J. Howlett.

LeVerrand:1985:EA

[Le 85] D. Le Verrand. Evaluating ADA. North Oxford Academic, Oxford, UK, 1985. ISBN 0-946536-15-5. vii + 288 pp. LCCN ???? Translation from French original [Le 82] by J. Howlett.

Ledgard:1981:AI

[Led81] Henry F. Ledgard. Ada, an introduction. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. ISBN 0-387-90568-5, 3-540-90568-5. 380 (est.) pp. LCCN QA76.73.A35 L42; QA76.73.A35L42.

Ledgard:1983:AI

[Led83] Henry F. Ledgard. Ada, an introduction. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1983. ISBN 0-387-90568-5. viii + 135 pp. LCCN QA76.73.A35 L42 1983.

Lee:1982:VLS

[Lee82] J. A. N. Lee. Validation and language standards. Computers and Standards, 1(4):219–223, December 1982. CO-DEN COSTEZ. ISSN 0167-8051 (print), 1878-3090 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0167805182900390.

Lee:1992:BRR

[Lee92] J. A. N. Lee. Book review: The Rationale for the Design of the Ada Programming Language, The Ada Companion Series by J. ichbiah, J. Barnes, R. Firth, and M. Woodger.: (Cambridge University Press, Cambridge, UK). ACM SIGPLAN Notices, 27(1):14, January 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Levy:1989:AFP

[Lev89] Stephen H. Levy. Ada, the Fortran Programmer’s Companion. Silicon Press, 25 Beverly Road, Summit, NJ 07901, USA, October 1989. ISBN 0-929306-02-3. various pp. LCCN

???? US$27.95. URL http://www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=0929306023.

Lyttle:1990:SDR

[LF90] Dan Lyttle and Ray Ford. A symbolic debugger for real-time embedded Ada software. Software—Practice and Experience, 20(5):499–514, May 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lamb:1983:SPV

[LH83] D. A. Lamb and P. N. Hilfinger. Simulation of procedure variables using Ada tasks. IEEE Transactions on Software Engineering, SE-9(1):13–15, January/February 1983. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=1703007.

Lopes:1994:VHL

[LHF94] Arthur V. Lopes, Rachelle S. Heller, and Michael B. Feldman. Very high-level debugging. Computers and Education, 22(3):231–238, April 1994. CODEN COMEDR. ISSN 0360-1315 (print), 1873-782X (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0360131594900043.

Li:1995:NAE

[Li95] Gansheng Li. A new approach for efficient implementation of Ada multi-tasking. ACM SIGPLAN Notices, 30(12):25–31, December 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lignelet:1990:SDA

[Lig90] Patrice Lignelet. Structures de données avec ADA: conception orientée objets. Techniques de l’informatique. Masson Editeur, Masson, France, 1990. ISBN 2-903607-78-8. ISSN 1146-5832. 160 pp. LCCN QA 76 .73 A35 L535 1990. Préface de J.-F. Perrot.

Lin:1993:REO

[Lin93] F. Lin. Re-engineering option analysis for managing software rejuvenation. Information and Software Technology, 35(8):462–467, August 1993. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0950584993900433.

Lapalme:1986:EUA

[LL86] Guy Lapalme and Jean-François Lamy. An experiment in the use of Ada in course in software engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 18(1):124–126, February 1986. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 17th SIGCSE Symposium on Computer Science Education.

Llaurado:1993:BRB

[Lla93] J. G. Llaurado. Book review: ADA, the enchantress of numbers: Betty A. Toole. Strawberry Press, Mill Valley, California, USA. 1992, 440 pp. International Journal of Bio-Medical Computing, 32(1):79–80, January 1993. CO-DEN IJBCBT. ISSN 0020-7101 (print), 1878-366X (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/002071019390008T.

Litvintchouk:1984:DAS

[LM84] S. D. Litvintchouk and A. S. Matsumoto. Design of Ada systems yielding reusable components: An approach using structured algebraic specification. IEEE Transactions on Software Engineering, SE-10(5):544–551, September/October 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=5010278.

Lander:1992:DAE

[LM92] Leslie C. Lander and Sandeep Mitra. Detection and avoidance of elaboration-time problems for multi-unit real-time Ada applications. Lecture Notes in Computer Science, 603: 165–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larsson:1993:AER

[LN93] Lars-Göran Larsson and Erik Nilsson. An Ada environment for relativistic cross-section calculations. Computer Physics Communications, 74(1):41–57, January 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/001046559390104K.

Loftus:1993:AY

[Lof93] C. Loftus, editor. Ada yearbook 1993. IOS Press, Postal Drawer 10558, Burke, VA 2209-0558, USA, 1993. xvi + 431 pp.

Luckham:1980:PMD

[LP80] David C. Luckham and Wolfgang Polak. A practical method of documenting and verifying Ada programs with packages. ACM SIGPLAN Notices, 15(11):113–122, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lewi:1986:DSP

[LP86] Johan Lewi and Jan Paredaens. Data structures of Pascal, Algol 68, PL. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 3-540-15121-4, 0-387-15121-4. xii + 395 pp. LCCN ????

Levy:1991:RAD

[LRT91] D. C. Levy, M. C. Randelhoff, and J. L. Tokar. Reconfigurable Ada distributed control system software. Lecture Notes in Computer Science, 499:279–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ledgard:1982:SAT

[LS82] Henry F. Ledgard and Andrew Singer. Scaling down Ada (or towards a Standard Ada Subset). Communications of the ACM, 25(2):121–125, 1982. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Llamosa:2004:RST

[LS04] Albert Llamosí and Alfred Strohmeier, editors. Reliable Software Technologies—Ada-Europe 2004: 9th ADA-Europe International Conference on Reliable Software Technologies, Palma de Mallorca, Spain, June 14–18, 2004, Proceedings, volume 3063 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNCSD9. ISBN 3-540-22011-

9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.R44 A33 2004. URL http://link.springer-

ny.com/link/service/series/0558/tocs/t3063.htm;
http://uclibs.org/PID/41011;   http://www.springerlink.
com/content/978-3-540-22011-4;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=
3063;   http://www.springerlink.com/openurl.asp?genre=
volume&id=doi:10.1007/b97913.

Lundberg:1989:PAS

[Lun89] Lars Lundberg. A parallel Ada system on an experimental multiprocessor. Software—Practice and Experience, 19(8): 787–800, August 1989. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lundberg:1990:PRG

[Lun90] Lars Lundberg. A protocol to reduce global communication in distributed Ada tasking. Journal of Parallel and Distributed Computing, 10(3):261–264, November 1990. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/074373159090017J.

Lundberg:1991:CHP

[Lun91] L. Lundberg. A coprocessor for high performance multiprocessor Ada tasking. Lecture Notes in Computer Science, 499: 147–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lundberg:1992:PSP

[Lun92] Lars Lundberg. Predicting the speedup of parallel Ada programs. Lecture Notes in Computer Science, 603:257–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luqi:1990:APD

[Luq90] Luqi. Automated prototyping and data translation. Data and Knowledge Engineering, 5(2):167–177, July 1990. CO-DEN DKENEW. ISSN 0169-023X (print), 1872-6933 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0169023X90900112.

Lutz:1998:NBPb

[Lut98] Michael J. Lutz. New books: Programming for high-end JavaBeans; Intel secrets; Ada and concurrency; data-modeling primer; technology and economics; dissecting a microprocessor; real-world ATM; when computers run amok. Computer, 31(10):109, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co1998/pdf/rx109.pdf.

Levy:1989:DSD

[LvdGvK89] A. M. Levy, A. J. van de Goor, and J. van Katwijk. Distributed system design using Ada as a tool for prototyping. Microprocessing and Microprogramming, 27(1–5):221–230, August 1989. CODEN MMICDT. ISSN 0165-6074 (print),

1878-7061 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0165607489900501.

Luckham:1984:AAB

[LvLS84] David C. Luckham, Friedrich W. von Henke, H. J. Larsen, and D. R. Stevenson. ADAM: an Ada-based language for multiprocessing. Software—Practice and Experience, 14(7): 605–642, July 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Liu:2004:PAM

[LX04] Yuan Liu and Baowen Xu. Process algebra model of Ada protected objects. ACM SIGPLAN Notices, 39(2):34–39, February 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liu:2003:DDA

[LXC03] Yuan Liu, Baowen Xu, and Zhenqiang Chen. Detecting deadlock in Ada rendezvous flow structure based on process algebra. Lecture Notes in Computer Science, 2495:262–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.de/link/
service/series/0558/bibs/2495/24950262.htm; http:/
/link.springer.de/link/service/series/0558/papers/
2495/24950262.pdf.

Lyons:1987:APS

[Lyo87] T. G. L. Lyons. Ada programming support environment (APSE). Computer Standards and Interfaces, 6(1):83–87, ???? 1987. CODEN CSTIEZ. ISSN 0920-5489 (print), 1872-

7018 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0920548987900481.

Lu:2004:CTO

[LZLX04] Hongmin Lu, Yuming Zhou, Jiangtao Lu, and Baowen Xu. A compile-time optimization framework for Ada rendezvous.

ACM SIGPLAN Notices, 39(2):18–25, February 2004. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McGarry:1989:MAS

[MA89] Frank E. McGarry and William W. Agresti. Measuring Ada for software development in the Software Engineering Laboratory. The Journal of Systems and Software, 9(2):149–159, February 1989. CODEN JSSODM. ISSN 0164-1212 (print),

1873-1228 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0164121289900174.

Miranda:1996:DAE

[MAAG96] F. J. Miranda, A. Alvarez, S. Arevalo, and F. J. Guerra. Drago: An Ada extension to program fault-tolerant distributed applications. Lecture Notes in Computer Science, 1088:235–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MacLaren:1980:ETA

[Mac80] Lee MacLaren. Evolving toward Ada in real time systems. ACM SIGPLAN Notices, 15(11):146–155, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Machanick:1983:NCW

[Mac83] Philip Machanick. A note on C. S. Wetherell’s ‘Problems with the Ada reference grammar’. ACM SIGPLAN Notices, 18(5): 44–45, May 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [Wet81].

MacLennan:1984:SMP

[Mac84]  Bruce J. MacLennan.  Simple metrics for program-

ming languages. Information Processing and Management, 20(1–2):209–221, ???? 1984. CODEN IPMADK. ISSN 0306-4573 (print), 1873-5371 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0306457384900517.

Madhav:1996:TAP

[Mad96] N. Madhav. Testing Ada 95 programs for conformance to rapide architectures. Lecture Notes in Computer Science, 1088:123–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Magel:2017:RIA

[Mag17] Kenneth Magel. Revisiting the impact of the Ada programming language. Computer, 50(9):10–11, September 2017. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL https://www.computer.org/csdl/mags/co/
2017/09/mco2017090010.html.

Mahjoub:1981:SCA

[Mah81] Ahmed Mahjoub. Some comments on Ada as a real-time programming language. ACM SIGPLAN Notices, 16(2):89–95, February 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mangold:1992:AMP

[Man92] Karlotto Mangold. AMPATS — a multi-processor Ada tool set. Lecture Notes in Computer Science, 603:300–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1995:AII

[Mar95] Richard E. Martin. Atlas II and IIA analyses and environments validation. Acta Astronautica, 35(12):771–791, June 1995. CODEN AASTCF. ISSN 0094-5765 (print), 1879-

2030 (electronic). URL http://www.sciencedirect.com/
science/article/pii/009457659500003I.

Mayoh:1982:PSA

[May82] Brian H. Mayoh. Problem solving with ADA. Wiley series in computing. Wiley, New York, NY, USA, 1982. ISBN 0-471-10025-0. viii + 233 pp. LCCN QA76.8.A15 M38 1982.

Mayoh:1983:PCL

[May83] Brian Mayoh. Programmare con il linguaggio Ada. (Italian) [Programming in the Ada language]. Tecniche nuove, Milano, Italia, 1983. ISBN 88-7081-120-4. viii + 239 pp. LCCN ????

Morrison:1986:DPI

[MB86] I. W. Morrison and A. Burns. The design and prototype implementation of a “structure attribute” model for tool interfacing within an IPSE. Microprocessing and Microprogramming, 18(1–5):233–240, December 1986. CODEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165607486900499.

Millan:1996:AOC

[MB96] T. Millan and P. Bazex. Ada/O2 coupling: a solution for an efficient management of persistence in Ada 83. Lecture Notes in Computer Science, 1088:396–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McCormick:1992:MRA

[McC92] John W. McCormick. A model railroad for Ada and software engineering. Communications of the ACM, 35(11):68– 70, November 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0001-0782/138848.html.

McCormick:1994:TAB

[MCD^{+}94] John W. McCormick, Fintan Culwin, Nicholas J. DeLillo, Michael B. Feldman, Richard Pattis, and Walter J. Savitch. Teaching Ada by the book: the pedagogy of Ada in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1):385–386, March 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McGettrick:1982:PVU

[McG82] Andrew D. McGettrick. Program verification using Ada, volume 13 of Cambridge computer science texts. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-24215-0, 0-521-28531-3 (paperback). xii + 345 pp. LCCN QA76.6 .M399 1982.

McGettrick:1983:CPL

[McG83] A. D. McGettrick. The current programming language standards scene VIIIA: ADA. Computers and Standards, 2(2–3):107–113, ???? 1983. CODEN COSTEZ. ISSN 0167-8051 (print), 1878-3090 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0167805183900104.

Maymir-Ducharme:1992:PER

[MD92] Fred A. Maymir-Ducharme. Porting embedded real-time Ada software. Lecture Notes in Computer Science, 603:92–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manas:1993:TSI

[MdMSA93] JoséA. Mañas, Tomás de Miguel, Joaquín Salvachúa, and Arturo Azcorra. Tool support to implement LOTOS formal specifications. Computer Networks and ISDN Systems, 25(7):815–839, February 1993. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
016975529390050E.

Martinez:2008:ACC

[MDPM08] Patricia López Martínez, José M. Drake, Pablo Pacheco, and Julio L. Medina. Ada-CCM: Component-based technology for distributed real-time systems. Lecture Notes in Computer Science, 5282:334–350, 2008. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/978-3-540-
87891-9_23.

Mearns:1987:BRB

[Mea87] Ian Mearns. Book review: Concurrent programming in Ada: by Alan Burns. The Ada companion series, Cambridge University Press, Cambridge, 1985, 256 pages, ISBN 0-521-30033-9, Price: £17.50 hardback. Science of Computer Programming, 9(3):309–310, December 1987. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016764238790013X.

Mearns:1988:BRB

[Mea88] Ian Mearns. Book review: A review of Ada tasking, by Alan Burns, Andrew M. Lister and Andrew J. Wellings. Lecture Notes in Computer Science 262, Springer, Berlin, 1987, 141 pp., Price £10.00 (soft cover), ISBN 3-540-18008-7. Science of Computer Programming, 11(1):90–91, October 1988. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0167642388900688.

Meek:1992:BRB

[Mee92] Brian L. Meek. Book review: ADA Yearbook 1992: by Fred Long. Chapman and Hall, London, UK. Pages 340. ISBN: 0-402-45630-3. Computer Standards and Interfaces, 14(4):355, August 1992. CODEN CSTIEZ. ISSN 0920-5489 (print),

1872-7018 (electronic). URL http://www.sciencedirect.
com/science/article/pii/092054899290046G.

Mercy:1984:BRB

[Mer84] Andy Mercy. Book review: The programming languages, Pascal, Modula, CHILL and ADA: C. H. Smedema, P. Medema and M. Boasson Prentice-Hall International, ISBN 0-13-729756-4. Advances in Engineering Software (1978), 6(2): 126, April 1984. CODEN AESODT. ISSN 0965-9978 (print),

0141-1195 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0141119584901475.

Metcalf:1985:FF

[Met85] Michael Metcalf. Has Fortran a future? Computer Physics Communications, 38(2):199–210, October/November 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900864.

Mellish:1991:CMA

[MG91] A. C. Mellish and J. E. Galletly. The concurrency models of Ada and occam — a practical comparison. Journal of Microcomputer Applications, 14(4):363–, October 1991. CO-DEN JMIADO, MIAPEZ. ISSN 0745-7138 (print), 1096-

374X (electronic). URL http://www.sciencedirect.com/
science/article/pii/074571389190021I.

Medina:2002:MSA

[MGDH02] Julio L. Medina, J. Javier Gutiérrez, José M. Drake, and Michael González Harbour. Modeling and schedulability analysis of hard real-time distributed systems based on Ada components. Lecture Notes in Computer Science, 2361:282– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2361/23610282.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2361/23610282.pdf.

Murphy:1991:EEL

[MGK91] Susan C. Murphy, Per Gunningberg, and John P. J. Kelly. Experiences with Estelle, LOTOS and SDL: a protocol implementation experiment. Computer Networks and ISDN Systems, 22(1):51–59, August 1991. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016975529190081M.

Miranda:2002:HUG

[MGM^{+}02] Javier Miranda, Francisco Guerra, Ernestina Martel, José Martín, and Alexis González. How to use GNAT to efficiently preprocess new Ada sentences. Lecture Notes in Computer Science, 2361:179–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2361/
23610179.htm;   http://link.springer-ny.com/link/
service/series/0558/papers/2361/23610179.pdf.

Montanari:1987:SDA

[MH87] U. Montanari and A. N. Habermann, editors. System development and Ada: CRAI Workshop on Software Factories and Ada, Capri, Italy, May 26–30, 1986: proceedings, volume 275 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 3-540-18341-8, 0-387-18341-8. LCCN ????

Moran:1997:LEA

[MH97] Tom Moran and Tom R. Halfhill. Letter to the Editor: Ada speaks Java. Byte Magazine, 22(4):20, April 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mitchell:1983:AASa

[Mit83a] Edward Mitchell. Augusta — an Ada subset for micros. Dr. Dobb’s Journal of Software Tools, 8(1):14, 16–23, January 1983. CODEN DDJOEB. ISSN 1044-789X.

Mitchell:1983:AASb

[Mit83b] Edward Mitchell. Augusta — an Ada subset for micros. Dr. Dobb’s Journal of Software Tools, 8(3):14–??, March 1983. CODEN DDJOEB. ISSN 1044-789X.

Mitchell:1983:AASc

[Mit83c] Edward Mitchell. Augusta — an Ada subset for micros. Dr. Dobb’s Journal of Software Tools, 8(5):14–??, May 1983. CO-DEN DDJOEB. ISSN 1044-789X.

Mitchell:1983:AASd

[Mit83d] Edward Mitchell. Augusta — an Ada subset for micros. Dr. Dobb’s Journal of Software Tools, 8(7):14–??, July 1983. CO-DEN DDJOEB. ISSN 1044-789X.

Mitchell:1987:EVA

[Mit87] Charles Z. Mitchell. Engineering VAX Ada for a multilanguage programming environment. ACM SIGPLAN Notices, 22(1):49–58, January 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

MacLean:1988:PIA

[MMH88] D. N. MacLean, R. H. Murray, and J. D. Hamilton. The performance implications of Ada in a Mascot 3 design: final report. Yard, Glasgow, Scotland, 1988. iii + 116 pp. LCCN ????

Miller:1987:SSC

[MMHS87] D. M. Miller, R. S. Maness, J. W. Howatt, and W. H. Shaw. A software science counting strategy for the full Ada language. ACM SIGPLAN Notices, 22(5):32–41, May 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McNamee:1990:CCR

[MO90] C. M. McNamee and R. A. Olsson. Comments on “Critical races in Ada programs”. IEEE Transactions on Software Engineering, 16(12):1439–??, December 1990. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=62452. See [KSB89].

Moore:1994:DTB

[MO94] Paul Moore and Peter G. O’Donoghue. Developing transputer-based systems using HOOD and Parallel C. Information and Software Technology, 36(6):353–360, June 1994.

CODEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0950584994900353.

Moffat:1981:EPA

[Mof81] David V. Moffat. Enumerations in Pascal, Ada, and beyond. ACM SIGPLAN Notices, 16(2):77–82, February 1981. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Molenmaker:1996:CPT

[Mol96] K. Molenmaker. Converting the part task nautical simulator to Ada 95. Lecture Notes in Computer Science, 1088:439–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moore:1995:OOF

[Moo95] David L. Moore. Object-oriented facilities in Ada 95. Dr. Dobb’s Journal of Software Tools, 20(10):28, 30, 32, 34–35, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Morris:1981:CAR

[Mor81] Alfred H. Morris, Jr. Can Ada replace FORTRAN for numerical computation? ACM SIGPLAN Notices, 16(12):10–13, December 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mossakowski:1986:BRP

[Mos86] Moss Mossakowski. Book review: Pulse operating system textbook casts doubts on ADA: D. Keeffe, G. M. Tomlinson, I. C. Wand and A. J. Wand PULSE: an ADAbased distributed operating system Academic Press, Orlando, FL, USA (1985) £15.50 pp xiii + 244. Microprocessors and Microsystems, 10(4):227–228, May 1986. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0141933186901547.

Moser:1990:DDG

[Mos90] L. E. Moser. Data dependency graphs for Ada programs. IEEE Transactions on Software Engineering, 16(5):498–509, May 1990. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=52773.

Miller:1990:FSA

[MP90] Nancy E. Miller and Charles G. Petersen. File structures with Ada. Benjamin/Cummings series in computer science. Benjamin/Cummings Pub. Co., Menlo Park, CA, USA, 1990. ISBN 0-8053-0440-1. various pp. LCCN ????

Masticola:1991:MAP

[MR91] Stephen P. Masticola and Barbara G. Ryder. A model of Ada programs for static deadlock detection in polynomial time. ACM SIGPLAN Notices, 26(12):97–107, December 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Michell:1998:GUA

[MS98] Steve Michell and Mark Saaltink. Guidance on the use of Ada95 in high integrity systems. Lecture Notes in Computer Science, 1411:276–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/1411/
14110276.htm;   http://link.springer-ny.com/link/
service/series/0558/papers/1411/14110276.pdf.

Melton:2002:ASR

[MS02a] Jim Melton and Alan R. Simon. Accessing SQL from the real world. SQL: 1999, ??(??):411–446, ???? 2002. CODEN

???? ISSN ???? URL http://www.sciencedirect.com/
science/article/pii/B9781558604568500138.

Melton:2002:C

[MS02b] Jim Melton and Alan R. Simon. Cursors, chapter 13, pages 447–467. Elsevier, Amsterdam, The Netherlands, 2002.

ISBN 1-55860-456-1. URL http://www.sciencedirect.
com/science/article/pii/B978155860456850014X.

McCormick:2011:BPE

[MSH11] John W. McCormick, Frank Singhoff, and Jérôme Hugues. Building Parallel, Embedded, and Real-Time Applications with Ada. Cambridge University Press, Cambridge, UK, 2011. ISBN 0-521-19716-3 (hardback), 0-511-97334-9 (ebook). xvii + 368 pp. LCCN QA76.73.A35 M375 2011. URL

http://assets.cambridge.org/97805211/97168/cover/
9780521197168.jpg;   http://library.books24x7.com/
library.asp?\^B&bookid=41447.

Murata:1989:DAS

[MSS89] T. Murata, B. Shenker, and S. M. Shatz. Detection of Ada static deadlocks using Petri net invariants. IEEE Transactions on Software Engineering, 15(3):314–326, March 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=21759.

Magnenat-Thalmann:1982:CIL

[MT82] Nadia Magnenat-Thalmann. Choosing an implementation language for Automatic Translation. Computer Languages, 7(3–4):161–170, ???? 1982. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0096055182900042.

Munro:1988:AP

[Mun88] Malcolm Munro. ADA for the professionals. Information and Software Technology, 30(7):452, September 1988. CO-DEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0950584988900456.

Murphy:1991:EAU

[Mur91] Michael G. Murphy. Evolution of an approach to undergraduate ADA and software engineering instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(4):55–57, December 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mitchell:1998:DAA

[MWR98] S. E. Mitchell, A. J. Wellings, and A. Romanovsky. Distributed atomic actions in Ada 95. The Computer Journal, 41(7):486–502, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/41/7/486.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/Volume_41/
Issue_07/Vol41_07.body.html#AbstractMitchell.

Mandrioli:1985:MAT

[MZGT85] Dino Mandrioli, Roberto Zicari, Carlo Ghezzi, and Francesco Tisato. Modeling the Ada task system by Petri nets. Computer Languages, 10(1):43–61, ???? 1985. CODEN CO-LADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0096055185900104.

Naiditch:1989:RAP

[Nai89] David Naiditch. Rendezvous with Ada: a programmer’s introduction. Wiley, New York, NY, USA, 1989. ISBN 0-471-61654-0. xvi + 477 pp. LCCN QA76.73.A35 N35 1989.

Narayana:1991:OTP

[Nar91] K. T. Narayana. Observing task preemption in Ada 9X. Lecture Notes in Computer Science, 571:107–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nedginn:1984:CAP

[NB84] Preet J. Nedginn and Trebor L. Bworn. CLOG: An Ada package for automatic footnote generation in UNIX (April 1984 special section). Communications of the ACM, 27(4): 351–??, 1984. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nielsen:1990:IPC

[NC90] Kjell Nielsen and Harald Carlsson. Inter-processor communication and Ada in distributed real-time systems. Computer Communications, 13(8):451–459, October 1990. CO-DEN COCOD7. ISSN 0140-3664 (print), 1873-703X (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0140366490901284.

Neelamkavil:1994:DDE

[ND94] Francis Neelamkavil and Joseph Drumgoole. Design and development of EASEL: a portable UIMS tool. Information and Software Technology, 36(12):743–749, December 1994. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0950584994900469.

Nebesh:1996:LUA

[NF96] B. Nebesh and M. Feldman. Learning to use Ada 95 components using HTML linking. In ACM [ACM96], pages 207–210. ISBN 0-89791-808-8. LCCN QA 76.73 A35 T75 1996.

Nicolescu:1980:SSC

[Nic80] Radu Nicolescu. Some short comments on the definition and the documentation of the Ada programming language. ACM SIGPLAN Notices, 15(7–8):64–71, July/August 1980. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nievergelt:1986:BRB

[Nie86] J. Nievergelt. Book review: An introduction to Ada: by S. J. Young, second (revised) edition, Ellis Horwood Ltd., Chichester, UK, 1984. Computer Compacts, 4(2– 3):70, March/June 1986. CODEN ???? ISSN ????

URL http://www.sciencedirect.com/science/article/
pii/016771368690065X.

Norris:2005:DAI

[NJ05] Sigrid Norris and Rodney H. Jones, editors. Discourse in action: introducing mediated discourse analysis. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2005. ISBN 0-415-36617-8 (paperback), 0-415-35429-3 (hardcover). xii + 229 pp. LCCN P302.84 .D574 2005. URL http://

www.loc.gov/catdir/enhancements/fy0653/2004023214-
d.html;   http://www.loc.gov/catdir/toc/ecip051/2004023214.
html.

Neelamkavil:1991:AGD

[NM91] F. Neelamkavil and O. Mullarney. Automatic graphics display management for abstract data types in Papillon prototype. Information and Software Technology, 33(4):265–272, May 1991. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-

6025 (electronic). URL http://www.sciencedirect.com/
science/article/pii/095058499190151Z.

Nishiyama:2002:SCA

[NMH^{+}02] A. Nishiyama, K. Miyoshi, T. Hikita, K. Tsukamoto, and M. Tsujigado. A study on CORBA applications for sequence control in Ada95 and Java. IECON Proceedings, 3(??):2397– 2402, 2002. CODEN ???? ISSN ????

Napier:1990:R

[NR90] John Napier and William Frank Richardson. Rabdology, volume 15 of Charles Babbage Institute reprint series for the history of computing. MIT Press, Cambridge, MA, USA, 1990. ISBN 0-262-14046-2. xxxviii + 135 pp. LCCN QA75 .N3613 1990. With an introduction by Robin E. Rider.

Nielsen:1987:CDL

[NS87a] Kjell W. Nielsen and Ken Shumate. Corrigenda: “Designing large real-time systems with Ada”. Communications of the ACM, 30(12):1073, December 1987. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
0001-0782/33454.html. See [NS87c].

Nielsen:1987:CLR

[NS87b] Kjell W. Nielsen and Ken Shumate. Corrigenda: “Designing large real-time systems with Ada”. Communications of the ACM, 30(12):1073, December 1987. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
0001-0782/33454.html. See [NS87c].

Nielsen:1987:DLR

[NS87c] Kjell W. Nielsen and Ken Shumate. Designing large realtime systems with Ada. Communications of the ACM, 30(8): 695–715, August 1987. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0001-0782/27655.html. See corri-

genda [NS87b].

Nielsen:1988:DLR

[NS88] Kjell Nielsen and Kenneth C. Shumate. Designing large realtime systems with Ada. Intertext Publications/Multiscience Press, New York, NY, 1988. ISBN 0-07-046536-3. xxix + 464 pp. LCCN QA76.73.A35 N54 1988.

Nyberg:1989:IAP

[NU89] K. Nyberg and J. Udell. IntegrAda (Ada programming support environment for the PC). Byte Magazine, 14(1):213– 214, 216, 218, 220, January 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Oest:1980:TFD

[OB80] O. N. Oest and Dines Bjørner, editors. Towards a formal description of Ada, volume 98 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. ISBN 3-540-10283-3, 0-387-10283-3. xii + 630 pp. LCCN ????

Oberndorf:1988:CAP

[Obe88] P. A. Oberndorf. The Common Ada Programming Support Environment (APSE) Interface Set (CAIS). IEEE Transactions on Software Engineering, 14(6):742–748, June 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=6154.

Oberndorf:1994:SIS

[Obe94] Patricia A. Oberndorf, editor. Proceedings of the Second International Symposium on Environments and Tools for Ada (SETA2), volume 14 (Special Issue) of ACM SIGADA Ada Letters. ACM Press, New York, NY 10036, USA, Fall 1994. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Oh:1996:GIP

[OBM96] D.-I. Oh, T. P. Baker, and S.-J Moon. The GNARL implementation of POSIX/Ada signal services. Lecture Notes in Computer Science, 1088:275–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oudshoorn:1996:BAA

[OC96] M. J. Oudshoorn and S. C. Crawley. Beyond Ada 95: The addition of persistence and its consequences. Lecture Notes in Computer Science, 1088:342–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oliveira:2008:ART

[OC08] M. V. M. Oliveira and A. L. C. Cavalcanti. ArcAngelC: a refinement tactic language for Circus. Electronic Notes in Theoretical Computer Science, 214(??):203– 229, June 28, 2008. CODEN ???? ISSN 1571-0661.

URL http://www.sciencedirect.com/science/article/
pii/S1571066108003538.

Organick:1984:TAP

[OCM^{+}84] E. I. Organick, T. M. Carter, M. P. Maloney, A. Davis, A. B. Hayes, D. Klass, G. Lindstrom, B. E. Nelson, and K. F. Smith. Transforming an Ada program unit to silicon and verifying its behavior in an Ada environment: a first experiment. IEEE Software, 1(1):31–38, 40, 42–49, January 1984. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Osterbye:1999:SAB

[ØK99] Kasper Østerbye and Wolfgang Kreutzer. Synchronization abstraction in the BETA programming language. Computer Languages, 25(3):165–187, October 1999. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0096055100000047.

Ortiz:2002:DCS

[OMÁ^{+}02] Francisco J. Ortiz, Alejandro S. Martínez, Barbara Álvarez, Andres Iborra, and José M. Fernández. Development of a control system for teleoperated robots using UML and Ada95. Lecture Notes in Computer Science, 2361:113–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2361/23610113.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2361/23610113.pdf.

Orme:1986:HSA

[Orm86] Tony Orme. How to succeed with Ada: guidelines for getting started. Data Processing, 28(5):236–240, June 1986. CODEN ???? ISSN 0011-684X (print), 1878-3058 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0011684X86902704.

Owen:1987:UAM

[Owe87] G. Scott Owen. Using Ada on microcomputers in the undergraduate curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 19(1):374–377, February 1987. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Owen:1989:ABS

[Owe89] G. Scott Owen. An Ada-based software engineering course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 21(1):213–216, February 1989. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Oliveira:2011:TLR

[OZC11] Marcel Oliveira, Frank Zeyda, and Ana Cavalcanti. A tactic language for refinement of state-rich concurrent specifications. Science of Computer Programming, 76 (9):792–833, September 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S016764231000211X.

Payne:1993:BRS

[Pay93] Alun Payne. Book review: Software design techniques for large Ada systems, W. E. Byme. Digital Press (Prentice Hall, Hemel Hempstead). ISBN 0-13-816174-7. Price £39.00 (hardback). Journal of Software Maintenance: Research and Practice, 5(1):61–62, ???? 1993. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Parrish:1996:ICI

[PCBE96] A. Parrish, D. Cordes, R. Borie, and S. Edara. Illustrating client and implementation readability tradeoffs in Ada and C++. Software—Practice and Experience, 26(7):799– 814, July 1996. CODEN SPEXBL. ISSN 0038-0644 (print),

1097-024X (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract?ID=16823.

Pollack:1982:SAMa

[PCH^{+}82a] Fred J. Pollack, George W. Cox, Dan W. Hammerstrom, Kevin C. Kahn, Konrad K. Lai, and Justin R. Rattner. Supporting Ada memory management in the iAPX-432. ACM SIGARCH Computer Architecture News, 10(2):117– 131, March 1982. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Pollack:1982:SAMb

[PCH^{+}82b] Fred J. Pollack, George W. Cox, Dan W. Hammerstrom, Kevin C. Kahn, Konrad K. Lai, and Justin R. Rattner. Supporting Ada memory management in the iAPX-432. ACM SIGPLAN Notices, 17(4):117–131, April 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pedersen:1988:BRB

[Ped88] Jan Storbank Pedersen. Book review: Ada language and methodology: By David A. Watt, Brian A. Wichmann, and William Findlay. Prentice-Hall, Englewood Cliffs, NJ, 1987, Price £16.95, ISBN 0-13-004078-9. Science of Computer Programming, 10(3):327, June 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0167642388900548.

Pyle:1980:APC

[PEGR80] I. C. Pyle, P. Elzer, S. J. Goldsack, and J. Robert. ADA in process control. Annual Review in Automatic Programming, 10(??):57–62, ???? 1980. CODEN AR-VAAM. ISSN 0066-4138 (print), 1878-545X (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0066413882900088.

Perrott:1987:PP

[Per87] R. H. Perrott. Parallel Programming. Addison-Wesley, Reading, MA, USA, 1987. ISBN 0-201-14231-7. LCCN QA76.6.P463 1987.

Pervin:1989:VAS

[Per89] William J. Pervin. Verification of Ada 1/0 statements. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 21(2):57–60, June 1989. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pfleeger:1991:MSE

[Pfl91] S. L. Pfleeger. Model of software effort and productivity. Information and Software Technology, 33(3):224–231, April 1991. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-

6025 (electronic). URL http://www.sciencedirect.com/
science/article/pii/095058499190137Z.

Pinho:2006:RST

[PH06] Luís Miguel Pinho and Michael González Harbour, editors. Reliable Software Technologies – Ada-Europe 2006: 11th Ada-Europe International Conference on Reliable Software Technologies, Porto, Portugal, June 5–9, 2006. Proceedings, volume 4006 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNCSD9. ISBN 3-540-34663-5 (print), 3-540-34664-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-540-34664-7.

Plodereder:2000:RST

[PK00] Erhard Plödereder and H. B. Keller, editors. Reliable software technologies, Ada-Europe 2000: 5th Ada-Europe International Conference, Potsdam, Germany, June 26–30, 2000: Proceedings, volume 1845 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNCSD9. ISBN 3-540-67669-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1

L43 no.1845. URL http://link.springer-ny.com/link/
service/series/0558/tocs/t1845.htm;   http://www.
springerlink.com/content/978-3-540-67669-0;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=
0302-9743&volume=1845.

Ploedereder:1992:BCA

[Plö92] Erhard Plödereder. Building consensus for Ada 9X. Communications of the ACM, 35(11):85–88, November 1992. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
0001-0782/138852.html.

Plaza:2007:EPL

[PM07] Inmaculada Plaza and Carlos Medrano. Exceptions in a programmable logic controller implementation based on ADA. Computers in industry, 58(4):347–354, May 2007. CO-DEN CINUD4. ISSN 0166-3615 (print), 1872-6194 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016636150600128X.

Porter:2001:BRM

[Por01] Roy Porter. Book review: Marina Frasca-Spada and Nick Jardine (eds.), Books and the Sciences in History. Cambridge: Cambridge University Press, 2000. Pp. xiv + 438.

ISBN 0-521-65939-6. £18.95, $29.95. British Journal for the History of Science, 34(2):233–250, June 2001. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

http://www.jstor.org/stable/4028021.

Papazoglou:1987:HMS

[PP87] Mike Papazoglou and Panayotis Pintelas. A hierarchical multimicroprocessor system for object oriented languages. Microprocessing and Microprogramming, 19(2):129–141, February 1987. CODEN MMICDT. ISSN 0165-6074 (print), 1878-

7061 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0165607487902651.

Price:1984:IA

[Pri84] David A. (David Andrew) Price. Introduction to Ada. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1984. ISBN 0-13-477653-4, 0-13-477646-1 (paperback). viii + 151 pp. LCCN QA76.73.A35 P74 1984.

Pinho:2002:TER

[PV02] Luís Miguel Pinho and Francisco Vasques. Transparent environment for replicated Ravenscar applications. Lecture Notes in Computer Science, 2361:297–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/2361/23610297.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
2361/23610297.pdf.

Panunzio:2012:ARC

[PV12] Marco Panunzio and Tullio Vardanega. Ada Ravenscar code archetypes for component-based development. Lecture Notes in Computer Science, 7308:1–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/chapter/10.1007/978-
3-642-30598-6_1/.

Purtilo:1992:FPA

[PW92] James M. Purtilo and Elizabeth L. White. A flexible program adaptation system: Case studies in Ada. The Journal of Systems and Software, 17(2):129–143, February 1992. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016412129290091W.

Pyle:1981:APL

[Pyl81] I. C. (Ian C.) Pyle. The Ada programming language: a guide for programmers. Prentice-Hall International, Upper Saddle River, NJ 07458, USA, 1981. ISBN 0-13-003921-7 (paperback). x + 293 pp. LCCN QA76.73.A35 P94.

Pyle:1985:PMA

[Pyl85] I. C. Pyle. Pascal, Modula and Ada. Computer Physics Communications, 38(2):191–197, October/November 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465585900852.

Pyle:1988:BRB

[Pyl88] Ian Pyle. Book review: Software development with Ada: Somerville, I, and Morrison, R. Addison-Wesley, Wokingham, UK £15.95 pp 364. Microprocessors and Microsystems, 12(4):236–237, May 1988. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0141933188901731.

Pyle:1986:A

[PZ86] Ian C. Pyle and Janusz. Tł. Zalewski. ADA. Biblioteka Inynierii Oprogramowania. Wydawnictwa Naukowo-Techniczne, Warszawa, Poland, 1986. ISBN 83-204-0755-9. 284 + 4 pp. LCCN ????

Radensky:1990:CAU

[Rad90] Atanas Radensky. Can Ada be used as a primary programming language?: major problems and their solutions by means of subsets. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(1):201–205, February 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Richardson:2001:LEO

[RAH^{+}01] Bruce Richardson, Anonymous, Nathan Hokanson, Ken O. Burtch, Jim V., Jerel Crosland, Paul Taylor, Sheldon Dubrowin, Paul Dale Roberts, Dean Provins, Kathy Lynn, and Andre Lessa. Letters to the editor: Offended; A real bastard; common misconception; Ada boy!; wacky names; penultimate Linux box?; SuSe too loosa; LJ interactive; sold on Soldier; groff is great; what’s up with Ogg?; changes to the Python Developer’s Handook. Linux Journal, 83:6, 141– 142, March 2001. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Raines:1992:RTA

[Rai92] Gary Raines. Real time Ada in the international space station freedom. Lecture Notes in Computer Science, 603:9–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramamritham:1987:VRC

[Ram87] Krithivasan Ramamritham. Verification of resource controller processes. Information systems, 12(1):57–67, ???? 1987. CODEN INSYD6. ISSN 0306-4379 (print), 1873-

6076 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0306437987900184.

Ramsey:1989:DFV

[Ram89] N. Ramsey. Developing formally verified Ada programs. ACM SIGSOFT Software Engineering Notes, 14(3):257–265, May 1989. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ramesh:1999:ICR

[Ram99] S. Ramesh. Implementation of communicating reactive processes. Parallel Computing, 25(6):703–727, June 1999. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167819199000137.

Rapkin:1998:AA

[Rap98] Marwin Rapkin. Ada’s attractions. Computers in Physics, 12(4):312–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822630.

Rising:1994:IHM

[RC94] Linda S. Rising and Frank W. Calliss. An information-hiding metric. The Journal of Systems and Software, 26(3):211–220, September 1994. CODEN JSSODM. ISSN 0164-1212 (print),

1873-1228 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0164121294900124.

Ruiz:2012:SCK

[RCM12] José F. Ruiz, Cyrille Comar, and Yannick Moy. Source code as the key artifact in requirement-based development: The case of Ada 2012. Lecture Notes in Computer Science, 7308: 49–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/
chapter/10.1007/978-3-642-30598-6_4/.

Real:1996:UAP

[REC96] J. Real, A. Espinosa, and A. Crespo. Using Ada 95 for prototyping real-time systems. Lecture Notes in Computer Science, 1088:262–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reed:1985:PSA

[Ree85] Joylyn Reed. A parser for a small Ada-like language — a student software engineering project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 17 (1):241–254, March 1985. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Refenes:1990:MPS

[Ref90]  Apostolos N. Refenes.  Message passing via singly-

buffered channels: an efficient & flexible communications control mechanism. Microprocessing and Microprogramming, 30(1–5):645–653, August 1990. CODEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
016560749090312W.

RTI:1989:IESa

[Rel89a] Relational Technology, Inc. INGRES/ embedded SQL companion guide for COBOL; INGRES/ embedded SQL companion guide for BASIC; INGRES/ embedded SQL companion guide for ADA; INGRES/ embedded SQL companion guide for FORTRAN; INGRES/ embedded SQL companion guide for PL/I. Relational Technology Inc., Alameda, CA, USA, 1989. 5 v. in 1 pp.

RTI:1989:IEC

[Rel89b] Relational Technology, Inc. INGRES/EQUEL companion guide for FORTRAN; INGRES/EQUEL companion guide for BASIC; INGRES/EQUEL companion guide for ADA; INGRES/EQUEL companion guide for PL/I; INGRES/ EQUEL companion guide for C; INGRES/EQUEL companion guide for COBOL; INGRES/EQUEL companion guide for PASCAL. Relational Technology Inc., Alameda, CA, USA, 1989. 7 v. in 1 pp.

Roberts:1981:TMA

[REMC81] Eric S. Roberts, Arthur Evans, Jr., C. Robert Morgan, and Edmund M. Clarke. Task management in Ada — a critical evaluation for real-time multiprocessors. Software—Practice and Experience, 11(10):1019–1051, October 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Reynolds:1985:PTM

[Rey85] Robert G. Reynolds. PARTIAL: a tool to monitor the stepwise refinement of Ada programs. ACM SIGSOFT Software Engineering Notes, 10(3):76–94, July 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reynolds:1987:MBR

[Rey87] Robert G. Reynolds. Metric-based reasoning about pseudocode design in the partial metrics system. Information and Software Technology, 29(9):497–502, November 1987. CO-DEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/095058498790005X.

Reynolds:1989:PMS

[Rey89] Robert G. Reynolds. The partial metrics system: a tool to support the metrics-driven design of pseudocode programs. The Journal of Systems and Software, 9(4):287–295, May 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-

1228 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0164121289900484.

Richard-Foy:1996:SAE

[RF96] M. Richard-Foy. Safe Ada executive: An executive for Ada safety critical applications. Lecture Notes in Computer Science, 1031:11–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ringstrom:1992:PMP

[RFF92] Johan Ringström, Peter Fritzson, and Johan Fagerström. PREDULA a multi-paradigm parallel programming and debugging environment. Microprocessing and Microprogramming, 34(1–5):45–48, February 1992. CODEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
016560749290099S.

Rivas:2001:MAK

[RH01] Mario Aldea Rivas and Michael González Harbour. MaRTE OS: An Ada kernel for real-time embedded applications. Lecture Notes in Computer Science, 2043:305–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2043/20430305.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2043/20430305.pdf.

Rivas:2002:PAI

[RH02] Mario Aldea Rivas and Michael González Harbour. A POSIX-Ada interface for application-defined scheduling. Lecture Notes in Computer Science, 2361:136–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/2361/23610136.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
2361/23610136.pdf.

Rosenberg:1980:CAC

[RLHS80] Jonathan Rosenberg, David Alex Lamb, Andy Hisgen, and Mark Sherman. The Charrette Ada compiler. ACM SIG-PLAN Notices, 15(11):72–81, November 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reynolds:1990:PMD

[RMP90] Robert G. Reynolds, Jonathan I. Maletic, and Stephen E. Porvin. PM: a metrics-driven plan compiler. Engineering Applications of Artificial Intelligence, 3(1):50–

61, March 1990. CODEN ????  ISSN ????  URL

http://www.sciencedirect.com/science/article/pii/
095219769090021D.

Rogers:1984:ALC

[Rog84] M. W. Rogers, editor. Ada: language, compilers, and bibliography. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-26464-2. v + 324 + 88 pp. LCCN QA76.73.A35 A29 1984. Published on behalf of the Commission of the European Communities.

Romanovsky:1996:ASC

[Rom96] Alexander Romanovsky. Application specific conversation schemes for ADA programs. Microprocessing and Microprogramming, 41(10):703–713, June 1996. CODEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0165607495000143.

Romanovsky:1997:PEH

[Rom97] Alexander Romanovsky. Practical exception handling and resolution in concurrent programs. Computer Languages, 23(1):43–58, April 1997. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0096055197000064.

Romanovsky:1998:SAA

[Rom98] A. Romanovsky. A study of atomic action schemes intended for standard Ada. The Journal of Systems and Software, 43(1):29–44, October 1998. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0164121298100201.

Romanovsky:1999:CDS

[Rom99] Alexander Romanovsky. Class diversity support in objectoriented languages. The Journal of Systems and Software, 48(1):43–57, August 1, 1999. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0164121299000448.

Romanovsky:2000:ECL

[Rom00] A. Romanovsky. Extending conventional languages by distributed/concurrent exception resolution. Journal of Systems Architecture, 46(1):79–95, January 1, 2000. CODEN JSARFB. ISSN 1383-7621 (print), 1873-6165 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1383762198000605.

Roos:1989:RTS

[Roo89] J. Roos. A real-time support processor for Ada tasking. ACM SIGARCH Computer Architecture News, 17(2):162– 171, April 1989. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Rosenblum:1985:MDA

[Ros85] David S. Rosenblum. Methodology for the design of Ada transformation tools in a Diana environment. IEEE Software, 2(2):24–33, March 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Rosskopf:1991:RIO

[Ros91] A. Rosskopf. Reusable input/output packages for Ada avionic applications. Lecture Notes in Computer Science, 499:338–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rosen:1992:WOS

[Ros92] J. P. Rosen. What orientation should Ada objects take? Communications of the ACM, 35(11):71–76, November 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
0001-0782/138849.html.

Rosskopf:1996:USA

[Ros96] A. Rosskopf. Use of a static analysis tool for safety-critical Ada applications: a critical assessment. Lecture Notes in Computer Science, 1088:183–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ralston:1993:ECS

[RR93] Anthony Ralston and Edwin D. Reilly, Jr., editors. Encyclopedia of Computer Science and Engineering. Van Nostrand Reinhold, New York, NY, USA, third edition, 1993. ISBN 0-442-27679-6. xxv + 1558 pp. LCCN QA76.15 .E48 1993.

Romanovsky:1997:IBC

[RRS^{+}97] A. Romanovsky, B. Randell, R. Stroud, J. Xu, and A. Zorzo. Implementation of blocking coordinated atomic actions based on forward error recovery. Journal of Systems Architecture, 43(10):687–699, ???? 1997. CODEN JSARFB. ISSN 1383-7621 (print), 1873-6165 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1383762197000295.

Rosen:2003:RST

[RS03] Jean-Pierre Rosen and Alfred Strohmeier, editors. Reliable software technologies — Ada-Europe 2003: 8th Ada-Europe International Conference on Reliable Software Technologies, Toulouse, France, June 16–20, 2003: Proceedings, volume 2655 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNCSD9. ISBN 3-540-40376-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A16 A23 2003; QA76.76.R44 A33 2003. URL http://link.springer-

ny.com/link/service/series/0558/tocs/t2655.htm;
http://uclibs.org/PID/35263;   http://www.springerlink.
com/content/978-3-540-40376-0;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=
2655.

Raduenz:1993:AAB

[RSC93] B. D. Raduenz, B. W. Suter, and E. R. Christensen. Analysis of an Ada based version of Glassman’s general Npoint Fast Fourier Transform. Computers and Mathematics with Applications, 26(2):61–65, July 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/089812219390322M.

Rosskopf:2000:AFC

[RT00] Alfred Roßkopf and Theodor Tempelmeier. Aspects of flight control software — a software engineering point of view. Control Engineering Practice, 8(6):675–680, June 2000. CO-DEN COEPEL. ISSN 0967-0661 (print), 1873-6939 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0967066100000125.

Rubine:1982:HAI

[Rub82] Dean Harris Rubine. A hybrid Ada interpreter. Thesis (M.S.), Department of Electrical Engineering and Computer Science, Massachusetts Institute of Technology, Cambridge, MA, USA, 1982. 112 pp.

Russell:1987:PLT

[Rus87] J. J. Russell. Programming languages: Time for a change? Computer Physics Communications, 45(1–3):269–273, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465587901639.

Real:2010:RST

[RV10] Jorge Real and Tullio Vardanega, editors. Reliable Software Technology — Ada-Europe 2010: 15th Ada-Europe International Conference on Reliable Software Technologies, Valencia, Spain, June 14–18, 2010. Proceedings, volume 6106 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNCSD9. ISBN 3-642-13549-8 (print), 3-642-13550-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.R44 A33 2010. URL http://hdl.

library.upenn.edu/1017.12/561474;   http://w2.vu.
edu.au/library/EBookSearch/files/Springer_E-Books_
Online.pdf;   http://www.springerlink.com/content/
978-3-642-13550-7.

Romanovsky:2011:RST

[RV11] Alexander Romanovsky and Tullio Vardanega, editors. Reliable Software Technologies — Ada-Europe 2011: 16th Ada-Europe International Conference on Reliable Software Technologies, Edinburgh, UK, June 20–24, 2011. Proceedings, volume 6652 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNCSD9. ISBN 3-642-21337-5 (paperback), 3-642-21338-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

????  URL http://www.springerlink.com/content/
978-3-642-21338-0;   http://www.springerlink.com/
openurl.asp?genre=book&isbn=978-3-642-21337-3.

Rogers:2000:SRA

[RW00] P. Rogers and A. J. Wellings. State restoration in Ada 95: a portable approach to supporting software fault tolerance. The Journal of Systems and Software, 50(3): 237–255, March 15, 2000. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/29/11/49/28/30/article.
pdf;   http://www.elsevier.nl/gej-ng/10/29/11/49/
28/abstract.html;   http://www.sciencedirect.com/
science/article/pii/S0164121299001004.

Radensky:1988:EAF

[RZP^{+}88] Atanas Radensky, Emilia Zivkova, Valia Petrova, Rumiana Lesseva, and Christina Zascheva. Experience with Ada as a first programming language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 20(4): 58–61, December 1988. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stanley:1985:APG

[S^{+}85] James Stanley et al. Ada, a programmer’s guide with microcomputer examples. Addison-Wesley, Reading, MA, USA, 1985. ISBN 0-201-16416-7. x + 358 pp. LCCN QA76.73.A35 S83 1985.

Sage:1987:WSI

[Sag87] Mike Sage. Workstation supports integration of Ada on embedded processors. Microprocessors and Microsystems, 11(5):277–279, June 1987. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0141933187902717.

Saib:1985:AI

[Sai85] Sabina Saib. Ada: an introduction. Holt, Rinehart and Winston, New York, NY, USA, 1985. ISBN 0-03-059487-1. 358 pp. LCCN ????

Sammet:1981:OHL

[Sam81] Jean E. Sammet. An overview of high-level languages. Advances in Computers, 20(??):199–259, ???? 1981. CODEN

???? ISSN ???? URL http://www.sciencedirect.com/
science/article/pii/S0065245808604982.

Sammet:1986:WAJ

[Sam86] Jean E. Sammet. Why Ada is not just another programming language. Communications of the ACM, 29(8):722–732, August 1986. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.org/pubs/

toc/Abstracts/0001-0782/6425.html.

Sandmayr:1981:CLC

[San81] H. Sandmayr. A comparison of languages: CORAL, PASCAL, PEARL, ADA and ESL. Computers in industry, 2(2):123–132, June 1981. CODEN CINUD4. ISSN 0166-3615 (print), 1872-6194 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0166361581900051.

Sankar:1989:AST

[San89a] Sriram Sankar. APE — a set of TEX macros to format Ada programs. TUGboat, 10(1):89–97, April 1989. ISSN 0896-3207.

Sankar:1989:NDA

[San89b] Sriram Sankar. A note on the detection of an Ada compiler bug while debugging an Anna program. ACM SIGPLAN Notices, 24(6):23–31, June 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sanden:1994:SSC

[San94] Bo Sandén. Software systems construction with examples in Ada. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-030834-X. xvi + 443 pp. LCCN QA76.76.D47 S26 1994.

Sanden:1995:DCS

[San95] Bo I. Sanden. Designing control systems with entity-life modeling. The Journal of Systems and Software, 28(3):225–237, March 1995. CODEN JSSODM. ISSN 0164-1212 (print),

1873-1228 (electronic). URL http://www.sciencedirect.
com/science/article/pii/016412129400058U.

Savoysky:1980:ADA

[Sav80] S. Savoysky. Analysis and description of automatic control system. Annual Review in Automatic Programming, 10(??):45–55, ???? 1980. CODEN ARVAAM.

ISSN 0066-4138 (print), 1878-545X (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0066413882900076.

Savoysky:1981:UAS

[Sav81] S. Savoysky. The use of Ada for the specification of automata in civil engineering. Annual Review in Automatic Programming, 11(??):129–138, ???? 1981. CODEN AR-VAAM. ISSN 0066-4138 (print), 1878-545X (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0066413881900197.

Schneeweiss:1996:TSM

[SAV96] H. Schneeweiss, V. Amiot, and O. Vix. Translating Shlaer/ Mellor object-oriented analysis models into Ada95. Lecture Notes in Computer Science, 1031:249–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stansifer:1994:MCP

[SBM94] Ryan Stansifer, Mike Beaven, and Dan C. Marinescu. Modeling concurrent programs with colored Petri nets. The Journal of Systems and Software, 26(2):129–148, August 1994. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0164121294900841.

Stratford-Collins:1982:APC

[SC82] M. J. (Michael J.) Stratford-Collins. ADA, a programmer’s conversion course, volume 20 of Ellis Horwood series in computers and their applications. E. Horwood, Chichester, West Sussex, UK, 1982. ISBN 0-85312-250-4, 0-85312-444-2, 0-470-27332-1. 170 pp. LCCN QA76.73.A35 S85 1982.

Shatz:1988:PNF

[SC88] S. M. Shatz and W. K. Cheng. A Petri net framework for automated static analysis of Ada tasking behavior. The Journal of Systems and Software, 8(5):343–359, December 1988. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0164121288900271.

Shen:1994:ACP

[SC94] Jun Shen and Gordon V. Cormack. Access control for private declarations in Ada. Computer Languages, 20(2):117–126, May 1994. CODEN COLADA. ISSN 0096-0551 (print), 1873-

6742 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0096055194900183.

Sutton:1997:AHI

[SC97] J. M. Sutton and B. A. Carré. Achieving high integrity at low cost: a constructive approach. Microprocessors and Microsystems, 20(8):455–461, April 1997. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0141933196011106.

Scarlato:1991:DAS

[Sca91] Philip P. Scarlato. Developing Ada software without target hardware. ACM SIGSOFT Software Engineering Notes, 16 (1):36–40, January 1991. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Scandura:1994:CLC

[Sca94] Joseph M. Scandura. Converting legacy code into Ada: a cognitive approach. Computer, 27(4):55–61, April 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Scheer:1982:AFA

[Sch82] Linda Sue Scheer. Ada, FORTRAN, ALGOL, JOVIAL, Pascal, PL/I, and LISP compared to Ada design requirements. Thesis (M.S.), Wright State University, Dayton, OH, USA, 1982. x + 121 pp.

Schrijver:1985:PDM

[Sch85] H. Schrijver. Profiling and debugging in modern programming languages. Computer Physics Communications, 38(2):289–293, October/November 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/001046558590092X.

Schefstrom:1986:RCT

[Sch86a] Dick Schefstrom. Revision control tools and the Ada program library. In USENIX Association [USE86b], pages 241–251.

Schiper:1986:PCI

[Sch86b] A. Schiper. Programmation concurrente: Illustré par des exemples en Portal, Modula-2 et Ada. Collection informatique. Pr. Polytechniques Romandes, Lausanne, Switzerland, 1986. ISBN 2-88074-093-2. 295 pp. LCCN

???? URL http://www.zentralblatt-math.org/zmath/
en/search/?an=0659.68032.

Schoitsch:1986:SEA

[Sch86c] Erwin Schoitsch. Software engineering aspects of realtime programming concepts. Computer Physics Communications, 41(2–3):327–361, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010465586900743.

Schoitsch:1988:SSS

[Sch88] Erwin Schoitsch. Software-safety and software quality assurance in real-time applications: Part 2: Real-time structures and languages. Computer Physics Communications, 50(1–2):189–211, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0010465588901257.

Scheiber:1999:NST

[Sch99] E. Scheiber. The numerical solution of Theodorsen integral equation. Computers and Mathematics with Applications, 38(9–10):221–231, November 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0898122199002771.

Stift:1998:SPA

[SD98] Martin J. Stift and Paul F. Dubois. Scientific programming with Ada 95: Object-oriented, parallel, and safe. Computers in Physics, 12(2):150–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168624.

Secord:1988:BRD

[Sec88] J. A. Secord. Book review: Dorothy Stein. Ada: A Life and a Legacy. Cambridge, Mass.: The MIT Press, 1985. Pp. xix + 321. ISBN 0-262-19242-X. £17.50 (hardcover), £8.95 (paperback). British Journal for the History of Science, 21(3):375– 377, September 1988. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.

org/stable/4026847.

Seidewitz:1989:GOO

[Sei89] Ed Seidewitz. General object-oriented software development: Background and experience. The Journal of Systems and Software, 9(2):95–108, February 1989. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0164121289900137.

Sennett:1992:DCA

[Sen92] C. T. Sennett. Demonstrating the compliance of Ada programs with Z specifications. In Cliff B. Jones, Roger C. Shaw, and Tim Denvir, editors, 5th Refinement Workshop: Proceedings of the 5th Refinement Workshop, organized by BCS-FACS, London, 8–10 January 1992, pages 70–87. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., January 1992. ISBN 3-540-19752-4. LCCN QA76.76.D47 R44 199.

Stenning:1981:AEP

[SFGT81] V. Stenning, R. Froggatt, R. Gilbert, and E. Thomas. The Ada environment: a perspective. Computer, 14(6):26–36, June 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Soupos:1991:UAI

[SG91] P. Soupos and S. Goutas. Using Ada to implement the transaction mechanism of a distributed object-oriented DBMS. Lecture Notes in Computer Science, 499:252–??, 1991. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schiper:1989:CPI

[SH89] André Schiper and J. (Jack) Howlett. Concurrent programming: illustrated with examples in Portal, Modula-2 and Ada. North Oxford Academic, London, UK, 1989. ISBN 0-946536-39-2. 234 pp. LCCN QA76.6. Translation to English by Jack Howlett of [Sch86b].

Shatz:1988:TCM

[Sha88] S. M. Shatz. Towards complexity metrics for Ada tasking. IEEE Transactions on Software Engineering, 14(8): 1122–1127, August 1988. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?arnumber=7623.

Sherman:1980:ACG

[SHLR80] Mark Sherman, Andy Hisgen, David Alex Lamb, and Jonathan Rosenberg. An Ada code generator for VAX 11/780 with Unix. ACM SIGPLAN Notices, 15(11):91–100, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shumate:1988:UCA

[Shu88] Kenneth C. Shumate. Understanding concurrency in Ada. Intertext Publications, New York, NY, USA, 1988. ISBN 0-07-057299-2. xii + 595 pp. LCCN QA76.73.A35 S483 1988.

Shultz:1989:DEO

[Shu89a] Roger K. Shultz. A decentralized embedded operating system supporting distributed execution of Ada tasks. In USENIX Association [USE89], pages 391–409. ISBN ???? LCCN ????

Shumate:1989:UAA

[Shu89b] Kenneth C. Shumate. Understanding ADA: with abstract data types. Wiley, New York, NY, USA, second edition, 1989. ISBN 0-471-60520-4. various pp. LCCN QA76.73.A15.

Silberschatz:1981:SMA

[Sil81] Abraham Silberschatz. On the synchronization mechanism of the Ada language. ACM SIGPLAN Notices, 16(2):96–103, February 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Silver:1991:UAS

[Sil91] James L. Silver. Using Ada to specify and evaluate projects in a data structures course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):337– 340, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Silberg:1992:CRV

[Sil92a] Steen Silberg. Considerations with regard to validation of Ada debuggers. Lecture Notes in Computer Science, 603: 197–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Silberg:1992:IIV

[Sil92b] Steen Silberg. Intel i860 versus digital signal processors (DSP). Microprocessing and Microprogramming, 35 (1–5):605–610, September 1992. CODEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
016560749290375H.

Skansholm:1988:AB

[Ska88] Jan Skansholm. Ada from the beginning. International computer science series. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-17522-3. xiii + 617 pp. LCCN QA76.73.A15.

Skansholm:1994:AB

[Ska94a] Jan Skansholm. Ada from the beginning. Addison-Wesley, Reading, MA, USA, second edition, 1994. ISBN 0-201-62448-6. xiii + 648 pp. LCCN QA76.73.A35 S4913 1994. Translation from Swedish to English by Shirley Booth.

Skazinski:1994:PAR

[Ska94b] Joseph G. Skazinski. Porting Ada: a report from the field. Computer, 27(10):58–64, October 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Skansholm:1995:AFB

[Ska95] Jan Skansholm. Ada från början: med Ada 95 standard. Studentlitteratur, Lund, Sweden, 3. upplagan edition, 1995. ISBN 91-44-25233-1. LCCN ????

Skansholm:1997:AB

[Ska97] Jan Skansholm. Ada 95: from the beginning. International computer science series. Addison-Wesley, Reading, MA, USA, third edition, 1997. ISBN 0-201-40376-5. xiii + 656 pp. LCCN QA76.73 .A28513 1994; QA76.73 .A28513 1997. Translation from Swedish to English by Shirley Booth.

Skansholm:2002:AFB

[Ska02] Jan Skansholm. Ada från början [Ljudupptagning]: med Ada 95 standard. (Swedish) [Ada from the beginning [sound recording]: with the Ada 95 standard]. TPB, Enskede, The Netherlands, 2002. 1 CD–R (34h 8m) pp. LCCN

????  URL http://katalog.tpb.se/wsHitList.Asp?
SCode1=TN&SearchStr1=C21792.

Skelly:1982:APS

[Ske82] Patrick G. Skelly. The ACM position on standardization of the Ada language. Communications of the ACM, 25(2):118– 120, 1982. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schwarz:1988:OAD

[SKL88] B. Schwarz, W. Kirchgässner, and R. Landwehr. An optimizer for Ada — design, experiences and results. ACM SIG-PLAN Notices, 23(7):175–184, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org:80/pubs/citations/
proceedings/pldi/53990/p175-schwarz/.

Spooner:1986:FSC

[SKW^{+}86] David L. Spooner, Arthur M. Keller, Gio Wiederhold, John Salasin, and Deborah Heystek. Framework for the security component of an Ada DBMS. In Kambayashi et al. [KCGO86], pages 347–354. ISBN 0-934613-18-4. LCCN QA

76.9 D3 I61 1986. URL http://www.vldb.org/dblp/db/
conf/vldb/SpoonerKWSH86.html.

Steigerwald:1991:CTR

[SLM91] R. Steigerwald, Luqi, and J. McDowell. CASE tool for reusable software component storage and retrieval in rapid prototyping. Information and Software Technology, 33(9):698–706, November 1991. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
095058499190043B.

Stansifer:1991:PNM

[SM91] Ryan Stansifer and Dan Marinescu. Petri net models of concurrent Ada programs. Microelectronics and Reliability, 31(4):577–594, ???? 1991. CODEN MCR-LAS. ISSN 0026-2714 (print), 1872-941X (electronic).

URL http://www.sciencedirect.com/science/article/
pii/002627149190004Q.

Smart:1996:DDD

[Sma96] J. D. Smart. A decade of development and deployment of distributed Ada systems. Lecture Notes in Computer Science, 1088:485–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smedema:1983:PLP

[SMB83] C. H. Smedema, P. (Piet) Medema, and M. (Maarten) Boasson. The programming languages: Pascal, Modula, CHILL, and Ada. Prentice/Hall International, Englewood Cliffs, NJ, USA, 1983. ISBN 0-13-729756-4 (paperback). 154 pp. LCCN QA76.7 .S6 1983. US$19.95.

Shatz:1990:DIP

[SMBT90] Sol M. Shatz, Khanh Mai, Christopher Black, and Shengru Tu. Design and implementation of a Petri net based toolkit for Ada tasking analysis. IEEE Transactions on Parallel and Distributed Systems, 1(4):424–441, October 1990. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sommerville:1995:PGA

[SMD95] I. Sommerville, L. Masera, and C. Demaria. Practical guidelines for Ada reuse in an industrial environment. Lecture Notes in Computer Science, 926:138–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smedema:1985:SII

[Sme85] C. H. Smedema. Some issues in the international standardization of CHILL and Ada. Computers and Standards, 4(2):95– 100, ???? 1985. CODEN COSTEZ. ISSN 0167-8051 (print),

1878-3090 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0167805185900221.

Smyth:1997:GGN

[Smy97] Gavin Smyth. GNAT: The GNU New York University Ada Translator: a high-quality, low-cost Ada 95 compiler. Dr. Dobb’s Journal of Software Tools, 22(12):86, 88, 89–90, 105– 107, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Staff:1985:PAA

[Sof85] Software Engineering Notes Staff. Proceedings of the ACM Ada TEC “Future Ada environment workshop”. ACM SIG-SOFT Software Engineering Notes, 10(2):77–117, April 1985. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sahraoui:1992:CSP

[SOK92] A. E. K. Sahraoui and N. Ould-Kaddour. Control software prototyping. Computers in industry, 20(3):327–334, October 1992. CODEN CINUD4. ISSN 0166-3615 (print), 1872-

6194 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016636159290081W.

Sommerville:1989:IPS

[Som89] Ian Sommerville. Integrated project support environments. Microprocessors and Microsystems, 13(4):254–262, May 1989. CODEN MIMID5. ISSN 0141-9331 (print), 1872-

9436 (electronic). URL http://www.sciencedirect.com/
science/article/pii/014193318990063X.

Souter:1990:PMA

[Sou90] John Souter. The position of MODULA-2 among programming languages. Microprocessors and Microsystems, 14(3):145–148, April 1990. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0141933190900643.

Scheer:1985:UAR

[SR85a] M. D. Scheer and S. Rajeev. A UNIX-based Ada runtime system. In USENIX Association [USE85b], pages 51–?? Abstract only.

Scheer:1985:UBA

[SR85b] M. D. Scheer and S. Rajeev. A UNIX-based Ada runtime system. In USENIX [USE85a], pages 51–?? Abstract only.

Srinivasan:1994:CLS

[Sri94] S. Srinivasan. A critical look at some Ada features. ACM SIGPLAN Notices, 29(3):18–22, March 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sridhar:2007:SDB

[Sri07] Nigamanth Sridhar. Serfs: Dynamically-bound parameterized components. The Journal of Systems and Software, 80(5):736–749, May 2007. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0164121206002056.

Sharma:2022:ECP

[SS22] Durgesh M. Sharma and Shishir Kumar Shandilya. An efficient cyber-physical system using hybridized enhanced support-vector machine with Ada-Boost classification algorithm. Concurrency and Computation: Practice and Experience, 34(21):e7134:1–e7134:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Standish:1984:APA

[ST84] Thomas A. Standish and Richard N. Taylor. Arcturus: a prototype advanced Ada programming environment. ACM SIGPLAN Notices, 19(5):57–64, May 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sebesta:1986:FIA

[ST86] Robert W. Sebesta and Mark A. Taylor. Fast identification of Ada and Modula-2 reserved words. Journal of Pascal, Ada and Modula-2, 5(2):36–39, March/April 1986. CODEN JPAME8, JOPAD5. ISSN 0747-1351, 0735-1232.

Shimojima:1987:VRT

[ST87] Takehiko Shimojima and Masanori Teramoto. V60 realtime operating system. Microprocessing and Microprogramming, 21(1–5):197–204, August 1987. CODEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
016560748790038X.

Stevenson:1980:ATA

[Ste80] D. R. Stevenson. Algorithms for translating Ada multitasking. ACM SIGPLAN Notices, 15(11):166–175, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stein:20xx:ALL

[Stexx] D. Stein. Ada: a Life and a Legacy. Wiley-IEEE Press, ????, 20xx. ISBN 0-262-25703-3. LCCN ???? URL http:

//0-ieeexplore.ieee.org.library.vu.edu.au/servlet/
opac?bknumber=6267341.

Stiff:1998:APS

[Sti98a] M. J. Stiff. (astro)physical supercomputing: Ada95 as a safe, object oriented alternative. Lecture Notes in Computer Science, 1411:128–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stift:1998:APS

[Sti98b] Martin J. Stift. (astro)physical supercomputing: Ada95 as a safe, object oriented alternative. Lecture Notes in Computer Science, 1411:128–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/1411/
14110128.htm;   http://link.springer-ny.com/link/
service/series/0558/papers/1411/14110128.pdf.

Shatz:1996:APN

[STMD96] Sol M. Shatz, Shengru Tu, Tadao Murata, and Sastry Duri. Application of Petri net reduction for Ada tasking deadlock analysis. IEEE Transactions on Parallel and Distributed Systems, 7(12):1307–1322, December 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Strohmeier:1996:RST

[Str96] Alfred Strohmeier, editor. Reliable software technologies, Ada-Europe ’96: 1996 Ada-Europe International Conference on Reliable Software Technologies, Montreux, Switzerland, June 10–14, 1996: proceedings, volume 1088 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNCSD9. ISBN 3-540-61317-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A16 A23 1996. URL http://link.springer-

ny.com/link/service/series/0558/tocs/t1088.htm;
http://uclibs.org/PID/108181;   http://www.springerlink.
com/content/978-3-540-61317-6;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=

Samuels:1998:LSA

[SvA^{+}98] Adam D. Samuels, Jerry van Dijk, Dawn Amore, Shlomi Fish, Scott Schwendinger, Arvid R. Hand, Jr., and Howard Mark. Letters: Something in the air; more on Ada; recycling PC’s; server-side scripting; stronger encryption; inner loops; Einstein kudos. Dr. Dobb’s Journal of Software Tools, 23(3): 8, 12, March 1998. CODEN DDJOEB. ISSN 1044-789X.

Slape:1983:CFA

[SW83] John K. Slape and Peter J. L. Wallis. Conversion of Fortran to Ada using an intermediate tree representation. The Computer Journal, 26(4):344–353, November 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/26/4/344.full.pdf+html.

Staalhane:1994:QRC

[SW94] Tor Stålhane and Kari J. Wedde. The quest for reliability: a case study. The Journal of Systems and Software, 26(1):69– 76, July 1994. CODEN JSSODM. ISSN 0164-1212 (print),

1873-1228 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0164121294900973.

Sward:2011:PAA

[Swa11] Ricky E. Sward, editor. Proceedings of the 2011 ACM annual international conference: Special interest group on the Ada programming language. ACM Press, New York, NY 10036, USA, 2011. ISBN 1-4503-1028-1. LCCN ???? URL http:/

/dl.acm.org/citation.cfm?id=2070337.

Taft:2000:CAR

[T^{+}00] S. Tucker Taft et al. Consolidated Ada reference manual: language and standard libraries: international standard ISO/IEC 8652/1995(E) with technical corrigendum 1, volume 2219 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-43038-5 (paperback). xix + 560 pp. LCCN QA76.73.A16 C65 2000.

Taft:1982:OBV

[Taf82] S. Tucker Taft. An object-based virtual operating system for the Ada programming support environment. Operating Systems Review, 16(1):14–25, January 1982. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tafvelin:1987:ACL

[Taf87] S. (Sven) Tafvelin, editor. Ada components: libraries and tools: proceedings of the Ada-Europe International Conference, Stockholm, 26–28 May 1987, The Ada companion series. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-34636-3. LCCN QA/76/.73/A35/A322/1987 IN; QA76.73.A35 A24 1987.

Taft:1996:PIAa

[Taf96a] S. T. Taft. Programming the Internet in Ada 95. Lecture Notes in Computer Science, 1088:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taft:1996:PIAb

[Taf96b] Tucker Taft. Programming the Internet in Ada 95. In Strohmeier [Str96], pages 1–16. CODEN LNCSD9. ISBN 3-540-61317-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A16 A23 1996. URL http://

link.springer-ny.com/link/service/series/0558/tocs/
t1088.htm;   http://uclibs.org/PID/108181;   http:
//www.springerlink.com/content/978-3-540-61317-6;
 http://www.springerlink.com/openurl.asp?genre=issue&
issn=0302-9743&volume=1088. Describes the Intermetrics

Ada 95 compiler, which generates Java Virtual Machine bytecodes directly.

Tang:1990:PGE

[Tan90] Ping Tak Peter Tang. A portable generic elementary function package in Ada and an accurate test suite. Technical report ANL-90/35, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, November 1990. iii

+ 35 pp. URL http://www.osti.gov/bridge/servlets/
purl/6310184-4n5sOR/6310184.PDF.

Tang:2004:AHR

[TC04] Xueyan Tang and Samuel T. Chanson. Adaptive hash routing for a cluster of client-side web proxies. Journal of Parallel and Distributed Computing, 64(10):1168–1184, October 2004. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731504000863.

Tai:1991:DCA

[TCO91] K. C. Tai, R. H. Carver, and E. E. Obaid. Debugging concurrent Ada programs by deterministic execution. IEEE Transactions on Software Engineering, 17(1):45–63, January 1991. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=67578.

Taft:1995:ARM

[TD95] S. Tucker Taft and Robert A. (Robert Anderson) Duff. Ada 95 reference manual: language and standard libraries, volume 1246 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-63144-5. xviii + 526 pp. LCCN QA76.73.A35 A22 1995. URL http:/

/link.springer.com/book/10.1007/BFb0034910/page/1;
 http://uclibs.org/PID/126664.

Taft:1997:ARM

[TD97] S. Tucker Taft and Robert A. (Robert Anderson) Duff. Ada 95 reference manual: language and standard libraries, volume 1246 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNCSD9. ISBN 3-540-63144-5. ISSN 0302-9743 (print), 1611-3349 (electronic). xviii + 526 pp. LCCN QA76.73.A35

A22 1997. URL http://link.springer-ny.com/link/
service/series/0558/tocs/t1246.htm;   http://www.
springerlink.com/content/978-3-540-63144-6;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=

0302-9743&volume=1246. International standard ISO/IEC 8652:1995(E).

Thomas:1992:EAS

[TDB92] W. M. Thomas, A. Delis, and V. R. Basili. An evaluation of Ada source code reuse. Lecture Notes in Computer Science, 603:80–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taft:2006:ARM

[TDB^{+}06] S. Tucker Taft, Robert A. Duff, Randall L. Brukardt, Erhard Plödereder, and Pascal Leroy, editors. Ada 2005 Reference Manual. Language and Standard Libraries: International Standard ISO/IEC 8652/1995 (E) with Technical Corrigendum 1 and Amendment 1, volume 4348 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNCSD9. ISBN 3-540-69335-1 (print), 3-540-69336-X (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). 618

(est.) pp. LCCN ???? URL http://www.springerlink.
com/content/978-3-540-69336-9.

Taft:2001:CAR

[TDBP01] S. Tucker Taft, R. A. Duff, R. L. Brukardt, and E. Plödereder, editors. Consolidated Ada reference manual: language and standard libraries: international standard ISO/IEC 8652/ 1995(E) with technical corrigendum 1, volume 2219 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CO-DEN LNCSD9. ISBN 3-540-43038-5. ISSN 0302-9743 (print), 1611-3349 (electronic). xix + 560 pp. LCCN QA76.73.A16

C65 2000. URL http://link.springer-ny.com/link/
service/series/0558/tocs/t2219.htm;   http://www.
springerlink.com/content/978-3-540-43038-4;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=
0302-9743&volume=2219.

Tam:1987:TAU

[TE87] Wing C. Tam and Michael A. Erlinger. On the teaching of Ada in an undergraduate computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 19(1):58–61, February 1987. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Teller:1984:PTJ

[Tel84] J. Teller, editor. Proceedings of the Third Joint Ada Europe/Ada TEC Conference: Brussels, 26–28 June 1984: industrial implications of Ada and Ada programming support environment, The Ada companion series. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-30102-5. LCCN ????

Temte:1986:ACU

[Tem86] Mark Temte. An Ada course for upper-level undergraduates. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 18(4):41–45, December 1986. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tempelmeier:1994:NHS

[Tem94] T. Tempelmeier. A note on hardware-software codesign. Annual Review in Automatic Programming, 18(??):121–126, ???? 1994. CODEN ARVAAM. ISSN 0066-4138 (print),

1878-545X (electronic). URL http://www.sciencedirect.
com/science/article/pii/0066413894900213.

Testi:1981:IRP

[Tes81] Maria Rossella Testi. Una interfaccia relazionale per il linguaggio ADA. (Italian) [A relational interface for the Ada language]. Technical report, ????, ????, 1981. ?? pp.

Texel:1982:ADA

[Tex82] Putnam P. Texel. Ada education : = Design concepts “+” Ada constructs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 14(1):201–204, February 1982. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 13th SIGCSE Symposium on Computer Science Education.

Tai:1980:CSI

[TG80] Kuo-Chung Tai and Ken Garrard. Comments on the suggested implementation of tasking facilities in the ’Rationale for the design of the Ada programming language’. ACM SIG-PLAN Notices, 15(10):76–84, October 1980. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Theron:1990:DCL

[Thé90] Jean-Louis Théron. Design and checking of a large ADA real-time system. Nuclear instruments and methods in physics research. Section A, Accelerators, spectrometers, detectors and associated equipment, 293(1–2):373–376, August 1, 1990. CODEN NIMAER. ISSN 0168-9002, 0167-5087.

URL http://www.sciencedirect.com/science/article/
pii/016890029091463L.

Tricas:1998:DCS

[TM98] Fernando Tricas and Javier Martínez. Distributed control systems simulation using high level Petri nets. Mathematics and Computers in Simulation, 46(1):47–55, April 1, 1998. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0378475497001572.

Thanh-Nu:1992:CIS

[TN92] Do Thanh-Nu. Compilation integration: a solution for the challenge of developing and reusing Ada software on different platforms. Lecture Notes in Computer Science, 603:230–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tolmach:1998:MAS

[TO98] Andrew Tolmach and Dino P. Oliva. From ML to Ada: Strongly-typed language interoperability via source translation. Journal of Functional Programming, 8(4):367–412, July 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-

7653 (electronic). URL https://www.cambridge.org/core/
product/9ACB018A7F71ECF8A0FEE0DFEDAC4E84.

Tokar:2001:NDA

[Tok01] Joyce L. Tokar. New developments in Ada 95 run-time profile definitions and language refinements. Lecture Notes in Computer Science, 2043:160–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/series/0558/bibs/
2043/20430160.htm;   http://link.springer-ny.com/
link/service/series/0558/papers/2043/20430160.pdf.

Tomayko:1989:LLT

[Tom89] James E. Tomayko. Lessons learned teaching Ada in the context of software engineering. The Journal of Systems and Software, 10(4):281–283, November 1989. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0164121289900757.

Tonndorf:1998:TYT

[Ton98] Michael Tonndorf. Ten years of tool based Ada compiler validations. an experience report. Lecture Notes in Computer Science, 1411:176–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/1411/
14110176.htm;   http://link.springer-ny.com/link/
service/series/0558/papers/1411/14110176.pdf.

Touati:1987:AOO

[Tou87] H. Touati. Is Ada an object oriented programming language? ACM SIGPLAN Notices, 22(5):23–26, May 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Toussaint:1994:AEF

[Tou94] M. Toussaint, editor. Ada in Europe: First International Eurospace-Ada-Europe Symposium, Copenhagen, Denmark, September 26–30, 1994: proceedings, volume 887 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNCSD9. ISBN 3-540-58822-1 (Berlin: paperback), 0-387-58822-1 (New York: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A35

I57 1994. DM60.00. URL http://link.springer-ny.com/
link/service/series/0558/tocs/t0887.htm;   http:/
/www.springerlink.com/content/978-3-540-58822-1;
http://www.springerlink.com/openurl.asp?genre=issue&
issn=0302-9743&volume=887.

Toussaint:1996:AES

[Tou96] Marcel Toussaint, editor. Ada in Europe: Second International Eurospace-Ada-Europe Symposium, Frankfurt/Main, Germany, October 2–6, 1995: proceedings, volume 1031 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNCSD9. ISBN 3-540-60757-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A35 I57 1995. URL http://link.springer-

ny.com/link/service/series/0558/tocs/t1031.htm;
http://www.springerlink.com/content/978-3-540-60757-
1;   http://www.springerlink.com/openurl.asp?genre=
issue&issn=0302-9743&volume=1031.

Taylor:1985:SAA

[TS85] R. N. Taylor and T. A. Standish. Steps to an Advanced Ada^{1} Programming Environment. IEEE Transactions on Software Engineering, SE-11(3):302–310, March 1985. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=1702006.

Tugwell:1983:BRB

[Tug83] Paul Tugwell. Book review: Programming embedded systems with Ada: V. A. Downes and S. J. Goldsack. Advances in Water Resources, 6(4):250, December 1983. CO-DEN AWREDI. ISSN 0309-1708 (print), 1872-9657 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0309170883900659.

Tugwell:1984:BRB

[Tug84] P. J. Tugwell. Book review: Programming embedded systems with Ada: Downes & Goldsack, Prentice-Hall International, 1984, pp., 00.00 / £00.00, ISBN 0-13-730010-7. Advances in Engineering Software (1978), 6(3):175, July 1984. CO-DEN AESODT. ISSN 0965-9978 (print), 0141-1195 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0141119584900470.

Twigg:1983:AGS

[Twi83] Tom Twigg. ADA gets support from UK. Microprocessors and Microsystems, 7(6):291, July/August 1983. CO-DEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0141933183904714.

USDOD:1982:MRL

[U. 82] U. S. Department of Defense. Manuel de référence du langage de programmation ADA. (French) [Ada Language Programming Reference Manual]. Eyrolles, Paris, France, 1982. xiii + 227 pp. LCCN QA/76/ .73/A35/E8312 IN; QA76.73.A35. Translation to French by Anne Kruchten and Philippe Kruchten.

USNRC:1997:ABS

[U. 97] U. S. National Research Council Committee on the Past and Present Contexts for the Use of Ada in the Department of Defense. Ada and beyond: software policies for the Department of Defense. National Academy Press, Washington, DC, USA, 1997. ISBN 0-585-00251-7 (e-book), 0-309-05597-0 (paperback). xi + 102 pp. LCCN QA76.73.A16 N38

1997. URL http://www.nap.edu/catalog.php?record_
id=5463;   http://www.nap.edu/openbook.php?record_
id=5463.

Uhl:1982:AGS

[U^{+}82] J. (Jürgen) Uhl et al. An Attribute Grammar for the Semantic Analysis of ADA, volume 139 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CODEN LNCSD9. ISBN 0-387-11571-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 511 pp. LCCN QA76.73.A35 A87 1982. URL http://link.springer-

ny.com/link/service/series/0558/tocs/t0139.htm;
http://www.springerlink.com/content/978-0-387-11571-
9;   http://www.springerlink.com/openurl.asp?genre=
issue&issn=0302-9743&volume=139.

US-DOD:1983:ANS

[UA83a] United States. Dept. of Defense and American National Standards Institute. American National Standard reference manual for the Ada programming language: ANSI/MIL-STD-1815A-1983. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, February 17, 1983. various pp.

USDOD:1983:PLA

[UA83b] United States. Dept. of Defense and American National Standards Institute. The Programming language Ada reference manual, volume 155 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNCSD9. ISBN 0-387-12328-8, 3-540-12328-8. ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 330 pp. LCCN QA76.73.A35

P72 1983. URL http://link.springer-ny.com/link/
service/series/0558/tocs/t0155.htm;   http://www.
springerlink.com/content/978-0-387-12328-8;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=

0302-9743&volume=155. “ANSI/MIL-STD-1815A-1983.” Ada was designed in accordance with requirements defined by the United States Dept. of Defense. “Approved 17 February 1983.”.

US-DOD:1983:RMA

[UA83c] United States. Dept. of Defense and American National Standards Institute. Reference manual for the Ada programming language: ANSI/MIL-STD-1815A-1983. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CODEN LNCSD9. ISBN 0-387-90887-0. ISSN 0302-9743 (print), 1611-3349 (electronic). 330 pp. LCCN QA267.A1 L43 no.155. Approved February 17, 1983.

USDOD:1981:PLA

[Uni81] United States. Dept. of Defense. The Programming Language Ada Reference Manual: Proposed Standard Document, volume 106 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. CODEN LNCSD9. ISBN 0-387-10693-6, 3-540-10693-6. ISSN 0302-9743 (print), 1611-3349 (electronic). x + 243 pp. LCCN QA76.73.A35

P763 1981. URL http://link.springer-ny.com/link/
service/series/0558/tocs/t0106.htm;   http://www.
springerlink.com/content/978-0-387-10693-9;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=

0302-9743&volume=106. A photographic reproduction of the official November 1980 printing (Honeywell, Minneapolis).

US-DOD-AJPO:1983:ANS

[Uni83] United States. Dept. of Defense. Ada Joint Program Office. American National Standard Reference Manual for the Ada Programming Language: ANSI/MIL-STD-1815A. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1983. various pp.

USNBS:1985:A

[Uni85] United States National Bureau of Standards. Ada, volume 119 of Federal information processing standards publication, FIPS PUB. U.S. Department of Commerce/National Bureau of Standards, Gaithersburg, MD, USA, 1985. 3 pp. LCCN ????

USENIX:1985:UCPb

[USE85a] USENIX, editor. USENIX Conference Proceedings, Winter, 1985. Dallas, TX. USENIX, Berkeley, CA, USA, Winter

USENIX:1985:PUA

[USE85b] USENIX Association, editor. Proceedings: USENIX Association Winter Conference, January 23–25, 1985, Dallas, Texas, USA. USENIX, P.O. Box 7, El Cerrito 94530, CA, USA, 1985.

USENIX:1986:SCP

[USE86a] USENIX Association, editor. Summer conference proceedings, Atlanta 1986: June 9–13, 1986, Atlanta, Georgia, USA. USENIX, P.O. Box 7, El Cerrito 94530, CA, USA, 1986.

USENIX:1986:UAW

[USE86b] USENIX Association, editor. USENIX Association Winter Conference proceedings: January 15–17, 1986, Denver, Colorado USA. USENIX, P.O. Box 7, El Cerrito 94530, CA, USA, 1986.

USENIX:1987:ECP

[USE87] USENIX Association, editor. EUUG Conference Proceedings, Autumn, 1987. Dublin, Ireland. EUUG, Buntingford, Herts, UK, Autumn 1987. ISBN ???? LCCN ????

USENIX:1989:DMS

[USE89] USENIX Association, editor. Distributed and Multiprocessor Systems Workshop Proceedings, October 5–6, 1989. Fort Lauderdale, FL. USENIX, Berkeley, CA, USA, October 5–6, 1989. ISBN ???? LCCN ????

Vajda:1986:CSP

[Vaj86] Ferenc Vajda. Concurrent systems, programming primitives and languages: a comparative study. Microprocessing and Microprogramming, 18(1–5):185–194, December 1986. CO-DEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165607486900438.

vandenBos:1980:CAP

[vdB80] Jan van den Bos. Comments on ADA process communication. ACM SIGPLAN Notices, 15(6):77–81, June 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanderLinden:1984:BRA

[vdL84] Peter van der Linden. Book review: ADA for experts: A. Nico Habermann and Dewayne E. Perry, ADA for experienced programmers, Addison-Wesley, Reading, MA, USA £16.10 pp 479. Microprocessors and Microsystems, 8(5):252, June 1984. CODEN MIMID5. ISSN 0141-9331 (print), 1872-

9436 (electronic). URL http://www.sciencedirect.com/
science/article/pii/014193318490440X.

vanderLinden:1981:MFA

[vdLN81] Peter van der Linden and Leonard K. Nicholson. Macro facilities in the Ada environment. ACM SIGPLAN Notices, 16 (8):67–68, August 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vardanega:2001:ACR

[VGdlP01] Tullio Vardanega, Rodrigo García, and Juan Antonio de la Puente. An application case for Ravenscar technology: Porting OBOSS to GNAT/ORK. Lecture Notes in Computer Science, 2043:392–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2043/
20430392.htm;   http://link.springer-ny.com/link/
service/series/0558/papers/2043/20430392.pdf.

Vignes:1993:SAR

[Vig93] J. Vignes. A stochastic arithmetic for reliable scientific computation. Mathematics and Computers in Simulation, 35(3):233–261, September 1993. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/037847549390003D.

Katwijk:1987:ACD

[vJK87] J. van (Jan) Katwijk. The Ada-compiler: on the design and implementation of an Ada language compiler. Thesis (doctoral), Technische Universiteit Delft, Delft, The Netherlands, 1987. 233 pp. Abstract in English and Dutch.

Vines:1988:GOO

[VK88] D. Vines and T. King. Gaia: an object-oriented framework for an Ada environment. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):206, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Katwijk:1992:AMT

[vK92] J. (Jan) van Katwijk, editor. Ada: moving towards 2000: 11th Ada-Europe International Conference, Zandvoort, The Netherlands, June 1–5, 1992: proceedings, volume 603 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. CODEN LNCSD9. ISBN 3-540-55585-4 (Berlin), 0-387-55585-4 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A35

A24 1992. URL http://link.springer-ny.com/link/
service/series/0558/tocs/t0603.htm;   http://www.
springerlink.com/content/978-3-540-55585-8;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=
0302-9743&volume=603.

Volz:1991:DAC

[VKT91] R. A. Volz, P. Krishnan, and R. Theriault. Distributed Ada: case study. Information and Software Technology, 33(4):292– 300, May 1991. CODEN ISOTE7. ISSN 0950-5849 (print),

1873-6025 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0950584991901544.

Volz:1987:TID

[VM87] R. A. Volz and T. N. Mudge. Timing issues in the distributed execution of Ada programs. IEEE Transactions on Computers, C-36(4):449–459, April 1987. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1676927.

vonMayrhauser:1993:IPS

[vMAW93] A. von Mayrhauser, K. Archie, and N. Weber. Incremental parsing for software maintenance tools. The Journal of Systems and Software, 23(3):235–243, December 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/016412129390098I.

Volz:1989:TED

[VMBK89] R. A. Volz, T. N. Mudge, G. D. Buzzard, and P. Krishnan. Translation and execution of distributed Ada programs: Is it still Ada? IEEE Transactions on Software Engineering, 15 (3):281–292, March 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?arnumber=21756.

vanRumste:1983:ING

[vR83] M. van Rumste. The iAPX432, a next generation microprocessor. Microprocessing and Microprogramming, 11(2): 69–106, February 1983. CODEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0165607483900753.

vanKatwijk:1984:DMR

[vv84] J. van Katwijk and J. van Someren. The doublet model: run-time model and implementation of Ada types. ACM SIGPLAN Notices, 19(1):78–92, January 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vardanega:2005:RST

[VW05] Tullio Vardanega and Andy Wellings, editors. Reliable Software Technology — Ada-Europe 2005: 10th Ada-Europe International Conference on Reliable Software Technologies, York, UK, June 20–24, 2005. Proceedings, volume 3555 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNCSD9. ISBN 3-540-26286-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.R44

A33 2005. URL http://www.springerlink.com/content/
978-3-540-26286-2;   http://www.springerlink.com/
openurl.asp?genre=issue&issn=0302-9743&volume=3555.

Ward:2002:LIC

[WA02] Michael Ward and Neil C. Audsley. Language issues of compiling Ada to hardware. Lecture Notes in Computer Science, 2361:88–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer-ny.

com/link/service/series/0558/bibs/2361/23610088.htm;
 http://link.springer-ny.com/link/service/series/
0558/papers/2361/23610088.pdf.

Wallis:1983:BRB

[Wal83] Peter Wallis. Book review: Problem solving with ADA, Brian Mayoh, Wiley, Chichester, 1982. no. of pages: viii + 233. Price: £10.75. Software—Practice and Experience, 13(2): 198, February 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wallis:1984:BRB

[Wal84a] Peter Wallis. Book review: Software Engineering with Ada, Grady Booch, Benjamin/Cummings, 1983. No. of pages: xix + 502. Price: £14.20. Software—Practice and Experience, 14(6):603, June 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wallis:1984:AST

[Wal84b] Peter J. L. Wallis, editor. Ada Software Tools Interfaces Workshop, Bath, July 13–15, 1983: proceedings, volume 180 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. CODEN LNCSD9. ISBN 0-387-13878-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A35 A34 1983. URL http://link.springer-

ny.com/link/service/series/0558/tocs/t0180.htm;
http://www.springerlink.com/content/978-0-387-13878-
7;   http://www.springerlink.com/openurl.asp?genre=
issue&issn=0302-9743&volume=180.

Wallich:1985:SAI

[Wal85] P. Wallich. Software: Artificial-intelligence software becomes available for minis and mainframes, and the U.S. Department of Defense converts to Ada and pushes to produce secure systems. IEEE Spectrum, 22(1):50–52, January 1985. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Walters:1991:RSA

[Wal91] Neal Walters. Requirements specification for Ada software under DoD-STD-2167A. The Journal of Systems and Software, 15(2):173–183, May 1991. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016412129190054A.

Warn:1986:LVA

[War86] Keith Warn. Lisp vs. Ada: Implications in diagnostics oriented expert systems. In IEEE [IEE86a], pages 409–415. CO-DEN AUPRDX. ISBN ???? ISSN 0734-7510. LCCN TK 7895 A8 A98 1986.

Waterman:1997:TTA

[Wat97] S. R. Waterman. Techniques for testing Ada 95. Lecture Notes in Computer Science, 1251:278–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wellings:1996:PRS

[WB96] A. J. Wellings and A. Burns. Programming replicated systems in Ada 95. The Computer Journal, 39(5):361–373, ???? 1996. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/39/5/361.full.pdf+html;   http://www.
oup.co.uk/jnls/list/comjnl/hdb/Volume_39/Issue_05/
390361.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_05/Vol39_05.body.html#AbstractWellings.

Wellings:1997:IAA

[WB97] A. Wellings and A. Burns. Implementing atomic actions in Ada 95. IEEE Transactions on Software Engineering, 23(2): 107–123, February 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?arnumber=585500.

Wellings:1997:TTA

[WBP97] A. J. Wellings, A. Burns, and O. Pazy. Task termination in Ada 95. Lecture Notes in Computer Science, 1251:149–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Williams:1985:MSL

[WCK85] Michael R. (Michael Roy) Williams and Martin Campbell-Kelly, editors. The Moore School lectures: theory and techniques for design of electronic digital computers, volume 9 of Charles Babbage Institute reprint series for the history of computing. MIT Press, Cambridge, MA, USA, 1985. ISBN 0-262-03109-4, 0-262-03109-3. LCCN TK7888.3. 49.95.

Wang:1996:ACC

[WCW96] Zhenyu Wang, Li Chen, and Zhijiang Wang. Ada concurrent complexity metrics based on rendezvous relations. Progr. Natur. Sci. (English Ed.), 6(3):359–368, 1996. ISSN 1002-0071.

Wearing:1992:SEA

[Wea92] Alison Wearing. Software engineering, Ada and metrics. Lecture Notes in Computer Science, 603:35–??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wegner:1979:PAI

[Weg79] P. Wegner. Programming with ADA: an introduction by means of graduated examples. ACM SIGPLAN Notices, 14 (12):1–46, December 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegner:1980:ALE

[Weg80a] Peter Wegner. The Ada language and environment. ACM SIGSOFT Software Engineering Notes, 5(2):8–14, April 1980. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wegner:1980:PAI

[Weg80b] Peter Wegner. Programming with Ada: An Introduction by Means of Graduated Examples. Prentice-Hall software series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1980. ISBN 0-13-730697-0. xi + 211 pp. LCCN QA76.73.A35 W46 1980.

Wegner:1990:TSC

[Weg90] Eberhard Wegner. Testing of software for conformity with standards. In J. L. Berg and H. Schumny, editors, An Analysis of the Information Technology Standardization Process, pages 289–297. Elsevier, Amsterdam, The Netherlands, 1990.

ISBN 0-444-87390-2. URL http://www.sciencedirect.
com/science/article/pii/B9780444873903500414.

Weiss:2003:DSA

[Wei03] Mark Allen Weiss. Data structures and algorithm analysis in Ada [sound recording]. TPB, Enskede, The Netherlands, 2003. 1 CD–R (47h 24m) pp. LCCN ???? URL http://

katalog.tpb.se/wsHitList.Asp?SCode1=TN&SearchStr1=
C22496.

Wetherell:1981:PAR

[Wet81] C. S. Wetherell. Problems with the Ada reference grammar. ACM SIGPLAN Notices, 16(9):90–104, September 1981. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See note [Mac83].

Wegener:1997:SUT

[WF97] J. Wegener and I. Fey. Systematic unit-testing of Ada programs. Lecture Notes in Computer Science, 1251:64–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wehrum:1986:SKF

[WHD86] R. P. Wehrum, W. Hoyer, and G. Dießl. On some key features of Ada: Language and programming environment. Computer Physics Communications, 41(2–3):271–283, August 1986. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465586900706.

Wheeler:1981:ESD

[Whe81] Thomas J. Wheeler. Embedded system design with Ada as the system design language. The Journal of Systems and Software, 2(1):11–21, February 1981. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/016412128190042X.

Whitaker:1981:CPA

[Whi81a] William A. Whitaker. Comments on portions of the ACM SIGPLAN conference on the ADA programming language not available in the proceedings. ACM SIGPLAN Notices, 16 (2):110–112, February 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Whitaker:1981:SAI

[Whi81b] William A. Whitaker. Summary of the ADA implementor’s meeting: December 1980. ACM SIGPLAN Notices, 16(2):

104–109, February 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

White:1989:CSP

[Whi89] Bebo White. The comparison and selection of programming languages for high energy physics applications. Computer Physics Communications, 57(1–3):538–542, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046558990283X.

Wichman:1984:BRB

[Wic84a] B. A. Wichman. Book review: Life cycle support in the Ada environment: J. McDermid and K. Ripken Cambridge University Press, Ada Companion Series, 1984, 247 pp., £12.50. Applied mathematical modelling, 8(5):370–371, October 1984. CODEN AMMODL. ISSN 0307-904X (print),

1872-8480 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0307904X84901525.

Wichmann:1984:BRA

[Wic84b] B. A. Wichmann. Book review: Ada in perspective: The Ada programming language: a tutorial S. H. Saib and R. E. Fritz, IEEE Computer Society (1983) 150 pp £15. Computer Aided Design, 16(3):178–179, May 1984. CODEN CAIDA5. ISSN 0010-4485 (print), 1879-2685 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0010448584900538.

Wichmann:1984:ATB

[Wic84c] Brian A. Wichmann. Is Ada too big? A designer answers the critics. Communications of the ACM, 27(2):98–103, February 1984. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wichmann:1988:BRB

[Wic88] Brian Wichmann. Book review: Good programming practice in Ada: Luker, P. A., Blackwell Scientific, Oxford, UK (1987) £14.95 (soft cover), pp 320. Microprocessors and Microsystems, 12(7):413, September 1988. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0141933188901500.

Williams:1987:ITI

[Wil87] David Thomas Williams. In-line translation of input/output statements of Fortran to ADA. Thesis (M.S.), Dept. of Computer Science, Wichita State University, Wichita, KS, USA, 1987. 43 pp.

Williams:2006:LRD

[Wil06a] Rob Williams. Languages for RTS development — C, Ada and Java. Real-Time Systems Development, ??(??): 341–357, ???? 2006. CODEN ???? ISSN ????

URL http://www.sciencedirect.com/science/article/
pii/B9780750664714500189.

Williams:2006:RTS

[Wil06b] Rob Williams. Real-time systems development. Elsevier Butterworth-Heinemann, Oxford, UK, 2006. ISBN 0-7506-6471-1 (paperback), 0-08-045640-5. URL http://site.

ebrary.com/lib/royallibrary/Doc?id=10138749.

Wimlett:1983:BRBa

[Wim83a] Chris Wimlett. Book review: Programming embedded systems with Ada: V. A. Dowries and S. J. Goldsack. Published by Prentice/Hall International. $10.95. pp. Data Processing, 25(5):42, June 1983. CODEN ???? ISSN 0011-684X (print),

1878-3058 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0011684X8390014X.

Wimlett:1983:BRBb

[Wim83b] Chris Wimlett. Book review: The Programming Languages Pascal, Modula, Chill and Ada: C. H. Smedema, P. Medema and M. Boasson. Published by Prentice/Hall International. £9.95. 154 pp. Data Processing, 25(9):42, November 1983. CODEN ???? ISSN 0011-684X (print), 1878-3058 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0011684X83901478.

Winter:1999:OSA

[Win99] Don C. Winter. Open systems Ada technology demonstration program. Microprocessors and Microsystems, 22(8):433–438, February 22, 1999. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0141933198001021.

Withrow:1990:EDS

[Wit90] Carol Withrow. Error density and size in Ada software. IEEE Software, 7(1):26–30, January 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wellings:2000:IOO

[WJS^{+}00] A. J. Wellings, B. Johnson, B. Sanden, J. Kienzle, T. Wolf, and S. Michell. Integrating object-oriented programming and protected objects in Ada 95. ACM Transactions on Programming Languages and Systems, 22(3): 506–539, 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://www.acm.

org/pubs/articles/journals/toplas/2000-22-3/p506-
wellings/p506-wellings.pdf;   http://www.acm.org/
pubs/citations/journals/toplas/2000-22-3/p506-wellings/

.

Weiss:1989:CQD

[WMS^{+}89] Eric A. Weiss, Rex Malik, H. E. Salzer, The Editor, and Herbert R. J. Grosch. Comments, queries, and debate: The Real Time Club; New York Mathematical Tables Project; early small computers; Ada’s First Stirring. Annals of the History of Computing, 11(1):51–54, January/March 1989. CODEN AHCOE5.

ISSN 0164-1239. URL http://dlib.computer.org/an/
books/an1989/pdf/a1051.pdf;   http://www.computer.
org/annals/an1989/a1051abs.htm.

Whittington:1997:LEC

[WN97] Paul Whittington and Nancy Nicolaisen. Letter to the editor: Could Ada95 outmuscle Java? Object Magazine, 6(12):6, 8, February 1997. CODEN OBMAFO. ISSN 1055-3614.

Wolcott:1991:APW

[Wol91] Peter Wolcott. Ada: progress in the West and the East. Mathematics and Computers in Simulation, 33(2):173–174, August 1991. CODEN MCSIDR. ISSN 0378-4754 (print),

1872-7166 (electronic). URL http://www.sciencedirect.
com/science/article/pii/037847549190162V.

Wolf:2008:BRC

[Wol08] Markus Wolf. Book review: Concurrent and Real-time Programming in Ada (3rd edition), Alan Burns, Andy Wellings, Cambridge University Press, 2007, $75.00, ISBN: 978-0-521-86697-2. ACM Queue: Tomorrow’s Computing Today, 6 (6):54, October 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Woodside:1989:TCB

[Woo89] C. Murray Woodside. Throughput calculation for basic stochastic rendezvous networks. Performance evaluation, 9(2):143–160, April 1989. CODEN PEEVD9. ISSN 0166-5316 (print), 1872-745X (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0166531689900394.

Wotawa:2000:DVD

[Wot00] F. Wotawa. Debugging VHDL designs using model-based reasoning. Artificial Intelligence in Engineering, 14(4): 331–351, October 2000. CODEN ???? ISSN ????

URL http://www.sciencedirect.com/science/article/
pii/S0954181000000212.

Wallis:1980:EIA

[WS80] Peter J. L. Wallis and Bernard W. Silverman. Efficient implementation of the Ada overloading rules. Information Processing Letters, 10(3):120–123, April 18, 1980. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0020019080900605.

Wiener:1983:PA

[WS83] Richard Wiener and Richard Sincovec. Programming in ADA. Wiley, New York, NY, USA, 1983. ISBN 0-471-87089-7. xvii + 345 pp. LCCN QA76.73.A35 W53 1983; QA/76/ .73/A35/W54/1983 IN.

Wiener:1984:SEM

[WS84] Richard S. Wiener and Richard Sincovec. Software engineering with Modula-2 and Ada. Wiley, New York, NY, USA, 1984. ISBN 0-471-89014-6. xix + 451 pp. LCCN ????

Waroquiers:2001:MLA

[WVC^{+}01] Philippe Waroquiers, Stef Van Vlierberghe, Dirk Craeynest, Andrew Hately, and Erik Duvinage. Migrating large applications from Ada83 to Ada95. Lecture Notes in Computer Science, 2043:380–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2043/
20430380.htm;   http://link.springer-ny.com/link/
service/series/0558/papers/2043/20430380.pdf.

Wallis:1984:RAA

[WW84] Peter J. L. Wallis and Brian A. Wichmann. Requirements analysis for Ada compilers. Communications of the ACM, 27 (1):37–41, 1984. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Watt:1987:ALM

[WWF87] David A. (David Anthony) Watt, Brian A. Wichmann, and William Findlay. Ada Language and Methodology. Prentice-Hall International series in computer science. Prentice-Hall International, Upper Saddle River, NJ 07458, USA, 1987. ISBN 0-13-004078-9 (paperback), 0-13-004086-X (paperback with disk). xvii + 518 pp. LCCN QA76.73.A35 W38 1987. US$28.67.

Wallach:1988:ULA

[WY88] Y. Wallach and E. Yaprak. Use of local-area networks as alternating sequential-parallel systems. Microprocessing and Microprogramming, 24(1–5):95–102, August 1988. CO-DEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0165607488900361.

Xu:1998:CSS

[Xu98] Baowen Xu. Comments on several syntax rules in Ada95. ACM SIGPLAN Notices, 33(2):65–67, February 1998. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yeung:1997:SBS

[Yeu97] W. L. Yeung. Semantics-based support tools for high integrity Ada software. Lecture Notes in Computer Science, 1251:100–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Young:1993:CTL

[YLT93] M. Young, D. L. Levine, and R. N. Taylor. Comments on “Temporal logic-based deadlock analysis for Ada” by G. M. Karam and R. J. A. Burh. IEEE Transactions on Software Engineering, 19(2):198–199, February 1993. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=214836. See [KB91].

Young:1982:BRBa

[You82a] S. J. Young. Book review: The ADA programming language: Pyle, I. C., Prentice-Hall International (1981) pp 293, £8.95. Microprocessors and Microsystems, 6(4):199, May 1982. CO-DEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0141933182905.

Young:1982:BRBb

[You82b] S. J. Young. Book review: The programming language ADA Reference Manual: Springer-Verlag (1981) pp 243, $7.90, DM 16.50. Microprocessors and Microsystems, 6(4):199, May 1982. CODEN MIMID5. ISSN 0141-9331 (print), 1872-

9436 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0141933182903787.

Young:1983:IA

[You83] Stephen J. Young. An Introduction to Ada. Ellis Horwood Series in Computers and their Applications, Editor: Brian Meek. Ellis Horwood, New York, NY, USA, 1983. ISBN 0-85312-535-X. 400 pp. LCCN QA76.73.A35 Y68 1983. UK£25.00.

Yeung:1990:IJD

[YT90] W. L. Yeung and G. Topping. Implementing JSD designs in Ada: a tutorial. ACM SIGSOFT Software Engineering Notes, 15(3):25–32, July 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Young:1995:CAT

[YTL^{+}95] Michal Young, Richard N. Taylor, David L. Levine, Kari A. Nies, and Debra Brodbeck. A concurrency analysis tool suite for Ada programs: rationale, design, and preliminary experience. ACM Transactions on Software Engineering and Methodology, 4(1):65–106, January 1995. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392

(electronic). URL http://www.acm.org/pubs/articles/
journals/tosem/1995-4-1/p65-young/p65-young.pdf;
http://www.acm.org/pubs/citations/journals/tosem/
1995-4-1/p65-young/.

Yang:2021:MTA

[YYB^{+}21] Zhibin Yang, Shenghao Yuan, Jean-Paul Bodeveix, Mamoun Filali, Tiexin Wang, and Yong Zhou. Multi-task Ada code generation from synchronous dataflow programs on multicore: Approach and industrial study. Science of Computer Programming, 207(??):??, July 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S016764232100037X.

Zamorano:1997:BSC

[ZAdlP97] J. Zamorano, A. Alonso, and J. A. de la Puente. Building safety-critical real-time systems with reusable cyclic executives. Control Engineering Practice, 5(7):999–1005, July 1997. CODEN COEPEL. ISSN 0967-0661 (print), 1873-

6939 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0967066197000889.

Zalewski:1988:STR

[Zal88] Janusz Zalewski. A step towards real time application of Ada. Computer Physics Communications, 50(1–2):273–279, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465588901373.

Zalewski:1992:RAD

[Zal92] Janusz Zalewski. Review of Ada in Distributed Real-Time Systems (Nielsen, K; 1990). Computer, 25(1):135, January 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zenil:2013:CUU

[Zen13] Hector Zenil, editor. A computable universe: understanding and exploring nature as computation. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2013. ISBN 981-4374-29-6. xliv + 810 pp. LCCN

QA267.7 .C676 2013. URL http://www.worldscientific.
com/worldscibooks/10.1142/8306. Foreword by Roger

Penrose.

Zipser:2007:CPM

[ZGMK07] S. Zipser, A. Gommlich, J. Matthes, and H. B. Keller. Combustion plant monitoring and control using infrared and video cameras. In Power Plants and Power Systems Control 2006: A Proceedings Volume from the IFAC Symposium on Power Plants and Power Systems Control, Kananaskis, Canada, 2006, pages 249–254. Elsevier, Amsterdam, The Netherlands, 2007. ISBN 0-08-046620-6.

URL http://www.sciencedirect.com/science/article/
pii/B9780080466200500426.

Zhu:1996:HPB

[ZLZ^{+}96] Jiang Zhu, T. G. Lewis, Lihua Zhao, Weldon Jackson, and Russel L. Wilson. HaRTS: Performance-based design of distributed hard real-time software. The Journal of Systems and Software, 32(2):143–156, February 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0164121294000662.

Zeigler:1991:RVB

[ZRC91] Bernard P. Zeigler, Jerzy W. Rozenblit, and Eric R. Christensen. Reducing the validation bottleneck with a knowledgebased, distributed simulation environment. Expert Systems with Applications, 3(3):329–342, ???? 1991. CODEN

ESAPEH. ISSN ???? URL http://www.sciencedirect.
com/science/article/pii/095741749190056K.

Zamorano:2001:IAR

[ZRdlP01] Juan Zamorano, José F. Ruiz, and Juan Antonio de la Puente. Implementing Ada.Real Time.Clock and absolute delays in real-time kernels. Lecture Notes in Computer Science, 2043:317–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2043/
20430317.htm;   http://link.springer-ny.com/link/
service/series/0558/papers/2043/20430317.pdf.

Zemanek:1986:RSA

[ZT86] Heinz Zemanek and James E. Tomayko. Reviews: Stan Augarten, Bit by Bit; Dorothy Stein, Ada, A Life and a Legacy; capsule reviews. Annals of the History of Computing, 8(4): 380, 382–383, October/December 1986. CODEN AHCOE5.

ISSN 0164-1239. URL http://dlib.computer.org/an/
books/an1986/pdf/a4380.pdf;   http://www.computer.
org/annals/an1986/a4380abs.htm.